A Complete Bibliography of Publications in the Proceedings of the American Philosophical

Society (1900–1949)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 August 2019 Version 1.00

Title word cross-reference

−183^{∘} [Bee39]. _{1} [Atk18c]. _{2} [Atk18c]. Σ [Doo03].

1189 [Dav39b]. 146 [SO37]. 1777 [Chi48]. 1867 [Col41a]. 1900 [Ano00e, Ano00j, Ano00i, Ano00k, Ano00m, Ano00l, Ano00o, Ano00n, Ano00h, Ano00q, Ano00p, Ano00s, Ano00r, Ano00u, Ano00t, Ano00w, Ano00v, Low01]. 1901 [Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01p, Ano01o, Ano01q, Ano01r, Ano01t, Ano01s, Ano01e]. 1902 [Ano02e, Ano02f, Ano02h, Ano02g, Ano02j, Ano02i, Ano02l, Ano02k, Ano02m, Ano02o, Ano02n, Ano02p, Ano02q]. 1903 [Ano03e, Ano03f, Ano03h, Ano03g, Ano03j, Ano03i, Ano03l, Ano03k, Ano03n, Ano03m, Ano03o, Ano03p, Ano03q, Ano03s, Ano03r]. 1904 [Ano04e, Ano04f, Ano04h, Ano04g, Ano04j, Ano04i, Ano04k, Ano04l, Ano04n, Ano04m, Ano04p, Ano04o, Ano04q, Ano04r]. 1905 [Ano05f, Ano05h, Ano05g, Ano05i, Ano05k, Ano05j, Ano05l, Ano05n, Ano05m, Ano05p, Ano05o]. 1906 [Ano06e, Ano06f, Ano06g, Ano06h, Ano06j, Ano06i, Ano06k, Ano06m, Ano06l, Ano06o, Ano06n, Ano06q, Ano06p, Ogb07]. 1907 [Ano07f, Ano07g, Ano07i, Ano07h, Ano07j, Ano07l, Ano07k, Ano07m, Ano07o, Ano07n, Ano07p, Ano07q]. 1908 [Ano08e, Ano08g, Ano08f, Ano08h, Ano08j, Ano08i, Ano08k, Ano08m, Ano08l, Ano08n]. 1918 [And19, Cam19, Lam19, Mil19a]. 1931 [Lin33]. 1937 [Con37a]. 1940 [Wie45]. 1941 [Bar45, Wad45b]. 1942 [Ang42].

2 [Kra47]. 200-Inch [Tho30].

61 [Str43].

A.D. [Dav39b]. A.M [Pet03]. Abbe [Mar17]. Abbey [Cro49]. Abbott [Cro43]. Abdominal [Str36]. Abilities [Tho41]. Abnormal [Fen48]. Aboard [Bau15]. Aboriginal [Dav41b, Mat01a, Mat03b]. Aborigines [Mat03a, Dav49b, Mat00a, Mat00c, Mat00d, Mat05a, Mat09a, Mat09c]. Absolute [Bur21, Gap38, Gap40]. Absorption [Her40, JA08, JS09, LL40, Mac25]. Abstract [Mil10b, Rus10]. Abstracting [Sch39]. Abstracts [Ano32a, Ano33a, Ano34a, Ano35a, Ano36a, Ste45a]. Abundance [Ble37]. Abuses [Joh39]. Academies [Lel37]. Academy [DM47, Jew41, Sin21a, Sin21b, Sin21c, Reh40]. Acceleration [Bru28]. Accessibility [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano06a, Ano06b, Ano06c, Ano07a, Ano07b, Ano07c, Ano08a, Ano08b, Ano08c, Ano09a, Ano09b, Ano09c, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano20a, Ano20b]. Accessibility [Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano25a, Ano25b, Ano25c, Ano26a, Ano26b, Ano26c, Ano26d, Ano26e, Ano26f, Ano27a, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano29b, Ano29c, Ano29d, Ano30a, Ano31a, Ano31b, Ano31c, Ano31d, Ano31e, Ano32b, Ano32c, Ano32d, Ano32e, Ano32f, Ano32g, Ano32h, Ano33b, Ano33c, Ano33d, Ano33e, Ano33f, Ano34b, Ano34c, Ano34d, Ano34e, Ano34f, Ano34g, Ano34h, Ano34i, Ano34j, Ano34k, Ano35b, Ano35c, Ano35d, Ano35e, Ano35f, Ano35g, Ano35h, Ano35i, Ano35j, Ano36b, Ano36c, Ano36d, Ano36e, Ano36f, Ano36g, Ano37a, Ano37b, Ano37c, Ano37d, Ano37e, Ano37f, Ano38a, Ano38b, Ano38c, Ano38d]. Accessibility [Ano38e, Ano38f, Ano39a, Ano39b, Ano39c, Ano39d, Ano39e, Ano39f, Ano39g, Ano39h, Ano39i, Ano40a, Ano40b, Ano40c, Ano40d, Ano40e, Ano40f, Ano40g, Ano40h, Ano40i, Ano40j, Ano41a, Ano41b, Ano41c, Ano41d, Ano41e, Ano41f, Ano41g, Ano42a, Ano42b, Ano42c, Ano42d, Ano42e, Ano43a, Ano43b, Ano43c, Ano43d, Ano43e, Ano44a, Ano44b, Ano44c, Ano44d, Ano44e, Ano44f, Ano44g, Ano44h, Ano45a, Ano45b, Ano45c, Ano45d, Ano46a, Ano46b, Ano46c, Ano46d, Ano46e, Ano47a, Ano47b, Ano47c, Ano47d, Ano47e, Ano48a, Ano48b, Ano48c, Ano48d, Ano48e, Ano48f, Ano49a, Ano49b, Ano49c, Ano49d, Ano49e, Ano49f, Ano49g]. Acclimatization [Fra45]. Accomplished [Ber45]. Accomplishments [Abb27, Shi45]. Account [Bar17, Bee39, MP44, Wil02, Bal02]. Accused [Kee34]. Achievement [McE38]. Achievements [Lin33]. Acid [KM04, KS06, Kra06, McC04, Mor26]. Acids [Ros47a]. Acoustic [KS27]. Acquired [Car49, Sto23]. Acquisition [Hay24b]. across [Hal41]. Act [Hue34]. Action [Dal25, Hei34, Hes26, Jen08, Kra05, Mac21a, Mac22, Mac24, Mas07, See15]. Activated [RE40]. Activating [Par34]. Activation [Par41, PS40]. Active [Dua22]. Activities [Der27, Hun32, McC45, Mye43, Sto46]. Activity [Bau10, BJ19, HM38, SM39, Wei41]. Acts [SO37]. Actually [Sto03]. Aculeata [Whe19]. Adaptation [Par15]. Adaptations [Bat49]. Added [Wil08b]. Addition [Ken29]. Additional [Tre28b]. Additions [Sin15]. Address [Bur23, Der27, Fle30, KW26, Smi03]. Addresses [Ano27b, Ano30b, Ano30c]. Adirondacks [Aga23]. Adjourned [Ano00e]. Adjustment [Bar09c, Kop43a, Pat31, Tol41]. Adjustments [Hob23, Wil46]. Administration [Bro16, Pea16, Poo33, Woo30]. Admiral [Haw45, Sau40]. Adrenal [Can11]. Adriatic [Joh20]. Adsorption [Bee39]. Adult [Heb42, Hrd36, Wis27]. Advance [Kae39]. Advancement [Con44]. Advances [Dye44, Osb04, Rit43, Whi44]. Advantages [Du 31]. Adventures [Blo13]. Aërology [Bla17]. Aëronautics [Bau17, Dur17]. Aëroplane [Wil17]. Aeshna [Lin40]. Æshnidæ [Cal34]. Æsthetic [Grè34a]. Æsthetics [Edg36]. Affairs [Bin48, Coh48, Cus48, Lin27, Poo33]. Affect [Blo44a]. Affecting [Bur12, Wet34]. Affinities [Ber14, Ort05, Pat35]. Afghanistan [Jac29]. Africa [Bar29]. After [KW44, TP08, Joh32a, New44, Rol30]. After-Effects [KW44]. After-Images [TP08]. against [Hau11, Mor35, Rav03]. Agallian [Bla44]. Agarics [Atk18b]. Agaricus [Atk15]. Agave [Tre10]. Age [Bro22, Cha22, Chi16, Cla22, Hay00a, Kra46, OC31, Pat31, WBAW39, Wis27]. Agencies [Fal46]. Agency [Hue34, Rod15]. Agent [Hau03, Mül36]. Agents [Par34, Par40]. Aggregates [Ric11]. Agricultural [Dou01]. Agriculture [Bai47, Lan06b, Tae39, Wil42b, dV06]. Aid [Hay20a, ZF47]. Aided [Sin21a, Sin21b, Sin21c]. Aims [Hun44, Sta48]. Air [Bar07b, Bar11c, Bid28, Bla09, Hob15, KS27, Mun10, Ros28, Squ17, Whi11, Whi20]. Aircraft [McC45]. al [Far39, Far39]. al-Dīn [Far39]. al-Ghazzāli [Far39]. Alaska [Har28, Hei47, Tit08]. Albert [RW07]. Albino [Han23]. Alcohol [Han23, JA08, WS07]. Alcoholic [AB02]. Alexander [Hir09, Lin33, All41, Ano41h, Ano41k, Bie49, Che41, Col41a, Con41, Jew41]. Alexandre [Chi49]. Algæ [Rod15, Set18]. Algebraic [Bro04, Lam04, Lam08]. Algodonite [Koe03]. Alkaline [Smy27]. Alleghenian [Ort13]. Alliances [Lin38]. Allied [GH39, Gro00b]. Alligator [Hay11]. Allotrophy [Mül26b]. Alloys [Sau27]. Almshouse [Hun32]. along [Ada10, Cop49]. Alphabet [Lym15]. Alps [Hyd35]. Also [OC31]. Alterations [Rog10]. Alternating [KV19, KN20]. Alternating-Current [KV19]. Altitude [Kra47]. Altitudes [Cou45a, Sen30]. Amazon [Eig21b]. Amblypod [Pat35]. Amblypoda [Pat34a]. Ambrosia [Hau22]. Ambystoma [Cra39]. Ameiurus [Par41]. Amending [Ame24, Ame33]. America [Cou45b, Fra45, Per45, Sip45b, AH13, AB18, Art25, Atk18a, Bal09a, Bal17, Bal19, Ber14, Bog38, Chi45, Chi47, Cob15, Col43b, Cri31, Eig17b, Eig21b, Fle41, For43, Fri45b, Har29b, Hob43, How36, Hrd43, Kno40, Kot45, Lut27, McC27, MP44, Mor43a, New48, Oli35, Pat28, Row16, Rus30, Rus47, SK36, Sim40, Sim42, Sim43, Spe26, Tay39, Tre35]. American [Ano00f, Ano30b, Far17, Jac42, May21, OM19, Ros03b, Adl27, Ano26j, Ano28e, Ano29e, Ano30c, Ano30f, Ano31f, Ano32a, Ano35l, Bal09b, Bal12, Bin48, Bro42a, Car27, Chi43a, Chi43b, Chi45, Chi49, Com40, Con40, Con41, Con47, Cur49, Cus48, Del47, Der27, Din43, Dri19, DM47, Edg43, EM49, Fer42, Fis12, Fis32, Ful36, Goo48, Gre29, Gro02a, Hei18, Hoo32, Hrd26a, Hum42, Hun32, Hun13, Jep33, Jep34, Kid48, Kyt49, Lel32, Lut48, Lyd48, May43, McG43, McK40, Mer36, Mit42, Nic47, Nol34, NW38, Odg43, Osb23, Pac46, Pac42, Phi04, Pri03, Ric42, Ros00, Ros03a, Ros07, Ros11, Sco27, Shr43, Shr46, Shr47, Sio49, Smi18, Tre15, Tre17, Tre30, Tre34, Vai43, Voe49, Vol38, Wer43, Wil42a, Wis42, vD43]. Americas [Her37]. Amino [Ros47a]. Amity [Dub10]. Ammonium [NP14]. among [Ano34l, Bro46, Cal29, Pen27, Set18, Whe18]. Amos [Sto18]. Amphibians [Cra38, Cra39, Cra40]. Amphipoda [Sho45]. Amplifications [Tre34]. Amplifiers [Eva37]. Amulets [Bon42]. Amylolytic [SM39]. Anæsthetics [Hei34]. Analogy [Rog12]. Analysis [Ble36, Dem36, HS46, Jon46, KU08, KS05, Mic06, New48, Sch39, Sei42, Shr43, Shu10]. Analyzed [Wee12]. Anasazi [Col43c]. Anatomical [Kie26]. Anatomy [Mar05, Rum16]. Anax [Cal34]. Ancestors [Gre28b, Moo08]. Ancestry [Osb28, Rob49b]. Ancient [Abb18, Hau11, Ort02, Noc42, Win40]. Ancients [See07b]. Andaman [Spi08b]. Andean [Ben43]. André [Tru37]. Anemias [HC40]. Anemones [Par16]. Anesthesia [Bas15]. Angeles [Sin10a]. Angle [Ste37]. Anglo [Car27, Kyt49]. Anglo-American [Car27]. Animal [Par22, Per40, Ste02]. Animals [Ano34l, Cal29, Car17, Di 03]. Anniversary [Ano30c]. Annotated [AB18]. Annual [Dou22, Ano30b]. Anode [KS05, KS06, WS06]. Anomalies [Hes38]. Anomalous [Ben38]. Anonymous [Hay24b]. Anoplophrya [KH38]. Ant [Sch44a, Whe18]. Antagonism [Ost16]. Antarctic [Bar45, Bla45b, Cla45, Cou45a, Cus45, Dei45, Ewi45, Fra45, Fri45a, Gla45, KBC45, Per45, Sch45a, Sch45b, Sho45, Sip45b, Wad45b, Boy45, Dav10c, Dor45, Loc45, Per40, Ste34b, Ste45a]. Antarctica [Bow30, Wie45, Ber45, Bry45, Cou45b, Gou40, Hob40, Joe40, Kno45a, Kno45b, McC45, Pas45, Ron45, Shi45, Ste45c, Ste45b, Ste47, Wad45a, War45]. Antecedents [Lel46, Wat39a]. Anthropoid [Gre27]. Anthropological [Tit16, Wis27]. Anthropologist [Sha39]. Anthropologists [Bro42b]. Anthropology [Col30, Hoo32, Hoo35a, Hrd43, Kid48]. Anticyclone [Hob15, Hob21, Hob43]. Antigens [MCPL37]. Antillean [Mat19]. Antioch [Mor36]. Antioch-on-the-Orontes [Mor36]. Antiquaries [Sch46]. Antiquity [Bar29, Col43b, Osb27]. Antiquorum [Cha38]. Antitoxin [Nor41]. Antonio [Eig19]. Apatemyidæ [Jep34]. Apparatus [JR40]. Apparent [Wat39b]. Apparently [Sto03]. Appear [Wri20]. Application [Ash19, Bas15, CF22, Kop43a, MHM31, MW25]. Applications [Ada23, Bai47, Dar47, Eva37, Lam03, McE38]. Applied [Ame33, Ber33]. Approximations [Lam11]. April [Ang42, Ano30b, Ano30c, Con37a, Wie45, Ano00j, Ano00i, Ano01f, Ano01g, Ano02e, Ano02f, Ano03e, Ano03f, Ano04e, Ano04f, Ano05f, Ano05h, Ano05g, Ano06e, Ano06f, Ano07f, Ano07g, Ano08e]. Arabia [Mon28]. Arabic [Bel45]. Arbitration [Bal07, Tow16]. Arc [Com22, Hes38]. Archaeological [Cre43, Abb18, But07, Dix30, Mor43a]. Archaeology [Col43b, Din43, Hen48, Kid48, Vai43, Din39]. Archeological [Lun33]. Archeology [Car49]. Arches [Lam19, Mil19a]. Architecture [Wat39a, Wer43]. Archives [Ros00]. Arctic [Bar40b, Bry02, Ekb40, Ell40, Har28, Hob10, Moz37, Ste40]. Area [Dal05, Lun33, Mor43a]. Areas [Liv13]. Areography [Low02]. Areopagitica [Osg45]. Areopagitica-1644 [Osg45]. Argentodomeykite [Koe03]. Arising [See08]. Arizona [Dou01, Mag10]. Armada [Bro41]. Army [Sch44a, Ash19, Mil19b]. Army-Ant [Sch44a]. Aromatic [Her40]. Arran’da [Mat07a]. Arrangement [Bre04, Mac24]. Arrive [Bra42]. Art [Bal08a, Bal18, Blo17, Mor32]. Arthropoda [Pac03]. Artifacts [Abb04a, MP44]. Artifical [Tit14b]. Artificial [KN20, Koe03, Mac23]. Arts [Kim43, Nic44b, Tay40]. Asellus [Nic02]. Ashm [Whi39b]. Asia [Det37, Dix30, Hrd21, Hrd32, Mac34]. Aspect [Tow16]. Aspects [Cal42, Coh14, Dur17, Fle41, Fra31a, Gow34, Grè34a, Hun17, Joh20, Kis39, Mil19b, Rad34, Rid35, Smi18, Squ17, Sve47, Wad45c, Web17, Wil16, Wol32, Wri20]. Asphaltum [Sin10a]. Assault [Com31]. Assembly [Bal02]. Asset [Com46]. Association [Cri12, Mag07, MP44]. Associations [Köh41]. Assyria [Jas08a]. Asteroids [See10b]. Astrapotheria [Sco37]. Astraspis [Bry36]. Astrology [Jas08a]. Astronomer [Mit38]. Astronomical [Bar07a, Neu48]. Astronomy [Abb27, Bro22, Din39, Mit42, Rus47, Sha47]. Atavus [Jep37]. Atlantic [Ano11f, Bal09b, Ber11, BBC^{+}38, Hal41, Sei40, Sei42]. Atlas [Bog49, Kur34]. Atmosphere [Mac36c, Men39]. Atmospheres [Str39]. Atmospheric [Bau15, KS14, Kop45a, SP45]. Atoll [May21]. Atom [Sny26c, Swa28, Wil11]. Atomic [Boh47, Com46, Hin10, Hin11, Hin13, Lan46, Opp46, Ric04, Sho46, SE04, Smy46, Sny26b, Sny26c, TE33, Tay41, Vin46, Wil46]. Atoms [Com31]. Attainment [Bar40b]. Attempt [HS04, War48]. Attempts [HS46]. Atterbury [Joh36]. Attitudes [Har48a, Noc42, Ste48, Sto48, Tho37b]. Atypical [Jon37]. August [Con15]. Augustus [Bar01]. Aurelia [Son38]. Aurora [Gar41]. Auroral [Wie45]. Australia [Dav36, Dav38, Dav49a, Mat00b, Mat05a, Mat07a, Mat07d, Mat09b, Mat10, Osb34]. Australian [Mat00a, Mat00d, Dav41b, Dav49b, Mat00c, Mat07b, Mat09a, Mat09c]. Authorship [Hay24b]. Autocollimating [Ric12]. Automatic [BM38, MW25, Ros13]. Autonomous [Wei41, WB43]. Autosynthetic [CGTR32]. Average [Wil10]. Avian [Rou12]. Aviary [BF27]. Aviation [Bla28, Bro44b, Wal26]. Avifauna [Bon34]. Axes [See11c]. Axiomatically [Tre16]. Aztecs [Nut26, Vai43].

B. [Lin48]. Bābur [Jac29]. Båråboedaer [Cas39]. Babylonia [Jas08a]. Babylonian [Bar17, Jas08b, Row29]. Bache [All41, Ano41h, Ano41k, Che41, Col41a, Con41, Jew41, Odg41]. Back [Ano03d, Ano05d, Ano26g, Ano27c, Ano30d, Ano47f, Ano47g, Ano47h, Ano47i, Ano47j, Ano48g, Ano48h, Ano48i, Ano48j, Ano48k, Ano48l, Ano49h, Ano49i, Ano49j, Ano49k, Ano49l, Ano49m, Ano49n]. Background [Abb36, Lel46]. Backward [Kid40]. Bacon [Bur27]. Bacteria [WB43]. Bacteria-Free [WB43]. Bacterial [MCPL37]. Bacteriolysis [AB02]. Badlands [Sin21a, Sin22a]. Bagatelles [Ros01]. Bahri [Win32]. Balance [Whi35]. Ball [Hum36]. Ballistics [WA19, Web19]. Balloons [Bla09]. Bands [Par35, SE42]. Bank [Ald49]. Banking [Nor32]. Barker [Tho11]. Barriers [Bla45c]. Bars [Hau01]. Barton [Pen42]. Basal [Ben24, BM32]. Base [Ber45, Bla45b, Bry45, Cou45a, Ekl45, McC45, Ron45, Shi45, Wie45, Per45, Sip45b]. Based [Art25, Mac36b]. Bases [Mac21a, Smi26a]. Basic [Sli48]. Basilica [Kra41]. Basin [Cre43, Jep30b, RW09]. Basins [Eig21b]. Basis [All19, Mon04b]. Bates [Pac04b]. Bay [SP46, NL39, New40]. Bays [Dal15]. Be [Bro31, Che27, Emm32, Joh28]. Beach [Kie26]. Bean [Fow04]. Bear [Bry32]. Bearing [Bro38, Col43b, Cri31, Det37, Dix30, Mat19, Ort02, Ric37, Sch15a, Sim40, Sim43, Smi06, Sto23]. Beartooth [Bry33, BR34, Bry35]. Beat [Coh14]. Beatty [San35]. Beauvois [Mer36]. Beds [RP35, Sin05, Sin21b, Sin24, Sin25a, Ste11a, Ste11b, Ste12, Ste13, Sin15, Wan23]. before [BLV47, Zir41]. Beginning [Bal15]. Beginnings [Bar40c, Qui45, Sim42, vD43]. Behavior [Cog26, Dav24, Ges41, KU08, MS20, Mac36b, PS40, Sch44a, Smi26a, Smi26b, Whi39b]. Behaviorist [Tit14a]. Belgium [Lin38]. Belgrade [Pri34]. Bellevue [Ano30b]. Belo [Few35]. belonging [Bry26]. Ben [Nol44]. Bench [Gil42, Jep40]. beneath [Abb04a, See08]. Benefactor [Jew41]. Benjamin [Abb06, But48, Dav06, Dua27, DM47, Har40, Pen42, Sio49]. Beowulf [Mal49]. Bermuda [Har08]. Bessemer [Boy43]. Bethlehem [Ogb07]. between [Bar09a, Bet44, Cop49, Cri13, HH41, Ken09, Mer05, Mil10b, Mil46, Rog12, Rus12, See11a, Spr48, Wen35, Wri20]. beyond [See11a]. Bias [Har48a]. Biases [Wil39c]. Biblical [Alb30]. Bibliography [Hay00b]. Bicentennial [Con43]. Biennis [Dav26]. Big [Sin21a, Sin22a]. Bighorn [Jep30b]. Bilingual [Bel45]. Bill [Lew42]. Binaries [Kop41, Kop43a, Kop44]. Binary [Kop43b, Kop45b]. Binocular [Wil39b]. Biochemical [Hep18, Sta37]. Biocolloids [Mac21a]. Biographical [Cha06b, RW07]. Biographies [Rol13]. Biological [BB30, Bro44a, Bro06, Mac36b, McC27, Moz37, Pac04b, Par39, Whe31]. Biology [Bry45, Per45]. Bioluminescence [Har32]. Bioluminescent [HS30]. Biophysical [CF22]. Bird [Rid24, Sin28]. Birdhawal [Mat07c]. Birds [BF27, Fri45a, Fri45b, RD45]. Black [Sch15b, Tre12, Tre21]. Blackhead [Tyz32]. Bladder [McM23]. Blaine [Vol38]. Blaise [Men40]. Blind [Eig19, ZF47]. Blindness [Hau21, Oli02]. Blood [Coh44, Cra38, HC40, Par38, Ste02]. Blue [Rod15]. Bocaina [Lut26]. Bodies [Lov09, See15]. Body [Blo17, Du 48, Fra45, Rid24, Rid27, Whi39b]. Bohemia [Few32]. Bomb [Vin46]. Bombshell [Cha08]. Bone [Dav41a]. Bones [Bel49, Sin10a]. Book [Tru36, War28]. Books [Cra17]. Boone [Nol44]. Borings [Dou01]. Bornholm [Pri24]. Botanical [Bai47, Mer33, Tru28]. Botanist [Mer36, Tru37]. Botanists [Fer42]. Botany [Bai15, Cou27, Mer43, Tre16]. Both [Her37]. Bottom [Pig38]. Bottoms [See08]. Boundary [Tit08]. Bowring [Col41b]. Box [Bel49]. Boyan [Pri17]. Brāhmana [Dum48]. Brachiators [Gre28b]. Brachiopod [Con02]. Bradbury [Tru29]. Brahman [Eme40]. Brain [Cas37, Cra38, Cri13, Heb42, Wil08a, Wil10]. Brains [Cra39, Cra40, Spi08b]. Branches [Lot39]. Brant [Sch36]. Brass [KS27]. Brazil [Ber35, Har29a, JB13, Sop16]. Brdo [Few35]. Brea [Sin10a]. Breeding [Bai03]. Bridge [Mer05]. Brief [Luc18, Lym18, Wil02]. Britain [Hay24b]. British [Bal09b]. Bromine [Hin13]. Bronze [Cal41]. Brooks [Con08]. Broom [Kun43]. Brown [Sto18]. Bruce [Bar12]. Bubonidæ [Shu00a]. Building [Dal25]. Bullfrog [RE40]. burchelli [Sch44a]. Burial [Mat09c, Mat10, McG43]. Burning [Hau09b]. Burr [Hol47]. Burrage [Hoo35b]. Bursahom [dT42]. Bursaria [Jen41]. Bush [Hau09b]. Business [Hue34, Joh28, Pat32]. Bust [Rob40]. Bütschlii [Wen37]. Butte [Bry32, Bry33, BR34, Bry35]. Butterflies [Ger43]. Byrd [Pas45, Sau40, Sio49, Ste45b, Haw45, Sau40].

C [Bee39]. C.E. [Pet03]. Calcium [Goo04, HMW14, RD45]. Calendar [Ano00f, Bar49b, Ken35b, Win40]. California [Cle35, HA04, Kra10, Sin10a, SP46, Smi06]. Called [Lym12, Mor32, Rob40]. Caloric [Bro49]. Calorimeter [BM38]. Calorimetry [Bee39]. Camarasaurus [OM19]. Cambarus [Ort05]. Cambrian [Dal15]. Cambridge [Tit16]. Camera [Bak40a, Bak40b, MHM31]. Cameras [Bak40a]. Campos [Har29a]. Can [Che27, Joh28]. Canada [Hay24b]. Canal [Joh38, Joh43, Tho07, Tow13]. Cancerous [WB43]. Cannon [Dan11].

Cannot [Emm32]. Canyons [Ste38]. Capientium [Dav39b]. Capitalist [War44]. Caps [Dav49b]. Captives [Kno40]. Capture [See10b]. Car [Dud04]. Carbon [BS07, Bog12, Hey10, Pre04]. Carboniferous [Hay00a]. Career [Ano41h, Sel48]. Carnegie [Bau15, Mor43a]. Carnivora [GH39]. Carnotite [Bla13]. Caroli [Str42]. Carolina [Hey13]. Carré [Men40]. Carried [Hoo26]. Carson [KW26]. Cartouches [Low03]. Casamayor [RP35]. Case [Lan06a]. Casks [Bry02]. Cassiopeiæ} [McD15]. Castaways [Gam44]. Cataloguing [Sch14]. Catalytic [Bee39]. Catastrophe [Car09]. Catawba [Few44]. Catfish [Par41]. Catfishes [Par40]. Cathode [KS05, KS06]. Cathodic [BS07]. Catlin [Bal18]. Cats [Jep33]. Caudal [Par35]. Cause [Lon48, See06]. Causes [Hov09]. Cave [Hib43]. Cayuga [RW09]. Celestial [MHM31]. Cell [Guy23, Ham29, Ham30, Har32, Mac21a, Mac24]. Cell-Masses [Mac21a]. Cell-Walls [Mac24]. Cells [Car29, Cri13, CGTR32, HC40, Hei30, Mac23, Mac28]. Cellular [Mon04b]. Celt [Pri20]. Cenozoic [Det37]. Central [Cre43, Dou02, Wei41, Har28, Kra10, Mat07a, SE42]. Centrechinoidea [Mac36b]. Centrifuged [How41]. Centuries [Dan11, Put47]. Century [Gre34b, Mun11, BS49, Bla36, Chi47, Hum42, Shr47]. Cepheid [SN48, Sha49]. Ceremonial [Mat09a, Mat09b]. Ceremonies [Mat00c, Mat00d]. Certain [AB02, AD41, Cle35, Cri21, Dav37b, Eas15, HC40, Har33, HHHN17, JS09, Kra05, Lov09, Mer37, Nic16, Pea16, Sco30, Wri20]. Certificate [San39]. Cetacea [Tru08]. Chain [Fer46]. Chamber [JR40]. Change [Not43]. Changed [Bec30]. Changes [Loe11, Par37, Par38, Par40, Rid27, Wis27, Zel31]. Changing [Mer34, Pat28]. Channels [Hau03]. Chapter [Car17, Dub10, Jas04]. Character [All26, Blo13, Met05]. Characteristic [dCW23]. Characteristics [Hob10, Rus12, Whi39a]. Characters [Dou22, Hel28, OM19, Sto23, Wis27]. Charles [Che11, Kee11, Bry09, Cha49, Chi49, Cop48, Mil35, Rob49b, Sel48, Swa36]. Charles-Alexandre [Chi49]. Chart [Bac36]. Charter [Ano30c, Sho46]. Charts [Bau12, Har07]. Chavin [Ben43]. Chemical [Cla06, Cri21, Cri30, Dem36, Ham29, Ham30, Ric03, Tay26]. Chemistry [Rog12, She18]. Cherokee [Few44]. Chesapeake [NL39, New40]. Chestnut [Rum16]. Chief [Lon48, Wes44]. Chiefly [Di 03, Mac48]. Chiffre [Men40]. Child [Dav31, Dav35]. Childhood [Wis27]. Childlessness [Kis39]. Chile [Kel09]. Chilean [Kel08]. Chili [Eig21a]. Chimæra [Tre33]. Chimpanzees [Fur16]. China [Osb23, RP06]. Chloride [JA08]. Chondrification [Cas33]. Choptank [Wes44]. Christian [Kra41, Eme39]. Chromatin [Nic02]. Chromatophore [Par44]. Chromic [Bee39]. Chromosome [Smi34]. Chromosphere [Mit30]. Chronology [Col43a, Cro43]. Chrysocolla [Kel09]. Church [Cro49, Cro49]. Chytridiaceous [Spa37]. Cicero [Ran32]. Ciliate [KH38]. Circle [Har28]. Circuital [Ken33]. Circular [Ada35a, Ada36a]. Circulating [Fen01]. Circulation [Cha06a, Hob15]. City [Ano30c, Grè34a, Car09]. Civets [GH39]. Civil [Cus48, Nic47]. Civilian [Spr48]. Civilization [Bar29, Hay40, Hou32, Hun13, dT42]. Civilizations [Row29]. Clark [Tru28]. Classic [Dic27]. Classical [Eme40]. Classics [Wri43]. Classification [Bog12, GH39, Hof39b, Jac31b, Kea02, Osb25, Pac03, Tre12, Tru08]. Clay [Abb04a]. Cleveland [Mar17]. Climates [Ber22, Cha06a, Dav10c, Sip45a, War25, War28]. Climatic [Dou22, Liv13]. Climatology [Hum42]. Clinical [Cri13, Swi23]. Clock [Tuc22]. Clones [Jen41]. Close [Kop42, Kop43b, Kop45b]. Clothing [Sip45a]. Cloud [JR40, SN48]. Clouds [Lan48c]. Cloudy [Bar09b]. Clovers [Har22]. Clustering [See12b]. Clusters [See12b, Sha19]. Coal [Ste11a, Ste11b, Ste12, Ste13]. Coalescence [Las41]. Coast [AH13, Col41a, Kra10, Bol47, Fri45b]. Coastal [Ano11f, Ber11, JW27]. Coccyges [Shu01]. Cochise [Hau43]. Cocoon [Gro02b]. Codes [Hal49]. Coiling [Few44]. Coit [And09]. Cold [Fra45, Sip45a]. Colias [Ger43]. Colleagues [EM49]. Collected [Sch45a]. Collecting [Cop48, Fen49a]. Collection [OM19, Ros13, San35]. Collections [Bre04, Ste34b, Art25]. College [Woo30]. Collembola [Gla45]. Collisions [SP42]. Colloidal [MS20]. Colombia [Eig24]. Colonial [Bra42, Jon42, Jon46, Qui45]. Colonic [Bas15]. Colonies [Hay24b]. Colonization [Jon46, Lut27]. Colony [Bra42]. Color [Ger43, Kop41, Kra04, Nic16, Oli04, Oli05, Par37, Par40, Ska46, Tit16]. Color-Index [Kop41]. Color-Photographs [Nic16]. Color-Signals [Oli04]. Colorado [Bla13]. Colored [JS09, TP08]. Colors [Whi39b]. Columbium [HS05, Smi05]. Columbus [Ols41]. Columns [KS27]. Combustion [Du 31]. Come [Dav06]. Comet [Bar10, Low11]. Comets [See11a]. Coming [Hrd32]. Commemorating [Ano30c]. Commemoration [KW26]. Comment [Woo30]. Commerce [Dub10]. Committee [Con36, Sch39, Con37a]. Commodities [Pea19]. Common [Mas04, Put47, Sch16]. Communication [Hov48, JA08, Kae31, Mac36c]. Commutative [Mil11]. Commutator [Mil07]. Company [Dav10b]. Comparative [Fox29, Har08, Har09, PS40, Tyz42, WB37]. Compared [Hob21]. Comparing [SK37]. Comparison [Ogb07, Row29]. Compass [Bau17, HB22]. Complete [Ber39a]. Complex [Smi38]. Components [MS20]. Composition [Com35, Mab03, Mer26, Sit39]. Compounds [Bog12, Her40]. Computation [Kop44]. Concave [Bar09c, Ric12]. Concentration [May19]. Concentrations [Fen48]. Concept [Lam05, Sch29]. Conception [Bec30, Fal14, Dal15]. Conceptions [Mul49]. Concerned [AB02]. concerning [BLV47, HM38, Tyz32]. Conclusions [Gro02a, Osb25]. Concretionary [Sin24]. Concretions [Rod15]. Condensation [Bar09b]. Condensed [Ekl45]. Condenser [Ada36b, Ada36c, Bar09a]. Condition [Mer26]. Conditions [Dal05, Hal41, Sch15b, SP46, Son38]. Conducting [Ken09, Ken24a]. Conductivity [Ken09]. Conductors [Ada37c, Bar09a]. Cones [LL40].

Confederacy [Fen49b, Wal49a]. Conference [Con37b]. Conflict [Blo44b]. Congenital [Oli02]. Congress [Coh48]. Congresses [Hun47]. Conic [Neu48]. Conjugation [Ros40]. Connected [Ada37b, Hob23, See06]. Connecticut [Pri03]. Connection [All41, Che41, Con40, Con41, Reh40, See11a]. Connections [Ken30]. Conservation [Com45, Lin49, Low45, Mer34, She18]. Consideration [Sch29, Whe18]. Considerations [Fal14, See10b, See13, Web19]. Considered [Hay24b]. Consolidation [Joh25, Joh29b, Joh30]. Conspiracy [Hol47]. Constant [Abb11]. Constantine [Bet44]. Constants [Ken28]. Constitution [Bec30, Cha39, Fen49b, Fox29, Gow34, Mor15, Rei15, Wil39a, Wil11, Ame24, Ame33]. Constitutional [Mon28]. Construction [Dav10b]. Consumer [Har48a, Ste48]. Consumption [Pea19, SS38]. Contact [Aga23, Lot39]. Containing [Bar09a]. Contemporary [Cla32, Sea48]. Content [BJ27, New40]. Contents [Mac23]. Continental [Cri31, Hob43]. Continents [See08]. Continual [Bru26, Bru28]. Continued [BM38, Det23, Pea16]. Continuity [HK37, Kra02]. Continuous [Har33, LL40, Mac30b]. Contraband [Moo12]. Contraction [See07b]. Contribution [Hay00b, Hrd43, KU08, Pat34a, Pat35, Pea19, RB08, Rid29, Sha19, Sin21c]. Contributions [Ber41, Dil47, Rei40]. Control [Dub44, Hei32, Kay45, Lan46, Lin27, Mac39, Rav13, Rid35, Sho46, Sta47, Sto12, Tau49, Wri21]. Controlling [Ang36, Par39, Pat34b]. Controls [BB30]. Convalescence [Rus46]. Convection [KS14]. Conversion [Hey10]. Cooling [See07b, SP45]. Cooperation [Boh47, Com45, Kyt49, Mac46, Sha47, Pat16]. Coördination [Wil32]. Cope [Gre34b, OM19]. Copernican [Fra44]. Copernicus [Bal08b]. Copper [Kel08, Kra05]. Copy [Bra42]. Cor [Str42]. Coral [Hof40, May20, Pal28, Set26]. Corals [Pal28]. Corbino [Ada15]. Cordyceps [Mai34]. Core [Pig38]. Core-Samples [Pig38]. Coregonus [Fow04]. Cores [BBC^{+}38]. Corinth [Cal41]. Cormophytaster [Tre16]. Corn [Ric39]. Cornelius [Pac35]. Cornwall [Tra06]. Corona [Whi11, Whi20]. Coronal [Lam19, Mil19a]. Coronas [Bar09b]. Corporations [Sel44]. Corpus [Cha38]. Correction [Mat07b]. Corrections [Ano34l]. Correlate [HS04]. Correlation [Bru28, Hoo35a, Ort20, Pat13, Ste37]. Correlations [Hal41]. Correspondence [Ano00f, Bet44, Roe48, Tru29, WPL47]. Corresponding [RB08]. Corridor [Bol47]. Corrigenda [Ano12f, Ano13f, Ano22e]. Corrigendum [Ano11f]. Corroded [Cal41]. Cosmic [Com35, Fle41, Kor41, KBC45, MN40, Sny26a, Sny26b, Ste36b]. Cosmic-Ray [KBC45, MN40]. Cosmic-Terrestrial [Ste36b]. Cosmical [See10c]. Cosmogony [See11b]. Cosmology [Rob49a]. Cosmopolitanism [Fra27]. Cost [Mun11, Sch39]. Cotton [Mun09]. Coudersport [Bal21]. Council [Lel32, Sch39]. Councils [Lel37]. Counsel [Ros13]. Count [Bro49]. Counter [Eva37]. County [Gil42, HA04, Jep30a, Jep40, Tra06]. Court [Ree36]. Courts [Put47]. Couthouy [Con02]. Cozumel [Ric37]. Crater [Mag10]. Crawford [Ros03b]. Creation [Bar17]. Creek [Sin15]. Crepidula [Con38]. Cretaceous [Dou02, McC26, Rus30, Sta03]. Criminal [Hoo32]. Crises [Cla32]. Crisis [Wil32]. Critical [Mon08, Par18, Shr43]. Criticism [Mar28]. Criticisms [Pac04b]. Critique [Gre44, New48]. Crocker [Cam19]. Crookes [Goo03]. Crop [Low45]. Crossability [Bla45c]. Crosses [Atk18c, Dav37b]. Crossopterygian [Gre35]. Crown [WB43]. Crown-Gall [WB43]. Crucifer [Rug41]. Crusades [Mun16]. Crust [See08]. Cruz [HA04, Sin05, Sin08]. Cryptaspis [Bry35]. Crystalline [Nor41, Sta37]. Crystallization [GG38]. Crystallized [Koe03]. Crystallographic [RB08, Wri03]. Crystallography [Rog12]. Crystals [Ric03]. Cuckoos [Shu01]. Culminating [Bar40b]. Cultural [Ken48, Noy47, Tay39, Vai43, Wir48]. Culture [Com40, Sau48, Sea48, Spe26]. Cultures [Hrd26a, Lew45]. Cure [Kun43]. Currency [Wes34]. Current [Eag30, KV19, KN20, Tae39]. Currents [May19]. Curse [Hau11]. Curtis [Emm03]. Curve [Pea24, Sha49]. Curves [Bro04, Kop42, Kop43b, Kop45b]. Customs [Mat04, Mat09c, Mat10, Met05]. Cutting [Kay45]. Cyaneum [KH38]. Cycle [HK37]. Cycles [Chi16, Det37]. Cyclic [Loe11]. Cygni [Str43]. Cylinders [Ada35a, Ada36a, Ada37a, Du 31, Ken09]. Cylindrical [Ada36b, Ada36c, Ada37b, Ada37c, KS27]. Cyto [Cle35]. Cyto-Taxonomic [Cle35].

D.C.L. [Doo10]. Dīn [Far39]. Dakota [Del44, Jep34, Sin21a, Sin21c, Sin22a, Wan23]. d’Alexandre [DH47]. Dallas [All41, Ano41h, Ano41k, Che41, Col41a, Con41, Jew41]. Dana [For36, Hof40]. Daniel [And09, Bar10, Nol44]. Danish [Pri24]. Danubia [Jás49]. Danubian [Few37]. Darkening [Kop41]. Darwin [Bry09, Con09, Ful09, Goo09]. Darwinian [Mul49]. Dating [Few35]. Datura [CBB36]. Davis [Ano30c]. Days [Car27]. Dead [Dav49a]. Death [CTR29, Lon48, Spi08a]. Debit [Bro35]. Debt [Con09]. Decade [Voe49]. Decapods [Ort02]. December [Lin33, Ano00k, Ano01h, Ano02h, Ano02g, Ano03h, Ano03g, Ano04h, Ano04g, Ano05i, Ano06g, Ano07i, Ano07h, Ano08g, Ano08f, Chi48, Low01]. Decimals [Fen01]. Decisions [Has00]. Declaration [Hay00b]. Decline [Lon48]. Decomposing [SS38]. Decora [Sin10b]. Decrease [See15]. Deep [BBC^{+}38, Cha06a, Wil20]. Deep-Sea [Cha06a]. Deeper [Hau03]. Defender [Row16]. Defense [Yer41]. Deficiency [Elv49]. Definitions [Ken33]. Deformans [Wil02]. Deformations [Eme17]. Degenerate [KH38]. Deglaciation [Sau48]. Degradation [Ehr16]. Degree [Nip19]. Degrees [Cop49]. Dehaven [Ekb40]. Deir [Win32]. Delaware [Pri02]. Demarcation [Tit08]. Democracies [Bro35]. Democracy [Lin27, Nic47]. Demonstration [Joe40]. Denis [Cro49]. Dental [Gre34b]. Dentition [GH39, Hel28]. Dependence [Sto03]. Deposit [Abb04a]. Deposition [KS06, Sch15b]. Deposits [Dal15, Hat02, Sin10a, Wil20, dTdC36]. Depression [Nor32, Pat32]. Depressions [Fis32]. Depth [Hay24a, Mer05, See12a]. Depths [See07a]. Dercum [Bru32]. Derivatives [Coh44, Dav37b, KM04]. Dermal [Sin10a]. Descent [Gro00a, Osb34]. Described [Tit16].

Describing [Mac36b]. Description [Fow04, Jep34, RP35, Sch17]. Descriptions [Eig17a, Hay00a, McC26]. Desert [Tre11]. Destruction [HC40, Ort09]. Detecting [May19]. Detection [Eva37, Hay20b]. Determination [Bla13, Dav08, Kop41, Len26, Ric04, See12a, Whi35]. Determined [Hin11, MPS20]. Determining [Ken28, Sei40]. Determinism [Hal43]. Detonation [Mun09, MH20]. Deux [DH47]. Development [Ame33, Atk15, Ber33, Bid28, Bru18, Cal34, Cog26, Con38, Cou27, Cri31, Dav35, Dav39a, Dav40, Dav44, Fal14, Fer46, Has00, Hoo35a, How41, McC27, Pac04a, Pat13, Ree38, RM43, Rus47, Sco27, Sea48, Sto12, Tay39, You48, RE40]. Developmental [Ges41, Rid29, Rug39, RE40]. Developments [Lut48, Mil27]. Devices [Rol22]. Devonian [Gre35, Bry32, Bry35]. Diabase [HA04]. Dialect [Pri24]. Dialogue [New18]. Diameters [Pea21]. Diaphragms [KT15, KT16]. Diary [Mim12]. Dickson [Car16]. Dicotyledons [Har07]. Dictatorships [Bro35]. Did [Ald49]. Diet [Ben18, Loc45, McC19]. Difference [Bar09a]. Different [Ano34l, Cal29, RB08, Son38]. Differential [Lam11]. Differentiation [Jen41]. Difficulties [Mac36b]. Diffraction [Bru13, LHYH36, Mic15]. Diffusion [BLV47, Jac31a]. Dilute [Kra06]. Dimensional [Ada35b]. Dimensions [Bro04, Gap38, Gap40]. Dimethyl [KM04]. Dinner [Ano30b, Ano30c]. Dinosaur [Rus30]. Diphtheria [Nor41]. Diploma [San39]. Direction [Dar44a, Tay44]. Directions [Wil39c]. Directly [Har28]. Disappearance [Par35]. Discharges [Nip11a]. Discover [Cha47]. Discoveries [Dix30, Ekb40, Hib43, Hrd32, Lun33, Mat19, Mit32, Osb27]. Discovery [Bal19, Bro44a, Foo48, Hob40]. Discrimination [Tit16]. Discussion [Bro40, Bru14, Bru21, Bru22a, Bru23, Bru26, Bru28, Dar44a, Hea40, Hof39b, LBD^{+}38, Mac36b, MS31, Tau32, Ano40k]. Discussions [Sch29]. Disease [Elv49, Sho48, Sta47, Wil02]. Diseases [Dub44, Riv47]. Disorientations [Con38]. Dispersal [Ort05]. Disposal [Dav49a]. Disruptive [Nip11a]. Dissociated [Com22]. Dissolution [SM39]. Distance [TP07]. Distances [Mit28, Rus10]. Distant [SW49]. Distinct [Osb34]. Distinguishing [Ste02]. Distortions [Nic48]. Distribution [Ada37a, Ber14, Bon34, Cal34, Eas40, Eig21b, Har11, Hel28, LHYH36, MN40, NL39, Ort02, Ort18, Rei20, Ros13, Wat39b, Wet47]. Distributions [Ada35a, Ada35b, Ada36a, Bar07b]. Districts [SE42]. Disturbances [Bau10, Ste37]. Diurnal [Tuc22]. Diverse [Son38]. Divide [Ort13]. Divisions [Mat00a, Mat07b]. Dixon [Cop48, Hoo35b]. Do [Cha47, JB13, Sop16]. Documentation [Sch39]. Documents [DH47]. Dodder [Kun43]. Dollar [Fis32]. Domestic [Kuz48, Pat34b, Pea16, Pea17]. Domeykite [Koe03]. Dominance [Dav08]. Double [Doo03]. Doublet [Sch14]. down [Swa35]. Dr. [Bur23, Chi48, EM49]. Dragonflies [Cal34]. Drainage [Ort18]. Drift [Bry02, Ell40, Tro10]. Drifts [Oli33]. Drosophila [How41]. Drosophyllum [Har25]. Drought [Wea44]. Dry [SP45]. Dryopithecus [Hel28]. Due [Pac04b]. Dune [Abb18, Har08]. Duration [HS30]. during [Ano26j, Ano28e, Ano29e, Ano30f, Ano31f, Ano32a, Ano33a, Ano34a, Ano35a, Ano36a, BJ19, Cou45b, HK37, Hrd36, McK40, Mit42, Mor43a, Sau48, Shr47]. Dust [Bar07b]. Dust-Free [Bar07b]. Duties [Tow15, Wri21]. Dwarf [Dav37b]. Dwarfism [Abb36]. Dwight [Hof40]. Dying [Pri03]. Dynamic [Der27, Sei40]. Dynamical [McE38, Ost16, Ric11, See12b, Web17].

Each [Mil10a]. Ear [WB37]. Earl [Ros03b]. Earliest [MP44]. Early [Bal17, Bro42a, Chi45, Edg43, Fer42, Fet43, Fle41, Fra31a, Heb42, Hib43, How36, How41, Hun32, Jac42, May43, McG43, Mit42, Ric42, Ros07, RM43, Din43, Kra41]. Earth [Hob23, Bro22, Bru26, Cha15, Cha22, Cla22, Hay15, HB22, Hrd26b, Pen27, Rei15, Rei20, Sch15a, See06, See07b, See08, See09]. Earthquakes [Hec38, Hov09, See06, See07a, See07b]. Earthworm [KH38]. East [Mac33a, Bla45b, Bry45, Ekl45, Ron45]. Eastern [Cro43, Cob15, Dix30, Kno40, dCW23]. Eccentricities [See11c]. Eccentricity [See15]. Echinoderms [Cla45]. Echoes [Pri03]. Eciton [Sch44a]. Eclipse [And19, Cam19, Fro19, KS19, Lam19, Mil19a, Mit26, Mit30, Mit32, Mit38]. Eclipses [Kop45a]. Eclipsing [Dug15, Gap38, Gap40, Kop41, Kop42, Kop43a, Kop44]. Ecologic [Har22]. Ecological [Kie26]. Ecology [Bow17, Mye43]. Economic [Bou43, Del34, Fra31a, Hue34, Kuz48, Mer05, Pat28, Pat31, Put47, Smi18]. Economics [Bon40, Lov43]. Economies [Kea02]. Economy [Fet43, Inn44, Kuz48, Lut48, Win45]. Ecuador [Eig21a]. Editors [Bla36]. Edsel [Pas45, War45]. Education [Ang27, Ang42, Ayd30, Gra28, Odg43, Ros13, Sta48, Woo40]. Educational [BB30, Gre29, Noy47, Sta48]. Educator [Odg41]. Edward [Ste45b, Wad45a, For36]. Edwin [Sch36]. Effect [Cro32, Cro43, Heb42, JS09, KS06, Pea16, RM43, TP07, Tuc22, Wil05, Rug39]. Effects [Ada15, AD41, Ben18, Cle49, Dav17, Det23, Fra45, Hov09, KW44, Kop43b, Nic10, Nip06, Ric11, Rug39, RE40, See11c, Spe40]. Efficacy [Ste49]. Efficient [Hau03]. Egg [RM43]. Eggs [How41, PS40, RE40]. Egyptian [Row29, Win40]. Eight [Mor43a]. Eighteen [Eig17b]. Eighteenth [Chi47]. Eighth [Dum48]. Eighty [Hin11]. Einstein [Ada20, EM49]. Elasmobranch [Par37]. Elderly [BM32]. Electric [Ada10, Ken24a, KS19]. Electrical [Ada35a, Ada36a, Hey10, Ken25, Ken26, WB37]. Electrically [KS14, Nip13]. Electricity [Ada35b, Ada37a, Bau15, Gis41, Nip11a, Spi08a]. Electro [Cri21, Cri30, KS05]. Electro-Analysis [KS05]. Electro-Chemical [Cri21, Cri30]. Electrolysis [KS07]. Electrolytic [Goo04, WS06]. Electromagnetic [Ken28, Ken31]. Electrometric [Bar09a]. Electron [Bar09b, Fal14, LHYH36, SP42]. Electron-Electron [SP42]. Electronic [ZF47]. Electrons [BRS38, Dav30, HH41, Hou39, Lor06, Ric11]. Electrostatic [Ada37b]. Elementary [Ste11b, Whe46, dV06]. Elements [Dem36, Kop43a, Kop44, Sei40]. Elephant [BL36a, OC31]. Elimination [Nip06, Sch13b]. Elliptic [Bar11b]. Ellsworth [Joe40]. Elwood [Swa36].

Emancipation [Hau05]. Embassy [Dub10]. Embryo [Rug39]. Embryology [Con02]. Embryonic [Sto12]. Embryos [How41]. Emotional [Can11]. Emotions [Cri12]. Emperor [Jac29]. Enactment [Mon28]. Enamel [OC31]. Encountering [Wil17]. Endings [Spe42]. Endocrine [dMS27]. Endowments [Kep37]. Enduring [Oli42, Oli47]. Energy [Aus13, BJ19, Com46, CGTR32, Hey10, HH41, Lan46, MN40, PGG43, Rus39, SP42, Sho46, Wil46]. Enforcement [Bur12]. Engineering [Com45, Hun17, Jac42, Ken26, Wol32]. Engines [Du 31]. England [Jon42, Kee34, Kur34, Mat03a, Bro41, Dav26, May43, Put47]. English [Hew04, Qui45, Wat39a]. Enquiry [Dal05]. Entagion [Bel45]. Entelodonts [Sin21a, Sin22b]. Entering [Blo17]. Enters [For43]. Entitled [Hay24b]. Entomology [Cal42]. Entropy [Dar44b]. Environment [Bat49, Bel48, Far17, Pac04b, Sau48, Tho39]. Environmental [Dor44, Jen08]. Eocene [Gaz36, Sin28, Ber14, Ber22, Jep30b, Jep32]. Eperythrozoa [Tyz42]. Epic [Hop10]. Epidemic [Riv47]. Epilepsy [Wee12]. Epoch [Jep33]. Epochs [Low00]. Equal [Gol24, Hal24]. Equation [Bro04]. Equations [Lam03, Lam08, Lam11, Mac02]. Equiangular [Ken26]. Equilibria [Mor49]. Equipment [Bry40]. Equivalent [Bak40a]. Eriptychius [Bry36]. Ernest [Wal48]. Errata [Ano00g, Ano07d]. Erroneously [Di 03]. Error [Ric03, Ric04]. Erythrina [Shu42]. Erythro [HC40]. Erythro-Stasis [HC40]. Eskimo [Hei49, Col43b]. Especial [Fle41, Nic02]. Especially [Par37, Put47, See08]. Essay [Sch46, CDG^{+}12, BLV47]. Essential [Ham29, Lew45]. Establish [Wri20]. Establishing [KN20]. Estuarine [Har11, New40]. Ether [Bas15, Mer03, Tro10]. Ethnic [Dav38]. Ethnological [Tit16]. Ethnology [Bal08a]. Ethyl [JA08]. Ethylene [Cro32]. Etymology [Edg38]. Euler [See15]. Euplotes [HK37]. Europe [Not43]. European [Dra46]. Eusmiloid [Jep33]. Eusmilus [SJ27]. Evans [Cha49, Dav06]. Evaporation [Bas15, Jen36, Jen38]. Eve [Hau11]. Evening [Ano30c, Dav27]. Event [Eme40]. Every [Mil11]. Evidence [Bru26]. Evolution [Bal19, Car09, Con34, Der27, Gre28a, Gre34b, GH39, Har32, Hob09, Jen08, Mer34, Osb04, Osb25, Pat16, See10c, Whe19, Wri49]. Evolving [Coh48]. Examination [Kee34]. Excavation [Mor36]. Excavations [Alb30, Cro49, Few32, Win32]. except [Osb34]. Exchange [Foo13, Hue34]. Excitement [Can11]. Excretory [Mon08]. Executive [Mac48]. Existence [See10a]. Existences [Sto03]. Existing [Hay03, See11a]. Expansion [Lam03]. Expansions [Lam04]. Expedition [And19, Bar45, But07, Con40, Coo40, Fro19, Mit38, Reh40, Sin15, Sin21c, Tit16, Wad45b, Bar40a, Bla45b, Bry40, Cam19, Cla45, Cou45a, Cox48, Cus45, Dei45, Ewi45, Fow40, Fra45, Fri45a, Gla45, KBC45, Mil19a, Per45, Sch45a, Sch45b, Sho45, Sip45b, Tru28]. Expeditions [Sco28]. Experiences [Dra46]. Experimental [Bru26, Cri13, CTR29, Rug41, Sto23]. Experiments [Bea36, Bru21, Bru22a, Bru23, Bru24, Bru25, Dem36, Dou01, Har33, HM38]. Explanation [Abb04b, Ber22, Low01]. Exploration [Bar40b, Bla09, MG49, Sau40, Sip45b]. Explorations [KT15, Osb23].

Exploratory [Haw45, Tho37b]. Explore [Bal09a]. Explorer [Tru37]. Exploring [Con40, Coo40, Reh40]. Explosions [Mun10]. Explosive [Nip13]. Exposure [Rug39, RE40]. Exposures [Fen48, MW25]. Expressed [Ken36]. Extended [Bar40b, Str39]. Extension [See11a]. Exterior [WA19]. External [Bin48, Dav37a]. Extract [Smi03]. Extremities [Dav44, Gre35]. Exudation [Mac25]. Exuviæ [Spa37]. Eye [FR18].

F [Bea46, But44, Atk18c]. F.R.S [Doo10]. Facade [Cro49]. Facial [NS37]. Facilitate [Kop44]. Factor [Ben37, Der27, Ehr16, Par44, Pat16, Row14, Tho39]. Factors [AB02, Ang36, Ben24, CBB36, DSA45, Fla39, Foo13, Gre36a, Not39, Pat32, Pat34b, Wet34, You48]. Facts [Mon04b]. Faeroe [Pri22]. Fairfield [Gre36b]. Falling [McC05]. Family [Bak40a, Boa43, Bry26, Hal48, Rob49b, SE42, WBAW39]. Farm [Nol44, Cro43]. Farmer [Wil43b]. Fatalism [Bro02]. Fatigue [Car48]. Fats [Par40]. Fauna [Bry33, BR34, Gro02a, Mat19, Ort09, Ort13, Sin05, Sin15, vI02]. Faunas [Gil42, Jep40, Rus30, Sin24]. Faunule [Sta03]. Faunules [SK37]. Fauré [KR40]. Features [Mac39]. February [Ano00e, Ano00m, Ano00l, Ano01i, Ano01j, Ano02j, Ano02i, Ano03j, Ano03i, Ano04j, Ano04i, Ano06h, Ano07j, Ano08h]. Federal [Bla45a, Mag49, New44, Ame24, Ame33]. Feeble [God12]. Feeble-Mindedness [God12]. Female [Mon04a, Rid27, RP06]. Fertility [Fla39]. Fertilization [Jen41, Whi35]. Fetus [Ges41]. Fever [Eig22, Eig24, Rav13, Swi23, WBAW39]. Fibers [Spe40]. Fibroma [Ber37]. Fiction [Blo17]. Field [Bak40a, Cou45a, Foo48, Goo03, Mac30a, War25]. Field-Work [War25]. Fifteen [Mil26]. Fifteenth [Dan11, Put47]. Fifth [Bru25]. Fiftieth [Ano30c]. Fifty [MPS20, Smy46]. Figural [KW44]. Figure [Fow04]. Figures [Dav41b, See12b]. Filipino [Met05]. Film [Joh39]. Filterable [Kne37]. Fin [Gre35]. Final [Osb25]. Finance [Whi40c]. Finances [Joh43]. Financial [Ano16j]. Finder [Hay24a]. Fine [Tay40]. Finished [Mag49]. Finley [Ano30b]. Fire [Hou32]. Fireball [Wat39c]. Fireballs [Oli35]. First [Bac36, Bal02, Bra42, Con40, Dav06, Fer46, JA08, Ree36, Reh40, Roy06, Jew41]. Fiscal [Blo44a, Hun44]. Fish [Bry33, BR34, Cas37, Eig19, Gre34b, Gre35, McC26, Par34, Sco30]. Fish-Like [Sco30]. Fisheries [Bal09b]. Fishes [Bro03, Bry26, Bry32, Bry34, Bry35, Eig17b, Eig21a, Eig21b, Eig22, Eig24, Fow40, Par18, Par37, Sch45b]. Five [Ben18, But44, Web19]. Fixed [Hob21, See10a]. Flames [Nip11a]. Flash [Mit19, Mit26]. Flat [Bak40a]. Flat-Field [Bak40a]. Flight [Joe40, Sau40]. Flightless [Sin28]. Flights [Haw45]. Flora [Ber14, Lut26]. Floras [Ano11f, Ber11]. Flower [Ogb07]. Flowering [Bro38, Eas40]. Fluctuation [LL40]. Fluid [Nip06]. Fluogermanates [Mül26a]. Fluorescence [Nic10]. Fluting [Bra13]. Foil [Kra05]. Folding [See08]. Folk [Eme40, Sch16]. Folk-Tale [Eme40]. Folklore [Tho49]. Following [Han23].

Folsom [How43]. Food [Bou43, Pea19, She18, Smi18]. Foods [Wil08b]. Foraminifera [Cus45, Mye42, Mye43]. Forced [KS14]. Ford [Pas45, War45]. Forecasting [War26]. Forehead [Hrd33]. Foreign [Coh48, Cus48, Kuz48, Lin27, Lut48, Poo33, Vol38, Wri21]. Forest [Bat49, Com45]. Forestry [Bla45a, Chi45]. Forests [Zon45]. Forgotten [Dub10]. Form [Dav27, Ges41, Jen08, MB15, Mic06, San39]. Formation [Gil42, Jep40, Mac33b, See06, See07a, See09, Sen30, Set26, Ste11a, Ste11b, Ste12, Ste13, dTdC36]. Formations [HS04]. Formed [Rod15]. Former [Bal12]. Forms [Sco30, TE33]. Formulæ [Ken36]. Fortress [Few32]. Forward [Bai03, Mum40]. Fossil [Bry26, Bry35, CL36, Gre36a, GH39, Hay11, Jep49, Mat19, McN41, MP44, OC31, Pal28, Ste16, Ste17, Ste18, Ste20]. Fossils [Jep30b, Ric33]. Foundation [dMS27]. Founder [Jew41]. Four [Lev26, Mil07]. Fourteenth [Dan11, Put47]. Fourth [Bru24]. Fowl [Pea16, Pea17]. Fragility [HC40]. Fragments [San35]. Fraley [Pat01]. France [BLV47, Jen48]. Francis [Bru32]. Franciscana [Atk18c, Dav37b]. Francisco [Cra07]. Francois [Tru37]. Franklin [Lin48, Sco32, Abb06, All41, Bac36, Chi48, Com46, Dav06, Dua27, DM47, Har40, Jen48, Nol44, Roe48, Ros01, Ros03a, Sio49]. Frederick [Tho11, Bar01, Pat01]. Free [Bar07b, Ros40, WB43]. Freedmen [Wes48]. Freedom [Bou43, Dal47]. Freeman [Cox48]. Fremiet [KR40]. French [McK40, Mur44, Ros07, Ros11]. Frequencies [Ada37c, KV19, Ste37]. Fresh [Ort09, Ort20, Sta03]. Fresh-Water [Ort09, Ort20, Sta03]. Freshwater [Ort02, Ort13, Ort18, Ric37]. Friends [Ros11]. Fringes [Bru13]. Frog [Rug39, RE40, RM43, Spe40, Spe42]. Front [Ano01d, Ano05e, Ano06d, Ano07e, Ano08d, Ano09d, Ano10e, Ano11g, Ano12g, Ano13g, Ano14d, Ano15f, Ano16k, Ano16l, Ano17h, Ano17i, Ano18h, Ano19h, Ano20g, Ano21e, Ano22f, Ano23g, Ano23h, Ano24d, Ano25d, Ano25e, Ano26h, Ano26i, Ano27d, Ano30e, Ano35k, Ano39j, Ano39k, Ano40l, Ano41i, Ano41j, Ano42f, Ano43f, Ano43g, Ano45e, Ano46f, Ano46g, Ano47k, Ano47l, Ano47m, Ano47n, Ano47o, Ano48m, Ano48n, Ano48o, Ano48p, Ano48q, Ano48r, Ano49o, Ano49p, Ano49q, Ano49r, Ano49s, Ano49t, Ano49u]. Frost [Sch36]. Fruited [Tre15]. Fuels [Ste16, Ste17, Ste18, Ste20]. Fullerton [Hir09, Lin33]. Fulrad [Cro49]. Fumes [Han23]. Function [HK37, NS37, Pat13, dMS28]. Functional [Cog26]. Functions [Cri13, Hue34, Lam03, Lam04, Nip19]. Fund [Sin21a, Sin21b, Sin21c]. Fundamental [Ehr16, Mon28, Rog12, Tre16]. Fundulus [Par35]. Fungicide [Kra06]. Furness [Sha35]. Further [Car08, CGTR32, Dem36, Fal14, Gre35, Har33, Mat09b, Mat10, Mül26b, MPG26, See08, See10b, See13, PS40]. Future [Hrd29, Sha39, Tho49].

Galaxies [SW49]. Galaxy [Pla30]. Galerula [Atk18a]. Gall [McM23, WB43]. Gallic [Car09]. Gamma [Eva37]. Garden [Tru36]. Garigue [Har26]. Garment [Sta48]. Gaseous [BF27, Low11, Wri20, Zel31]. Gases [Com22, Fen48, Tro02]. Gaston [Jas16]. Gastric [Reh16]. Gates [Boy43]. Geminid [Whi47]. Geminorum} [Bar12]. Gene [Bea46]. Genealogy [Ano41k]. Genera [Eig21b, Gro00b, Gro02a, Hay03, Mai34]. General [Ano00f, Ano02e, Ano03e, Ano04e, Ano05f, Ano06e, Ano07f, Bau15, Com11, Hun32, Sip45a, Ano00f]. Generalizations [Lov09]. Generally [Mer42]. Generated [Mil07, Mil10a]. Generation [Bru15, Bru17, Bru26, Bru27, Bru28, Con34, Har33]. Generations [Dav24]. Generic [Tre29]. Genesis [Ges41, Hal49, Smy27, Jas04]. Genetic [Ger43, Gow34]. Genetics [Jen41]. Genic [Whi35]. Genth [Bar01]. Genus [Atk18a, Bro03, Cal34, Jep34, McC26, NW38, Ort05, OM19, Tre28a]. Geochemical [Cla12]. Geoffrey [Tat38]. Geographic [Far17, Joh20]. Geographical [Bla45b, Ort02, Sip45b]. Geography [Dav02, Ort02, Pen27, Tre30]. Geologic [Cha06a]. Geological [Cha22, Ric37, BBC^{+}38, Sin21a]. Geology [Ber33, Ber39b, Cha15, Cla06, Cra07, Dav10c, Fie38, Hec38, Hob09, JB13, Kno45a, Sop16, Wad45a, Wil42a]. Geomagnetic [McC41]. Geomagnetism [Ber41, Fle41, Gar41, Gis41]. Geophysical [Hay15]. Geophysics [Fie38, Mac47]. George [Jas21, Bal18, Tho11]. Georgian [Wer43]. Germ [Guy23, Sto23]. Germ-Cell [Guy23]. Germ-Plasm [Sto23]. German [Fra08, Fra27, Ros00]. Germanic [Fra08, Mül26b]. Germanium [MPG26]. Germans [Pri31]. Germany [Sch15b]. Germinal [Shu10]. Ghazzāli [Far39]. Gibbs [Mor10]. Gilman [And09]. Gippsland [Mat07c]. Girdle [Gre35]. Girls [Ben24]. Given [Mil11]. Gives [Mer05]. Giving [AB18]. Glacial [Hob15, Hob21, Hob43, Lev26, Low00]. Glaciers [Gou40, Hob43]. Glaciology [Kno45b]. Glass [Bak40b]. Glaucophane [Smi06]. Glaucophane-Bearing [Smi06]. Globe [Hob15]. Globular [See12b]. Glozel [Rie30]. God [Wes48]. Gold [Kem32]. Golgotha [Hau20]. Goodwin [Smy13]. Gortyna [Gro00b]. Gosiute [Cha13]. Got [Ros47b]. Gottschee [Pri31]. Governing [Sip45a]. Government [Dod33, Joh34, Lik48, Tho46, You33]. Governmental [Fal46, Mac48, Mil46, Oli05, Rad34]. Governments [Bla45a]. Governor [Bal15]. Grahamellae [Tyz42]. Grande [JB13, Sop16]. Grandiflora [Dav26, Dav27]. Granites [Bra13]. Grant [Sin21a, Sin21b, Sin21c]. Granting [Ano30c]. Graphical [Nip19]. Grasping [Hoo38]. Grass [Art25, Har06]. Grass-Killing [Har06]. Grating [Bar09c, Bar11b, Mic15, Ric12]. Grave [Ash03, Bon41]. Grave-Stelae [Bon41]. Gravimetric [Ewi38]. Gravitation [Bru14, Bru21, Bru22a, Bru23, Bru24, Bru25, Bru26, Bru28, Bru29]. Gravitational [Bru28]. Gravity [Hes38, Rei40]. Gray [Dav34]. Great [Cre43, Hay24b, Ash19, Bie49, Dub10, Hat02]. Greatest [Mer05, Sch48]. Greek [Luc18, Mar28]. Greeks [Noc42]. Green [Cra39, Rod15]. Greene [Ano00f, Roe48]. Greenland [Lee40]. Gregorian [Ken35b]. Gresham [Bal08b]. Ground [Ber45]. Groundplans [Bet46]. Group [Bon41, CGTR32, Du 48, Mil10b, Mil11, Pac03, Tre11, Tre21]. Groups [Mil07, Mil10a, Mil10b]. Growing [You33]. Growth [Ano34l, Cal29, Cal34, Car27, Cog26, Dar47, Dav31, Dav37a, Ges41, Ham29, Ham30, Hob23, Hob44, Hrd36, Jen08, Jon37, Lan48c, Lin40, Liv13, Loe08, MS17, Mac19, Mac21b, Mac22, Mac30b, Mac39, Not39, Pea24, Ree38, Tho39]. Growth-Control [Mac39]. Guadaloupe [Hay24b]. Guatemala [Hun13]. Gudermannians [Ken29]. Guide [War28]. Guide-Book [War28]. Gulf [Ano11f, Bac36, Ber11]. Gun [Mun09, Web19]. Gundungurra [Mat01b]. Gusts [Wil17].

H [Ano30b]. Habit [Whe18]. Habitat [Chi47]. Habitus [Gre36a]. Habrobracon [Whi39b]. Haemobartonellae [Tyz42]. Hæmolysis [AB02]. Hair [Bro42b]. Half [Gre34b, Wil10]. Hall [Ada15, Ekb40, Gre35]. Halley [Low11]. Hamites [Jas04]. Hampton [KW26]. Harbors [Hau03]. Hard [Bog49]. Hardened [Bru15, Bru17]. Hardy [DH47]. Harmonization [Hun44]. Harmony [Blo44b]. Harrison [Vol38]. Hartshorne [Mor00]. Harvard [Fra08, Hoo26, Hoo35a]. Harvest [Jor06]. Hays [KW26]. Haze [Bar11a, Bar19]. Head [Dav40]. Health [Dra46, Fal46, Fox33, Hei26, Mil46, Mud47, Row42, Shr46, Sto46, Wil08b]. Hearing [Par18]. Heart [BL36a, Coh14]. Heart-Beat [Coh14]. Heat [AD41, BM38, BM32, Bro49, Bru15, Bru17, Bru18, Bru26, Bru27, Bru28, CBB36, Har33, Kun43, Mac02, McE38]. Heated [KS14]. Heavenly [See15]. Heavier [Bid28]. Heavier-Than-Air [Bid28]. Heaviside [Ada23]. Heights [Mit30, Oli33]. Held [Ano30b, Ano30c, See07b]. Hemoglobins [RB08]. Hemolysis [HC40, Jac31b]. Henderson [Bur23]. Henneguyi [KR40, Ros40]. Henri [Veb12]. Henricus [Jon11]. Henry [Fle35, RW07, Sch46, Che11, CDG^{+}12, Gre36b, Kee11, Mor00, Wal48]. Hepatoscopy [Jas08a]. Her [Hay24b]. Herculis [Bob36, Bob35]. Hereditary [Mor15]. Heredity [Bro06, God12, Hrd35, Jen08, Mon04b, Wee12, Whi17, Jen08]. Heroes [Kra46]. Heroic [Kra46]. Herschel [See12b]. Heterocaryotic [DSA45]. Heteroploidy [CH49]. Hewettite [HMW14]. Hierosolymam [Dav39b]. High [Ada37c, Bea36, Har48b, Kra47, See12b, Sen30, Tro02, Whi11, Whi20]. Higher [Ayd30, Ros13]. Highest [Hun13]. Highly [Bar09a]. Himalaya [See13]. Hindu [Blo17, Edg36, Hop10]. Hints [Pac03]. Hiss [Tit14b]. Histological [Abb36, Swi23]. Historia [Tat38]. Historical [Bar13, Gro00b, Hew04, Sco23, Shr46, Ste11a, Vai43]. Histories [Fis12, Spe42]. Historiography [Lov38, Sch44b, Sch46, Shr43]. History [Alb30, Bro42a, Che27, Chi45, Cob15, Col43c, Cri31, Dav26, Dub10, Fen49a, Fen49b, Fet43, Hay40, Hrd43, Hun32, Hyd35, Jac42, Kra46, Luc18, Nic02, Put47, Ric37, Rob11, Ros00, Ros03b, See09, Sho27, Sta48, OM19, Osb23]. Hittite [Bar26]. Hoff [Jon11]. Hohokam [Hau43]. Holway [Art25]. Homes [Ree36]. Homolka [Few32]. Honduras [Ric38]. Honest [Bog49]. Honorable [Ano30c]. Hookeri [Dav34]. Hopes [Whe31]. Hoplophoneus [Sin21b]. Horatio [Boy43]. Horizons [Hib43, Lew43]. Hormonal [Rid35]. Hormones [RD45]. Horse [Dum48, Osb04]. Horse-Sacrifice [Dum48]. Hospital [Hun32]. Host [Ree38, Wen35]. Host-Parasite [Wen35]. Hosts [Wen35]. Hotel [Ano30b, Ano30c]. Hour [Ste37]. House [Wri21]. Household [BJ19]. Howard [Mag37, Tre37]. Howard-Smithiï [Tre37]. Huacos [Ash03]. Hudson [Dav10b]. Huebnerian [Gro02a]. Hughes [Cha49]. Hugo [Cle36]. Hulbert [MG49]. Human [Ben37, BB30, BLV47, Bro44b, Cal42, Chi47, Com46, Dav39a, Dav44, Det37, Fra45, Hel28, Hou32, Hrd35, Hrd37, Jor06, Lew45, Mar48, MP44, Osb28, Rid48, Ros47a, Ste02]. Humanism [Ran32]. Humanistic [Lel32]. Humanitarian [Moo47]. Humanities [Lan48a]. Hume [WPL47]. Humming [KU08]. Hundred [Ano30c]. Hungarian [Pri35]. Hunting [SE42]. Huron [Bar49a]. Hyænodontidæ [Mat05b]. Hybrid [Dav37b, Tre17]. Hybridization [Shu10]. Hybrids [Atk18c]. Hydration [Mac19]. Hydrogen [May19, MPG26, TE33]. Hydrogen-Ion [May19]. Hydrographic [Sei40]. Hydrography [SP46]. Hydroid [Bro03]. Hydrometric [Har11]. Hydrophone [Hay20a]. Hydrostatic [RM43]. Hydrous [HMW14]. Hydroxide [Mül36]. Hygiene [Ove46]. Hyla [Rug41]. Hymenoptera [Whi35]. Hypochlorites [Mas09]. Hypotheses [Pac04b]. Hypothesis [Dav35, Par35]. Hypotonic [HC40]. Hyracodons [Sin22a]. Hyrax [Fox33]. Hysteria [Hau21].

I. [KW26]. Ice [Bal21, Hob10, Wad45c]. Idea [Del47, Jon42]. Ideas [Lov38]. Identical [BB30, Ken25, Ken30]. Identification [Bel49]. Idiom [Pri27]. Idvorsky [Ken35a]. Igneous [Bru26, Dal25]. Ignorance [Bog49]. Ihyā’ [Far39]. II [Ada36a, Ada36c, Bru21, Bru22a, Bru23, BR34, Cra40, Gil42, Jen08, Jen41, RE40, SP46, Ste11b, Ste17, Whi40a, Whi17, Wil17]. III [Cou45b, Bru17, Bru26, BR34, Fra45, Per45, Sip45b, Ste12, Ste18, Whi40b]. Illusion [Sch48]. Illusions [Mor37]. Illustrated [Has00, Par15, Sch15b, See07a]. Image [Par22]. Images [TP08]. Imbibition [MS17]. Immunization [Dye44]. Impact [Bin48, Cur49, Kuz48]. Impairment [Bru28]. Impatiens [Smi34]. Impedance [KS27]. Imperceptible [TP07]. Implications [Bou43, Del34, Not43, Pat28, Rid35, Vin46]. Importance [Fie38, War25]. Important [Gal42, MW37]. Impressed [KT15]. Improvement [Hau03]. Improvements [Goo48, Nor32, Poo33]. Improving [Hau01]. Impulse [Bea36, Jon46]. Inch [Mic15, Tho30]. Incidence [CH49]. Including [Bai47, Mac36b, Tru29]. Inclusion [Ric03]. Income [Blo44a, Kuz44]. Incoming [MN40]. Increase [Ham29, Ham30, Mac23, RD45, See11c]. Increased [CBB36, HC40]. Independence [Hay00b]. Index [Ano35l, Kop41, SK37]. Index-Method [SK37]. India [Bro31, Bro47, dTdC36, dT42]. Indian [Bau12, Bon34, Eme40, Far17, Voe49, dCW23, Wes44, Wis42]. Indians [Cha13, Kno40, Smi38]. Indicated [LL40, Rei15]. Indicating [Ric33, Rol22]. Indices [Rog12]. Indies [Hes38, Mat19, Ols41]. Indirect [Edg36]. Individual [Boa43, Dav31, Spe40]. Inductor [HB22]. Industrial [Bak48, Wet34]. Industries [OC31]. Industry [Dil47]. Inédits [DH47].

Inexhaustible [Dut44]. Infant [Ges41]. Infection [Kne37, Tyz32]. Infections [Kee44, Kee48]. Infectious [Ber37, Dub44]. Inferred [See09, See12b]. Infinite [Lam08, Mac02]. Infirmities [Cri33]. Infliction [Spi08a]. Influence [AB02, Ben10, BLV47, Cha06a, Ful09, Goo09, Har11, Hei18, KS14, Lut48, Ort13, Ree38, Row29, Tay39, Wil08b]. Influences [Guy23]. Information [Ano00x, Ano00y, Ano01u, Ano01v, Ano02r, Ano02s, Ano03t, Ano03u, Ano04s, Ano04t, Ano05q, Ano05r, Ano06r, Ano06s, Ano07r, Ano07s, Ano08o, Ano08p, Ano09f, Ano09g, Ano10g, Ano10h, Ano11i, Ano11j, Ano12i, Ano12j, Ano13i, Ano13j, Ano14f, Ano14g, Ano15h, Ano15i, Ano16n, Ano16o, Ano17k, Ano17l, Ano18j, Ano18k, Ano19j, Ano19k, Ano20i, Ano20j, Ano21g, Ano21h, Ano22g, Ano22h, Ano23j, Ano23k, Ano24f, Ano24g, Ano25g, Ano25h, Ano26k, Ano26l, Ano26m, Ano27f, Ano28f, Ano28g, Ano29f, Ano29g, Ano31g, Ano31h, Ano32i, Ano32j, Ano33g, Ano33h, Ano34n, Ano34o, Ano34p, Ano34q, Ano35m, Ano35n, Ano36h, Ano36i, Ano37h, Ano37i, Ano37j, Ano38g, Ano38h, Ano38i, Ano39l, Ano39m, Ano39n, Ano39o, Ano40n, Ano40o, Ano40p, Ano40q]. Information [Ano41l, Ano41m, Ano41n, Ano42g, Ano42h, Ano43h, Ano43i, Ano44i, Ano44j, Ano44k, Ano45f, Ano45g, Ano46h, Ano46i, Ano47p, Ano47q, Ano48s, Ano49v]. Infra [MG49, Ran23]. Infra-Red [MG49, Ran23]. Infrared [Bad36]. Inhabitants [Spa37]. Inheritance [All19, All24, Cou16, Dav08, DSA45, Han23, Har18, RP06, Sto23, Whi17]. Inherited [Det23]. Initiation [Mat00d]. Injected [Rum16]. Injurious [Spe42]. Injury [Heb42, Spe40]. Inland [Hob10]. Inland-Ice [Hob10]. Insect [Spa37]. Insectivorous [Har25]. Insects [Car17, Whe18]. Insidious [Ric03]. Instinct [Rid35]. Institute [All41]. Institution [Mor43a, WPL47]. Institutions [Ano27b, Joh39]. Instrument [Sch17]. Instruments [McC41]. Insurance [Hue32, Lin18]. Integumentary [Par37]. Intellectual [Mac46]. Intelligence [Heb42, Lan48b]. Intelligent [Wil10]. Intensity [LHYH36]. Interchange [Mil47]. Intercontinental [Eme17]. Interest [Hay24b]. Interesting [Smi26b]. Interests [Tho37b]. Interference [Bar11b]. Interferometer [Pea21]. Interglacial [Ric33]. Interior [Cha15, Rei15, Sch15a]. Intermingling [Dav17]. Internal [Bru26, Cha39, Du 31, dMS28]. International [Bal07, Bal16, Boh47, Bro46, Bro16, But44, Con47, Coo46, Dub10, Eag30, Fle47, Hun47, Ken48, Ken24b, Lei47, Lin38, Mac48, Nic48, Pat32, Pat34b, Pic06, Sha47, Sho38, Sta47, Sve47, Tow16, Vin46, Whi40c, Wil43a, Lew42, Sta48]. Interplay [Bar29]. Interpretation [Bar40c, Bro44a, Cri30, Eas15, Edg38, Few35]. Interrelations [Son38, Ste16, Ste17, Ste18, Ste20]. Interstellar [Dun39]. Interviewing [Har48a]. Intestinal [Kra05]. Intoxication [AB02]. Intra [MW25]. Intra-Thoracic [MW25]. Intrinsic [Not39]. Introduced [Mer37]. Introduction [Con43, Dar40a, Wad45b]. Introductory [Cob15, Con42, Mil30]. Intrusive [Ste45b]. Intrusives [Ste45c]. Invariance [Wig49]. Investigation [CTR29, Har11, KW44, Ric03, Tho37b, Sin21a, Sin21b, Sin21c]. Investigations [Mer03, Mor43a, Rei15, Sta37]. Involuntary [Kis39]. Involved [DSA45]. Involving [Dav37b]. Iodamœba [Wen37]. Ion [May19]. Ionic [Mül36]. Ionization [Com22]. Ionized [Bar09a]. Ionosphere [Dar40a]. Ionospheric [Ber41]. Ions [Bea36, Tru30, Zel31]. Irish [BS49]. Iron [Hey10, Sau27]. Iroquois [Fen49b, Wal49a]. Irradiation [Rug39]. Irrigator [LL40]. Island [Hei47, Hes38, Ric38, Ric39, Ric37]. Islands [Hof40, Ste45c, Ste47, Par15, Spi08b]. Isostasy [Rei11]. Isotope [Ble36]. Isotopes [Ble37, TE33, Tay48, Ure46]. Issue [Bro47]. Issues [Sli48]. Isthmian [Tho07]. Italy [Pac46]. Itinere [Dav39b]. Itself [Mil10a]. IV [Bru28, Ste13, Ste20, Whi47].

J. [Cha06b]. Jacobus [Jon11]. James [Bra42, Hof40, Pac35]. January [Ano00e, Ano00o, Ano00n, Ano01l, Ano01k, Ano02l, Ano02k, Ano03l, Ano03k, Ano04k, Ano04l, Ano05k, Ano05j, Ano06j, Ano06i, Ano07l, Ano07k, Ano08j, Ano08i]. Japanese [Dub10, Lym12]. Jastrow [CM21]. Jeannette [Ell40]. Jefferson [Bec43, Bet44, Chi43b, Con43, Kim43, Kim44, Mor43b, Sha43, Tru29, Tru36, Wil43b, Wri43]. Jelly [Bro03]. Jelly-Fishes [Bro03]. Jeremiah [Cop48]. Jersey [Har09, Ric33]. Jews [Ros47b]. Jivaro [Sti33]. John [Ano30b, Ano30c, Col41b, Mil35, Tru29]. Jointing [Ehr16]. Jonah [Hau07b]. Jordão [Har29a]. Joseph [Bro42a, Jas21, Mor09, Pet03]. Journey [Ron45]. Jr [Bea46, But44, Lin33]. Jr. [CM21]. Judaism [Hau09b]. Judges [Cha47]. Judgment [TP07]. Judicial [Tow16]. Judiciary [Wri21]. juglandis [Whi39b]. Jules [Veb12]. July [Lin33]. June [And19, Ano00h, Cam19, Lam19, Mil19a]. Junto [Car27]. Jurassic [Cri31]. Juries [Dub14]. Jurisprudence [Car17]. Jurors [Dub14]. Justices [Put47]. Jutland [Pri27].

K.C.B. [Ric17]. Kānda [Dum48]. Kane [Ekb40]. Kansan [Wil20]. Kansas [McC26]. Kashmir [dT42]. Keen [Dar33]. Keep [Joh32a]. Keith [Con08]. Killing [Har06]. Kindred [See06]. Kinetic [Bru14, Bru21, Bru22a, Bru23, Bru26, Bru28, Cri14]. Kinetics [HS30]. King [Ste45b, Wad45a]. Kingdom [Moo08]. Kites [Bla09]. Knowledge [But48, Dar33, Dut44, Hes26, Kep37, Mer26, Osb04, Sch15a, Sha19, Whi17]. Known [Bry36, Car17, Wil02]. Kodiak [Hei47]. Kupferschiefer [Sch15b].

L [KW26]. Labial [NS37]. Labile [MCPL37]. Labor [But44]. Laboratory [Hin11, KN20]. Labrador [SE42]. Ladie [Sta48]. Lafayette [Ben36, Nol34]. Lagrange [Sar44]. Lake [Jen36, Jen38, RW09, Ber39b]. Lakes [BJ27]. Lamarck [Dav27]. Lamarckiana [Atk18c, Dav24, Dav26, Dav27]. Lambertians [Ken29]. Land [Hob40, Joe40, Low45, Pas45, Rei20, Ric37, Ric38, Ric39, Ron45, Wad45a, Bal09a, Ekl45, Sau40, Ste34a]. Landing [Ols41]. Lands [Bar40b, Ste45b]. Langley [Wal26, WPL47]. Language [Bar49a, Ken24b, Mat01b, Mat07a, Mat07c, Pri14, Pri22]. Languages [Mat03a, Mat03b, Mat04, Mat07d]. Laplace [See15, Bro04]. Large [BF27, Tay41, Tre15]. Larvæ [Whe18]. Larval [Cal34]. Last [Mil26, Sau48]. Late [Heb42, Lew43, Few32]. Later [dTdC36]. Latin [Rol22, Bog38]. Latitude [Ogb07, Sch15a]. Lava [See08]. Law [Ben38, Car27, Cha47, Eag30, Fra31a, Has00, Mon28, Rog12, Tow16, Bal08b, Put47]. Lawn [LL40]. Laws [Rog12, WPL47]. Lawyer [Mor43b]. Layer [Par39, Sin21c]. Layers [Mac24]. Lea [Che11, Kee11]. Leaf [Har08, Har09, Mac33b]. Leaf-Surface [Mac33b]. Leafhoppers [Bla44]. League [Hei26]. Leakage [See08]. Learned [Lel32]. Learning [Ano27b, Con44, Tol41]. Least [Kop43a, Mer03]. Leaves [SM39, Tre26]. Lebanon [Tra06]. Lecture [Ang42, Bea46, But44, Com46, Con34, Swa35]. Ledgers [Bro35]. Left [Rid24]. Legal [Bal16, Ful36, Ste02]. Legendary [Fis12]. Legislative [Wil16]. Length [Rol22]. Lengthened [Mac30b]. Lengthening [Fra31b]. Lens [Bar07a]. Leonid [Oli33]. Leonis [McD15]. Leopard [RE40]. Lepidopterous [Ano00g]. Leprosy [Hei32]. Leptauchenia [Sin10b]. Lesions [Swi23]. Lesley [Cha06b]. Leslie [Goo31]. Lessening [Du 31]. Lessons [But44]. Lesueur [Chi49]. Letters [Mil11, Nic44b]. Lettres [BLV47]. Level [JW27]. Lewis [Dav06, Tru28]. Lias [Sch15b]. Liberation [SS38]. Liberty [Cus48, You34]. Libraries [Sio49]. Library [Ano00f, Ros03b]. Lick [Cam19]. Life [BF27, Che11, Fra31b, HK37, Hen23, Hrd36, Kim44, Lin33, Mye43, Per40, Sin25b, Tru29, Wis27]. Light [Alb30, AD41, BJ19, Dic27, Her40, Hrd29, Hrd32, Kop42, Kop43b, Mac34, Nic44b, Rea49, Sha49, SW49, Wri19]. Lighter [Ros28]. Lighter-Than-Air [Ros28]. Lighting [FR18]. Lightning [Hum36]. Like [Sco30]. Limb [Kop41]. Limitation [Der29]. Limitations [Par41]. Limiting [You48]. Lindeniopiper [Tre29]. Line [Dav24, KN20, Lam05, RP06, Cop49]. Linear [Ken09, Lam11, Whi40a]. Lines [Osb34]. Linguistic [Dut44, Kur34, Mac36b]. Linguistics [Edg43, Lyd48, Voe49]. Linn [Nic02]. Linn. [Gro02a]. Linnæus [Dav26]. Linnean [Di 03]. Lipoids [Mac22, Mac28]. Liquor [Hep18]. List [MPS20]. Liston [Kyt49]. Literary [Edg36]. Literature [Hew04, Jon46, Kra42, Kra46, Sch29, Sch39, Ska46]. Lithology [Wan22]. Lithosphere [Ehr16]. Litter [RP06]. Little [Car17, Cou45b, Fra45, Per45, Sip45b]. Liver [HM38]. Living [Bec43, Cro32, Der31, Hei30, Hei34, Jac31a, Mac23, Moo08, Mun11, Spe40, Spe42]. Lizards [Gil42]. LL.D [Doo10]. LL.D. [And09, Hir09, Mor09, Pat01, Tho11]. Loan [Pri35]. Local [Bla45a, Dod33, New44]. Locality [Dou01]. Locke [BLV47]. Logan [Bra42]. Logic [Sin25b]. Logical [Bro06]. London [Hay24b, Rea49]. Long [Ada10, BM38, Det23, Dix34, KN20, Kuz48, Oli42, Oli47, War26]. Long-Continued [Det23]. Long-Range [War26]. Longitude [Tuc22]. Looking [Kid40, Mum40]. Loss [BJ19]. Louis [Hen23]. Loureiro [Mer33]. Low [Con38, GG38]. Lowell [Lam19, Sli38a]. Lower [Ber14, Bry32, Dou02, Fox29, Hel28, Bry35, Few37, Gaz36, Jep30b, Sin15, Sin28]. Luciopimelodinæ [Dri19]. Lumbrici [KH38]. Luminous [Bar11a, Bar19]. Lusitanicum [Har25]. Lytechinus [Mac36b].

M [CDG^{+}12]. M.D [Hir09, Tho11]. M.K.S. [Ken36]. Macchia [Har26]. Machine [Bid28, Pat31]. Machines [Ros28]. Maclaurin [Lam03]. Maclure [Moo47]. Macrochelys [Hay11]. Macroscopic [RM43]. Maculosus [Cra40]. Made [Ano30b, Ano30c, Che27, Joh28, KBC45]. Magellanic [SN48]. Magic [Hop10]. Magician [McG43]. Magnetic [Bau10, Bau12, Bru18, Hal14, Hal41, Hec41, Joh41, Ken33, Ken36]. Magnetism [Fle38, Hey41, McN41, Rei40]. Magnitude [Tuc33]. Main [Hrd26a, Mon04b, Ron45]. Maine [Pri14]. Maintenance [Fox33]. Major [GH39, See11c, Ano00f]. Major-General [Ano00f]. Make [Cha47]. Making [Bur12, Few44, Fis12]. Male [Rug41]. Malformation [Oli02]. Malignant [Car29]. Mammal [Sim40, Sim43]. Mammal-Bearing [Sim40, Sim43]. Mammalian [Bro42b, Loe11, PS40]. Mammals [Fox29, Gre36a, MP44, RD45, Tyz42]. Man [Bal17, Bar17, Cri33, Fox29, Gre27, Gre28a, Gre28b, Gre34b, Hal24, Hel28, Hoo38, How36, Mac36a, Osb27, OC31, Rid29, Wil10, Col43b, Hrd29, Hrd32, Mac34, Sch48]. Manganese [Bru18, Dal15]. Mangrove [Bow17]. Manhattan [Dav10b]. Manna [Hau22]. Manual [Pri16]. Manuscript [WPL47]. Manuscripts [Car49]. Map [Bog38, Dav38]. Maracaibo [Eig21b]. Marble [Rob40]. March [Ano00q, Ano00p, Ano01m, Ano01n, Ano02m, Ano03n, Ano03m, Ano04n, Ano04m, Ano05l, Ano06k, Ano07m, Ano08k]. Marie [Pas45, Sau40, Ste45b]. Marine [AH13, Ewi38, HS04, HC25, Her37, Mye43, New48, Oli04, Ric33, Rud38, SK36]. Market [Hue34]. Markings [Emm32, Pac04a, Pac04b]. Marriage [San39]. Mars [Low00, Low01, Low03]. Marsh [Har11, Har16, Sin21a, Sin21b, Sin21c]. Marsupial [Sin05]. Martinichthys [McC26]. Maryland [CL36, Cop49]. Marylandica [Ber39a]. Mason [Cop48, Rob49b]. Maspero [Jas16]. Mass [Ble36, Dav31, Dem35, Dem36, See11c]. Mass-Spectrograph [Ble36]. Masses [Gap40, Mac21a]. Mastodon [Bel49]. Mastodons [Cas39]. Mateo [HA04]. Material [Cop48, Dun39, Mac24, Mac28, Mor15, OM19, Pri19, Pri35, Sch45a, Whi31]. Materials [Bee39, Fen49a, Lin48, Tau49]. Maternal [Rid35]. Mathematical [Bra42, Wey49]. Mating [Son38]. Matter [Ano01d, Ano03d, Ano05d, Ano05e, Ano06d, Ano07e, Ano08d, Ano09d, Ano10e, Ano11g, Ano12g, Ano13g, Ano14d, Ano15f, Ano16k, Ano16l, Ano17h, Ano17i, Ano18h, Ano19h, Ano20g, Ano21e, Ano22f, Ano23g, Ano23h, Ano24d, Ano25d, Ano25e, Ano26g, Ano26h, Ano26i, Ano27c, Ano27d, Ano30d, Ano30e, Ano35k, Ano39j, Ano39k, Ano40l, Ano41i, Ano41j, Ano42f, Ano43f, Ano43g, Ano45e, Ano46f, Ano46g, Ano47f, Ano47g, Ano47h, Ano47i, Ano47j, Ano47k, Ano47l, Ano47m, Ano47n, Ano47o, Ano48g, Ano48h, Ano48i, Ano48j, Ano48k, Ano48l, Ano48m, Ano48n, Ano48o, Ano48p, Ano48q, Ano48r, Ano49h, Ano49i, Ano49j, Ano49k, Ano49l, Ano49m, Ano49n, Ano49o, Ano49p, Ano49q, Ano49r, Ano49s, Ano49t, Ano49u, Der31, Nip13]. May [Blo44a, Ano00s, Ano00r, Ano01p, Ano01o, Ano02o, Ano02n, Ano03o, Ano03p, Ano04p, Ano04o, Ano05n, Ano05m, Ano06m, Ano06l, Ano07o, Ano07n].

Maya [Lun33, Mor43a]. Mays [Abb36]. McClenahan [Mag37]. McCrae [Ste36a]. McMath [MG49]. Mean [Bur21, Joh29a, See11c]. Meaning [Fra44, Whe18]. Means [Bla09, Sch13a]. Measles [WBAW39]. Measurement [BM38, Bar09a, Mas07, Pea21]. Measurements [Bee39, KV19, KN20, SP45]. Measuring [JR40, Nip06, OC31]. Mechanical [Dur17]. Mechanics [McE38, Mer03, Rid29]. Mechanism [Coh14, HC40, Har48b, Moo36, Par16, Par39, Sho48, Sin25b, Smi17, Tay26]. Medal [Com46]. Mediæval [Cra17, Car17]. Median [Pac04a]. Medical [CF22, Gre44, Shr47]. Medicine [Dye44, Mud47, Pac42, dMS27]. Medico [Ste02]. Medico-Legal [Ste02]. Medieval [Fra08]. Mediterranean [Din43, Har26]. Medium [Bar09a, Ken09, Mac02, See15]. Meeting [Ano00e, Ano00j, Ano00i, Ano00k, Ano00m, Ano00l, Ano00o, Ano00n, Ano00h, Ano00q, Ano00p, Ano00s, Ano00r, Ano00u, Ano00t, Ano00w, Ano00v, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01p, Ano01o, Ano01q, Ano01r, Ano01t, Ano01s, Ano01e, Ano02e, Ano02f, Ano02h, Ano02g, Ano02j, Ano02i, Ano02l, Ano02k, Ano02m, Ano02o, Ano02n, Ano02p, Ano02q, Ano03e, Ano03f, Ano03h, Ano03g, Ano03j, Ano03i, Ano03l, Ano03k, Ano03n, Ano03m, Ano03o, Ano03p, Ano03q, Ano03s, Ano03r, Ano04e, Ano04f, Ano04h, Ano04g, Ano04j, Ano04i, Ano04k, Ano04l, Ano04n, Ano04m, Ano04p, Ano04o, Ano04q, Ano04r, Ano05f, Ano05h, Ano05g, Ano05i, Ano05k, Ano05j, Ano05l, Ano05n, Ano05m, Ano05p]. Meeting [Ano05o, Ano06e, Ano06f, Ano06g, Ano06h, Ano06j, Ano06i, Ano06k, Ano06m, Ano06l, Ano06o, Ano06n, Ano06q, Ano06p, Ano07f, Ano07g, Ano07i, Ano07h, Ano07j, Ano07l, Ano07k, Ano07m, Ano07o, Ano07n, Ano07p, Ano07q, Ano08e, Ano08g, Ano08f, Ano08h, Ano08j, Ano08i, Ano08k, Ano08m, Ano08l, Ano08n, Ano37g, Kot45]. Meetings [Ano26j, Ano28e, Ano29e, Ano30f, Ano31f, Ano32a, Ano33a, Ano34a, Ano35a, Ano36a]. Megaliths [dT42]. Melanogaster [How41]. Melanophore [Par38, Par41]. Melanophores [Par34, Par41]. Melchior [Ste45c, Ste47]. Member [Mat05b]. Members [Bal12, Hun32, Ros07]. Memorial [Bea46, But44, Con34, Lin33, Swa35]. Memory [Cri30]. Men [Ben18, Mil47]. Menander [Rob40]. Mendel [Ben36]. Mendelian [Dav08, Wee12]. Mendenhall [Swa36]. Mental [Ben10, Ful09, Ove46, Tho41]. Mentality [Fur16]. Merchants [Nic44b]. Mercury [KS05, KS06]. Mery [Mim12, Ros11]. Mesohippus [Sin25a]. Mesopotamia [Bar29]. Meta [Smi26b]. Meta-Tungstates [Smi26b]. Metabolic [Rid27]. Metabolism [Ben24, BF27, Ben37, Mor26, Wil05, Wil08b]. Metahewettite [HMW14]. Metal [MPG26]. Metallic [MPG26, Nic16]. Metals [Hou39, KS06, Mas07, Mül26a, Wri19]. Metamorphism [Aga23]. Metazoa [Mon08]. Meteor [Oli36, Oli42, Oli47, Whi38, Whi40a, Whi40b, Whi47, Wil39c, Mag10]. Meteorites [Bre04, Foo13, Mer26]. Meteoroid [Wat39b]. Meteorological [Mil19b]. Meteorologist [Abb06]. Meteorology [Bla28, Hum42, Rei40]. Meteors [Oli33, Wat39d, Wil39b]. Meter [McC05]. Method [Bar09c, Bar09b, Cla06, Hec38, Ken28, Kop43a, Lan41, OC31, Par41, Pea21, SK37, Sny02, WA19, Wee12]. Methodology [Mar48]. Methods [Ada23, Bar07b, Bea36, Bur12, Coh14, Dem35, Ewi38, Hau01, Lel37, MW37, MCPL37, Rud38, Ure46]. Methyl [Du 48, JA08]. Metrazol [Spe40]. MEV [HH41]. Mexican [Tre37]. Mexico [Nut26, Pal28, Ric37]. Michael [Ken35a]. Michaux [Tru37]. Michigan [Mai34, San35, SO37, Ste34b]. Micrometers [Ewe07]. Microscopy [Bro42b]. Middle [Col43a, HS04, Mor43a, Vai43]. Midwestern [Wea44]. Migrating [Hob21]. Migration [Du 48, Osb34]. Mild [Ric33]. Mile [Web19]. Military [Bro44b, Spr48]. Milky [See12a]. Millennium [Alb30]. Miller [Pet03, Goo31]. Millikan [Sny26b]. Million [SP42]. Mind [Ges41]. Mindedness [God12]. Mine [Bal21]. Mineral [Lei47]. Minerals [Bru26, Bru27, Kel08, Rog13, Smi06, Ste34b]. Minimum [Par39]. Minis [KW26]. Minnesota [Jen36, Jen38]. Minoris [McD15]. Minute [Bro02]. Minutes [Ano09e, Ano10f, Ano11h, Ano12h, Ano13h, Ano14e, Ano15g, Ano16m, Ano17j, Ano18i, Ano19i, Ano20h, Ano21f, Ano23i, Ano24e, Ano25f, Ano26j, Ano27e, Ano28e, Ano29e, Ano30f, Ano31f, Ano32a, Ano33a, Ano34a, Ano35a, Ano36a, Bru22b]. Miocene [BL36b, HA04, Hat02]. Miscellaneous [Sch45a]. Misnomers [Bal23]. Missiles [Har48b]. Mississippi [Hau04]. Missouri [Cas37]. Mites [Ewi45]. Mixtures [Hoo26, JA08, Mun10]. Mode [See09]. Modern [Abb27, Eme40, Ewe07, Inn44, McE38, Mor32, Mul49, Pri02, Pri34, Sig46, Wal26]. Modification [Sto23]. Modifications [Han23, Hel28]. Moghul [Jac29]. Mogollon [Hau43]. Moisture [CBB36, LL40]. Molar [Gre34b, Hel28]. Molecular [Smy36, TE33]. Molecules [Low11, Tay41, Tru30]. Mollusca [Moz37]. Molluscan [New48, SK36, SK37, Sta03, vI02]. Mollusks [Bar45, Her37, Ric37, Ric38, Ric39]. Molybdenum [WS06]. Monetary [Nic48, Wil43a]. Money [War43]. Mongolia [Osb23]. Mongoose [Eme40]. Monmouth [Tat38]. Monocotyledons [Har07]. Monsoon [Dal05]. Montagnais [SE42]. Montana [Dou02, Sta03]. Monterey [SP46]. Montgomery [Hol47]. Moon [Dal46, Emm32, Ste37]. Moral [Ful09]. Moreau [Mim12, Ros11]. Morocco [Hen48]. Morphologic [Har22]. Morphological [Mon04a]. Morphology [Atk15, Ges41, Mon08, Pri14]. Morris [CM21]. Mortality [Spi39]. Mortgage [Nol44]. Mosaic [Sta37]. Mosaics [Whi39b]. Mosquitoes [Bat49]. Most [MW37, Moo08]. Motif [Blo17]. Motion [Cle49, Joh39, MHM31, Str43, vdK44]. Motions [See11c]. Motivation [Tol41]. Motives [Tat38]. Motorized [Boy45]. Mould [Har06]. Mount [And19]. Mountain [Dal25, Dor45, Hob44, Rei11, See06, See07a, See08]. Mountain-Building [Dal25]. Mountains [Hob23, HA04, See13, Wad45a]. Mounted [Sin25a, SJ29]. Mounting [Ric12]. Mourning [Dav49b]. Mourning-Caps [Dav49b]. Movement [Bai03, Hob23, Hoo38]. Movements [MW25, See08]. MS [SO37]. Ms. [Ros03b]. Muhlenberg [Wal48]. Muhlenbergs [Wal49b]. Muladeva [Blo13]. Mule [Bis41]. Müller [Pac04b]. Multiple [BRS38, Whi39b]. Mummy [Ash03]. Mummy-Grave [Ash03]. Murderers [Del44]. Muscular [Ben10, BJ19]. Museum [Fra08, OM19, Osb23, Sin21a]. Music [Lue40]. Mussels [Ort18, Ort20]. Mustelus [Par37]. mut [Shu42]. Mutant [Whi39b]. Mutation [CBB36]. Mutational [Shu42]. Mutual [Ort05]. MV [Hay20a]. Mysteries [Sha30]. Mystery [Bal19, Rie30, Str42]. Myth [Fis12]. Myth-Making [Fis12]. Myxomatosis [Ber37].

n [Mil11]. N.Y. [RW09]. Naiades [Ort20]. Name [Jas08b]. Names [Cha13, Di 03, Mer37, Pri17]. Naming [Tre17]. Narratio [Dav39b]. Narratives [Bar13, Lam07]. Naskapi [SE42]. Natal [Dav39a, Dav40, JB13]. Nathanael [Ano00f]. Nation [Bro31, Hau09a, Joh24, Wal49a]. National [Con40, Del34, Fal46, Hue34, Kuz44, Mil46, Mud47, Pea19, Reh40, Sin21a, Sin21b, Sin21c, Win45, Jew41, Sch39]. Nations [Fen49a, Hei26, Koh43, Noy47]. Native [Mat00b, Mat04, Hun13]. Natives [Spi08b]. Natural [CH49, Goo09, Ham30, Mul49, OM19, Osb23, Reh40, Rid48, Sha43, Tre21, Zir41]. Naturalism [Bro02]. Naturalist [Pen42]. Nature [CTR29, Der29, Eig21a, Fox33, Heb42, Köh41, Kra04, Lym12, Mor49, Sch29, Sto03, WPL47, Hop31]. Naval [Oli04, Rea49]. Navali [Dav39b]. Navigable [Hau03]. Navigation [Hay20a]. Navy [Hay20a, Dua16]. Nayades [Ort18]. near [Mac33a, Sin10a]. Nearby [vdK44]. Nebraska [Sin15]. Nebulæ [Cur18, Sli17, Wri20]. Nebular [Bak40b]. Necessitating [Oli05]. Necessity [Hof39b]. Nectar [Hau22]. Nectaries [Bro38]. Necturus [Cra40]. Needed [Bog49]. Needs [Bai15, Kot45, Low45]. Negative [Lor06]. Neglected [Tru28]. Negotiates [Chi48]. Nelsonii [Fow04]. Neo [Dav24]. Neo-Lamarckiana [Dav24]. Neodymium [JA08, JS09]. Neolithic [Few32, Few37]. Neoplasm [WB43]. Nepenthes [Hep18]. Nephelometer [MB15]. Neptune [See11a]. Neptunian [Sli38b]. Nerve [Spe40, Spe42]. Nerves [Spe40, Spe42]. Nervous [Par11, Pat13, Wei41]. Nets [Ken30]. Networks [Ken24a, Ken25]. Neurohumoral [Par35]. Neurohumors [Par34, Par40]. Neurology [Las41]. Neuromotor [HK37, KH38, KR40]. Neuromuscular [Par16]. Neurones [Cog26]. Neuropsychiatry [Str44]. Neutral [Dar38, Row16]. Neutrality [Lin38]. Neutralization [Sch13b]. Neutralized [Tow15]. Neutrons [Kor41]. Never [Mag49]. Newcomb [Doo10]. Newer [Coh14, Hes26]. Newfoundland [Dal15]. Newhouse [Fen49b]. Newt [Fan38]. Newton [Bra42]. Next [Joh25]. Nicaragua [Ric39]. Nicobar [Spi08b]. Night [Bar11a, Bar19]. Nine [Osb34]. Nineteenth [BS49, Shr47]. Ninth [Dum48]. Ninus [Lev44]. Nitro [CGTR32]. Nitrogen [NP14, Ric04]. Noctua [Gro02a]. Noctuid [Gro02a]. Noise [Wol32]. Nolineæ [Tre11]. Nomadism [Sch44a]. Nomenclature [HC25, Rog13]. Non [Der31, HS04, Jac31a, Whi40a]. Non-Linear [Whi40a]. Non-Living [Der31, Jac31a]. Non-Marine [HS04]. Normal [Cri21, Heb42, Hrd34, Oli05, Reh16]. Norte [JB13, Sop16]. North [Ber14, BBC^{+}38, Cob15, Har29b, Kno40, New48, SK36, Sei40, Sei42, Spe26, AH13, AB18, Atk18a, Bar40b, Cri31, Gro02a, Hey13, Hob43, Lee40, Mat07b, MP44, Rus30, Sim42, Tre34]. Northeast [Dav06, Rit43]. Northeastern [Jep30a, Spe26]. Northern [Cre43]. Northwest [Jen36, Jen38]. Northwestern [Eig17b]. Norway [Sco30]. Nose [Dav37a, Dav39a]. Note [AB02, Bel49, Her37, Mas07, Rob49b, Spi08b]. Notes [Atk18b, Bar12, Dan11, Edg43, Fen01, Few44, Fox33, Har25, Kel08, Lym18, Mag10, Mat09b, Mat10, Mat05b, Ort18, Phe39, Phe40, Rei09, Rog10, Sha43, Sim43]. Notice [Cha06b, RW07]. Notostylops [RP35]. Notoungulates [RP35]. Nova [Bar12, Bob35, Bob36]. Novæ [Bar22]. Novel [Lev44]. Novelty [Hay40]. November [Ano00u, Ano00t, Ano01q, Ano01r, Ano02p, Ano03q, Ano04q, Ano06o, Ano06n, Ano07p, Ano08m, Ano08l]. Nth [Nip19]. Nuclear [Dar40b, Wen37]. Nuclei [Bar07b]. Nullipore [Set26]. Number [Cur18, Ham29, Ham30, Son38]. Numbers [Ben38, Dic27, SS45]. Nürnberg [Bid47]. Nutrition [Ros47a, She18]. Nutritional [Car29]. Nutritive [Ben10].

Oaks [Cob15, Tre12, Tre15, Tre17, Tre21]. Oat [Ree38]. Objective [Pic12, Sch13a]. Objectives [Ure46]. Obligations [Tow13]. Oblivion [Bar49a]. Observances [Hop10]. Observation [Bar07b, Dug15]. Observations [Bar10, Cou45b, Dub14, Fox29, Fro19, Fur16, Gre35, HS05, Hof39a, Kop41, KBC45, Lam19, McE38, Mud47, Ogb07, RM43, Sli38a, Smi05, Spi08a, Tyz32, Wie45, Wil39b, Wri20, dTdC36]. Observatories [McC41]. Observatory [Mil19a, Ogb07, And19, Bar12, Cam19, Fro19, Lam19, MG49, MPS20, Sli38a]. Observed [Fra45, Mit26, Oli35, See12b]. Observing [May19]. Obtained [Fow40, Gro02a]. Obtaining [Sch13a]. Occlusion [Ric03]. Occurrence [Abb04a, Bry26, Mab03, Par39, Son38]. Occurring [Her37]. Occurs [Sto03]. Ocean [Hau01, May19, McE38, Pig38, See08, Bau12, Bry02]. Oceanic [Fle38]. Oceanography [McE38, Sve47]. October [Ano00w, Ano00v, Ano01t, Ano01s, Ano02q, Ano03s, Ano03r, Ano04r, Ano05p, Ano05o, Ano06q, Ano06p, Ano07q, Ano08n]. Octocoral [Dei45]. Octolasium [KH38]. Odonata [Ano34l, Cal29, Cal34, Lin40]. Œ. [Atk18c]. Oenothera [Dav27, Dav37b, Shu42, Atk18c, Dav15, Dav24, Dav26, Dav34]. Oenotheras [Cle35]. Officers [Ano34m]. Offspring [MA36]. Oil [Bas15, Foo48]. Oil-Ether [Bas15]. Oils [Bas15, Par40]. Old [Bel49, Dic43, Jás49, Lin49, Mac29, Mac30a, Mar28, Ash03, Mon28, Pri16, Pri17, Pri19]. Old-World [Mac29]. Oldest [Bry36, Kra42, San35]. Oligocene [Jep34, Hat02, Jep33, Sin21c, Sin22b, Sin24]. Oligodynamic [Kra05]. Oliver [Mor10]. Omorhamphus [Sin28]. Onagra [Cle37]. One [Abb04b, Ano30c, Bro02, DSA45, Hop31, Pac03, Wil10, Wil20]. One-Half [Wil10]. One-Minute [Bro02]. Oniscus [Nic02]. Operation [Joh34, LL40]. Operational [Ada23]. Operations [Bla45b, Dra46]. Operators [Mil07, Mil10a]. Ophiocordyceps [Mai34]. Ophiura [Ber39a]. Opinion [Bin48, Cla32, Gal42, Lik48, Ste11a]. Opinions [Har48a, Ste48]. Opisthenogenesis [Pac04a]. Opisthonecta [KR40, Ros40]. Optical [Nip11b]. Orang [Fur16, Str36]. Orang-Utan [Str36]. Orang-Utans [Fur16]. Orange [Ger43]. Orbit [Doo03]. Orbital [Str43]. Orbits [See11c, See15, Wat39b]. Order [Mil07, See12b]. Orders [Mil07]. Ore [Cha08, Phi04]. Oregon [Cre43, HC25]. Oreodon [Sin21c, Sin24]. Ores [Bla13, Wri19]. Oresme [Bal08b]. Organic [BJ27, Smi26a]. Organisms [Chi16, Cro32, Kra05, Pac04b]. Organization [Dar44a, Har18, Hil16, Mat00c, Tay44, But44, Noy47]. Organs [Car08, Mon08, dMS27]. Orient [Hei18]. Oriental [Mer37]. Orientation [Shu44]. Origin [Bon34, Car09, Cha08, Dal46, Der27, Eig21a, Eig21b, Emm32, Gre27, Gre28a, Gre35, Har16, Hat02, Hau09b, Hoo38, Kra04, Mat00c, Mat19, Neu48, Osb27, Pac04b, Par11, See10b, See13, Whe18, Whi39a, Win40, Zir41, Bry09]. Origins [Bar29, Dix20, Jon42]. Orinoco [Eig21b]. Ornithology [Ekl45]. Orontes [Mor36]. Orthic [Bro04]. Orthodoxy [Sch31]. Orthogenesis [Cri33, Jep49]. Osborn [Gre34b, Gre36b]. Oscillations [Sei42]. Oscillograph [Ken28, McC05]. Osmotic [Jac31b]. Osteitis [Wil02]. Osteology [Mat05b, Pat34a, Pat35, Shu00a, Shu00b, Shu01]. Other [Ano27b, BB30, Bry35, Chi16, Gap40, Mil10a, Rus12, Wil08b]. Our [Con40, Dav06, Fis32, Jac42, Lov27, Mer27, Osb04, Reh40, Sch15a, Sha39]. Outer [Dav39a]. Outline [Kop45a, Moo16]. Outlook [Ayd30, Hob09]. Ovary [Loe11]. Overgrowth [Smi17]. Overhead [You33]. Overlooked [Mer36, Mer42]. Ovulation [Rid27]. Ownership [Joh34]. Oxide [Bee39, Mül26b]. Oxygen [Hin10, Par39, SS38, Tay48].

Pacific [AH13, Hob44, Coo40, Eig21a]. Paine [Ald49]. Painter [Sel48]. Painters [May43]. Painting [Luc18]. Paired [Di 03]. Pakistan [Bro47]. Palæaspis [Bry26]. Palæolithic [Bar40c]. Palaeonictis [SJ29]. Palæontology [Sco27]. Paleocene [Gil42, Jep30a, Jep37, Jep40, Pat35]. Paleoniscid [Cas37]. Paleontological [Cla22]. Paleontology [Ber39b, Jep30a, Sim42]. Palestine [Mac36a, Row29]. Palisot [Mer36]. Palladium [Wha04]. Palmer [Bry45, Ekl45, Kno45a, Kno45b, Ron45, Joe40]. Pamphlet [Hay24b]. Pamunkey [Few44]. Panama [Joh38, Joh43, Tow13]. Panels [Laz48]. Paper [Bro02, Bru22a, Bru23, Bru24, Bru25, Bru27, Mer42, Oli47]. Papers [AB18, Hal48, Hum42, Ros03a]. Papilionidæ [Ano00g]. Papyrus [San35]. Paradox [Tho35]. Paragenesis [Smi06]. Parahyba [Sop16]. Parallaxes [Bur21, MPS20]. Parallel [Cop49, Ken09, Ken30]. Paramecium [Jen08, Jen41, Son38]. Paramylodon [Sin10a]. Paramys [Jep37]. Parasite [Wen35]. Parasitic [Wen35, Whe19, Whi39b]. Parasitism [Set18]. Park [Gil42, Jep30a, Jep40]. Parnassi [Ano00g]. Parnassi-Papilionidæ [Ano00g]. Part [AB18, Jep40, Gil42, Hel28, SP46, Wil20]. Parthenogenesis [Phi03]. Particle [Whe46]. Particles [Dar38, Lan48c, MN40]. Particular [Hes38]. Parties [Cou45a]. Partition [Bar11c]. Parts [Osb34, BR34]. Pascoite [HMW14]. Passage [Mac28]. Passamaquoddy [Pri14]. Passenger [Dud04]. Past [Cla32, Hrd29, Mor43a, See09]. Pasteur [Bur23, Hen23]. Patagonia [RP35, Sco28]. Patagonian [vI02]. Patella [HK37]. Pathological [Cri21, Rum16]. Pathology [Gow34, Moo36]. Patriarchal [Bar13]. Pattern [Hel28, Sch44a]. Patterns [LHYH36, Tay39]. Pattersoni [Eig19]. Pd [Wha04]. Peace [Kae31, Sho38, Put47]. Peale [Sel48]. Peary [Bar40b, Lee40]. Peas [Whi17]. Peaslee [Sto18]. Pectoral [Gre35]. Peculiarities [Bar22]. Pediastrum [Har18]. Pedicellate [Tre35]. Peeper [Rug41]. Peirce [WPL47]. Pelagic [Fri45b]. Pellagra [Elv49, McC19]. Penalty [Spi08a]. Penicillin [Kee44]. Peninsula [Bry45, Cra07, Kno45a, Kno45b, SE42]. Peninsularity [Joe40]. Pennsylvania [All41, Bal15, Bry34, Car49, Che41, Cop49, Gre35, Mur44, Ort09, Tra06, Wil20]. Pennsylvanian [Cas37]. Penrose [Bea46, But44, Hir09, Lin33, Ang42, Con34, Swa35]. Pentadactylate [Gre35]. People [Han44, Hrd26a, Shr46]. Peopling [Hrd21, Hrd26b]. Pepper [Tre26]. Peppers [Tre30, Tre35]. Perception [Oli05, Tit16]. Peregrinorum [Dav39b]. Perfection [Ros13]. Performance [Bak40a, Mic15]. Periodical [Sch39]. Periodicities [Chi16]. Periods [BM38, Mac30b]. Peritresius [BL36b]. Permeability [Mac23]. Persistent [Bru27, Lan06a]. Personal [Blo44a, Bry09, Cha13]. Personality [Sar44, Sea48]. Personnel [Bry40]. Persons [Kee34]. Perspective [War48]. Pertaining [Bar07a, Mac02]. Peru [Ash03, Eig21a]. Peruvians [Mea26]. Peter [Cha06b]. Petrifactions [Rog10]. Petroglyphs [Hei47]. Petrography [Ste34a, Ste45a, Ste45b, War45]. Petroleum [Mab03, Mun10]. Petrology [Idd11]. Phalænoptilus [Mar05]. Phallic [Mat00d]. Phase [McC05]. Phases [Joh30]. Pheasant [Bis41]. Phenomena [AB02, Bro44a, Bru13, CGTR32, Hal14, Mac36b, McM38, See06]. Phenomenon [Nip11b]. Philadelphia [Ano30b, Hun32, Bal02, Bel48, Hrd43, Hun32, Lin49, Ogb07, Pau36, Reh40]. Philippines [Dix30, Hei18, Kie26]. Phillips [CDG^{+}12]. Philosopher [Sco32, Sha43]. Philosophical [Adl27, Ano00f, Ano26j, Ano28e, Ano29e, Ano30b, Ano30c, Ano30f, Ano31f, Ano32a, Ano35l, Bal12, Bro42a, Chi43a, Chi43b, Chi45, Con40, Con41, Con47, Der27, DM47, Fer42, Fra44, Hum42, Hun32, Jac42, Mit42, Nol34, Odg43, Pac46, Pac42, Ros03b, Ros03a, Ros07, Ros11, Sio49, Wis42, vD43, Sio49]. Philosophiques [BLV47]. Philosophy [Bec43, Dic43, Koh40, New18, Sch29, Sta48]. Phosphates [NL39]. Phosphorescence [Har40, HHHN17, Nic10, Nic16, Par20]. Phosphorus [New40, SS38]. Photo [KS19]. Photo-Electric [KS19]. Photographic [Bad36, Bar10, Sch14, Sch17, Shi45, Whi38, Whi40a, Whi40b, Whi47, vdK44]. Photographs [JR40, Nic16]. Photography [Bar07a, MHM31]. Photometric [Kop43a, Tuc33]. Photometry [KS19]. Phylogenetic [Cob15, Cri12, GH39, HS46]. Phylogeny [Bro38, Osb25]. Physical [Ben24, Cri13, CTR29, Dar47, Hoo35a, Hrd43, Mag10, Pac04b, Squ17, Tay44, Wad45c, Whi39a, Wig49]. Physician [Sig46]. Physicists [Ric42]. Physics [Boh47, Dar38, Mil26, See06, See08, Smy46, Ste30]. Physiological [AD41, Ben18, Hes26, Lec44]. Physiology [Bow17, Dil47, Fen48, Rug41]. Picture [Joh39, MHM31]. Pictures [Mat01a]. Pierids [Gro00a]. Pierpont [Wal26]. Pietro [Kra41]. Pigeons [Rid27]. Piggot [BBC^{+}38]. Pike [Hol47]. Pile [Fer46]. Pine [Mac33b]. Pines [Dou22]. Piper [Tre29]. Piperaceae [Tre28a, Tre34]. Pipiens [RM43]. Pisum [Whi17]. Pitcher [Hep18]. Pits [NS37]. Pitting [Bra13]. Pittsburgh [TW24]. Pituitary [RD45]. Place [Cha13, Com47, Sig46, Sto03]. Plain [Ano11f, Ber11]. Plains [Hat02, Smi38]. Plan [PG39, Ros13, Row42]. Plana [Con38]. Plane [Bar09c]. Planet [Jas08b, Sli38b]. Planetary [Cle49]. Planets [See10a, See11a, See11c, Sli38a, Wil39a]. Planevector [KV19]. Plankton [SS38]. Planning [Del34, Grè34a, Lou43, Rad34]. Plant [Bai03, Har25, Liv13, MS20, Mac28, Mer37, Moo08, Sta47]. Plant-Breeding [Bai03]. Plants [Ber35, Bla45c, Bro38, Di 03, Eas15, Eas40, Har08, Har09, Har11, Kra04, Mac24, Mac25, Rod15, Smi17, Ste49, WB43]. Plasm [Sto23]. Plasma [RD45]. Plasmatic [Mac24]. Plate [McC05]. Plateau [See13]. Plateaus [See08]. Plates [Bar12]. Platinum [Hey13, KS14]. Platonism [Sho27]. Plaza [Ano30c]. Plea [Oli05]. Pleistocene [Lev26, OC31, dTdC36]. Pliocene [Sin15]. Plutonium [Sto46]. Pneumatic [Mac32, Mac36c]. Pneumonia [TW24]. Pœcilogenesis [Pac04b]. Poète [DH47]. Poetry [Bro41, Col41b, Edg36]. Poincaré [Veb12]. Point [Bro22, Dua22, Lel37]. Points [Lam04, Lot39]. Poison [Vau12]. Poisoning [Eva37]. Poisons [Pea16]. Poland [RP06]. Polar [Dav10c]. Polarized [Wri19]. Pole [Bar40b, Sau40]. Polecat [Gil42, Jep40]. Police [Bur12, Has00]. Policy [Blo44a, Lik48, Mac48, Vol38]. Political [Bal16, Bec43, Dic43, Fen49a, Fet43, Inn44, Nic48, Sco32]. Politics [Lin38, Mar40, Row14]. Polygons [Ken26]. Polynesia [Her37]. Polynesian [Dix20]. Polynesians [Dix34]. Polyphemus [Gro02b]. Polyphyletic [Pac03]. Pondings [Wil20]. Pools [Har16]. Popes [Mun16]. Popularity [McK40]. Population [Boa43, CH49, Kae39, Lot39, Not39, Not43, Pea24, Sha39, Tae39, Tay39, Tho39, Wri49]. Porcelain [LL40]. Porous [LL40]. Porous-Porcelain [LL40]. Portion [Few37]. Portrait [Bar07a]. Portraits [Bie49, Kim44]. Portuguese [Har25]. Position [Ben43, Vai43]. Positions [Sch14, Wat39c]. Positive [Lor06]. Possibility [Par39]. Possible [Ber22, Cha06a, CGTR32, Hau43, Tay39]. Post [Con44, Dav39a, Dav40, Koh43, Lou43, Sel44, Shu44, Wil43a, Not43]. Post-Natal [Dav39a, Dav40]. Post-War [Con44, Koh43, Lou43, Sel44, Shu44, Wil43a, Not43]. Postnatal [Dav37a, Dav44]. Posts [Oli05]. Posture [Gre28a]. Potassium [JS09]. Potato [Kun43]. Potential [Bar09a]. Potentialities [Blo44a]. Potentially [Gol24, Hal24]. Potentials [Com22]. Potentiometer [KV19]. Potteries [Ash03]. Pottery [Few44, Hei49]. Pottery-Making [Few44]. Power [Ame33, Bur12, Cri33, Has00, Mil10a, See12b]. Powers [Bur12]. Practical [KU08, Ken31, Lym15, Sto12, TE33]. Practice [Ame24, Nor32, Wri21]. Practices [Kay45]. Prairies [Wea44]. Prapāthakas [Dum48].

Praseodymium [JA08]. Pre [BS49, Bro41]. Pre-Armada [Bro41]. Pre-Nineteenth [BS49]. Precipitation [BS07, Hei30, WS06]. Precipitins [Ste02]. Precursor [Boy43]. Predecessors [Nut26]. Predicting [Ben24]. Prediction [Sny26a]. Preface [Byr45]. Prehistoric [Bal23, Bal25, Hen48, Mac33a, Mac34, Mac36a, Mea26, dT42]. Prehistory [Det37, Dix30, Mac29, Mac30a]. Preliminary [AB02, Atk18b, Ewe07, Few32, Hof39b, Hoo32, Kee34, Kra42, MHM31, RB08, Spi08b, Ste45c, Sch39]. Preparation [MCPL37, MPG26]. Present [Bai15, Bar26, Gra28, Hrd29, Kee44, Mer26, Nor32, Osb28, Rob49a, Roy06, Ste38, Ste11a, War26, Whi17]. Presented [Oli35]. Presenting [Ken26]. Preservation [Joh39, MCPL37]. Preservatives [Wil05, Wil08b]. President [Jew41, Row42, Smi03, Wri21]. Presiding [Ano30b, Ano30c]. Pressure [Dal05, KS14, RM43]. Pressures [Mac25, Nip06]. Presumably [Bru26]. Preventive [Sau49]. Pribilof [Par15]. Prices [Ang36, Pat34b]. Priestley [Bro42a]. Primary [Par35, SP42]. Primates [SS45, Str36]. Primitive [Gre28b, Mat05b, Moo08, Par11]. Primrose [Dav27]. Princeton [Sin21a, But07, Sco28, Sin15, Sin21c]. Principal [Lew45]. Principia [Bra42]. Principle [Mer03]. Principles [Com11, McE38, Roy06, Sip45a, Woo40]. Printer [Lin48]. Printing [Sch39]. Prints [Bet46]. Printz [Bal15]. Prior [Hum42]. Prism [Pic12, Sch13a]. Private [Com45, Kay45, Mil46]. Prize [CDG^{+}12]. Probable [Bru27, Mac22]. Problem [Bar26, Col43b, Hal43, Han23, Hau04, Joh20, Joh24, JW27, Joh38, Lan48b, Lov09, Osb28, Par15, Pat03, Rou12, Sch35, Ste38, Sto12]. Problems [Ada37b, BB30, Bro44b, Cam19, Coo46, CF22, Dic27, Dil47, How43, Idd11, Joh29b, Joh34, Lut27, Men39, Oli35, Spe26, Sta47, Ste11b, Tyz32, Whe46, Wil43a]. Proboscidea [Osb25]. Proboscidean [Osb34]. Proceedings [Ano35l]. Process [Fis12, KN20]. Processes [Ben10, Cri21, Jac31a, KW44, See07a]. Prochromosomes [Smi34]. Procuring [Hau03]. Produced [Con38, See08, WB43]. Production [BM38, Bea36, BM32, HH41, Koe03, Kor41, Lan48c, Tit14b, WS07]. Productivity [Mye42]. Products [MH20]. Progeny [Pea16]. Program [Com47, Fal46, Mer34, Sel44, Shr43, Ste40, Con43]. Programs [Mud47]. Progress [Con34, Del47, Fis32, Hei32, Hoo26, Mac30a, Ste36b, Tho07, Tit08, Whi31]. Progressive [Cas33]. Project [Sto46]. Prolonged [Ben18, Bru18]. Prominences [Lam19]. Promoter [But48, Dar33]. Promoting [Bla45a, Lel37, MW37]. Promotion [Jon46, Kep37]. Pronouns [Hew04, Pri15]. Pronunciation [BS49]. Proof [Low11]. Propagandist [Bro41]. Propagation [Ada10, Mac02, Mun10]. Proper [vdK44]. Properties [AD41, Car29, Dav30, Goo03, KT16, Ken24a, Mil10b, Wri03]. Property [You34]. Proposed [Bal07, Hue34, Ken35b, War28]. Prospect [Mac29, Sch13a]. Prospective [Tay39]. Prospects [Whe46]. Protection [Dor44, Sto46]. Protective [Par40]. Protein [Vau12]. Proteins [Sta37, Wyc37]. Protest [Hau11]. Protoceras [Sin25a]. Protophyta [GG38]. Protoplasm [CGTR32, Hei34, Kra02, MS20]. Protoplasmic [Loe49]. Protozoa [Wen35, Jen08, Wen37]. Providence [WBAW39]. Provinces [New48, SK36]. Provision [Ame24]. Provisional [Dav35]. Pseudomorphs [Rog10]. Psychiatry [Sau49]. Psychoanalysis [Lan41]. Psychological [Fla39]. Psychology [Dub14, Las41, Tit14a, Yer41]. Pteraspidæ [Bry26]. Public [Bin48, Com45, Dra46, Gal42, Goo48, Kay45, Mal39]. Publication [Sch35, Wet47]. Publications [Mer43, Wet47]. Published [Hay24b, Hum42]. Pupin [Ken35a]. Pure [Dav15, Fen01]. Purification [Mas09]. Purpose [Bry40, Rid29]. Puzzle [Tre37]. Puzzles [Hop31]. Puzzling [Tre26]. Pycnocarpa [Atk18c]. Pygidiidæ [Eig17a]. Pyrite [Tra06].

Quaker [Bar49b]. Quantitative [Len26, Mye42, Ric03]. Quantum [Eva37]. Quantum-Counter [Eva37]. Quarante [DH47]. Quarante-Deux [DH47]. Quartermaster [Ano00f]. Quartermaster-General [Ano00f]. Queensland [Mat01a, Mat03b]. Question [Bal09b, Roy06]. Questions [Bal16].

R. [Bea46, But44]. R.I. [WBAW39]. Rôle [CGTR32, Hob15, Pre04]. Rabbit [Ber37, PS40]. Race [Ben37, Boa43, Dav17, Hoo26, You47]. Racemic [KM04]. Races [Gol24, Hal24]. Racial [Hel28, Wir48]. Radial [Sch13a]. Radiant [Wat39c]. Radiated [Aus13]. Radiation [Abb11, Du 31, Kor41]. Radii [Gap40]. Radio [Dua22, Hal41]. Radio-Active [Dua22]. Radioactivity [Bol11]. Radiotelegraphy [Aus13]. Radium [Eva37, Phi04]. Rafinesque [Bet44, Cra40, Mer42, Mer43]. Raider [Sch44a]. Railroad [Dav10b, Joh25, Joh29b, Joh30, Joh32b]. Railroads [Joh34, War31]. Railways [Joh32a]. Rain [Bat49]. Rainfall [Dal05, Jen36, Jen38]. Raised [Sli48]. Ramsay [Ric17]. Rana [RM43]. Rancho [Sin10a]. Range [SP42, War26, Web19]. Ranges [Kra10, Rei11, See08, War45, Pas45]. Rare [Pri16]. Rat [Han23]. Rate [Bas15, BL36a, CBB36, Hob23]. Rates [Ano34l, Cal29, Cal34, Tuc22]. Rather [Pac04b]. Ratio [Pea17]. Rational [Lym15, Rog12]. Rationalised [Ken31]. Rats [Det23]. Ray [Eva37, KBC45, MW25, MN40]. Rays [Com35, Han23, Hes26, Rug39, RE40, Sny26a, Sny26b]. Re [Bal09a]. Re-Explore [Bal09a]. Reacting [Fer46]. Reaction [Hau03, Hei30]. Reactions [HS30, Par22, Tay26, Wig46]. Reading [Car48, ZF47]. Real [Sto03]. Realism [Ful36]. Reality [Ste49]. Received [Aus13]. Recently [Bru15, Bru17, Car49, Oli35]. Reception [Bry09]. Recessives [Whi39b]. Recognize [Dav06]. Reconstruction [Col43c]. Record [Dou01, Jep49]. Records [Put47]. Recovery [Spe40, Wea44, Wet34, You34]. Recurrence [Bal16]. Recurrent [Eme17]. Red [Bow17, Cox48, HC40, MG49, Ran23, Rus10, Set18, Sin21c]. Redefinition [HS46]. Redevelopment [Lin49]. Reduction [Ben18]. Reef [May20, Set26]. Reef-Formation [Set26]. Reefs [May20, Pal28]. Reference [Ano34l, Cal29, CGTR32, Fle41, Gre35, Hes38, Hoo26, Hrd43, Mac36b, Nic02, Pea16, Spe40, Str36, Tit16, Wen37]. Reflecting [Bar11b]. Reflection [Kop43b]. Reflections [Hrd35, Sha39]. Reflexogenous [Wei41]. Reform [Blo44a]. Reforms [Ken35b]. Refugees [Mur44]. Regard [Hay24b, Mon04b]. regarding [Hrd35, Mer03, Roy06, Spi08a]. Region [Bob35, Cre43, JB13, Ric37, Hei49, Hob44, Jen36, Jen38]. Regional [Grè34a, Rad34]. Regions [Hob10]. Registrants [Row42]. Regulating [Mac28]. Regulation [Ham30, Oli04]. Rehabilitation [Row42, Rus46]. Related [Fla39, Liv13, Lot39, Rod15]. Relating [Car49, Has00, Osb27, Tow13]. Relation [Cog26, Cri12, Cri13, Dal25, Far17, FR18, Gis41, Gow34, HC40, Har32, Jon37, Kae39, Lel32, Mac33b, May20, McC19, Mer37, Mer05, Mye43, Par35, Rou12, Sin25b, Spi39, Spr48, Whi40c, Zon45, Bru26]. Relations [Bak48, Bro46, Con47, Ken48, Kuz48, Lei47, Mag12, Mar48, Mil10b, Nic48, Par22, Rid48, Rus12, Ste36b, Vin46, Wen35, Wir48, Wri21, You47]. Relationship [SS38, Wri20]. Relationships [Cle37, Cob15, Mil46, Pat34a]. Relative [Bar29, Ble37, Bur21, Hew04, Pea19, Wri21]. Relativistic [Rob49a]. Relativity [Cle49, Wey49]. Religion [Bar40c, Ros47b]. Religious [Noc42, Pri16]. Remains [Ber39a, Pri34]. Remarks [Ano40m, Bau17, Bur23, Con42, Hoo32]. Reminiscences [Bry09]. Removing [Bla45c]. Renilla [Par20]. Reorganization [Dod33]. Repercussions [Cus48]. Replicas [Kim44]. Report [Con36, Con37a, Ekl45, Few32, KBC45, McC45, RB08, Sch39, Ste36b, Ste45c]. Reported [Abb04b, Eme40]. Reports [Bar40a, Sco28, Tit16]. Representation [Nip19]. Representative [Moo08]. Representatives [Lew45, Ano37g]. Represented [Gro02a]. Reproduction [Har18]. Reproductive [Rug41]. Repulsion [Low11]. Requirements [Gam44]. Research [Bak48, Bea39, Ber41, Con36, Con37a, Cre43, CF22, Dar44a, Fle41, Gra28, Gre44, Hoo35a, Kra47, Laz48, Lel37, Mac33a, McC27, MW37, Nic44a, Oli36, RC37, Sch39, Shr47, Ska46, Ste36b, Tau49, Tay44, Tho49, Wey49, Whe46, Wir48]. Researches [Bad36, See08, See10c]. Resistance [Ada37c, Ken09]. Resisting [See15]. Resonance [Wig46]. Resources [Lei47]. Respect [WBAW39]. Respective [Ken29]. Respectively [Mil07]. Respiration [Har32, dMS28]. Respiratory [Fen48, Kne37]. Responses [Par38, Par41, Par44, WB37]. Responsibilities [Mil46, Wri21]. Responsibility [Kep37]. Responsive [Dav41a]. Restoration [OM19]. Restored [Sin10b]. Resulting [Rug39, RE40]. Results [Bar07a, Bau15, Cre43, Dug15, Fro19, Gro02a, Hof39b, KS07, MHM31, See10c, Sin15, Sin21c, Tru28, Wad45b, Wie45]. Résumé [Mab03]. Retinal [Par22]. Retrospect [Cla32, Mac29]. Return [Sch31]. Returns [Jen48]. Revenues [Joh43]. Reversal [Cha06a]. Reversion [Dav10a]. Review [Gre28a, Hum42, Kuz48, Mon08, Phi03, Sim40]. Revised [Gre34b]. Revision [Jep34, Mag49]. Revolt [Mor35]. Revolution [Lin38, Fis12, Fra44, Gre29, McK40]. Revolutionary [Wal49b]. Revolutions [Cur49]. Rheumatic [Swi23]. Rhinochimaera [Wil08a].

Richard [Hir09, Lin33]. Riddle [Bea39]. Ridgway [Mar05]. Right [Rid24]. Rights [Lew45, Row16, Tow15, Wri21, Lew42]. Rings [Dou22]. Rio [JB13, Sop16]. Ripening [Mas04]. Rites [Mat00d]. River [Cox48, Jep34, Sin21c, Sin22b, Sin24, Wan23, Hau04, Jep32, Wan22]. Rivers [Hau03]. Roatan [Ric38]. Robert [Kyt49]. Rock [Mat01a]. Rockefeller [Wad45a]. Rockets [Kra47]. Rocks [Bru26, Bru27, Pas45, Smi06, Smy27, Ste34a, Ste34b, Ste45b, Ste47]. Rodent [Jep37]. Rodmani [Atk15]. Roger [Bur27]. Roi [DH47]. Roland [Hoo35b]. Role [Bla45a, Coh48, Dye44, Fal46, Hec38, Kot45, Ros47a, Rus47, Tho46]. Romantic [Fra27]. Rome [Car09, Fra31a]. Rose [May21]. Rosengarten [Jas21]. Ross [Wad45c]. Rotating [KS05, KS06, WS06]. Rotation [Det23, Pla30, Smy36]. Round [Con37b, LBD^{+}38, MS31]. Rowland [Bar09c]. Royal [Sio49]. Ruling [Mic15]. Rumford [Bro49]. Running [Swa35]. Rural [Tay39]. Rush [But48]. Russian [DM47, Pri17, Pri19]. Rusts [AB18, Art25].

S [Hay20a]. S. [Kra41]. Sabre [Jep33]. Sabre-Tooth [Jep33]. Sacrifice [Dum48]. Safeguard [Hay20a]. Sailor [Lin18]. Saint [Ros11, Mim12]. Saint-Mery [Mim12]. Sales [Bue44]. Salisbury [For36]. Salt [Har11, Har16]. Salts [JS09, Mac21a, NP14]. Same [Dou01, Mil11]. Samoa [May20, May21]. Samples [Pig38]. Sampling [Ste48]. Samuel [Bet44, Car16, Wal26]. San [Cra07, Eig19, HA04]. Sanarelli [Ber37]. Sanctioned [dMS28]. Sand [Har08]. Sandia [Hib43]. Sanitary [Ash19]. Sanitation [Hei18]. Santa [HA04, Sin05, Sin08]. Sap [Mac25]. Sara [Hau21]. Sargasso [Far14]. Satellites [See10b]. Satisfy [Bro04]. Saturday [Ano30b]. Sauripterus [Gre35]. Sauropod [OM19]. Saved [Bar49a]. Saw [Ols41]. Sayre [Ogb07]. Sc.D [Mor09, Tho11]. Scales [Tuc33]. Scaliger [Hal48]. Scarlet [WBAW39]. Scattering [BRS38]. Scene [For43]. Schedule [Ano37g]. Schmidt [Bak40a, Bak40b]. Scholar [Bra42, Mal39]. Scholarly [Sch39]. Scholarship [Lan48b, Nic44b]. Schweinitz [AB18]. Science [Ash19, Bar40a, Ber33, Che27, Chi43a, Com40, Com47, Dal47, Joh28, Lot39, Mil47, Nic44a, Roy06, dMS28, Sho27, Tho35, DM47]. Sciences [Bai47, Dar47, Ful09, Goo09, Lan48a, Mac46, Pen27, Reh40, Rid48, Tay44, Whe31, You48, Jew41, Sin21a, Sin21b, Sin21c]. Scientific [Bel48, Bro02, Cal42, Fle47, Hun47, Joh39, Lan41, Mar48, Mil19b, Noy47, dMS27, Sch35, Sch39, Wad45b, Wil42b]. Scientist [Ken24b, Moo47]. Scientists [Bro46]. Scores [Heb42]. Screes [Har29b]. Sculpture [Fra08]. Sea [BBC^{+}38, Cha06a, Har11, Har40, JW27, Joh41, Lam07, Mye42, Par16, Par39, See07a, SS38, Ell40, Far14]. Sea-Anemones [Par16]. Sea-Level [JW27]. Sealevel [Joh29a]. Seals [Par15]. Search [Gro02a, Sli38b]. Seasonal [Cal34, Ger43]. Second [Bru27, MPS20, Oli47, Pat35, Ste36b, Bru22a]. Secondary [HH41]. Secretion [Can11, Reh16]. Secretions [dMS28]. Secretory [HM38]. Section [Dou02]. Sections [Neu48]. Secular [See07b, See08, See11c, See15].

Securing [Pig38]. Securities [Hue34]. Sedimentary [Pas45]. Sediments [Wan22]. Seeds [CBB36, Lan06a]. Seen [Sch49]. Segments [Pac04a]. Segregate [Tre29]. Segregation [Dav37b, Jon37]. Segregations [Shu42]. Seismic [Hec38, Hob09, Ste37]. Seismological [Rei09, Rei15]. Selection [Jen08, Jep49, May49, Mul49, Sip45a, Ste49, Zir41]. Selective [Whi35]. Selenium [Len26]. Self [Bar11a, Bar19, Eas40, Hal43, Jen41, Jen41]. Self-Determinism [Hal43]. Self-Differentiation [Jen41]. Self-Fertilization [Jen41]. Self-Luminous [Bar11a, Bar19]. Self-Sterility [Eas40]. Selfed [Dav24]. Semiramis [Lev44]. Semites [Jas04]. Senate [Wri21]. Sensation [Tit16]. Sensations [Baz26]. Sense [Par18]. Separated [Tay48]. Separation [Ure46]. September [Ogb07, Wie45, Ano01e, Ogb07]. Septentrionalis [Con02]. Sequence [Hau43]. Sequences [Ben43]. Sequoias [Dou22]. Series [Ken25, Lam03, Lam08]. Seringe [Dav27]. Sermon [Cra17]. Sermon-Books [Cra17]. Serological [Guy23]. Serpentines [Kra10]. Serra [Lut26]. Service [Bar45, Bla45b, Cla45, Cou45a, Cus45, Dei45, Ewi45, Fra45, Fri45a, Gla45, KBC45, Oli04, Per45, Sch45a, Sch45b, Sho45, Sip45b, Wad45b]. Services [Hen23]. Sessions [Put47]. Seth [Fen49b]. Settlement [Lut27]. Seven [Dav24]. Seventeenth [Bla36]. Seventy [Web19]. Seventy-Five-Mile [Web19]. Sex [All19, Mon04a, Pea17, Rid24, RD45, Whi35]. Sex-Determination [Whi35]. Shadows [TP07]. Shakespeare [Bla36, Pau36, Phe39, Phe40, Sch16, Sha35]. Shakespearean [Sch31]. Shale [Sch15b]. Shalt [Bec31]. Shamanism [Sti33]. Shape [Ort20]. Shaping [Com40]. Shelf [Wad45c]. Shells [Ada37b]. Ship [Bon41]. Shock [Moo36]. Shope [Ber37]. Short [Hal41, HS30, Mac36b]. Should [Bal09a]. Shower [Whi40b, Whi47]. Showers [Abb04b]. Sibs [All24]. Side [Bro35]. Sidereal [See12b]. Sides [Rid24]. Sight [Dua16]. Sign [Jas08b]. Signalling [Dua16]. Signals [Low01, Oli04]. Significance [Abb18, Bru27, McM23, McN41, Mer37, Par11, Swi23, Woo31]. Silicates [Har33]. Silurian [Sco30]. Siluridæ [Dri19]. Silver [Hin10, Mül36]. Silviam [Dav39b]. Similar [Bar09c]. Simon [Doo10]. Simple [Her40, KT15, Ken24a]. Since [Cra17, Car27]. Sinclairella [Jep34]. Singular [Lam04]. Sinking [SS38]. Sinopa [Mat05b]. Sir [Jas16, Qui45, Ric17, Sch46]. Sister [Ano27b, Mer27]. Site [Few35]. Sites [Few37]. Situation [Joh32b, Sei40]. Siwalik [dTdC36]. Six [Fen49a, McC26]. Sixteen [Eig17a]. Sixth [Wal49a]. Size [Jen08, RP06, WBAW39]. Skeleton [Gre36a, Sin10b, Sin25a, SJ29]. Sketch [Cob15, Gro00b, Sco23, Tru29]. Sketchbooks [Chi49]. Skull [Cas37, Gaz36, GH39, Oli02, SJ27]. Slav [Pri20]. Slaves [Wes48]. Slavonic [Pri16, Pri35]. Slavs [Col41b]. Sledge [Ron45]. Slime [Har06]. Slope [Eig21a]. Slopes [Har29b]. Slovenia [Pri31]. Small [BJ27, KS14, Koh43, Sin22b, Tyz42, SN48]. Smith [Pen42, Con35, Qui45]. Smithiï [Tre37]. Smithsonian [WPL47]. Smokes [Lan48c]. Smut [Ree38]. Smyth [RW07]. Snake [Sin15]. Snakes [NS37]. Snapper [Hay11]. Snow [Lan48c]. So-Called [Lym12]. Social [Blo44a, Bou43, Hov48, Hun44, Lan48a, Laz48, Mac46, Mer34, Nic44a, Shu44, Tho39, War44, Whe18, Wil46, Woo31, You48]. Societies [Ano27b, Lel37, Mer27, Lel32]. Society [Ano30b, Ano30c, Ano35l, Sig46, Adl27, Ano00f, Ano26j, Ano28e, Ano29e, Ano30f, Ano31f, Ano32a, Bal12, Bro42a, Chi43a, Chi43b, Chi45, Con40, Con41, Con47, Der27, DM47, Fer42, Hum42, Hun32, Jac42, Mit42, Nol34, Odg43, Pac46, Pac42, Ros03b, Ros03a, Ros07, Ros11, Sio49, Wis42, vD43]. Sociological [Grè34a]. Sociology [Mat05a]. Socrates [New18]. Soil [LL40, Sch13b, Sen30, Zon45]. Soil-Moisture [LL40]. Solander [Dav26, Dav27]. Solar [Abb11, And19, Bau10, Hal14, McM38, MG49, Men39, See09, See11a]. Soldier [Lin18, San39]. Soldiers [Sto48]. Solid [Bak40b]. Solid-Glass [Bak40b]. Solids [Smy36]. Soluble [Smi26a, Smi26b]. Solution [Hey10, Joh38, JS09, Lam03, Lam08, Lam11]. Solutions [Mag07, Mag12]. Solve [Ros47b]. Solvent [Ric03]. Solvents [JA08]. Somatic [Jon37]. Some [Ada23, Ada35b, Atk18b, Bai15, Bal12, Bar07a, Bar22, Bee39, Bla44, Bla45c, Bro41, Bru13, Bru21, Bru22a, Bru23, Bru24, Bru25, Bru26, Bru27, Bru28, Cla12, Cra38, Di 03, Dou22, Dub14, Dug15, Ewe07, Fal14, Fer42, Fox29, Fox33, Fra31a, Gow34, Gre29, HS46, HS05, Hay00a, Her37, Hrd35, Joh29b, Joh30, Joh32b, Kel08, KT16, Ken24a, Kop44, Lou43, Mac02, Mat03b, Mat07d, Mat09c, Mil10b, Mil19b, Not43, Pat28, Rad34, Ric42, RP35, Rog10, San35, Sha39, Sli48, Smi18, Spa37, Sta37, Ste34a, Ste45c, Ste45b, Tre26, Tru28, Web19, Wet34, Whe18, Bru26, Cam19, Ste11b]. Songs [Eme37]. Sonic [Hay24a]. Soul [Bon41]. Source [Cop48, Dut44, Lin48, Ric03]. Sources [PGG43, Ric04, Put47]. South [Cre43, Eig17b, Jep34, Mat03a, Mon28, Ogb07, Sin21a, Sin21c, Art25, Dou02, Dri19, Eig21b, Far17, Fri45b, Lut27, Mat00a, Mat00d, Sau40, Sim40, Sim43, Sin22a, Ste34a, Tre35, Wan23]. Southeast [Lew43]. Southeastern [SE42, Ber14]. Southern [Ron45, Hei49, Hof39b, Kno45a, Kno45b, Pas45, Pic06, Sei42, War45]. Southwest [Dav06, Dou22, Hib43]. Southwestern [Hei47, Hob44]. Sovereignty [War43]. Sows [RP06]. Space [Dun39, Mer03]. Spafford [Boy43]. Span [Fra31b]. Spearthrower [Dav36]. Special [Ano00h, Ano01e, Ano34l, Cal29, CGTR32, Gre35, Hoo26, Hrd43, Mac36b, Pea16, Rid29, Spe40, Str36, Tit16, Wen37]. Specialists [Bar40a]. Specializations [Ano00g]. Specialized [Gro02b]. Specially [Bar07a]. Speciation [May49]. Species [Atk18b, Cle37, Dav15, Eig17a, Eig17b, Her37, McC26, Mer37, NW38, Ort05, Pal28, Tre10, Tre28b, Zir41, dV06, Bry09]. Specific [Ste02]. Specificity [RB08]. Specimen [Hay11]. Spectacles [Tay41]. Spectra [Com22, JA08, JS09, Mit30, Ran23, Rus12, Tro02, WPG41]. Spectral [Hof39b, HHHN17]. Spectrograph [Bak40b, Bar12, Ble36]. Spectroscopic [Low11, Mit32, Wri20]. Spectroscopy [Dem35, Mil27]. Spectrum [Bob35, Bob36, MG49, Mit19, Mit26, Sny26a]. Speech [Pri03]. Speed [Hal49]. Spelman [Sch46]. Sperm [Rug39, RE40]. Spermatogenesis [Nic02]. Sphaerocarpos [All26, All19, All24]. Spiral [Cur18, Ken26]. Spirit [Goo48]. Split [Ada36b, Ada36c]. Spontaneous [Bru15, Bru17, CH49]. Spores [Cou16]. Sporophyte [MA36]. Spring [Dou01, Rug41]. Sproul [Mil19a, MPS20]. Square [Mil10a]. Squares [Kop43a]. St [Mil35]. St. [Cro49]. St.-Denis [Cro49]. Stabilization [Wes34, Wil43a]. Stable [Mor49, Tay48]. Stage [BS49, McK40, Ric33]. Stages [Lev26, OC31]. Standard [Kem32, Sha49, Tuc33]. Standardizing [Bar09b]. Standpoint [Ges41, Hay15, Ken33, Mer43, She18]. Staple [Pea19]. Star [Doo03, Pea21, Sha19, Ste30]. Starch [SM39]. Stars [Cha39, Hof39b, McD15, MPS20, Mit28, PG39, PGG43, Rus10, Rus12, Sch14, See10a, SN48, Sny02, Str39, Wri20, vdK44]. Stasis [HC40]. State [All41, Bla45a, Cri13, Has00, Inn44, New44, Nip13, Rob49a, Ros13, Ros47b, Roy06, Shu44, Whi17, Rit43]. Stated [Ano00e, Ano00j, Ano00i, Ano00k, Ano00m, Ano00l, Ano00o, Ano00n, Ano00q, Ano00p, Ano00s, Ano00r, Ano00u, Ano00t, Ano00w, Ano00v, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01p, Ano01o, Ano01q, Ano01r, Ano01t, Ano01s, Ano02f, Ano02h, Ano02g, Ano02j, Ano02i, Ano02l, Ano02k, Ano02m, Ano02o, Ano02n, Ano02p, Ano02q, Ano03f, Ano03h, Ano03g, Ano03j, Ano03i, Ano03l, Ano03k, Ano03n, Ano03m, Ano03o, Ano03p, Ano03q, Ano03s, Ano03r, Ano04f, Ano04h, Ano04g, Ano04j, Ano04i, Ano04k, Ano04l, Ano04n, Ano04m, Ano04p, Ano04o, Ano04q, Ano04r, Ano05h, Ano05g, Ano05i, Ano05k, Ano05j, Ano05l, Ano05n, Ano05m, Ano05p, Ano05o, Ano06f, Ano06g, Ano06h, Ano06j, Ano06i]. Stated [Ano06k, Ano06m, Ano06l, Ano06o, Ano06n, Ano06q, Ano06p, Ano07g, Ano07i, Ano07h, Ano07j, Ano07l, Ano07k, Ano07m, Ano07o, Ano07n, Ano07p, Ano07q, Ano08e, Ano08g, Ano08f, Ano08h, Ano08j, Ano08i, Ano08k, Ano08m, Ano08l, Ano08n]. Statement [Ano16j, Lew45, Mil30]. States [Bar45, Bla45b, Bon40, Cla45, Col41a, Cou45a, Cus45, Dei45, Ewi45, Fra45, Fri45a, Gla45, Hrd43, KBC45, Mil19b, Per45, Sch45a, Sch45b, Sho45, Sip45b, Wad45b, Ang27, Bry26, Bur12, Con44, Cou27, Dar47, Fet43, Hau07a, Hec41, Kuz48, Liv13, Mim12, Ort05, Ree36, Row14, Tho49, Tow13, dCW23, Wri21]. Station [Dor45, Ort20]. Statistical [Cla06]. Statistics [Cla12]. Status [Bar26, Dar40b, Eag30, Gra28, Kee44, Osb28, Pat28, Ste38, War26]. Steadily [KN20]. Steam [Emm03]. Steel [Bru15, Bru17, Bru18]. Stegocephalia [Cas33]. Stelae [Bon41]. Stellar [PGG43, Rus39, Sha30]. Stem [Gre27]. Stenostomum [NW38]. Steps [Joh25]. Stereoscopic [JR40]. Sterility [Eas15, Eas40]. Stibiodomeykite [Koe03]. Stiffness [Mer05]. Still [Bec43]. Stimulation [Can11]. Stimuli [TP08]. Stimulus [Bog49, Ham29, Wey49]. Stock [Hue34]. Stockholders [Sel44]. Stocks [Son38]. Stone [OC31]. Stones [Mat09a, Mat09b]. Stories [Cra17]. Storms [Dav06]. Straight [KS27, Lam05]. Straits [Tit16]. Strand [Har09]. Stratford [Ano30b]. Stratigraphy [AH13, HC25, Jep30a, Sin21c, Wan23]. Stream [Nip06, Bac36]. Streamers [Mil19a]. Streaming [Loe49]. Streams [Rod15]. Streptomycin [Kee48]. Striges [Shu00a]. Strigidæ [Shu00a]. String [Dav41b]. Structural [Poo33]. Structure [Bry26, Cri31, Fle38, HK37, Har08, Har09, Hes38, HHHN17, How41, Mac48, New44, NS37, Pat13, Pla30, Sau27, Smi34, War45, War44, Wen37, Wri49]. Studies [Bon42, Car08, Cle35, CGTR32, Dav31, Din43, Eva37, HC40, Har48a, Hep18, Hof40, Jen08, Jen36, Jen38, Joh29a, Lel32, Lik48, McM38, Mül26b, MPG26, New40, PS40, Rug41, Sch44a, Sen30, Spe40, Spe42, Ste48, TE33, Voe49, Wen37, Whi38, Whi40a, Whi40b, Whi47, Whi17, Wol32]. Study [Ano11f, Ber11, Bry36, Coh14, Con34, Cra17, Dub44, Ewe07, Fla39, Gap38, GH39, Har22, Hob44, Hoo26, Kie26, Lot39, Luc18, Mac36b, Mar05, MW25, Moz37, Mye42, New18, Oli35, PG39, RB08, Ste40, Tyz42, War25, Wat39c, Wat39d, WB37, Whe18, Whe19, Wri19, Wyc37, vdK44]. Studying [Sto48]. Sub [Hof39b, Moz37]. Sub-Arctic [Moz37]. Sub-Classification [Hof39b]. Subfamily [Dri19]. Subfreezing [SP45]. Subjected [Spe42]. Subjective [Wet34]. Subliminally [TP08]. Submarine [Fie38, Hec38, Ste38]. Submarines [Hay20b]. Submerged [Spa37]. Subsidence [JW27]. Substance [Cra38]. Substances [Bru28, Mac28, Par40, RB08, Sch13b, Wil08b]. Substitutes [Coh44]. Substitution [Mil10b, Mil11]. Substitutions [Mil11]. Succession [OC31]. Successive [Lam11]. Such [Par40]. Suetonius [Rol13]. Suggested [Nor32]. Suggestion [Edg36]. Suggestions [Joh32b]. Sulfur [Dav37b]. Sulphides [HHHN17, Nic16]. Sulphur [Dou01]. Sulphuric [KS06, Kra06]. Sumer [Kra46]. Sumerian [Kra42, Pri15]. Summary [Gre34b, HC25, Ste11a]. Summer [dCW23]. Sun [Bru26, Low11, See11c, Sit39, SK43]. Supercooled [Lan48c]. Superintendent [Col41a]. Superiority [Mon04a]. Supernova [WPG41]. Supernovæ [Hof39a, Whi39a]. Superposed [Bru13]. Supervision [Oli05]. Supplies [Mas09]. Supply [Par38, Smi18]. Support [Dar44a, RC37, Tay44]. Supporting [Bru26]. Supposed [Low01]. Supreme [Ree36]. Surface [Boy45, Emm32, Hei30, Mac33b]. Surfaces [JW27, KT15]. Surrounding [Goo03]. Survey [Col41a, Hec41, Kra42, Par18, Shr46, Ska46, Wil42b, BBC^{+}38]. Surveying [Rud38]. Surveys [Ber45, Ewi38]. Susceptibility [Bru18]. Swarm [Sch44a]. Swarm-Raider [Sch44a]. Swedes [Bal15]. Swimming [Ros40]. Symmetry [Mac36b]. Symposium [Dar44a, Mil30, Wad45b]. Symptoms [Tho37b]. Synchronization [MW25]. Syncretistic [Bon42]. Synonymy [Di 03, Ort18]. Synthesis [KM04, WPG41]. Synthetic [WS07]. Syria [But07]. Syriac [New18]. Syrian [New18]. System [Cri14, Dud04, HK37, Ken36, KH38, KR40, Mac32, Par11, Pat13, See09, See11a, Wei41]. Systematic [Bai15, Dav02, Wil39c]. Systems [Jac31a, Kop42, Kop43b, Kop45b, Mac36c, See12b, Shu44].

T [McD15]. Table [Con37b, LBD^{+}38, MS31]. Tables [Kop44]. Tacarigua [Ber39b]. Tactile [Tit16]. Tadpoles [Spe40, Spe42]. Tæniodont [Gaz36]. Tail [Low11, Par35]. Tailed [Cra39, Cra40]. Taittirīya [Dum48]. Taittirīya-Brāhmana [Dum48]. Taken [Bar12]. Takes [Sto03]. Tale [Eme40, Pri02]. Talus [Har29b]. Tangier [Hen48]. Tantalum [Smi05]. Task [Com45]. Tasks [Mag49]. Taste [Mor35]. Tatar [Pri19]. Taurid [Whi40b]. Tax [Blo44a, Bue44, Mag49, New44, Ros13, Shu44]. Taxation [Blo44b, Sel44]. Taxes [Kuz44]. Taxonomic [Cle35, Har07]. Taylori [Gre35, Jep32]. Technique [Pig38, Shi45]. Techniques [You47]. Technological [Kae39]. Technology [Whi31]. Teeth [Hel28]. Telea [Gro02b]. Telegraph [Hal49]. Telephone [KU08, KT16]. Telephonic [KT15, KV19]. Telescope [Pic06, Tho30]. Telescopic [Wat39d]. Tellurium [Len26]. Temminckii [Hay11]. Temperature [Baz26, JS09, Nic10, See07b]. Temperatures [Con38, GG38, SP45, Tro02]. Temple [Win32]. Ten [Mic15, Sha30, Ste40]. Ten-Year [Ste40]. Tendentium [Dav39b]. Tennessee [Ort18]. Tenth [Jas04]. Terebratulina [Con02]. Terenuthis [Bon41]. Tergo [Pac04a]. Term [Kuz48]. Terminology [Bal25]. Terms [HS46, Tre16]. Terrestrial [Bau10, Fle38, Gis41, Rei40, Ste36b]. Territory [Tow15]. Tertiary [Ano11f, AH13, Ber11, Ber35, Dou02, HC25, Sim40, Sim43, vI02]. Test [Dav15, Heb42]. Testament [Mar28]. Testimony [Ash03]. Tests [Tit16]. Tetrad [All24]. Tetrahedral [Eme17]. Tetrapoda [Gre35]. Texas [Eig19, Hay11]. Text [Mal49, SO37]. Textual [Mar28]. Theatre [Dra46]. Their [Bai47, Bar29, Bro35, Bru28, Cob15, Cra17, Dar47, Det37, Dix30, Han23, Hov09, Ken29, Kim44, Kuz48, Lei47, Mat04, Mat19, May19, May20, MCPL37, Nut26, Ort05, Osb34, Ric37, Row29, Sch15a, Sel44, Sha19, Ste02, Tay39, Vai43, Wen35, Whi39b]. Them [Wri20]. Themselves [Ken26]. Theobald [Con35]. Theorem [See15]. Theorems [Ken29]. Theoretical [KU08, Kop42, Kop45b, Roy06, TE33, WPG41]. Theories [Chi47, May20]. Theory [Ada20, Bro49, Bru14, Bru21, Bru22a, Bru23, Bru26, Bru28, Cri21, Dar40b, Dic27, Dix20, Edg36, Gre34b, Her40, Kop45a, Loe49, Mag07, Neu48, Ost16, Qui45, See07a, See07b, See12b, Wey49, Wig49, Wil17, Wri21]. Therapeutic [Mül36]. There [Lev26]. Therein [Wil20]. Thermal [KS14, Mag12, SP46]. These [JA08]. Thick [Ada37b]. Thinking [Bog49]. Third [Bru23, Dum48]. Thirty [Osb34]. Thirty-Nine [Osb34]. Thomas [Qui45, Ald49, Bec43, Bet44, Sha43, Ste36a, Tru29, Tru36, Wil43b, Wri43]. Thoracic [MW25, Str36]. Thorium [SK43]. Thou [Bec31]. Thought [Cur49, Der29, Fra27, Rid29, Tho37a, Tho37b]. Thoughts [Lou43, Mas04]. Thousand [Rol30]. Three [Lov09, Pal28]. Tibet [See13]. tigrinum [Cra39]. Timber [Kot45]. Timberlands [Kay45]. Time [BM38, Gal42, Lec44, OC31, Par44, Sei40, Ste11a, Zel31]. Tissue [Loe08, dMS28, WB43]. Tissues [Spe40]. Titanoides [Pat34a, Pat35]. Titanotherium [Sin21b]. Titicaca [Eig21b]. TNT [MH20]. Toads [Abb04b]. Tobacco [Sta37]. Tobacco-Mosaic [Sta37]. Tobit [Hau21]. Todas [Eme37, Eme39]. Today [All40]. Tolls [Joh38]. Tomb [Jac29]. Tomorrow [Rid29]. Tones [KT15]. Tools [Tau49]. Tooth [Bry32, Jep33]. Topographic [Dou22]. Topography [Dal46]. torosus [CH49]. Torres [Tit16]. Torrey [Dav34]. Torsions [Eme17]. Torture [Kno40]. Total [Mit32]. Totalitarian [Bon40, Koh40, Mar40, Woo40]. Totalitarianism [Hay40, Hea40, Whi40c]. Totality [Mil11]. Toxic [Sch13b]. Toxicity [Tru30].

Trade [Sli48, Whi40c]. Trades [Dal05]. Trades-Monsoon [Dal05]. Tradition [Cus48]. Traditional [Fen49b, Spr48]. Trail [Loc45]. Trails [Whi40a]. Train [Oli33]. Train-Drifts [Oli33]. Trains [Oli36, Oli42, Oli47, Tho37a, Tho37b]. Trans [Sli38b]. Transactions [Sio49]. Transections [Lee40]. Transept [Kra41]. Transformation [Ber37]. Transforms [Mil10a]. Transient [KN20]. Transit [Sch17]. Transiting [Sny02]. Transition [Lov43]. Translation [AB18, Dum48]. Transmission [All26, Hal41, Kun43, Sho48, Tyz32]. Transmitted [Bla44]. Transpiration [Bar11c]. Transplantation [Car08]. Transportation [Boy45, Hau07a, Joh24, Wil32]. Treatment [Bru18, Del44, Han23, Kee44, Kee48, Whi44]. Treatments [Spe42]. Treaty [Bog38, Bur12, Tow13, Rea49]. Treaty-Making [Bur12]. Trees [Mac21b, Mac32, Mac33b, Mac36c, Mac39, Rum16]. Trends [Col30, Dub44, Kae39, Kuz48, Shr47, Spr48, Tae39, Tay39]. Trial [Car17, Bid47]. Trianaeopiper [Tre28b, Tre28a]. Triassic [Bry34]. Tribe [Mat07c]. Tribes [Mat00b, Mat04, Mat07b, Mat07d]. Tribunal [Bal07]. Tribute [Lin33]. Trifolium [Har22]. Trinity [Dal15]. Trionychidæ [Hay03]. Tripartite [Kra41]. Triploid [MA36]. Triploidy [Fan38]. Trisulphoxyarsenic [McC04]. Trituberculy [Gre34b]. Triturus [CH49, Fan38]. Trogloglanis [Eig19]. Tropical [Bat49, Lut27, Sen30]. Tropics [Bra13]. Troyan [Pri17]. True [Hin10, Hin13]. Trunks [Rum16]. Truss [Mer05]. Trusts [Pat03]. Tube [Goo03]. Tubercles [Pac04a]. Tuberculifera [Lin40]. Tuberculosis [Lon48, Rav03]. Tubes [KS27, Mun10]. Tubulidentate [Jep32]. Tubulodon [Jep32]. Tumor [Loe08, Rou12]. Tundra [Har28]. Tungstates [Smi26a, Smi26b]. Tungsten [SE04]. Tunnel [Dav10b]. Turbine [Emm03]. Turkeys [Tyz32]. Turkish [Pri34]. Turtle [BL36b, Sin21c]. Turtle-Oreodon [Sin21c]. Turtles [CL36]. Tuscaroras [Wal49a]. Tutuila [May20]. TV [McD15]. TW [McD15]. Twelfth [Mun11]. Twentieth [Hum42]. Twenty [Ben18, But44, Mor43a, JA08]. Twenty-Eight [Mor43a]. Twenty-First [JA08]. Twenty-Five [Ben18, But44]. Twin [Atk18c]. Twins [BB30]. Two [Ada35b, Ada37a, AB18, Bar09a, Bon42, Bro04, Mil07, Mil10a, Rol30, Tre16, Tre28b]. Two-Dimensional [Ada35b]. TX [McD15]. Type [Bak40b, DSA45, Gro02a, Hay20a, OM19, dCW23]. Types [Gro02a, Par16, Son38]. Typhoid [Rav13]. Typogeny [Hrd37]. Typotheria [Sin08].

U. [Hay20a]. U.S.A. [Ano00f]. ’Ulūm [Far39]. Ultimate [Sny26c]. Ultra [Hes26]. Ultra-Violet [Hes26]. Ultracentrifugal [Wyc37]. Ultraviolet [AD41]. Undergo [Zel31]. Underground [Wal49b]. Understanding [Coo46, Ken48, BLV47]. Undisturbed [Pig38]. Unemployment [Hue32, Woo31]. UNESCO [Com47, Lel46, Tho46]. Unfamiliar [Sho48]. Uniform [Ken09]. Union [Sta48, Sta48]. Unions [Fle47, Sli48]. United [Bla45b, Col41a, Ree36, Wad45b, dCW23, Ang27, Bar45, Bry26, Bur12, Cla45, Con44, Cou27, Cou45a, Cus45, Dar47, Dei45, Ewi45, Fet43, Fra45, Fri45a, Gla45, Hau07a, Hec41, Hrd43, KBC45, Kuz48, Liv13, Mil19b, Mim12, Noy47, Ort05, Per45, Row14, Sch45a, Sch45b, Sho45, Sip45b, Tho49, Tow13, Wri21]. Units [Ken31, Ken36]. Univalent [Mül26a]. Universal [Sny26b, Sny26c]. Universe [Sha19, Sha30, Sto03, Swa35]. Universities [Lov27, RC37]. University [Ang27, But07, Sco28, Sin15, Sin21c, Bet46, Che41, Fra08, Hoo26, Hoo35a, SO37, Sin21a, Ste34b]. Unpublished [Tru29]. Unrationalised [Ken31]. Unsolved [Sha30]. Unstable [Mor49]. Untangling [Hop31]. Unusual [Tre21, Tre33]. Uplift [See08]. upon [AB02, Bin48, Cro32, Fox33, Heb42, KS14, Kop43b, Kuz48, Mas07, Ort02, Ort13, RB08, RM43, Sch15a, See11c, Sto03, Tuc22, Wil08b]. Upper [Bla09, Kne37, Rus30, Ort18, Ber22, Gre35, dTdC36]. Upright [Gre28a]. Uranium [Bla13, WS06]. Uranous [JS09]. Uranyl [JS09]. Uric [Mor26]. Use [Hew04, KS05, Laz48, Mas09, McC05, Mer03, Rol22, Sch14, Ska46, Tau49, WS06]. Used [Mat09a]. Useful [But48, Dar33]. Uses [Hey41, Tay48]. Using [KS06]. USSR [Mud47]. Utah [Bla13, Cha13]. Utan [Str36]. Utans [Fur16].

V [Kra47]. V-2 [Kra47]. Vaccination [Rav13]. Vacuo [MPG26]. Valence [Fal14, NP14, Noy27]. Validity [Rog12]. Valley [Dou01, Nut26, Few37]. Value [Bar13, Foo13, Ste02]. Values [Gre29]. Vanadates [HMW14]. Vanadium [Bla13, Hin11]. Van’t [Jon11]. Vapor [Mun10]. Variability [Sei40]. Variable [McD15, PG39, PGG43, SN48]. Variables [Dug15, Gap38, Gap40]. Variation [Bro06, Hrd34, Jen08, Tuc22, Whi17]. Variations [CH49, Ger43, Sch15a]. Variegatus [Mac36b]. Variety [Kel09]. Variorum [Sha35]. Various [Hal24, JS09, Spe42]. Vascularity [Cra39, Cra40]. Vase [Luc18]. Vase-Painting [Luc18]. Vasomotor [Par38]. Vasorum [Cha38]. Vegetation [Far14, Har16, Har28, Har29a, Har29b, Kie26]. Vegetative [All26]. Velocities [Sch13a]. Velocity [Bea36, Har48b, Kop43b, Kop45b, Nip06]. Venezuela [Ber39b]. Ventilation [Dud04]. Verb [Lym12]. Verbs [Pri15]. Vergil [Rol30, Rob40]. versus [Ken31, Lin38, Set26, War31, Wei41]. Vertebrae [SS45]. Vertebrate [Jep30b, Sim42]. Vertebrates [Bry36, Hay00a, Mat19, RW09, RB08]. Vertical [Hob23, Sei42]. Vessels [Car08, Cra38]. VI [Spe40]. Vibrating [KT15, KT16]. Victoria [Ste34a, Mat03b, Mat04, Mat07c]. Vie [DH47]. View [Bro22, Dua22, Lel37]. Viewing [JR40]. Viewpoint [Cha15, Cha22, Cla22]. Views [Dav10a, Har40, Tit14a]. Vigenère [Men40]. Vigor [DSA45]. VII [Wad45a, PS40, Ste45b]. VIII [Spe42]. Vinca [Few35]. Vincoli [Kra41]. Violet [Hes26]. Virginia [BL36b, Bet46, Wat39a]. Virginianus [BL36b]. Viridescens [Fan38]. Virus [Ber37, Sta37, Wyc37]. Viruses [Bla44, Kne37]. Viscera [Str36]. Vision [Tit16]. Visual [Bob35, Car48, KW44, Tuc33]. Vital [RB08]. Vitality [Lan06a]. Vitrification [GG38]. Vitro [PS40]. Viverridae [GH39]. Vivo [PS40]. Vol [Ano35l]. Volatility [MPG26].

Volcanic [Emm32]. Volcanism [Sny26b, Sny26c]. Volcanoes [Hof40]. Voltage [Whi11, Whi20]. Voltaic [Bar09a]. Voltaire [BLV47]. Volts [SP42]. Volume [Ano00x, Ano00y, Ano01u, Ano01v, Ano02r, Ano02s, Ano03t, Ano03u, Ano04s, Ano04t, Ano05q, Ano05r, Ano06r, Ano06s, Ano07r, Ano07s, Ano08o, Ano08p, Ano09f, Ano09g, Ano10g, Ano10h, Ano11i, Ano11j, Ano12i, Ano12j, Ano13i, Ano13j, Ano14f, Ano14g, Ano15h, Ano15i, Ano16n, Ano16o, Ano17k, Ano17l, Ano18j, Ano18k, Ano19j, Ano19k, Ano20i, Ano20j, Ano21g, Ano21h, Ano22g, Ano22h, Ano23j, Ano23k, Ano24f, Ano24g, Ano25g, Ano25h, Ano26k, Ano26l, Ano26m, Ano27f, Ano28f, Ano28g, Ano29f, Ano29g, Ano31g, Ano31h, Ano32i, Ano32j, Ano33g, Ano33h, Ano34n, Ano34o, Ano34p, Ano34q, Ano35m, Ano35n, Ano36h, Ano36i, Ano37h, Ano37i, Ano37j, Ano38g, Ano38h, Ano38i, Ano39l, Ano39m, Ano39n, Ano39o, Ano40n, Ano40o, Ano40p, Ano40q]. Volume [Ano41l, Ano41m, Ano41n, Ano42g, Ano42h, Ano43h, Ano43i, Ano44i, Ano44j, Ano44k, Ano45f, Ano45g, Ano46h, Ano46i, Ano47p, Ano47q, Ano48s, Ano49v, Mac23]. Voluntary [Kis39]. Vorticellid [Ros40]. Vowel [Rol22]. Voyage [Mim12]. Voyages [Dix34, Ekb40]. Vries [Cle36, Dav26, Shu42]. vs [Dav31].

W [Ano30c]. Wales [Mat03a]. Walking [Lam07]. Wallace [Joh36]. Walls [Mac24]. Want [Bou43]. Wants [Har48a, Ste48]. War [Ang42, Bal16, Con44, Gal42, Koh40, Koh43, Lou43, Moo12, Nic44a, Sch44b, Sel44, Shu44, Smi38, Wil43a, Ash19, New44, Nic47, Not43]. Warfare [Rav03]. Wartime [Dye44]. Was [Bar49a, Dav06, Dav27]. Washington [vD32, HC25, Mor43a]. Wasp [Whi39b]. Water [Bar11c, BJ27, Gam44, Har11, JA08, Mas07, Mas09, Ort09, Ort20, Rei20, SS38, Sta03, Zon45]. Waters [Joh29a, New40]. Waterways [Hau05, Hau09a]. Watson [Smy13]. Wave [Dav30, Hal41]. Waves [Ada10]. Way [Joh32b, Joh32a, Wil32, See12a]. Way-Out [Wil32]. Weapons [Opp46]. Weather [Cou45b, Dor45, dCW23, War26, War28, War31]. Weathering [Sen30]. Wednesday [Ano30c]. Weight [Hin11, Hin13, Ric04, SE04, Wil10]. Weights [Hin10]. Weismann [Con15]. Welch [Fle35]. Welsh [BS49]. Were [Gre28b, Lev26]. West [Eig21b, Ber45, Bol47, Bon34, Col43a, Cop49, Cou45a, Fri45b, HS04, Hes38, Mat19, McC45, Ols41, Per45, Shi45, Sip45b, Wie45]. Western [Mat05a, Aga23, Cas37, Dav49a, Har29b, Hay40, Her37, Mat00b, Mat07d, New48, Ort09, Pal28, SK36, Sei40]. Westover [Sio49]. Wet [Bar07b]. Whale [Hau07b]. Wharton [Mor09]. Where [Bac36, Gre27]. Which [Bro04, Mil07, Mil10a, Mil11, RD45, dMS28, Wri20, Zel31]. White [Cob15, Sin24, Wil10, Jep34, Sin21c, Sin22b, Wan22, Wan23]. Whose [Mil07]. Wide [Fle41, Low45]. Widmanstättian [Sau27]. Widowhood [Spi39]. Wild [BF27]. Wilkes [Bal09a, Bar40a, Bry40, Fow40, Hob40, Rei40]. Wilkinson [Hol47]. Will [Bro31]. William [Goo31, Ric17, Con08, Dar33, Fle35, Joh36, Moo47, Sio49, Smy13]. Williams [Dar33]. Willson [Sel48]. Wilson [And19, Fin24, Pet03, Pac35]. Wind [Jep32]. Wing [Ano00g, Ger43]. Wing-Color [Ger43]. Wires [Ada10, KS14]. Witches [Kun43]. Witches’-Broom [Kun43]. Wolcott [Mor10]. Women [BJ19, BM32]. Wood [Mac33b, Win45]. Woodpeckers [Shu00b]. Woodrow [Fin24]. Woods [Low45]. Work [Bar07a, Bar40a, Bau15, Ben10, BJ19, Bog49, BBC^{+}38, Edg43, Hei26, Hin11, Hoo26, Joh41, Mer33, Mer03, Mil19b, See07a, Sta48, Tuc22, War25]. Worker [Sta48]. Works [Che11, Ste45a]. World [Ang42, Chi43a, Con44, Con09, Fle41, Hil16, Kae31, Koh43, Kot45, Kra42, Kra46, Lan46, Lew45, Low45, Mac29, Mac30a, Mer43, Osb34, Pat28, Row14, War28]. World-Wide [Fle41, Low45]. Wounding [Har48b]. Wounds [Whi44]. Write [Ald49]. Writing [Tat38]. Wrongly [Rob40]. Wyandot [Bar49a]. Wynicaco [Wes44]. Wyoming [Jep40, Bry32, Bry33, BR34, Bry35, Gaz36, Gil42, Jep30b, Jep30a, Jep32, Sin28].

X [Bru32, Han23, MW25, Rug39, RE40]. X-Ray [MW25]. X-Rays [Han23, Rug39, RE40]. Xeniophyte [Tre16].

Year [Ste40]. Years [But44, Hin11, Mil26, Mit42, Mor43a, Rol30, Smy46]. Yeast [AD41]. Yellow [Dou22, Eig22, Eig24]. Yerkes [Bar12, Fro19]. York [Ano30c, Rit43, Joh29a]. Young [BJ19]. Yucca [Tre37]. Yugoslavia [Few35]. Yugoslavian [Few37]. Yuma [How43].

Zea [Abb36]. Zebulon [Hol47]. Zone [Ash19, Hen48, See10b]. Zones [Sin24]. Zoological [Sch45a, Jac31b].

References

Abbott:1902:IAI

[AB02] A. C. Abbott and D. H. Bergey. The influence of alcoholic intoxication upon certain factors concerned in the phenomena of hæmolysis and bacteriolysis. A preliminary note. Proceedings of the American Philosophical Society, 41(169):140–142, April 1902. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

         //www.jstor.org/stable/983530.

Arthur:1918:ATP

[AB18] J. C. Arthur and G. R. Bisby. An annotated translation of the part of Schweinitz’s two papers giving the rusts of North America. Proceedings of the American Philosophical Society, 57(3):173–292, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

         URL https://www.jstor.org/stable/983925.

Abbott:1904:OAB

[Abb04a] Charles Conrad Abbott. On the occurrence of artifacts beneath a deposit of clay. Proceedings of the American Philosophical Society, 43(176):161–162, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983501.

Abbott:1904:OER

[Abb04b] Charles Conrad Abbott. One explanation of reported showers of toads. Proceedings of the American Philosophical Society, 43 (176):163–164, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983502.

Abbe:1906:BFM

[Abb06] Cleveland Abbe. Benjamin Franklin as meteorologist. Proceedings of the American Philosophical Society, 45(183):117–128, May/ September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983821.

Abbot:1911:SCR

[Abb11] C. G. Abbot. The solar constant of radiation. Proceedings of the American Philosophical Society, 50(199):235–245, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984038.

Abbott:1918:AOS

[Abb18] Charles C. Abbott. The archæological significance of an ancient dune. Proceedings of the American Philosophical Society, 57(1):49–59, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983881.

Abbot:1927:AMA

[Abb27] C. G. Abbot. Accomplishments of modern astronomy. Proceedings of the American Philosophical Society, 66(??):231–250, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301066.

Abbe:1936:HBD

[Abb36] Lucy B. Abbe. The histological background for dwarfism in Zea Mays. Proceedings of the American Philosophical Society, 76(6):743– 747, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984563.

Anderson:1941:EHU

[AD41] Thomas F. Anderson and B. M. Duggar. The effects of heat and ultraviolet light on certain physiological properties of yeast. Proceedings of the American Philosophical Society, 84(5):661–688, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984845.

Adams:1910:PLE

[Ada10] E. P. Adams. The propagation of long electric waves along wires. Proceedings of the American Philosophical Society, 49(197):364– 369, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984091.

Adams:1915:HCE

[Ada15] E. P. Adams. The Hall and Corbino effects. Proceedings of the American Philosophical Society, 54(216):47–51, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983911.

Adams:1920:ET

[Ada20] E. P. Adams. The Einstein theory. Proceedings of the American Philosophical Society, 59(3):176–183, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984497.

Adams:1923:SAH

[Ada23] E. P. Adams. Some applications of Heaviside’s operational methods. Proceedings of the American Philosophical Society, 62(2):26–47, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984508.

Adams:1935:EDC

[Ada35a] E. P. Adams. Electrical distributions on circular cylinders. Proceedings of the American Philosophical Society, 75(1):11–70, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984722.

Adams:1935:STD

[Ada35b] E. P. Adams. Some two-dimensional distributions of electricity. Proceedings of the American Philosophical Society, 75(6):549–563, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984616.

Adams:1936:EDC

[Ada36a] E. P. Adams. Electrical distributions on circular cylinders. II. Proceedings of the American Philosophical Society, 76(2):125–149, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984664.

Adams:1936:SCCa

[Ada36b] E. P. Adams. The split cylindrical condenser I. Proceedings of the American Philosophical Society, 76(3):251–267, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984544.

Adams:1936:SCCb

[Ada36c] E. P. Adams. The split cylindrical condenser II. Proceedings of the American Philosophical Society, 76(3):269–302, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984545.

Adams:1937:DET

[Ada37a] E. P. Adams. The distribution of electricity on two cylinders. Proceedings of the American Philosophical Society, 78(1):191–243, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984663.

Adams:1937:EPC

[Ada37b] E. P. Adams. Electrostatic problems connected with thick cylindrical shells. Proceedings of the American Philosophical Society, 78 (2):245–270, December 10, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984535.

Adams:1937:RCC

[Ada37c] E. P. Adams. The resistance of cylindrical conductors at high frequencies. Proceedings of the American Philosophical Society, 78 (2):271–285, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984536.

Adler:1927:APS

[Adl27] Cyrus Adler. The American Philosophical Society. Proceedings of the American Philosophical Society, 66(??):717–721, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301089.

Agar:1923:CMW

[Aga23] William M. Agar. Contact metamorphism in the Western Adirondacks. Proceedings of the American Philosophical Society, 62(3):95– 174, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984504.

Arnold:1913:MTS

[AH13] Ralph Arnold and Harold Hannibal. The marine tertiary stratigraphy of the North Pacific Coast of America. Proceedings of the American Philosophical Society, 52(212):559–605, November/December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984146.

Albright:1930:MBH

[Alb30] William F. Albright. A millennium of biblical history in the light of recent excavations. Proceedings of the American Philosophical Society, 69(1):441–461, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984378.

Aldridge:1949:WDT

[Ald49] Alfred Owen Aldridge. Why did Thomas Paine write on the bank? Proceedings of the American Philosophical Society, 93(4):309–315, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143156.

Allen:1919:BSI

[All19] Charles E. Allen. The basis of sex inheritance in Sphærocarpos. Proceedings of the American Philosophical Society, 58(5):289–316, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984196.

Allen:1924:ITS

[All24] Charles E. Allen. Inheritance by tetrad sibs in “Sphærocarpos”. Proceedings of the American Philosophical Society, 63(2):222–235, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984476.

Allen:1926:TVC

[All26] Charles E. Allen. The transmission of the vegetative character in “Sphaerocarpos”. Proceedings of the American Philosophical Society, 65(2):83–89, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984415.

Allen:1940:T

[All40] Frederick Lewis Allen. Today. Proceedings of the American Philosophical Society, 83(4):517–527, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984832.

Allen:1941:ADB

[All41] Henry Butler Allen. Alexander Dallas Bache and his connection with the Franklin Institute of the State of Pennsylvania. Proceedings of the American Philosophical Society, 84(2):145–149, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984913.

Ames:1924:APF

[Ame24] Herman V. Ames. The amending provision of the Federal Constitution in practice. Proceedings of the American Philosophical Society, 63(1):62–75, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984442.

Ames:1933:RDA

[Ame33] Herman V. Ames. Recent development of the amending power as applied to the Federal Constitution. Proceedings of the American Philosophical Society, 72(2):87–100, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984745.

Andrews:1909:DCG

[And09] Charles M. Andrews. Daniel Coit Gilman, LL.D. Proceedings of the American Philosophical Society, 48(193):lxii–lxx, September/ December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984068.

Anderson:1919:EMW

[And19] J. A. Anderson. The expedition of the Mount Wilson Observatory to the solar eclipse of June 8, 1918. Proceedings of the American Philosophical Society, 58(4):255–258, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984401.

Angell:1927:UEU

[Ang27] James R. Angell. University education in the United States. Proceedings of the American Philosophical Society, 66(??):645–654, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301086.

Angell:1936:FCP

[Ang36] James W. Angell. The factors controlling prices. Proceedings of the American Philosophical Society, 76(5):627–635, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984751.

Angell:1942:EWW

[Ang42] James R. Angell. Education in a world at war. Penrose Lecture, April 24, 1942. Proceedings of the American Philosophical Society, 85(5):433–439, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985062.

Anonymous:1900:Aa

[Ano00a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 39(161):??, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1900:Ab

[Ano00b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 39(162):??, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1900:Ac

[Ano00c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 39(163):??, September 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1900:Ad

[Ano00d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 39(164):??, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1900:ASM

[Ano00e] Anonymous. Adjourned stated meeting, January 19, 1900, February 2, 1900. Proceedings of the American Philosophical Society, 39(161):

69, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983543.

Anonymous:1900:CCM

[Ano00f] Anonymous. Calendar of the correspondence of Major-General Nathanael Greene, Quartermaster-General U.S.A., in the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 39(162):154–344, April 1900. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983662.

Anonymous:1900:ESL

[Ano00g] Anonymous. Errata: Specializations of the lepidopterous wing; Parnassi-Papilionidæ. Proceedings of the American Philosophical Society, 39(161):126, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983552.

Anonymous:1900:SMJc

[Ano00h] Anonymous. Special meeting, June 19, 1900. Proceedings of the American Philosophical Society, 39(162):358, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983667.

Anonymous:1900:SMAb

[Ano00i] Anonymous. Stated meeting, April 20, 1900. Proceedings of the American Philosophical Society, 39(162):345, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983663.

Anonymous:1900:SMAa

[Ano00j] Anonymous. Stated meeting, April 6, 1900. Proceedings of the American Philosophical Society, 39(162):127–129, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983660.

Anonymous:1900:SMD

[Ano00k] Anonymous. Stated meeting, December 7, 1900. Proceedings of the American Philosophical Society, 39(164):664–665, October/ December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983783.

Anonymous:1900:SMFb

[Ano00l] Anonymous. Stated meeting, February 16, 1900. Proceedings of the American Philosophical Society, 39(161):94, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983547.

Anonymous:1900:SMFa

[Ano00m] Anonymous. Stated meeting, February 2, 1900. Proceedings of the American Philosophical Society, 39(161):94, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983546.

Anonymous:1900:SMJb

[Ano00n] Anonymous. Stated meeting, January 19, 1900. Proceedings of the American Philosophical Society, 39(161):68, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983542.

Anonymous:1900:SMJa

[Ano00o] Anonymous. Stated meeting, January 5, 1900. Proceedings of the American Philosophical Society, 39(161):3–4, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983540.

Anonymous:1900:SMMb

[Ano00p] Anonymous. Stated meeting, March 16, 1900. Proceedings of the American Philosophical Society, 39(161):95–96, January 1900. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983549.

Anonymous:1900:SMMa

[Ano00q] Anonymous. Stated meeting, March 2, 1900. Proceedings of the American Philosophical Society, 39(161):95, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983548.

Anonymous:1900:SMMd

[Ano00r] Anonymous. Stated meeting, May 18, 1900. Proceedings of the American Philosophical Society, 39(162):357, April 1900. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983666.

Anonymous:1900:SMMc

[Ano00s] Anonymous. Stated meeting, May 4, 1900. Proceedings of the American Philosophical Society, 39(162):345–346, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983664.

Anonymous:1900:SMNb

[Ano00t] Anonymous. Stated meeting, November 16, 1900. Proceedings of the American Philosophical Society, 39(164):640, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983781.

Anonymous:1900:SMNa

[Ano00u] Anonymous. Stated meeting, November 2, 1900. Proceedings of the American Philosophical Society, 39(164):640, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983780.

Anonymous:1900:SMOb

[Ano00v] Anonymous. Stated meeting, October 19, 1900. Proceedings of the American Philosophical Society, 39(164):638–639, October/ December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983779.

Anonymous:1900:SMOa

[Ano00w] Anonymous. Stated meeting, October 5, 1900. Proceedings of the American Philosophical Society, 39(164):555, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983775.

Anonymous:1900:VIa

[Ano00x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 39(161):1, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983539.

Anonymous:1900:VIb

[Ano00y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 39(164):xiii–xv, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983774.

Anonymous:1901:Aa

[Ano01a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 40(165):??, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1901:Ab

[Ano01b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 40(166):??, July 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1901:Ac

[Ano01c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 40(167):??, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1901:FM

[Ano01d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 40(165):iii–xvi, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983420.

Anonymous:1901:SMS

[Ano01e] Anonymous. Special meeting, September 26, 1901. Proceedings of the American Philosophical Society, 40(167):137–138, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983753.

Anonymous:1901:SMAa

[Ano01f] Anonymous. Stated meeting, April 12, 1901. Proceedings of the American Philosophical Society, 40(165):56, January 1901. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983428.

Anonymous:1901:SMAb

[Ano01g] Anonymous. Stated meeting, April 19, 1901. Proceedings of the American Philosophical Society, 40(165):59, January 1901. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983430.

Anonymous:1901:SMD

[Ano01h] Anonymous. Stated meeting, December 6, 1901. Proceedings of the American Philosophical Society, 40(167):165–166, December 1901.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983761.

Anonymous:1901:SMFa

[Ano01i] Anonymous. Stated meeting, February 1, 1901. Proceedings of the American Philosophical Society, 40(165):52, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983424.

Anonymous:1901:SMFb

[Ano01j] Anonymous. Stated meeting February 15, 1901. Proceedings of the American Philosophical Society, 40(165):53–54, January 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983425.

Anonymous:1901:SMJb

[Ano01k] Anonymous. Stated meeting, January 18, 1901. Proceedings of the American Philosophical Society, 40(165):51–52, January 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983423.

Anonymous:1901:SMJa

[Ano01l] Anonymous. Stated meeting, January 4, 1901. Proceedings of the American Philosophical Society, 40(165):3–4, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983421.

Anonymous:1901:SMMa

[Ano01m] Anonymous. Stated meeting, March 1, 1901. Proceedings of the American Philosophical Society, 40(165):54–55, January 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983426.

Anonymous:1901:SMMb

[Ano01n] Anonymous. Stated meeting, March 15, 1901. Proceedings of the American Philosophical Society, 40(165):55, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983427.

Anonymous:1901:SMMd

[Ano01o] Anonymous. Stated meeting, May 17, 1901. Proceedings of the American Philosophical Society, 40(166):85–86, July 1901. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983573.

Anonymous:1901:SMMc

[Ano01p] Anonymous. Stated meeting, May 3, 1901. Proceedings of the American Philosophical Society, 40(166):61, July 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983571.

Anonymous:1901:SMNa

[Ano01q] Anonymous. Stated meeting, November 1, 1901. Proceedings of the American Philosophical Society, 40(167):160, December 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983758.

Anonymous:1901:SMNb

[Ano01r] Anonymous. Stated meeting, November 15, 1901. Proceedings of the American Philosophical Society, 40(167):165, December 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983760.

Anonymous:1901:SMOb

[Ano01s] Anonymous. Stated meeting, October 18, 1901. Proceedings of the American Philosophical Society, 40(167):159–160, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983757.

Anonymous:1901:SMOa

[Ano01t] Anonymous. Stated meeting, October 4, 1901. Proceedings of the American Philosophical Society, 40(167):138–140, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983754.

Anonymous:1901:VIa

[Ano01u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 40(165):i, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983419.

Anonymous:1901:VIb

[Ano01v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 40(167):xxiii–xxiv, December 1901. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983752.

Anonymous:1902:Aa

[Ano02a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 41(168):??, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1902:Ab

[Ano02b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 41(169):??, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1902:Ac

[Ano02c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 41(170):??, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1902:Ad

[Ano02d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 41(171):??, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1902:GMA

[Ano02e] Anonymous. General meeting, April 3, 4 and 5, 1902. Proceedings of the American Philosophical Society, 41(168):36–40, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983524.

Anonymous:1902:SMA

[Ano02f] Anonymous. Stated meeting, April 18, 1902. Proceedings of the American Philosophical Society, 41(170):259, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983556.

Anonymous:1902:SMDb

[Ano02g] Anonymous. Stated meeting, December 19, 1902. Proceedings of the American Philosophical Society, 41(171):414, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983491.

Anonymous:1902:SMDa

[Ano02h] Anonymous. Stated meeting, December 5, 1902. Proceedings of the American Philosophical Society, 41(171):406–407, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983489.

Anonymous:1902:SMFb

[Ano02i] Anonymous. Stated meeting, February 21, 1902. Proceedings of the American Philosophical Society, 41(168):35, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983522.

Anonymous:1902:SMFa

[Ano02j] Anonymous. Stated meeting, February 7, 1902. Proceedings of the American Philosophical Society, 41(168):35, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983521.

Anonymous:1902:SMJb

[Ano02k] Anonymous. Stated meeting, January 17, 1902. Proceedings of the American Philosophical Society, 41(168):34, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983520.

Anonymous:1902:SMJa

[Ano02l] Anonymous. Stated meeting, January 3, 1902. Proceedings of the American Philosophical Society, 41(168):3–4, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983517.

Anonymous:1902:SMMa

[Ano02m] Anonymous. Stated meeting, March 7, 1902. Proceedings of the American Philosophical Society, 41(168):36, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983523.

Anonymous:1902:SMMc

[Ano02n] Anonymous. Stated meeting, May 16, 1902. Proceedings of the American Philosophical Society, 41(170):264–265, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983559.

Anonymous:1902:SMMb

[Ano02o] Anonymous. Stated meeting, May 2, 1902. Proceedings of the American Philosophical Society, 41(170):264, April 1902. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983558.

Anonymous:1902:SMN

[Ano02p] Anonymous. Stated meeting, November 21, 1902. Proceedings of the American Philosophical Society, 41(171):406, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983488.

Anonymous:1902:SMO

[Ano02q] Anonymous. Stated meeting, October 3, 1902. Proceedings of the American Philosophical Society, 41(171):400, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983486.

Anonymous:1902:VIa

[Ano02r] Anonymous. Volume information. Proceedings of the American Philosophical Society, 41(168):1, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983516.

Anonymous:1902:VIb

[Ano02s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 41(171):415–417, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983484.

Anonymous:1903:Aa

[Ano03a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 42(172):??, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1903:Ab

[Ano03b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 42(173):??, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1903:Ac

[Ano03c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 42(174):??, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1903:BM

[Ano03d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 42(174):i–xvi, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983725.

Anonymous:1903:GMA

[Ano03e] Anonymous. General meeting, April 2, 3 and 4, 1903. Proceedings of the American Philosophical Society, 42(172):6–11, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983637.

Anonymous:1903:SMA

[Ano03f] Anonymous. Stated meeting, April 17, 1903. Proceedings of the American Philosophical Society, 42(174):265, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983709.

Anonymous:1903:SMDb

[Ano03g] Anonymous. Stated meeting, December 18, 1903. Proceedings of the American Philosophical Society, 42(174):396, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983721.

Anonymous:1903:SMDa

[Ano03h] Anonymous. Stated meeting, December 4, 1903. Proceedings of the American Philosophical Society, 42(174):352, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983718.

Anonymous:1903:SMFb

[Ano03i] Anonymous. Stated meeting, February 20, 1903. Proceedings of the American Philosophical Society, 42(172):5, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983634.

Anonymous:1903:SMFa

[Ano03j] Anonymous. Stated meeting, February 6, 1903. Proceedings of the American Philosophical Society, 42(172):4–5, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983633.

Anonymous:1903:SMJb

[Ano03k] Anonymous. Stated meeting, January 16, 1903. Proceedings of the American Philosophical Society, 42(172):4, January/April 1903. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983632.

Anonymous:1903:SMJa

[Ano03l] Anonymous. Stated meeting, January 2, 1903. Proceedings of the American Philosophical Society, 42(172):3–4, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983631.

Anonymous:1903:SMMb

[Ano03m] Anonymous. Stated meeting, March 20, 1903. Proceedings of the American Philosophical Society, 42(172):6, January/April 1903. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983636.

Anonymous:1903:SMMa

[Ano03n] Anonymous. Stated meeting, March 6, 1903. Proceedings of the American Philosophical Society, 42(172):5, January/April 1903. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983635.

Anonymous:1903:SMMc

[Ano03o] Anonymous. Stated meeting, May 1, 1903. Proceedings of the American Philosophical Society, 42(174):266, May/December 1903. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983710.

Anonymous:1903:SMMd

[Ano03p] Anonymous. Stated meeting, May 15, 1903. Proceedings of the American Philosophical Society, 42(174):266, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983711.

Anonymous:1903:SMN

[Ano03q] Anonymous. Stated meeting, November 6, 1903. Proceedings of the American Philosophical Society, 42(174):345–346, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983716.

Anonymous:1903:SMOb

[Ano03r] Anonymous. Stated meeting, October 16, 1903. Proceedings of the American Philosophical Society, 42(174):274, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983714.

Anonymous:1903:SMOa

[Ano03s] Anonymous. Stated meeting, October 2, 1903. Proceedings of the American Philosophical Society, 42(174):267, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983712.

Anonymous:1903:VIa

[Ano03t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 42(172):1, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983630.

Anonymous:1903:VIb

[Ano03u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 42(174):xix–xxii, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983708.

Anonymous:1904:Aa

[Ano04a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 43(175):??, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1904:Ab

[Ano04b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 43(176):??, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1904:Ac

[Ano04c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 43(177):??, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1904:Ad

[Ano04d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 43(178):??, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1904:GMA

[Ano04e] Anonymous. General meeting, April 7, 8 and 9, 1904. Proceedings of the American Philosophical Society, 43(176):97–102, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983492.

Anonymous:1904:SMA

[Ano04f] Anonymous. Stated meeting, April 15, 1904. Proceedings of the American Philosophical Society, 43(177):338, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983770.

Anonymous:1904:SMDb

[Ano04g] Anonymous. Stated meeting, December 16, 1904. Proceedings of the American Philosophical Society, 43(178):450, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983582.

Anonymous:1904:SMDa

[Ano04h] Anonymous. Stated meeting, December 2, 1904. Proceedings of the American Philosophical Society, 43(178):393, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983580.

Anonymous:1904:SMFb

[Ano04i] Anonymous. Stated meeting, February 19, 1904. Proceedings of the American Philosophical Society, 43(175):53, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983414.

Anonymous:1904:SMFa

[Ano04j] Anonymous. Stated meeting, February 5, 1904. Proceedings of the American Philosophical Society, 43(175):14, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983412.

Anonymous:1904:SMJa

[Ano04k] Anonymous. Stated meeting, January 1, 1904. Proceedings of the American Philosophical Society, 43(175):3–4, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983409.

Anonymous:1904:SMJb

[Ano04l] Anonymous. Stated meeting, January 15, 1904. Proceedings of the American Philosophical Society, 43(175):4, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983410.

Anonymous:1904:SMMb

[Ano04m] Anonymous. Stated meeting, March 18, 1904. Proceedings of the American Philosophical Society, 43(175):96, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983418.

Anonymous:1904:SMMa

[Ano04n] Anonymous. Stated meeting, March 4, 1904. Proceedings of the American Philosophical Society, 43(175):54, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983415.

Anonymous:1904:SMMd

[Ano04o] Anonymous. Stated meeting, May 20, 1904. Proceedings of the American Philosophical Society, 43(177):339, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983772.

Anonymous:1904:SMMc

[Ano04p] Anonymous. Stated meeting, May 6, 1904. Proceedings of the American Philosophical Society, 43(177):339, April/September 1904. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983771.

Anonymous:1904:SMN

[Ano04q] Anonymous. Stated meeting, November 4, 1904. Proceedings of the American Philosophical Society, 43(178):380, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983578.

Anonymous:1904:SMO

[Ano04r] Anonymous. Stated meeting, October 7, 1904. Proceedings of the American Philosophical Society, 43(177):340, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983773.

Anonymous:1904:VIa

[Ano04s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 43(175):1, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983408.

Anonymous:1904:VIb

[Ano04t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 43(178):457–460, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983575.

Anonymous:1905:Aa

[Ano05a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 44(179):??, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1905:Ab

[Ano05b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 44(180):??, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1905:Ac

[Ano05c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 44(181):??, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1905:BM

[Ano05d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 44(181):249–267, August/December 1905. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983751.

Anonymous:1905:FM

[Ano05e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 44(180):??, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983508.

Anonymous:1905:GMA

[Ano05f] Anonymous. General meeting, April 12, 13, and 14, 1905. Proceedings of the American Philosophical Society, 44(179):35–39, January/ April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983602.

Anonymous:1905:SMAb

[Ano05g] Anonymous. Stated meeting, April 28, 1905. Proceedings of the American Philosophical Society, 44(181):241, August/December 1905. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983745.

Anonymous:1905:SMAa

[Ano05h] Anonymous. Stated meeting, April 7, 1905. Proceedings of the American Philosophical Society, 44(179):35, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983601.

Anonymous:1905:SMD

[Ano05i] Anonymous. Stated meeting, December 15, 1905. Proceedings of the American Philosophical Society, 44(181):244, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983750.

Anonymous:1905:SMJb

[Ano05j] Anonymous. Stated meeting, January 20, 1905. Proceedings of the American Philosophical Society, 44(179):4–5, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983597.

Anonymous:1905:SMJa

[Ano05k] Anonymous. Stated meeting, January 6, 1905. Proceedings of the American Philosophical Society, 44(179):3–4, January/April 1905.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983596.

Anonymous:1905:SMMa

[Ano05l] Anonymous. Stated meeting, March 17, 1905. Proceedings of the American Philosophical Society, 44(179):5, January/April 1905. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983598.

Anonymous:1905:SMMc

[Ano05m] Anonymous. Stated meeting, May 19, 1905. Proceedings of the American Philosophical Society, 44(181):241–242, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983747.

Anonymous:1905:SMMb

[Ano05n] Anonymous. Stated meeting, May 5, 1905. Proceedings of the American Philosophical Society, 44(181):241, August/December 1905. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983746.

Anonymous:1905:SMOb

[Ano05o] Anonymous. Stated meeting, October 20, 1905. Proceedings of the American Philosophical Society, 44(181):243, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983749.

Anonymous:1905:SMOa

[Ano05p] Anonymous. Stated meeting, October 6, 1905. Proceedings of the American Philosophical Society, 44(181):242, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983748.

Anonymous:1905:VIa

[Ano05q] Anonymous. Volume information. Proceedings of the American Philosophical Society, 44(179):1–2, January/April 1905. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983595.

Anonymous:1905:VIb

[Ano05r] Anonymous. Volume information. Proceedings of the American Philosophical Society, 44(181):245–247, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983743.

Anonymous:1906:Aa

[Ano06a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 45(182):??, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1906:Ab

[Ano06b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 45(183):??, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1906:Ac

[Ano06c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 45(184):??, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1906:FM

[Ano06d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 45(184):420, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983373.

Anonymous:1906:GMA

[Ano06e] Anonymous. General meeting April 17, 18, 19 and 20, 1906. Proceedings of the American Philosophical Society, 45(182):10–13, January/ April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983675.

Anonymous:1906:SMA

[Ano06f] Anonymous. Stated meeting April 6, 1906. Proceedings of the American Philosophical Society, 45(182):10, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983674.

Anonymous:1906:SMD

[Ano06g] Anonymous. Stated meeting December 7, 1906. Proceedings of the American Philosophical Society, 45(184):415, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983384.

Anonymous:1906:SMF

[Ano06h] Anonymous. Stated meeting February 2, 1906. Proceedings of the American Philosophical Society, 45(182):9, January/April 1906. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983672.

Anonymous:1906:SMJb

[Ano06i] Anonymous. Stated meeting January 19, 1906. Proceedings of the American Philosophical Society, 45(182):4, January/April 1906. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983670.

Anonymous:1906:SMJa

[Ano06j] Anonymous. Stated meeting January 5, 1906. Proceedings of the American Philosophical Society, 45(182):3–4, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983669.

Anonymous:1906:SMMa

[Ano06k] Anonymous. Stated meeting March 16, 1906. Proceedings of the American Philosophical Society, 45(182):10, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983673.

Anonymous:1906:SMMc

[Ano06l] Anonymous. Stated meeting May 18, 1906. Proceedings of the American Philosophical Society, 45(183):178–182, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983828.

Anonymous:1906:SMMb

[Ano06m] Anonymous. Stated meeting May 4, 1906. Proceedings of the American Philosophical Society, 45(183):164–165, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983826.

Anonymous:1906:SMNb

[Ano06n] Anonymous. Stated meeting November 16, 1906. Proceedings of the American Philosophical Society, 45(184):415, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983383.

Anonymous:1906:SMNa

[Ano06o] Anonymous. Stated meeting November 2, 1906. Proceedings of the American Philosophical Society, 45(184):273, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983381.

Anonymous:1906:SMOb

[Ano06p] Anonymous. Stated meeting October 19, 1906. Proceedings of the American Philosophical Society, 45(184):244, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983376.

Anonymous:1906:SMOa

[Ano06q] Anonymous. Stated meeting October 5, 1906. Proceedings of the American Philosophical Society, 45(184):243–244, October/ December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983375.

Anonymous:1906:VIa

[Ano06r] Anonymous. Volume information. Proceedings of the American Philosophical Society, 45(182):1–2, January/April 1906. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983668.

Anonymous:1906:VIb

[Ano06s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 45(184):417–419, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983372.

Anonymous:1907:Aa

[Ano07a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 46(185):??, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1907:Ab

[Ano07b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 46(186):??, April/September 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1907:Ac

[Ano07c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 46(187):??, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1907:E

[Ano07d] Anonymous. Errata. Proceedings of the American Philosophical Society, 46(187):416, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983481.

Anonymous:1907:FM

[Ano07e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 46(187):435–454, October/December 1907. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983469.

Anonymous:1907:GMA

[Ano07f] Anonymous. General meeting April 18, 19, 20, 1907. Proceedings of the American Philosophical Society, 46(185):65–69, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983439.

Anonymous:1907:SMA

[Ano07g] Anonymous. Stated meeting April 5, 1907. Proceedings of the American Philosophical Society, 46(185):65, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983438.

Anonymous:1907:SMDb

[Ano07h] Anonymous. Stated meeting December 20, 1907. Proceedings of the American Philosophical Society, 46(187):360, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983477.

Anonymous:1907:SMDa

[Ano07i] Anonymous. Stated meeting December 6, 1907. Proceedings of the American Philosophical Society, 46(187):360, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983476.

Anonymous:1907:SMF

[Ano07j] Anonymous. Stated meeting February 15, 1907. Proceedings of the American Philosophical Society, 46(185):63–64, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983436.

Anonymous:1907:SMJb

[Ano07k] Anonymous. Stated meeting January 18, 1907. Proceedings of the American Philosophical Society, 46(185):62–63, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983435.

Anonymous:1907:SMJa

[Ano07l] Anonymous. Stated meeting January 4, 1907. Proceedings of the American Philosophical Society, 46(185):1–2, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983432.

Anonymous:1907:SMMa

[Ano07m] Anonymous. Stated meeting March 1, 1907. Proceedings of the American Philosophical Society, 46(185):64–65, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983437.

Anonymous:1907:SMMc

[Ano07n] Anonymous. Stated meeting May 17, 1907. Proceedings of the American Philosophical Society, 46(186):301, April/September 1907. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983741.

Anonymous:1907:SMMb

[Ano07o] Anonymous. Stated meeting May 3, 1907. Proceedings of the American Philosophical Society, 46(186):300, April/September 1907. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983740.

Anonymous:1907:SMN

[Ano07p] Anonymous. Stated meeting November 15, 1907. Proceedings of the American Philosophical Society, 46(187):359, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983475.

Anonymous:1907:SMO

[Ano07q] Anonymous. Stated meeting October 4, 1907. Proceedings of the American Philosophical Society, 46(187):340, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983472.

Anonymous:1907:VIa

[Ano07r] Anonymous. Volume information. Proceedings of the American Philosophical Society, 46(185):??, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983431.

Anonymous:1907:VIb

[Ano07s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 46(187):431–434, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983468.

Anonymous:1908:Aa

[Ano08a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 47(188):??, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1908:Ab

[Ano08b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 47(189):??, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1908:Ac

[Ano08c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 47(190):??, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1908:FM

[Ano08d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 47(190):i, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983830.

Anonymous:1908:SMA

[Ano08e] Anonymous. Stated meeting April 3, 1908. Proceedings of the American Philosophical Society, 47(188):13–17, January/April 1908. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983792.

Anonymous:1908:SMDb

[Ano08f] Anonymous. Stated meeting December 18, 1908. Proceedings of the American Philosophical Society, 47(190):697–698, September/ December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983840.

Anonymous:1908:SMDa

[Ano08g] Anonymous. Stated meeting December 4, 1908. Proceedings of the American Philosophical Society, 47(190):697, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983839.

Anonymous:1908:SMF

[Ano08h] Anonymous. Stated meeting February 7, 1908. Proceedings of the American Philosophical Society, 47(188):12, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983790.

Anonymous:1908:SMJb

[Ano08i] Anonymous. Stated meeting January 17, 1908. Proceedings of the American Philosophical Society, 47(188):2, January/April 1908. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983788.

Anonymous:1908:SMJa

[Ano08j] Anonymous. Stated meeting January 3, 1908. Proceedings of the American Philosophical Society, 47(188):1–2, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983787.

Anonymous:1908:SMM

[Ano08k] Anonymous. Stated meeting March 6, 1908. Proceedings of the American Philosophical Society, 47(188):13, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983791.

Anonymous:1908:SMNb

[Ano08l] Anonymous. Stated meeting November 20, 1908. Proceedings of the American Philosophical Society, 47(190):697, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983838.

Anonymous:1908:SMNa

[Ano08m] Anonymous. Stated meeting November 6, 1908. Proceedings of the American Philosophical Society, 47(190):645, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983835.

Anonymous:1908:SMO

[Ano08n] Anonymous. Stated meeting October 2, 1908. Proceedings of the American Philosophical Society, 47(190):644–645, September/ December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983834.

Anonymous:1908:VIa

[Ano08o] Anonymous. Volume information. Proceedings of the American Philosophical Society, 47(188):??, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983786.

Anonymous:1908:VIb

[Ano08p] Anonymous. Volume information. Proceedings of the American Philosophical Society, 47(190):699–702, September/December 1908. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983829.

Anonymous:1909:Aa

[Ano09a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 48(191):??, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1909:Ab

[Ano09b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 48(192):??, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1909:Ac

[Ano09c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 48(193):??, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1909:FM

[Ano09d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 48(193):i, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984058.

Anonymous:1909:M

[Ano09e] Anonymous. Minutes. Proceedings of the American Philosophical Society, 48(193):iii–xvi, September/December 1909. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984070.

Anonymous:1909:VIa

[Ano09f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 48(191):??, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983806.

Anonymous:1909:VIb

[Ano09g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 48(193):xvii–xx, September/December 1909. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984057.

Anonymous:1910:Aa

[Ano10a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 49(194):??, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1910:Ab

[Ano10b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 49(195):??, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1910:Ac

[Ano10c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 49(196):??, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1910:Ad

[Ano10d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 49(197):??, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1910:FM

[Ano10e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 49(197):i, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984088.

Anonymous:1910:M

[Ano10f] Anonymous. Minutes. Proceedings of the American Philosophical Society, 49(197):iii–xviii, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984094.

Anonymous:1910:VIa

[Ano10g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 49(194):1–2, January/June 1910. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983985.

Anonymous:1910:VIb

[Ano10h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 49(197):371–374, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984087.

Anonymous:1911:Aa

[Ano11a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 50(198):??, January/April 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1911:Ab

[Ano11b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 50(199):??, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1911:Ac

[Ano11c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 50(200):??, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1911:Ad

[Ano11d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 50(201):??, September/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1911:Ae

[Ano11e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 50(202):??, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1911:CST

[Ano11f] Anonymous. Corrigendum: A study of the tertiary floras of the Atlantic and Gulf Coastal Plain. Proceedings of the American Philosophical Society, 50(202):xv, October/December 1911. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984007. See [Ber11].

Anonymous:1911:FM

[Ano11g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 50(202):i, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984000.

Anonymous:1911:M

[Ano11h] Anonymous. Minutes. Proceedings of the American Philosophical Society, 50(202):iii–xiv, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984006.

Anonymous:1911:VIa

[Ano11i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 50(198):i–iv, January/April 1911. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983953.

Anonymous:1911:VIb

[Ano11j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 50(202):xvii–xx, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983999.

Anonymous:1912:Aa

[Ano12a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 51(203):??, January/March 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1912:Ab

[Ano12b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 51(204):??, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1912:Ac

[Ano12c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 51(205):??, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1912:Ad

[Ano12d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 51(206):??, August/September 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1912:Ae

[Ano12e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 51(207):??, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1912:C

[Ano12f] Anonymous. Corrigenda. Proceedings of the American Philosophical Society, 51(207):603, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984025.

Anonymous:1912:FM

[Ano12g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 51(207):i, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984016.

Anonymous:1912:M

[Ano12h] Anonymous. Minutes. Proceedings of the American Philosophical Society, 51(207):iii–xix, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984023.

Anonymous:1912:VIa

[Ano12i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 51(203):i–iv, January/March 1912. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983976.

Anonymous:1912:VIb

[Ano12j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 51(207):xxi–xxiii, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984015.

Anonymous:1913:Aa

[Ano13a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 52(208):??, January/April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1913:Ab

[Ano13b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 52(209):??, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1913:Ac

[Ano13c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 52(210):??, May/August 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1913:Ad

[Ano13d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 52(211):??, September/October 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1913:Ae

[Ano13e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 52(212):??, November/December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1913:C

[Ano13f] Anonymous. Corrigenda. Proceedings of the American Philosophical Society, 52(208):iv, January/April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983998.

Anonymous:1913:FM

[Ano13g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 52(212):i, November/December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984145.

Anonymous:1913:M

[Ano13h] Anonymous. Minutes. Proceedings of the American Philosophical Society, 52(212):iii–xviii, November/December 1913. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984150.

Anonymous:1913:VIa

[Ano13i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 52(208):i–iv, January/April 1913. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983994.

Anonymous:1913:VIb

[Ano13j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 52(212):xix–xxi, November/December 1913. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984144.

Anonymous:1914:Aa

[Ano14a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 53(213):??, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1914:Ab

[Ano14b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 53(214):??, June/July 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1914:Ac

[Ano14c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 53(215):??, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1914:FM

[Ano14d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 53(215):i, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984075.

Anonymous:1914:M

[Ano14e] Anonymous. Minutes. Proceedings of the American Philosophical Society, 53(215):iii–xxi, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984082.

Anonymous:1914:VIa

[Ano14f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 53(213):i–iii, January/May 1914. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984125.

Anonymous:1914:VIb

[Ano14g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 53(215):xxiii–xxvi, August/December 1914. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984074.

Anonymous:1915:Aa

[Ano15a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 54(216):??, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1915:Ab

[Ano15b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 54(217):??, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1915:Ac

[Ano15c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 54(218):??, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1915:Ad

[Ano15d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 54(219):??, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1915:Ae

[Ano15e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 54(220):??, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1915:FM

[Ano15f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 54(220):i, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984107.

Anonymous:1915:M

[Ano15g] Anonymous. Minutes. Proceedings of the American Philosophical Society, 54(220):iii–xv, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984112.

Anonymous:1915:VIa

[Ano15h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 54(216):i–iv, January/April 1915. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983904.

Anonymous:1915:VIb

[Ano15i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 54(220):xvii–xxi, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984106.

Anonymous:1916:Aa

[Ano16a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(??):??, ???? 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ab

[Ano16b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(1):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ac

[Ano16c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(2):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ad

[Ano16d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(3):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ae

[Ano16e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(4):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Af

[Ano16f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(5):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ag

[Ano16g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(6):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ah

[Ano16h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(7):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:Ai

[Ano16i] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 55(8):??, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1916:FS

[Ano16j] Anonymous. [Financial statement. 1915]. Proceedings of the American Philosophical Society, 55(??):3–16, ???? 1916. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984168.

Anonymous:1916:FMa

[Ano16k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 55(??):1–2, ???? 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984167.

Anonymous:1916:FMb

[Ano16l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 55(8):I–i, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984011.

Anonymous:1916:M

[Ano16m] Anonymous. Minutes. Proceedings of the American Philosophical Society, 55(8):iii–xv, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984014.

Anonymous:1916:VIa

[Ano16n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 55(1):i–iv, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984169.

Anonymous:1916:VIb

[Ano16o] Anonymous. Volume information. Proceedings of the American Philosophical Society, 55(8):xvii–xx, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984010.

Anonymous:1917:Aa

[Ano17a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(1):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Ab

[Ano17b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(2):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Ac

[Ano17c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(3):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Ad

[Ano17d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(4):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Ae

[Ano17e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(5):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Af

[Ano17f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(6):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:Ag

[Ano17g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 56(7):??, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1917:FMa

[Ano17h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 56(1):i, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983959.

Anonymous:1917:FMb

[Ano17i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 56(7):i, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983884.

Anonymous:1917:M

[Ano17j] Anonymous. Minutes. Proceedings of the American Philosophical Society, 56(7):iii–xix, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983890.

Anonymous:1917:VIa

[Ano17k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 56(1):i–iv, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983958.

Anonymous:1917:VIb

[Ano17l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 56(7):xxi–xxiii, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983883.

Anonymous:1918:Aa

[Ano18a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(1):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Ab

[Ano18b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(2):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Ac

[Ano18c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(3):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Ad

[Ano18d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(4):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Ae

[Ano18e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(5):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Af

[Ano18f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(6):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:Ag

[Ano18g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 57(7):??, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1918:FM

[Ano18h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 57(7):i, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984525.

Anonymous:1918:M

[Ano18i] Anonymous. Minutes. Proceedings of the American Philosophical Society, 57(7):iii–xiii, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984528.

Anonymous:1918:VIa

[Ano18j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 57(1):i–iv, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983879.

Anonymous:1918:VIb

[Ano18k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 57(7):xiv–xvi, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984524.

Anonymous:1919:Aa

[Ano19a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(1):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Ab

[Ano19b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(2):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Ac

[Ano19c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(3):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Ad

[Ano19d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(4):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Ae

[Ano19e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(5):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Af

[Ano19f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(6):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:Ag

[Ano19g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 58(7):??, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1919:FM

[Ano19h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 58(7):i, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984212.

Anonymous:1919:M

[Ano19i] Anonymous. Minutes. Proceedings of the American Philosophical Society, 58(7):iii–xiv, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984215.

Anonymous:1919:VIa

[Ano19j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 58(1):i–iv, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984477.

Anonymous:1919:VIb

[Ano19k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 58(7):xv–xviii, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984211.

Anonymous:1920:Aa

[Ano20a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(1):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:Ab

[Ano20b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(2):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:Ac

[Ano20c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(3):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:Ad

[Ano20d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(4):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:Ae

[Ano20e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(5):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:Af

[Ano20f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 59(6):??, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1920:FM

[Ano20g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 59(6):i, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984388.

Anonymous:1920:M

[Ano20h] Anonymous. Minutes. Proceedings of the American Philosophical Society, 59(6):iii–xv, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984392.

Anonymous:1920:VIa

[Ano20i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 59(1):i–iii, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984217.

Anonymous:1920:VIb

[Ano20j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 59(6):xvi–xix, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984387.

Anonymous:1921:Aa

[Ano21a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 60(1):??, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1921:Ab

[Ano21b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 60(2):??, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1921:Ac

[Ano21c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 60(3):??, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1921:Ad

[Ano21d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 60(4):??, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1921:FM

[Ano21e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 60(4):i, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984513.

Anonymous:1921:M

[Ano21f] Anonymous. Minutes. Proceedings of the American Philosophical Society, 60(4):iii–xxiv, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984523.

Anonymous:1921:VIa

[Ano21g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 60(1):i–iv, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984220.

Anonymous:1921:VIb

[Ano21h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 60(4):xxv–xxvii, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984512.

Anonymous:1922:Aa

[Ano22a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 61(1):??, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1922:Ab

[Ano22b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 61(2):??, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1922:Ac

[Ano22c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 61(3):??, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1922:Ad

[Ano22d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 61(4):??, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1922:C

[Ano22e] Anonymous. Corrigenda. Proceedings of the American Philosophical Society, 61(1):iv, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984429.

Anonymous:1922:FM

[Ano22f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 61(4):i, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984266.

Anonymous:1922:VIa

[Ano22g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 61(1):i–iv, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984428.

Anonymous:1922:VIb

[Ano22h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 61(4):289–291, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984265.

Anonymous:1923:Aa

[Ano23a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(1):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:Ab

[Ano23b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(2):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:Ac

[Ano23c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(3):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:Ad

[Ano23d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(4):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:Ae

[Ano23e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(5):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:Af

[Ano23f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 62(6):??, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1923:FMa

[Ano23g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 62(1):i, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984176.

Anonymous:1923:FMb

[Ano23h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 62(6):i, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984187.

Anonymous:1923:M

[Ano23i] Anonymous. Minutes. Proceedings of the American Philosophical Society, 62(6):iii–xii, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984189.

Anonymous:1923:VIa

[Ano23j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 62(1):i–iv, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984175.

Anonymous:1923:VIb

[Ano23k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 62(6):xiii–xv, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984186.

Anonymous:1924:Aa

[Ano24a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 63(1):??, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1924:Ab

[Ano24b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 63(2):??, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1924:Ac

[Ano24c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 63(3):??, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1924:FM

[Ano24d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 63(3):i, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984325.

Anonymous:1924:M

[Ano24e] Anonymous. Minutes. Proceedings of the American Philosophical Society, 63(3):iii–xiv, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984330.

Anonymous:1924:VIa

[Ano24f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 63(1):i–iv, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984436.

Anonymous:1924:VIb

[Ano24g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 63(3):317–320, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984324.

Anonymous:1925:Aa

[Ano25a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 64(1):??, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1925:Ab

[Ano25b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 64(2):??, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1925:Ac

[Ano25c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 64(3):??, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1925:FMa

[Ano25d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 64(1):??, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984298.

Anonymous:1925:FMb

[Ano25e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 64(3):i, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984449.

Anonymous:1925:M

[Ano25f] Anonymous. Minutes. Proceedings of the American Philosophical Society, 64(3):iii–xiii, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984450.

Anonymous:1925:VIa

[Ano25g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 64(1):i–iii, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984297.

Anonymous:1925:VIb

[Ano25h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 64(3):309–312, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984448.

Anonymous:1926:Aa

[Ano26a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(1):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:Ab

[Ano26b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(2):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:Ac

[Ano26c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(3):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:Ad

[Ano26d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(4):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:Ae

[Ano26e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(5):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:Af

[Ano26f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 65(5):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1926:BM

[Ano26g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 65(5):??, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984245.

Anonymous:1926:FMa

[Ano26h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 65(1):i, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984332.

Anonymous:1926:FMb

[Ano26i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 65(5):iii, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984281.

Anonymous:1926:MMA

[Ano26j] Anonymous. Minutes of the meetings of the American Philosophical Society during 1926. Proceedings of the American Philosophical Society, 65(5):v–xv, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984244.

Anonymous:1926:VIa

[Ano26k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 65(1):i–ii, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984331.

Anonymous:1926:VIb

[Ano26l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 65(5):i–103, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984280.

Anonymous:1926:VIc

[Ano26m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 65(5):iii–385, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984243.

Anonymous:1927:A

[Ano27a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 66(??):??, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1927:ASS

[Ano27b] Anonymous. Addresses from sister societies and other institutions of learning. Proceedings of the American Philosophical Society, 66 (??):681 + 683–714, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301088.

Anonymous:1927:BM

[Ano27c] Anonymous. Back matter. Proceedings of the American Philosophical Society, 66(??):715–ii, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301094.

Anonymous:1927:FM

[Ano27d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 66(??):i–17, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301055.

Anonymous:1927:M

[Ano27e] Anonymous. Minutes. Proceedings of the American Philosophical Society, 66(??):iii–xiii, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301093.

Anonymous:1927:VI

[Ano27f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 66(??):739–750, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301054.

Anonymous:1928:Aa

[Ano28a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 67(1):??, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1928:Ab

[Ano28b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 67(2):??, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1928:Ac

[Ano28c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 67(3):??, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1928:Ad

[Ano28d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 67(4):??, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1928:MMA

[Ano28e] Anonymous. Minutes of the meetings of the American Philosophical Society during 1928. Proceedings of the American Philosophical Society, 67(1):ix–xx, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984251.

Anonymous:1928:VIa

[Ano28f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 67(1):i–vii, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984250.

Anonymous:1928:VIb

[Ano28g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 67(4):379–384, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984246.

Anonymous:1929:Aa

[Ano29a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 68(1):??, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1929:Ab

[Ano29b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 68(2):??, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1929:Ac

[Ano29c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 68(3):??, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1929:Ad

[Ano29d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 68(4):??, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1929:MMA

[Ano29e] Anonymous. Minutes of the meetings of the American Philosophical Society during 1929. Proceedings of the American Philosophical Society, 68(1):ix–xix, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984394.

Anonymous:1929:VIa

[Ano29f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 68(1):i–vii, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984393.

Anonymous:1929:VIb

[Ano29g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 68(4):321–326, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984340.

Anonymous:1930:A

[Ano30a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 69(1):??, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1930:AMA

[Ano30b] Anonymous. Addresses made at the Annual Dinner of the American Philosophical Society Held at the Bellevue–Stratford Hotel, Philadelphia, Saturday, April 26, 1930, John H. Finley, Presiding. Proceedings of the American Philosophical Society, 69(1):554–573, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984381.

Anonymous:1930:AMD

[Ano30c] Anonymous. Addresses made at the dinner commemorating the one hundred and fiftieth anniversary of the granting of a charter to the American Philosophical Society Held at the Hotel Plaza, New York City, Wednesday Evening, April 2, 1930, the Honorable John W. Davis Presiding. Proceedings of the American Philosophical Society, 69(1):529–553, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984380.

Anonymous:1930:BM

[Ano30d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 69(1):575–580, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984382.

Anonymous:1930:FM

[Ano30e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 69(1):i–vi, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984344.

Anonymous:1930:MMA

[Ano30f] Anonymous. Minutes of the meetings of the American Philosophical Society during 1930. Proceedings of the American Philosophical Society, 69(1):vii–xix, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984345.

Anonymous:1931:Aa

[Ano31a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 70(1):??, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1931:Ab

[Ano31b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 70(2):??, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1931:Ac

[Ano31c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 70(3):??, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1931:Ad

[Ano31d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 70(4):??, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1931:Ae

[Ano31e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 70(5):??, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1931:MMA

[Ano31f] Anonymous. Minutes of the meetings of the American Philosophical Society during 1931. Proceedings of the American Philosophical Society, 70(1):vii–xviii, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984233.

Anonymous:1931:VIa

[Ano31g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 70(1):i–vi, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984232.

Anonymous:1931:VIb

[Ano31h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 70(5):403–410, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984493.

Anonymous:1932:AMM

[Ano32a] Anonymous. Abstracts of the minutes of the meetings of the American Philosophical Society during 1932. Proceedings of the American Philosophical Society, 71(1):ix–xx, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984815.

Anonymous:1932:Aa

[Ano32b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(1):??, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Ab

[Ano32c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(2):??, March 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Ac

[Ano32d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(3):??, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Ad

[Ano32e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(4):??, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Ae

[Ano32f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(5):??, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Af

[Ano32g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(6):??, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:Ag

[Ano32h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 71(7):??, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1932:VIa

[Ano32i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 71(1):i–vii, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984814.

Anonymous:1932:VIb

[Ano32j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 71(7):421–427, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984761.

Anonymous:1933:AMM

[Ano33a] Anonymous. Abstracts from the minutes of the meetings during 1933. Proceedings of the American Philosophical Society, 72(1):xi–xxii, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984604.

Anonymous:1933:Aa

[Ano33b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 72(1):??, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1933:Ab

[Ano33c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 72(2):??, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1933:Ac

[Ano33d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 72(3):??, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1933:Ad

[Ano33e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 72(4):??, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1933:Ae

[Ano33f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 72(5):??, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1933:VIa

[Ano33g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 72(1):i–ix, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984603.

Anonymous:1933:VIb

[Ano33h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 72(5):371–377, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984693.

Anonymous:1934:AMM

[Ano34a] Anonymous. Abstracts from the minutes of the meetings during 1934. Proceedings of the American Philosophical Society, 74(1):vii–xxiii, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984579.

Anonymous:1934:Aa

[Ano34b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(1):??, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ab

[Ano34c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(2):??, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ac

[Ano34d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(3):??, February 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ad

[Ano34e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(4):??, April 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ae

[Ano34f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(5):??, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Af

[Ano34g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 73(6):??, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ag

[Ano34h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 74(1):??, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ah

[Ano34i] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 74(2):??, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Ai

[Ano34j] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 74(3):??, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:Aj

[Ano34k] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 74(4):??, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1934:CDR

[Ano34l] Anonymous. [Corrections]: Different rates of growth among animals with special reference to the Odonata. Proceedings of the American Philosophical Society, 73(1):70, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984741. See [Cal29].

Anonymous:1934:O

[Ano34m] Anonymous. Officers. Proceedings of the American Philosophical Society, 73(6):371–397, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984652.

Anonymous:1934:VIa

[Ano34n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 73(1):i–viii, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984739.

Anonymous:1934:VIb

[Ano34o] Anonymous. Volume information. Proceedings of the American Philosophical Society, 73(6):399–400, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984651.

Anonymous:1934:VIc

[Ano34p] Anonymous. Volume information. Proceedings of the American Philosophical Society, 74(1):i–vi, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984578.

Anonymous:1934:VId

[Ano34q] Anonymous. Volume information. Proceedings of the American Philosophical Society, 74(4):319–323, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984793.

Anonymous:1935:AMM

[Ano35a] Anonymous. Abstracts from the minutes of the meetings during 1935. Proceedings of the American Philosophical Society, 75(1):ix– xxvi, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984720.

Anonymous:1935:Aa

[Ano35b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75:??, ???? 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ab

[Ano35c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(1):??, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ac

[Ano35d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(2):??, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ad

[Ano35e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(3):??, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ae

[Ano35f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(4):??, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Af

[Ano35g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(5):??, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ag

[Ano35h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(6):??, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ah

[Ano35i] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(7):??, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:Ai

[Ano35j] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 75(8):??, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1935:FM

[Ano35k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 75:??, ???? 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984746.

Anonymous:1935:IPA

[Ano35l] Anonymous. Index. Proceedings of the American Philosophical Society. Vol. 51 (1912) to Vol. 75 (1935). Proceedings of the American Philosophical Society, 75:1–71, ???? 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984747.

Anonymous:1935:VIa

[Ano35m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 75(1):i–viii, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984719.

Anonymous:1935:VIb

[Ano35n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 75(8):775–781, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984589.

Anonymous:1936:AMM

[Ano36a] Anonymous. Abstracts from the minutes of the meetings during 1936. Proceedings of the American Philosophical Society, 76(1):xv–xxxiii, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984707.

Anonymous:1936:Aa

[Ano36b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(1):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:Ab

[Ano36c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(2):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:Ac

[Ano36d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(3):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:Ad

[Ano36e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(4):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:Ae

[Ano36f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(5):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:Af

[Ano36g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 76(6):??, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1936:VIa

[Ano36h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 76(1):i–xiii, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984706.

Anonymous:1936:VIb

[Ano36i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 76(6):947–955, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984560.

Anonymous:1937:Aa

[Ano37a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 77(1):??, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:Ab

[Ano37b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 77(2):??, February 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:Ac

[Ano37c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 77(3):??, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:Ad

[Ano37d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 77(4):??, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:Ae

[Ano37e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 78(1):??, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:Af

[Ano37f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 78(2):??, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1937:RMS

[Ano37g] Anonymous. [Representatives and meeting schedule]. Proceedings of the American Philosophical Society, 77(4):561–565, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984777.

Anonymous:1937:VIa

[Ano37h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 77(1):i–vii, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984767.

Anonymous:1937:VIb

[Ano37i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 77(4):631–633, April 22, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984769.

Anonymous:1937:VIc

[Ano37j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 78(1):i–iv, October 22, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984653.

Anonymous:1938:Aa

[Ano38a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 78(3):??, January 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:Ab

[Ano38b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 78(4):??, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:Ac

[Ano38c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 79(1):??, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:Ad

[Ano38d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 79(2):??, June 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:Ae

[Ano38e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 79(3):??, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:Af

[Ano38f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 79(4):??, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1938:VIa

[Ano38g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 78(4):675–677, March 31, 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984799.

Anonymous:1938:VIb

[Ano38h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 79(1):i–iv, April 21, 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984673.

Anonymous:1938:VIc

[Ano38i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 79(4):741–742, November 15, 1938. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984937.

Anonymous:1939:Aa

[Ano39a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 80(1):??, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ab

[Ano39b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 80(2):??, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ac

[Ano39c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 80(3):??, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ad

[Ano39d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 80(4):??, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ae

[Ano39e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 81(1):??, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Af

[Ano39f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 81(2):??, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ag

[Ano39g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 81(3):??, August 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ah

[Ano39h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 81(4):??, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:Ai

[Ano39i] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 81(5):??, December 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1939:FMa

[Ano39j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 80(4):??, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985050.

Anonymous:1939:FMb

[Ano39k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 81(2):??, June 30, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984977.

Anonymous:1939:VIa

[Ano39l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 80(1):i–iv, January 20, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984968.

Anonymous:1939:VIb

[Ano39m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 80(4):627, February 15, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985049.

Anonymous:1939:VIc

[Ano39n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 81(1):i–iv, May 31, 1939. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984960.

Anonymous:1939:VId

[Ano39o] Anonymous. Volume information. Proceedings of the American Philosophical Society, 81(5):677–678, December 31, 1939. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985009.

Anonymous:1940:Aa

[Ano40a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 82(1):??, February 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ab

[Ano40b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 82(2):??, March 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ac

[Ano40c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 82(3):??, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ad

[Ano40d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 82(4):??, May 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ae

[Ano40e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 82(5):??, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Af

[Ano40f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 83(1):??, July 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ag

[Ano40g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 83(2):??, August 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ah

[Ano40h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 83(3):??, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Ai

[Ano40i] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 83(4):??, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:Aj

[Ano40j] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 83(5):??, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1940:D

[Ano40k] Anonymous. [Discussion]. Proceedings of the American Philosophical Society, 83(4):583–588, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984839.

Anonymous:1940:FM

[Ano40l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 83(4):??, September 20, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984830.

Anonymous:1940:R

[Ano40m] Anonymous. Remarks. Proceedings of the American Philosophical Society, 83(4):515–516, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984831.

Anonymous:1940:VIa

[Ano40n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 82(1):i–v, February 23, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984901.

Anonymous:1940:VIb

[Ano40o] Anonymous. Volume information. Proceedings of the American Philosophical Society, 82(5):949–950, June 29, 1940. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984882.

Anonymous:1940:VIc

[Ano40p] Anonymous. Volume information. Proceedings of the American Philosophical Society, 83(1):i–iv, July 20, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985118.

Anonymous:1940:VId

[Ano40q] Anonymous. Volume information. Proceedings of the American Philosophical Society, 83(5):747–748, October 31, 1940. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985129.

Anonymous:1941:Aa

[Ano41a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(1):??, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Ab

[Ano41b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(2):??, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Ac

[Ano41c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(3):??, May 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Ad

[Ano41d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(4):??, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Ae

[Ano41e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(5):??, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Af

[Ano41f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 84(6):??, August 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:Ag

[Ano41g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 85(1):??, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1941:CAD

[Ano41h] Anonymous. Career of Alexander Dallas Bache. Proceedings of the American Philosophical Society, 84(2):??, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984911.

Anonymous:1941:FMa

[Ano41i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 84(2):??, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984909.

Anonymous:1941:FMb

[Ano41j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 84(4):??, June 30, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984950.

Anonymous:1941:GAD

[Ano41k] Anonymous. Genealogy of Alexander Dallas Bache. Proceedings of the American Philosophical Society, 84(2):??, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984910.

Anonymous:1941:VIa

[Ano41l] Anonymous. Volume information. Proceedings of the American Philosophical Society, 84(1):i–iv, April 25, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984847.

Anonymous:1941:VIb

[Ano41m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 84(6):903–904, August 26, 1941. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984875.

Anonymous:1941:VIc

[Ano41n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 85(1):??, November 17, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985120.

Anonymous:1942:Aa

[Ano42a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 85(2):??, January 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1942:Ab

[Ano42b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 85(3):??, February 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1942:Ac

[Ano42c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 85(4):??, June 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1942:Ad

[Ano42d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 85(5):??, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1942:Ae

[Ano42e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 86(1):??, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1942:FM

[Ano42f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 86(1):i–iv, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985070.

Anonymous:1942:VIa

[Ano42g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 85(5):505–506, September 30, 1942. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985061.

Anonymous:1942:VIb

[Ano42h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 86(1):i–iv, September 25, 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985069.

Anonymous:1943:Aa

[Ano43a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 86(2):??, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1943:Ab

[Ano43b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 86(3):??, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1943:Ac

[Ano43c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 87(1):??, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1943:Ad

[Ano43d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 87(2):??, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1943:Ae

[Ano43e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 87(3):??, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1943:FMa

[Ano43f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 86(2):i–iii, February 10, 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985098.

Anonymous:1943:FMb

[Ano43g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 86(3):i–iii, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985021.

Anonymous:1943:VIa

[Ano43h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 86(3):477–478, July 8, 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985020.

Anonymous:1943:VIb

[Ano43i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 87(1):i–iv, July 14, 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984993.

Anonymous:1944:Aa

[Ano44a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 87(4):??, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ab

[Ano44b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 87(5):??, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ac

[Ano44c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(1):??, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ad

[Ano44d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(2):??, July 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ae

[Ano44e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(3):??, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Af

[Ano44f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(4):??, October 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ag

[Ano44g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(5):??, November 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:Ah

[Ano44h] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 88(6):??, December 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1944:VIa

[Ano44i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 87(5):459–460, May 5, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985286.

Anonymous:1944:VIb

[Ano44j] Anonymous. Volume information. Proceedings of the American Philosophical Society, 88(1):i–iii, June 16, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985182.

Anonymous:1944:VIc

[Ano44k] Anonymous. Volume information. Proceedings of the American Philosophical Society, 88(6):535–536, December 28, 1944. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985238.

Anonymous:1945:Aa

[Ano45a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 89(1):??, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1945:Ab

[Ano45b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 89(2):??, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1945:Ac

[Ano45c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 89(3):??, October 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1945:Ad

[Ano45d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 89(4):??, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1945:FM

[Ano45e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 89(1):i–vi, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985308.

Anonymous:1945:VIa

[Ano45f] Anonymous. Volume information. Proceedings of the American Philosophical Society, 89(1):i–v, April 30, 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985307.

Anonymous:1945:VIb

[Ano45g] Anonymous. Volume information. Proceedings of the American Philosophical Society, 89(4):627–628, December 10, 1945. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985359.

Anonymous:1946:Aa

[Ano46a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 90(1):??, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1946:Ab

[Ano46b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 90(2):??, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1946:Ac

[Ano46c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 90(3):??, July 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1946:Ad

[Ano46d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 90(4):??, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1946:Ae

[Ano46e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 90(5):??, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1946:FMa

[Ano46f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 90(1):i–iii, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301030.

Anonymous:1946:FMb

[Ano46g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 90(4):??, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301016.

Anonymous:1946:VIa

[Ano46h] Anonymous. Volume information. Proceedings of the American Philosophical Society, 90(1):i–iv, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301029.

Anonymous:1946:VIb

[Ano46i] Anonymous. Volume information. Proceedings of the American Philosophical Society, 90(5):432–433, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301044.

Anonymous:1947:Aa

[Ano47a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 91(1):??, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1947:Ab

[Ano47b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 91(2):??, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1947:Ac

[Ano47c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 91(3):??, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1947:Ad

[Ano47d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 91(4):??, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1947:Ae

[Ano47e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 91(5):??, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1947:BMa

[Ano47f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 91(1):??, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143137.

Anonymous:1947:BMb

[Ano47g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 91(2):??, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143606.

Anonymous:1947:BMc

[Ano47h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 91(3):??, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143403.

Anonymous:1947:BMd

[Ano47i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 91(4):??, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143347.

Anonymous:1947:BMe

[Ano47j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 91(5):??, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143426.

Anonymous:1947:FMa

[Ano47k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 91(1):??, February 25, 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143117.

Anonymous:1947:FMb

[Ano47l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 91(2):??, April 5, 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143597.

Anonymous:1947:FMc

[Ano47m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 91(3):??, August 29, 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143394.

Anonymous:1947:FMd

[Ano47n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 91(4):??, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143342.

Anonymous:1947:FMe

[Ano47o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 91(5):??, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143420.

Anonymous:1947:VIa

[Ano47p] Anonymous. Volume information. Proceedings of the American Philosophical Society, 91(4):402–404, October 24, 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143341.

Anonymous:1947:VIb

[Ano47q] Anonymous. Volume information. Proceedings of the American Philosophical Society, 91(5):457–iv, December 3, 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143419.

Anonymous:1948:Aa

[Ano48a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(1):??, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:Ab

[Ano48b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(2):??, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:Ac

[Ano48c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(3):??, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:Ad

[Ano48d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(4):??, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:Ae

[Ano48e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(5):??, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:Af

[Ano48f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 92(6):??, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1948:BMa

[Ano48g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(1):??, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143630.

Anonymous:1948:BMb

[Ano48h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(2):??, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143414.

Anonymous:1948:BMc

[Ano48i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(3):??, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143589.

Anonymous:1948:BMd

[Ano48j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(4):??, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143362.

Anonymous:1948:BMe

[Ano48k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(5):??, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143054.

Anonymous:1948:BMf

[Ano48l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 92(6):??, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143200.

Anonymous:1948:FMa

[Ano48m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(1):??, March 8, 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143618.

Anonymous:1948:FMb

[Ano48n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(2):??, May 5, 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143404.

Anonymous:1948:FMc

[Ano48o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(3):??, July 19, 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143579.

Anonymous:1948:FMd

[Ano48p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(4):??, October 25, 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143348.

Anonymous:1948:FMe

[Ano48q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(5):??, November 12, 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143041.

Anonymous:1948:FMf

[Ano48r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 92(6):??, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143195.

Anonymous:1948:VI

[Ano48s] Anonymous. Volume information. Proceedings of the American Philosophical Society, 92(6):504–iv, December 27, 1948. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143194.

Anonymous:1949:Aa

[Ano49a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(1):??, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Ab

[Ano49b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(2):??, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Ac

[Ano49c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(3):??, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Ad

[Ano49d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(4):??, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Ae

[Ano49e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(5):??, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Af

[Ano49f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(6):??, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:Ag

[Ano49g] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 93(7):??, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1949:BMa

[Ano49h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(1):??, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143274.

Anonymous:1949:BMb

[Ano49i] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(2):??, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143439.

Anonymous:1949:BMc

[Ano49j] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(3):??, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143478.

Anonymous:1949:BMd

[Ano49k] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(4):??, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143162.

Anonymous:1949:BMe

[Ano49l] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(5):??, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143393.

Anonymous:1949:BMf

[Ano49m] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(6):??, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143340.

Anonymous:1949:BMg

[Ano49n] Anonymous. Back matter. Proceedings of the American Philosophical Society, 93(7):??, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143145.

Anonymous:1949:FMa

[Ano49o] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(1):??, April 18, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143266.

Anonymous:1949:FMb

[Ano49p] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(2):??, May 16, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143427.

Anonymous:1949:FMc

[Ano49q] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(3):??, June 10, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143466.

Anonymous:1949:FMd

[Ano49r] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(4):??, September 9, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143153.

Anonymous:1949:FMe

[Ano49s] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(5):??, November 30, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143386.

Anonymous:1949:FMf

[Ano49t] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(6):??, December 29, 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143334.

Anonymous:1949:FMg

[Ano49u] Anonymous. Front matter. Proceedings of the American Philosophical Society, 93(7):??, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143139.

Anonymous:1949:VI

[Ano49v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 93(7):546–iv, December 30, 1949. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143138.

Arthur:1925:GRS

[Art25] J. C. Arthur. The grass rusts of South America; based on the Holway Collections. Proceedings of the American Philosophical Society, 64(2): 131–223, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984193.

Ashmead:1903:THM

[Ash03] Albert S. Ashmead. Testimony of the Huacos (mummy-grave) potteries of Old Peru. Proceedings of the American Philosophical Society, 42(174):378–396, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983720.

Ashford:1919:ASS

[Ash19] Bailey K. Ashford. The application of sanitary science to the Great War in the zone of the Army. Proceedings of the American Philosophical Society, 58(5):317–336, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984197.

Atkinson:1915:MDA

[Atk15] Geo. F. Atkinson. Morphology and development of Agaricus Rodmani. Proceedings of the American Philosophical Society, 54(219): 309–343, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983877.

Atkinson:1918:GGN

[Atk18a] Geo. F. Atkinson. The genus Galerula in North America. Proceedings of the American Philosophical Society, 57(5):357–374, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984179.

Atkinson:1918:PNS

[Atk18b] Geo. F. Atkinson. Preliminary notes on some new species of agarics. Proceedings of the American Philosophical Society, 57(4):354–356, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983942.

Atkinson:1918:THC

[Atk18c] George F. Atkinson. Twin hybrids from crosses of Œnothera Lamarckiana and Franciscana with Œ. Pycnocarpa, in the F_{1} and F_{2}. Proceedings of the American Philosophical Society, 57(2):130–143, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984164.

Austin:1913:RRE

[Aus13] L. W. Austin. Radiated and received energy in radiotelegraphy. Proceedings of the American Philosophical Society, 52(210): 407–419, May/August 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984085.

Aydelotte:1930:OHE

[Ayd30] Frank Aydelotte. The outlook for higher education. Proceedings of the American Philosophical Society, 69(1):271–280, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984362.

Bache:1936:WFF

[Bac36] Franklin Bache. Where is Franklin’s first chart of the Gulf Stream? Proceedings of the American Philosophical Society, 76(6):731–741, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984562.

Badger:1936:RPI

[Bad36] Richard M. Badger. Researches in the photographic infrared. Proceedings of the American Philosophical Society, 76(6):776–779, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984570.

Bailey:1903:FMP

[Bai03] L. H. Bailey. The forward movement in plant-breeding. Proceedings of the American Philosophical Society, 42(172):54–68, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983642.

Bailey:1915:SPN

[Bai15] L. H. Bailey. Some present needs in systematic botany. Proceedings of the American Philosophical Society, 54(216):58–65, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983913.

Bailey:1947:BST

[Bai47] Liberty Hyde Bailey. Botanical sciences and their applications, including agriculture. Proceedings of the American Philosophical Society, 91(1):22–26, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143121.

Baker:1940:FFF

[Bak40a] James G. Baker. A family of flat-field cameras, equivalent in performance to the Schmidt camera. Proceedings of the American Philosophical Society, 82(3):339–349, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984933.

Baker:1940:SGS

[Bak40b] James G. Baker. The solid-glass Schmidt camera and a new type nebular spectrograph. Proceedings of the American Philosophical Society, 82(3):323–338, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984932.

Bakke:1948:IRR

[Bak48] E. Wight Bakke. Industrial relations research. Proceedings of the American Philosophical Society, 92(5):376–380, November 1948. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143049.

Balch:1902:FAA

[Bal02] Thomas Willing Balch. The first “Assembly Account” — Philadelphia, 1748. Proceedings of the American Philosophical Society, 41 (170):260–264, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983557.

Balch:1907:PIT

[Bal07] Thomas Willing Balch. The proposed international tribunal of arbitration of 1623. Proceedings of the American Philosophical Society, 46(186):302–311, April/September 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983742.

Balch:1908:AE

[Bal08a] Edwin Swift Balch. Art and ethnology. Proceedings of the American Philosophical Society, 47(188):30–36, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983794.

Balch:1908:LOC

[Bal08b] Thomas Willing Balch. The Law of Oresme, Copernicus and Gresham. Proceedings of the American Philosophical Society, 47 (188):18–29, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983793.

Balch:1909:WAS

[Bal09a] Edwin Swift Balch. Why America should re-explore Wilkes Land. Proceedings of the American Philosophical Society, 48(191):34–50, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
983809.

Balch:1909:ABA

[Bal09b] Thomas Willing Balch. The American–British Atlantic fisheries question. Proceedings of the American Philosophical Society, 48(193): 319–353, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984059.

Balch:1912:SFM

[Bal12] Thomas Willing Balch. Some former members of the American Philosophical Society. Proceedings of the American Philosophical Society, 51(207):580–601, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984022.

Balch:1915:SGP

[Bal15] Thomas Willing Balch. The Swedes, Governor Printz and the beginning of Pennsylvania. Proceedings of the American Philosophical Society, 54(216):12–13, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983907.

Balch:1916:LPI

[Bal16] Thomas Willing Balch. Legal and political international questions and the recurrence of war. Proceedings of the American Philosophical Society, 55(3):267–280, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984117.

Balch:1917:EMA

[Bal17] Edwin Swift Balch. Early man in America. Proceedings of the American Philosophical Society, 56(6):473–483, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983969.

Balch:1918:AGC

[Bal18] Edwin Swift Balch. The art of George Catlin. Proceedings of the American Philosophical Society, 57(2):144–154, 1918. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984165.

Balch:1919:EMD

[Bal19] Edwin Swift Balch. Evolution and mystery in the discovery of America. Proceedings of the American Philosophical Society, 58(1):55–73, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984480.

Balch:1921:CIM

[Bal21] Edwin Swift Balch. The Coudersport Ice Mine. Proceedings of the American Philosophical Society, 60(4):553–559, 1921. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984520.

Balch:1923:PM

[Bal23] Edwin Swift Balch. Prehistoric misnomers. Proceedings of the American Philosophical Society, 62(2):57–62, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984510.

Balch:1925:PT

[Bal25] Edwin Swift Balch. Prehistoric terminology. Proceedings of the American Philosophical Society, 64(1):83–88, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984306.

Barker:1901:FAG

[Bar01] George F. Barker. Frederick Augustus Genth. Proceedings of the American Philosophical Society, 40(167):x–xxii, December 1901. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983764.

Barnard:1907:SRA

[Bar07a] E. E. Barnard. Some of the results of astronomical photography pertaining specially to the work with a portrait lens. Proceedings of the American Philosophical Society, 46(187):417–429, October/ December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983482.

Barus:1907:DND

[Bar07b] C. Barus. On distributions of nuclei in dust-free wet air, and on methods of observation. Proceedings of the American Philosophical Society, 46(185):70–79, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983440.

Barus:1909:EMV

[Bar09a] Carl Barus. The electrometric measurement of the voltaic potential difference, between the two conductors of a condenser, containing a highly ionized medium. Proceedings of the American Philosophical Society, 48(192):189–193, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984155.

Barus:1909:EMS

[Bar09b] Carl Barus. The electron method of standardizing the coronas of cloudy condensation. Proceedings of the American Philosophical Society, 48(192):177–188, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984154.

Barus:1909:APG

[Bar09c] Carl Barus. On an adjustment for the plane grating similar to Rowland’s method for the concave grating. Proceedings of the American Philosophical Society, 48(192):166–176, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984153.

Barnard:1910:POD

[Bar10] E. E. Barnard. Photographic observations of Daniel’s Comet. Proceedings of the American Philosophical Society, 49(194):3–16, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983986.

Barnard:1911:SLN

[Bar11a] E. E. Barnard. Self-luminous night haze. Proceedings of the American Philosophical Society, 50(199):246–253, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984039.

Barus:1911:EIR

[Bar11b] C. Barus. Elliptic interference with reflecting grating. Proceedings of the American Philosophical Society, 50(198):125–138, January/April 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983956.

Barus:1911:TAT

[Bar11c] C. Barus. The transpiration of air through a partition of water. Proceedings of the American Philosophical Society, 50(198):117–124, January/April 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983955.

Barrett:1912:NPN

[Bar12] Storrs B. Barrett. Notes on plates of Nova Geminorum of 1912 taken with the Bruce spectrograph of the Yerkes Observatory. Proceedings of the American Philosophical Society, 51(207):568, October/ December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984020.

Barton:1913:HVP

[Bar13] George A. Barton. The historical value of the patriarchal narratives. Proceedings of the American Philosophical Society, 52(209):184–200, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983865.

Barton:1917:NBA

[Bar17] George A. Barton. A new Babylonian account of the creation of man. Proceedings of the American Philosophical Society, 56(3):275– 280, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983951.

Barnard:1919:SLN

[Bar19] E. E. Barnard. Self-luminous night haze. Proceedings of the American Philosophical Society, 58(3):223–235, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984470.

Barnard:1922:SPN

[Bar22] E. E. Barnard. Some peculiarities of the novæ. Proceedings of the American Philosophical Society, 61(2):99–106, 1922. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984407.

Barton:1926:PSH

[Bar26] George A. Barton. The present status of the Hittite problem. Proceedings of the American Philosophical Society, 65(3):232–243, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984492.

Barton:1929:OCA

[Bar29] George A. Barton. The origins of civilization in Africa and Mesopotamia, their relative antiquity and interplay. Proceedings of the American Philosophical Society, 68(4):303–312, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984342.

Bartlett:1940:RWE

[Bar40a] Harley Harris Bartlett. The reports of the Wilkes Expedition, and the work of the specialists in science. Proceedings of the American Philosophical Society, 82(5):601–705, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984888.

Bartlett:1940:PEE

[Bar40b] Robert A. Bartlett. Peary’s extended exploration of Arctic lands culminating in the attainment of the North Pole. Proceedings of the American Philosophical Society, 82(5):935–947, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984900.

Barton:1940:POB

[Bar40c] George A. Barton. The palæolithic beginnings of religion — an interpretation. Proceedings of the American Philosophical Society, 82(2):131–149, March 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985012.

Bartsch:1945:MUS

[Bar45] Paul Bartsch. Mollusks of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):294, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985332.

Barbeau:1949:HHW

[Bar49a] Marius Barbeau. How the Huron–Wyandot language was saved from oblivion. Proceedings of the American Philosophical Society, 93(3):226–232, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143470.

Barton:1949:QC

[Bar49b] Samuel G. Barton. The Quaker calendar. Proceedings of the American Philosophical Society, 93(1):32–39, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143268.

Baskerville:1915:REE

[Bas15] Chas. Baskerville. On the rate of evaporation of ether from oils and its application in oil-ether colonic anesthesia. Proceedings of the American Philosophical Society, 54(218):270–278, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983939.

Bates:1949:AMT

[Bat49] Marston Bates. The adaptations of mosquitoes to the tropical rain forest environment. Proceedings of the American Philosophical Society, 93(4):340–346, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143161.

Bauer:1910:SAT

[Bau10] L. A. Bauer. Solar activity and terrestrial magnetic disturbances. Proceedings of the American Philosophical Society, 49(194):130– 144, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983993.

Bauer:1912:NMC

[Bau12] L. A. Bauer. New magnetic charts of the Indian Ocean. Proceedings of the American Philosophical Society, 51(205):240–241, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984141.

Bauer:1915:GRW

[Bau15] L. A. Bauer. General results of the work in atmospheric electricity aboard the Carnegie, 1909–1914. Proceedings of the American Philosophical Society, 54(216):14–17, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983908.

Bauer:1917:RCA

[Bau17] Louis A. Bauer. Remarks on the compass in aëronautics. Proceedings of the American Philosophical Society, 56(3):255–257, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983949.

Bazett:1926:TS

[Baz26] Henry C. Bazett. Temperature sensations. Proceedings of the American Philosophical Society, 65(2):131–135, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984420.

Blakeslee:1930:ITB

[BB30] Albert F. Blakeslee and Howard J. Banker. Identical twins as biological controls in educational and other human problems. Proceedings of the American Philosophical Society, 69(1):379–384, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984371.

Bradley:1938:GSW

[BBC^{+}38] W. H. Bradley, M. N. Bramlette, J. A. Cushman, L. G. Henbest, K. E. Lohman, and P. D. Trask. The Geological Survey’s work on the Piggot North Atlantic deep sea cores. Proceedings of the American Philosophical Society, 79(1):41–46, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984678.

Beams:1936:EPH

[Bea36] J. W. Beams. Experiments on the production of high velocity ions by impulse methods. Proceedings of the American Philosophical Society, 76(6):771–772, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984567.

Bean:1939:RR

[Bea39] Donald Bean. The riddle in research. Proceedings of the American Philosophical Society, 80(1):37–47, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984971.

Beadle:1946:RFP

[Bea46] George Wells Beadle. R. A. F. Penrose, Jr., Memorial Lecture. The Gene. Proceedings of the American Philosophical Society, 90 (5):422–431, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301048.

Beck:1930:CCC

[Bec30] James M. Beck. The changed conception of the constitution. Proceedings of the American Philosophical Society, 69(1):99–115, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984350.

Beck:1931:TS

[Bec31] James M. Beck. Thou shalt not. Proceedings of the American Philosophical Society, 70(3):297–308, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984458.

Becker:1943:WSL

[Bec43] Carl Becker. What is still living in the political philosophy of Thomas Jefferson? Proceedings of the American Philosophical Society, 87(3):201–210, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984866.

Beebe:1939:ACC

[Bee39] Ralph A. Beebe. Adsorption calorimetry on catalytic materials and an account of some measurements on chromic oxide at −183^{∘}C. Proceedings of the American Philosophical Society, 81(1):1–14, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984961.

Bell:1945:ABE

[Bel45] Harold Idris Bell. The Arabic bilingual entagion. Proceedings of the American Philosophical Society, 89(3):531–542, October 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985433.

Bell:1948:SEP

[Bel48] Whitfield J. Bell, Jr. The scientific environment of Philadelphia, 1775–1790. Proceedings of the American Philosophical Society, 92(1): 6–14, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143620.

Bell:1949:BOB

[Bel49] Whitfield J. Bell, Jr. A box of old bones: A note on the identification of the mastodon, 1766–1806. Proceedings of the American Philosophical Society, 93(2):169–177, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143437.

Benedict:1910:IMM

[Ben10] Francis G. Benedict. The influence of mental and muscular work on nutritive processes. Proceedings of the American Philosophical Society, 49(195):145–163, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983891.

Benedict:1918:PEP

[Ben18] Francis G. Benedict. Physiological effects of a prolonged reduction in diet on twenty-five men. Proceedings of the American Philosophical Society, 57(5):479–490, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984182.

Benedict:1924:PFP

[Ben24] Francis G. Benedict. Physical factors in predicting the basal metabolism of girls. Proceedings of the American Philosophical Society, 63(1):25–56, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984440.

Benedict:1936:LBM

[Ben36] Francis G. Benedict. Lafayette Benedict Mendel. Proceedings of the American Philosophical Society, 76(3):394–395, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984552.

Benedict:1937:RFH

[Ben37] Francis G. Benedict. Race: A factor in human metabolism. Proceedings of the American Philosophical Society, 78(1):101–110, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984660.

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society, 78(4):551–572, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984802.

Bennett:1943:PCA

[Ben43] Wendell C. Bennett. The position of Chavin in Andean sequences. Proceedings of the American Philosophical Society, 86(2):323–327, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985112.

Berry:1911:STF

[Ber11] Edward W. Berry. A study of the tertiary floras of the Atlantic and Gulf Coastal Plain. Proceedings of the American Philosophical Society, 50(199):301–315, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www. jstor.org/stable/984046. See corrigendum [Ano11f].

Berry:1914:ADL

[Ber14] Edward Wilber Berry. The affinities and distribution of the lower Eocene flora of Southeastern North America. Proceedings of the American Philosophical Society, 53(214):129–250, June/July 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984055.

Berry:1922:PEU

[Ber22] Edward W. Berry. A possible explanation of Upper Eocene climates. Proceedings of the American Philosophical Society, 61(1):1–14, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984430.

Berkey:1933:RDG

[Ber33] Charles P. Berkey. Recent development of geology as an applied science. Proceedings of the American Philosophical Society, 72(1):25– 37, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984606.

Berry:1935:TPB

[Ber35] Edward W. Berry. Tertiary plants from Brazil. Proceedings of the American Philosophical Society, 75(7):565–590, 1935. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984809.

Berry:1937:TVR

[Ber37] George Packer Berry. The transformation of the virus of rabbit fibroma (shope) into that of infectious myxomatosis (Sanarelli). Proceedings of the American Philosophical Society, 77(4):473–476, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984774.

Berry:1939:MCR

[Ber39a] Charles T. Berry. More complete remains of Ophiura Marylandica. Proceedings of the American Philosophical Society, 80(1):87–94, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984975.

Berry:1939:GPL

[Ber39b] Edward W. Berry. Geology and paleontology of Lake Tacarigua, Venezuela. Proceedings of the American Philosophical Society, 81 (4):547–568, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984826.

Berkner:1941:CIR

[Ber41] L. V. Berkner. Contributions of ionospheric research to geomagnetism. Proceedings of the American Philosophical Society, 84(2): 309–321, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984925.

Berlin:1945:GSA

[Ber45] Leonard M. Berlin. Ground surveys accomplished at West Base, Antarctica. Proceedings of the American Philosophical Society, 89(1): 386, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985348.

Betts:1944:CBC

[Bet44] Edwin M. Betts. The correspondence between Constantine Samuel Rafinesque and Thomas Jefferson. Proceedings of the American Philosophical Society, 87(5):368–380, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985288.

Betts:1946:GPU

[Bet46] Edwin M. Betts. Groundplans and prints of the University of Virginia, 1822–1826. Proceedings of the American Philosophical Society, 90(2): 81–90, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301049.

Benedict:1927:GML

[BF27] Francis G. Benedict and Edward L. Fox. The gaseous metabolism of large wild birds under aviary life. Proceedings of the American Philosophical Society, 66(??):511–534, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301081.

Biddlecombe:1928:DHT

[Bid28] C. H. Biddlecombe. The development of the heavier-than-air machine. Proceedings of the American Philosophical Society, 67(3):297– 305, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984295.

Biddle:1947:NT

[Bid47] Francis Biddle. The Nürnberg Trial. Proceedings of the American Philosophical Society, 91(3):294–302, August 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143402.

Bieber:1949:PAG

[Bie49] Margarete Bieber. The portraits of Alexander the Great. Proceedings of the American Philosophical Society, 93(5):373–421 + 423–427, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143389.

Binder:1948:IEA

[Bin48] Carroll Binder. The impact of external affairs upon American public opinion. Proceedings of the American Philosophical Society, 92(4):249–256, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143353.

Bissonnette:1941:MP

[Bis41] Thomas Hume Bissonnette. The “mule” pheasant. Proceedings of the American Philosophical Society, 85(1):49–70, November 1941. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985125.

Benedict:1919:ELY

[BJ19] Francis G. Benedict and Alice Johnson. The energy loss of young women during the muscular activity of light household work. Proceedings of the American Philosophical Society, 58(1):89–96, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984482.

Birge:1927:OCW

[BJ27] Edward A. Birge and Chancey Juday. The organic content of the water of small lakes. Proceedings of the American Philosophical Society, 66(??):357–372, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301073.

Benedict:1936:HRE

[BL36a] Francis G. Benedict and Robert C. Lee. The heart rate of the elephant. Proceedings of the American Philosophical Society, 76(3):335– 341, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984548.

Berry:1936:NTP

[BL36b] Charles T. Berry and W. Gardner Lynn. A new turtle, Peritresius Virginianus, from the Miocene of Virginia. Proceedings of the American Philosophical Society, 76(2):175–190, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984666.

Blair:1909:EUA

[Bla09] William R. Blair. The exploration of the upper air by means of kites and balloons. Proceedings of the American Philosophical Society, 48 (191):8–33, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983808.

Blair:1913:DUV

[Bla13] Andrew A. Blair. The determination of uranium and vanadium in the carnotite ores of Colorado and Utah. Proceedings of the American Philosophical Society, 52(209):201–205, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983866.

Blair:1917:A

[Bla17] William R. Blair. Aërology. Proceedings of the American Philosophical Society, 56(3):189–211, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983946.

Blair:1928:MA

[Bla28] William R. Blair. Meteorology for aviation. Proceedings of the American Philosophical Society, 67(3):287–296, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984294.

Black:1936:SSC

[Bla36] Matthew W. Black. Shakespeare’s Seventeenth Century editors. Proceedings of the American Philosophical Society, 76(5):707–717, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984756.

Black:1944:SVT

[Bla44] L. M. Black. Some viruses transmitted by agallian leafhoppers. Proceedings of the American Philosophical Society, 88(2):132–144, July 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985667.

Black:1945:RFS

[Bla45a] John D. Black. The role of federal, state, and local governments in promoting forestry. Proceedings of the American Philosophical Society, 89(2):434–443, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985640.

Black:1945:GOE

[Bla45b] Richard B. Black. Geographical operations from East Base, United States Antarctic Service Expedition, 1939–1941. Proceedings of the American Philosophical Society, 89(1):4–12, April 1945. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985311.

Blakeslee:1945:RSB

[Bla45c] Albert F. Blakeslee. Removing some of the barriers to crossability in plants. Proceedings of the American Philosophical Society, 89(4):561–574, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985361.

Bleakney:1936:IAM

[Ble36] Walker Bleakney. Isotope analysis with the mass-spectrograph. Proceedings of the American Philosophical Society, 76(6):774–776, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984569.

Bleakney:1937:RAI

[Ble37] Walker Bleakney. The relative abundance of isotopes. Proceedings of the American Philosophical Society, 77(3):395–409, March 1937.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984734.

Bloomfield:1913:CAM

[Blo13] Maurice Bloomfield. The character and adventures of Muladeva. Proceedings of the American Philosophical Society, 52(212):616– 650, November/December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984148.

Bloomfield:1917:AEA

[Blo17] Maurice Bloomfield. On the art of entering another’s body: A Hindu fiction motif. Proceedings of the American Philosophical Society, 56 (1):1–43, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983960.

Bloch:1944:FPS

[Blo44a] Henry Simon Bloch. Fiscal policy and social reform as it may affect the potentialities of the personal income tax. Proceedings of the American Philosophical Society, 88(1):36–42, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985187.

Blough:1944:CHT

[Blo44b] Roy Blough. Conflict and harmony in taxation. Proceedings of the American Philosophical Society, 88(1):22–28, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985185.

Bonno:1947:DIL

[BLV47] Gabriel Bonno, Locke, and Voltaire. The diffusion and influence of Locke’s “Essay concerning Human Understanding” in France before Voltaire’s “Lettres Philosophiques”. Proceedings of the American Philosophical Society, 91(5):421–425, December 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143422.

Benedict:1932:BHP

[BM32] Francis G. Benedict and Mary Henderson Meyer. The basal heat production of elderly women. Proceedings of the American Philosophical Society, 71(4):143–165, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984786.

Barrows:1938:NAC

[BM38] W. M. Barrows, Jr. and J. R. Murlin. New automatic calorimeter for the continued measurement of heat production over long periods of time. Proceedings of the American Philosophical Society, 78(3):483–512, January 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984765.

Boas:1943:IFP

[Boa43] Franz Boas. Individual, family, population, and race. Proceedings of the American Philosophical Society, 87(2):161–164, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985093.

Bobrovnikoff:1935:SNH

[Bob35] N. T. Bobrovnikoff. The spectrum of Nova Herculis in the visual region. Proceedings of the American Philosophical Society, 75(8):717– 753, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984591.

Bobrovnikoff:1936:SNH

[Bob36] N. T. Bobrovnikoff. On the spectrum of nova herculis. Proceedings of the American Philosophical Society, 76(4):497–521, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984600.

Bogert:1912:CCC

[Bog12] Marston Taylor Bogert. The classification of carbon compounds. Proceedings of the American Philosophical Society, 51(205):252–268, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984143.

Boggs:1938:MLA

[Bog38] S. Whittemore Boggs. The map of Latin America by treaty. Proceedings of the American Philosophical Society, 79(3):399–410, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984857.

Boggs:1949:AIN

[Bog49] S. Whittemore Boggs. An atlas of ignorance: A needed stimulus to honest thinking and hard work. Proceedings of the American Philosophical Society, 93(3):253–258, June 1949. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143475.

Bohr:1947:API

[Boh47] Niels H. D. Bohr. Atomic physics and international cooperation. Proceedings of the American Philosophical Society, 91(2):137–138, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143598.

Boltwood:1911:R

[Bol11] Bertram B. Boltwood. Radioactivity. Proceedings of the American Philosophical Society, 50(200):333–346, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983927.

Bolton:1947:WCC

[Bol47] Herbert E. Bolton. The West Coast corridor. Proceedings of the American Philosophical Society, 91(5):426–429, December 1947. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143423.

Bond:1934:DOW

[Bon34] James Bond. The distribution and origin of the West Indian avifauna. Proceedings of the American Philosophical Society, 73(5):341– 349, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984625.

Bonn:1940:ETS

[Bon40] Moritz J. Bonn. The economics of the totalitarian states. Proceedings of the American Philosophical Society, 82(1):77–87, February 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984906.

Bonner:1941:SSG

[Bon41] Campbell Bonner. The ship of the soul on a group of grave-stelae from Terenuthis. Proceedings of the American Philosophical Society, 85(1):84–91, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985127.

Bonner:1942:TSS

[Bon42] Campbell Bonner. Two studies in syncretistic amulets. Proceedings of the American Philosophical Society, 85(5):466–471, September 1942.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985066.

Boudreau:1943:SEI

[Bou43] Frank G. Boudreau. Social and economic implications of freedom from want of food. Proceedings of the American Philosophical Society, 87(2):126–132, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985088.

Bowman:1917:EPR

[Bow17] H. H. M. Bowman. Ecology and physiology of the red mangrove. Proceedings of the American Philosophical Society, 56(7):589–672, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983887.

Bowman:1930:A

[Bow30] Isaiah Bowman. Antarctica. Proceedings of the American Philosophical Society, 69(1):19–43, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984348.

Boyd:1943:HGS

[Boy43] Julian P. Boyd. Horatio Gates Spafford, precursor of Bessemer. Proceedings of the American Philosophical Society, 87(1):47–50, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984998.

Boyd:1945:MST

[Boy45] Vernon D. Boyd. Motorized surface transportation in the Antarctic. Proceedings of the American Philosophical Society, 89(1):379–381, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985346.

Bryant:1934:FFB

[BR34] William L. Bryant and Rudolph Ruedemann. The fish fauna of Beartooth Butte, Wyoming. Parts II and III. Proceedings of the American Philosophical Society, 73(3):127–167, February 1934. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984757.

Branner:1913:FPG

[Bra13] J. C. Branner. The fluting and pitting of granites in the tropics. Proceedings of the American Philosophical Society, 52(209):163–174, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983863.

Brasch:1942:JLC

[Bra42] Frederick E. Brasch. James Logan, a colonial mathematical scholar, and the first copy of Newton’s “Principia” to arrive in the Colony. Proceedings of the American Philosophical Society, 86(1):3–12, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985072.

Brezina:1904:ACM

[Bre04] Aristides Brezina. The arrangement of collections of meteorites. Proceedings of the American Philosophical Society, 43(176):211–247, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983506.

Brooks:1902:SNF

[Bro02] William Keith Brooks. Is scientific naturalism fatalism? A oneminute paper. Proceedings of the American Philosophical Society, 41(169):145–146, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983532.

Brooks:1903:NGH

[Bro03] William Keith Brooks. On a new genus of hydroid jelly-fishes. Proceedings of the American Philosophical Society, 42(172):11–14, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983638.

Brooks:1904:OCA

[Bro04] Charles Edward Brooks. Orthic curves; or, algebraic curves which satisfy Laplace’s equation in two dimensions. Proceedings of the American Philosophical Society, 43(177):294–331, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983768.

Brooks:1906:HVL

[Bro06] William Keith Brooks. Heredity and variation; logical and biological. Proceedings of the American Philosophical Society, 45 (182):70–76, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983680.

Brown:1916:IA

[Bro16] Philip Marshall Brown. International administration. Proceedings of the American Philosophical Society, 55(4):312–322, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984104.

Brown:1922:AEP

[Bro22] Ernest W. Brown. The age of the Earth from the point of view of astronomy. Proceedings of the American Philosophical Society, 61(4): 283–285, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984269.

Brown:1931:IWN

[Bro31] W. Norman Brown. India’s will to be a nation. Proceedings of the American Philosophical Society, 70(4):317–343, 1931. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984198.

Brooks:1935:DDD

[Bro35] Robert C. Brooks. Democracies and dictatorships: The debit side of their ledgers. Proceedings of the American Philosophical Society, 75 (6):443–481, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984609.

Brown:1938:BNP

[Bro38] William H. Brown. The bearing of nectaries on the phylogeny of flowering plants. Proceedings of the American Philosophical Society, 79(4):549–570 + 572–595, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984940.

Brooks:1940:D

[Bro40] Van Wyck Brooks. Discussion. Proceedings of the American Philosophical Society, 83(4):549–558, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984835.

Brooke:1941:SPA

[Bro41] Tucker Brooke. Some pre-Armada propagandist poetry in England (1585–1586). Proceedings of the American Philosophical Society, 85(1):71–83, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985126.

Bronk:1942:JPE

[Bro42a] Detlev W. Bronk. Joseph Priestley and the early history of the American Philosophical Society. Proceedings of the American Philosophical Society, 86(1):103–107, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985082.

Brown:1942:MMH

[Bro42b] F. Martin Brown. The microscopy of mammalian hair for anthropologists. Proceedings of the American Philosophical Society, 85(3):250–274, February 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985006.

Bronk:1944:DIB

[Bro44a] Detlev W. Bronk. The discovery and interpretation of biological phenomena. Proceedings of the American Philosophical Society, 87(4):307–312, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985042.

Bronk:1944:HPM

[Bro44b] Detlev W. Bronk. Human problems in military aviation. Proceedings of the American Philosophical Society, 88(3):189–195, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985413.

Bronk:1946:IRA

[Bro46] Detlev W. Bronk. International relations among scientists. Proceedings of the American Philosophical Society, 90(4):304–308, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301026.

Brown:1947:IPI

[Bro47] W. Norman Brown. India’s Pakistan issue. Proceedings of the American Philosophical Society, 91(2):162–180, April 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143602.

Brown:1949:CRC

[Bro49] Sanborn C. Brown. Count Rumford and the caloric theory of heat. Proceedings of the American Philosophical Society, 93(4):316–325, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143157.

Bethe:1938:MSE

[BRS38] H. A. Bethe, M. E. Rose, and L. P. Smith. The multiple scattering of electrons. Proceedings of the American Philosophical Society, 78(4):573–585, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984803.

Brush:1913:SDP

[Bru13] Charles F. Brush. Some diffraction phenomena; superposed fringes. Proceedings of the American Philosophical Society, 52(209):276–282, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983872.

Brush:1914:DKT

[Bru14] Charles F. Brush. Discussion of “a kinetic theory of gravitation”. Proceedings of the American Philosophical Society, 53(213):118–128, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984133.

Brush:1915:SGH

[Bru15] Charles F. Brush. Spontaneous generation of heat in recently hardened steel. Proceedings of the American Philosophical Society, 54 (217):154–164, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983920.

Brush:1917:SGH

[Bru17] Charles F. Brush. Spontaneous generation of heat in recently hardened steel. III. Proceedings of the American Philosophical Society, 56 (4):353–363, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983965.

Brush:1918:DMS

[Bru18] Charles F. Brush. Development of magnetic susceptibility in manganese steel by prolonged heat treatment. Proceedings of the American Philosophical Society, 57(4):344–353, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983941.

Brush:1921:DKT

[Bru21] Charles F. Brush. Discussion of a kinetic theory of gravitation, II, and some new experiments in gravitation. Proceedings of the American Philosophical Society, 60(2):43–61, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984383.

Brush:1922:DKT

[Bru22a] Charles F. Brush. Discussion of a kinetic theory of gravitation, II; and some new experiments in gravitation. Second paper. Proceedings of the American Philosophical Society, 61(3):167–183, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984226.

Brush:1922:M

[Bru22b] Charles F. Brush. Minutes. Proceedings of the American Philosophical Society, 61(4):iii–xviii, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984271.

Brush:1923:DKT

[Bru23] Charles F. Brush. Discussion of a kinetic theory of gravitation II.; and some new experiments in gravitation. Third paper. Proceedings of the American Philosophical Society, 62(3):75–89, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984502.

Brush:1924:SNE

[Bru24] Charles F. Brush. Some new experiments in gravitation. Fourth paper. Proceedings of the American Philosophical Society, 63(1):57–61, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984441.

Brush:1925:SNE

[Bru25] Charles F. Brush. Some new experiments in gravitation. Fifth paper. Proceedings of the American Philosophical Society, 64(1):36–50, 1925.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984301.

Brush:1926:DKT

[Bru26] Charles F. Brush. Discussion of the kinetic theory of gravitation. III. Some experimental evidence supporting theory; continual generation of heat in some igneous rocks and minerals. Relation of this to the internal heat of the Earth and presumably of the Sun. Proceedings of the American Philosophical Society, 65(3):207–231, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984491.

Brush:1927:PGH

[Bru27] Charles F. Brush. Persistent generation of heat in some rocks and minerals, and its probable significance: Second paper. Proceedings of the American Philosophical Society, 66(??):251–266, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301067.

Brush:1928:DKT

[Bru28] Charles F. Brush. Discussion of the kinetic theory of gravitation IV: Correlation of continual generation of heat in some substances, and impairment of their gravitational acceleration. Proceedings of the American Philosophical Society, 67(2):105–117, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984272.

Brush:1929:G

[Bru29] Charles F. Brush. Gravitation. Proceedings of the American Philosophical Society, 68(1):55–68, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984399.

Brubaker:1932:FXD

[Bru32] Albert P. Brubaker. Francis X. Dercum. Proceedings of the American Philosophical Society, 71(2):39–48, March 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984684.

Bryant:1902:DCA

[Bry02] Henry G. Bryant. Drift casks in the Arctic Ocean. Proceedings of the American Philosophical Society, 41(169):154–161, April 1902. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983534.

Bryce:1909:PRC

[Bry09] James Bryce. Personal reminiscences of Charles Darwin and of the reception of the “Origin of Species”. Proceedings of the American Philosophical Society, 48(193):iii–xiv, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984063.

Bryant:1926:SPA

[Bry26] William L. Bryant. On the structure of “Palæaspis” and on the occurrence in the United States of fossil fishes belonging to the family Pteraspidæ. Proceedings of the American Philosophical Society, 65(4): 256–271, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984237.

Bryant:1932:LDF

[Bry32] William L. Bryant. Lower Devonian fishes of Bear Tooth Butte, Wyoming. Proceedings of the American Philosophical Society, 71 (5):225–254, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984700.

Bryant:1933:FFB

[Bry33] William L. Bryant. The fish fauna of Beartooth Butte, Wyoming. Proceedings of the American Philosophical Society, 72(5):285–314, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984694.

Bryant:1934:NFT

[Bry34] William L. Bryant. New fishes from the Triassic of Pennsylvania. Proceedings of the American Philosophical Society, 73(5):319–326, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984622.

Bryant:1935:COL

[Bry35] William L. Bryant. Cryptaspis and other Lower Devonian fossil fishes from Beartooth Butte, Wyoming. Proceedings of the American Philosophical Society, 75(2):111–141, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984532.

Bryant:1936:SOK

[Bry36] William L. Bryant. A study of the oldest known vertebrates, Astraspis and Eriptychius. Proceedings of the American Philosophical Society, 76(4):409–427, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984595.

Bryan:1940:PEP

[Bry40] Captain G. S. Bryan. The purpose, equipment and personnel of the Wilkes Expedition. Proceedings of the American Philosophical Society, 82(5):551–560, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984885.

Bryant:1945:BEB

[Bry45] Herwil M. Bryant. Biology at East Base, Palmer Peninsula, Antarctica. Proceedings of the American Philosophical Society, 89(1):256– 269, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985328.

Blum:1907:CPC

[BS07] William Blum and Edgar F. Smith. The cathodic precipitation of carbon. Proceedings of the American Philosophical Society, 46 (185):59–62, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983434.

Bartley:1949:PNC

[BS49] J. O. Bartley and D. L. Sims. Pre-Nineteenth Century stage Irish and Welsh pronunciation. Proceedings of the American Philosophical Society, 93(5):439–447, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143391.

Buehler:1944:ST

[Bue44] Alfred G. Buehler. The sales tax. Proceedings of the American Philosophical Society, 88(1):29–35, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985186.

Burr:1912:TMP

[Bur12] Charles H. Burr. The treaty-making power of the United States and the methods of its enforcement as affecting the police powers of the States. Proceedings of the American Philosophical Society, 51(206): 270–417 + 419–422, August/September 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984009.

Burns:1921:MRA

[Bur21] Keivin Burns. On mean relative and absolute parallaxes. Proceedings of the American Philosophical Society, 60(4):496–502, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984515.

Burnet:1923:RDH

[Bur23] Etienne Burnet. Remarks on Dr. Henderson’s address on Pasteur. Proceedings of the American Philosophical Society, 62(1):xiv–xvi, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984178.

Burke:1927:RB

[Bur27] R. B. Burke. Roger Bacon. Proceedings of the American Philosophical Society, 66(??):79–88, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301059.

Butler:1907:PUA

[But07] Howard Crosby Butler. The Princeton University Archæological Expedition to Syria. Proceedings of the American Philosophical Society, 46(185):182–186, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983452.

Butler:1944:RFP

[But44] Harold B. Butler. R. A. F. Penrose, Jr., Memorial Lecture. The International Labor Organization lessons for twenty-five years. Proceedings of the American Philosophical Society, 88(3):214–219, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985417.

Butterfield:1948:BRP

[But48] Lyman H. Butterfield. Benjamin Rush as a promoter of useful knowledge. Proceedings of the American Philosophical Society, 92(1):26–36, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143622.

Byrd:1945:P

[Byr45] Richard E. Byrd. Preface. Proceedings of the American Philosophical Society, 89(1):iii–iv, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985309.

Calvert:1929:DRG

[Cal29] Philip P. Calvert. Different rates of growth among animals with special reference to the Odonata. Proceedings of the American Philosophical Society, 68(3):227–274, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/984210. See corrections [Ano34l].

Calvert:1934:RGL

[Cal34] Philip P. Calvert. The rates of growth, larval development and seasonal distribution of dragonflies of the genus Anax (Odonata: Æshnidæ). Proceedings of the American Philosophical Society, 73(1): 1–70, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984740.

Caley:1941:CBC

[Cal41] Earle R. Caley. The corroded bronze of Corinth. Proceedings of the American Philosophical Society, 84(5):689–761, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984846.

Calvert:1942:ESH

[Cal42] Philip P. Calvert. Entomology, scientific and human aspects. Proceedings of the American Philosophical Society, 86(1):123–129, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985084.

Campbell:1919:CEE

[Cam19] W. W. Campbell. The Crocker Eclipse Expedition from the Lick Observatory, June 8, 1918. Some eclipse problems. Proceedings of the American Philosophical Society, 58(4):241–254, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984400.

Cannon:1911:SAS

[Can11] W. B. Cannon. The stimulation of adrenal secretion by emotional excitement. Proceedings of the American Philosophical Society, 50 (199):226–227, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984036.

Carrel:1908:FST

[Car08] Alexis Carrel. Further studies on transplantation of vessels and organs. Proceedings of the American Philosophical Society, 47(190): 677–696, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983837.

Carter:1909:ECR

[Car09] Jesse Benedict Carter. The evolution of the City of Rome from its origin to the Gallic catastrophe. Proceedings of the American Philosophical Society, 48(192):129–141, April/August 1909. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984151.

Carson:1916:SD

[Car16] Hampton L. Carson. Samuel Dickson. Proceedings of the American Philosophical Society, 55(8):XIV–XXII, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984013.

Carson:1917:TAI

[Car17] Hampton L. Carson. The trial of animals and insects. A little known chapter of Mediæval jurisprudence. Proceedings of the American Philosophical Society, 56(5):410–415, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984029.

Carson:1927:GAA

[Car27] Hampton L. Carson. The growth of Anglo-American law since the days of the Junto. Proceedings of the American Philosophical Society, 66(??):593–619, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301084.

Carrel:1929:NPM

[Car29] Alexis Carrel. The nutritional properties of malignant cells. Proceedings of the American Philosophical Society, 68(2):129–132, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984313.

Carmichael:1948:RVF

[Car48] Leonard Carmichael. Reading and visual fatigue. Proceedings of the American Philosophical Society, 92(1):41–42, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143624.

Carpenter:1949:RAM

[Car49] Edmund S. Carpenter. Recently acquired manuscripts relating to the archeology of Pennsylvania. Proceedings of the American Philosophical Society, 93(2):166–168, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143436.

Case:1933:PCS

[Cas33] Ermine C. Case. Progressive chondrification in the Stegocephalia. Proceedings of the American Philosophical Society, 72(4):265–283, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984691.

Case:1937:BSP

[Cas37] E. C. Case. The brain and skull of a Paleoniscid fish from the Pennsylvanian of Western Missouri. Proceedings of the American Philosophical Society, 78(1):1–10, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984654.

Case:1939:MBR

[Cas39] E. C. Case. The mastodons of Båråboedaer. Proceedings of the American Philosophical Society, 81(4):569–572, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984827.

Cartledge:1936:HMF

[CBB36] J. L. Cartledge, L. V. Barton, and A. F. Blakeslee. Heat and moisture as factors in the increased mutation rate from Datura seeds. Proceedings of the American Philosophical Society, 76(5):663–685, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984754.

Choate:1912:HMP

[CDG^{+}12] Joseph H. Choate, J. M. Dickinson, John C. Gray, Henry Wade Rogers, and James Brown Scott. The Henry M. Phillips Prize Essay. Proceedings of the American Philosophical Society, 51(206):

269, August/September 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984008.

Crile:1922:ABR

[CF22] G. W. Crile and Hugo Fricke. Application of biophysical research to medical problems. Proceedings of the American Philosophical Society, 61(3):237–245, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984231.

Crile:1932:FSA

[CGTR32] George Crile, Otto Glasser, Maria Telkes, and Amy Rowland. Further studies of autosynthetic cells with special reference to the possible rôle of the nitro group in the energy phenomena of protoplasm. Proceedings of the American Philosophical Society, 71(7):411–420, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984762.

Costello:1949:HTT

[CH49] Donald P. Costello and Catherine Henley. Heteroploidy in Triturus torosus. I. The incidence of spontaneous variations in a “natural population”. Proceedings of the American Philosophical Society, 93(5):428–438, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143390.

Chamberlin:1906:PRD

[Cha06a] Thomas C. Chamberlin. On a possible reversal of deep-sea circulation and its influence on geologic climates. Proceedings of the American Philosophical Society, 45(182):33–43, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983677.

Chance:1906:BNJ

[Cha06b] Henry M. Chance. A biographical notice of J. Peter Lesley. Proceedings of the American Philosophical Society, 45(184):i–xiv, October/ December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983385.

Chance:1908:OBO

[Cha08] H. M. Chance. The origin of bombshell ore. Proceedings of the American Philosophical Society, 47(188):136–140, January/April 1908. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983805.

Chamberlin:1913:PPN

[Cha13] Ralph V. Chamberlin. Place and personal names of the Gosiute Indians of Utah. Proceedings of the American Philosophical Society, 52(208):1–20, January/April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983995.

Chamberlin:1915:IEV

[Cha15] T. C. Chamberlin. The interior of the Earth from the viewpoint of geology. Proceedings of the American Philosophical Society, 54 (219):279–289, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983874.

Chamberlin:1922:AEG

[Cha22] T. C. Chamberlin. The age of the Earth from the geological viewpoint. Proceedings of the American Philosophical Society, 61(4):247– 271, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984267.

Chase:1938:CVA

[Cha38] George H. Chase. The Corpus Vasorum Antiquorum. Proceedings of the American Philosophical Society, 79(4):681–693, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984945.

Chandrasekhar:1939:ICS

[Cha39] S. Chandrasekhar. The internal constitution of the stars. Proceedings of the American Philosophical Society, 81(2):153–187, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984981.

Chafee:1947:DJM

[Cha47] Zechariah Chafee, Jr. Do judges make or discover law? Proceedings of the American Philosophical Society, 91(5):405–420, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143421.

Chafee:1949:CEH

[Cha49] Zechariah Chafee, Jr. Charles Evans Hughes. Proceedings of the American Philosophical Society, 93(3):267–281, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143477.

Cheyney:1911:LWH

[Che11] Edward Potts Cheyney. On the life and works of Henry Charles Lea. Proceedings of the American Philosophical Society, 50(202): v–xli, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984003.

Cheyney:1927:HHC

[Che27] Edward P. Cheyney. How history can be made a science. Proceedings of the American Philosophical Society, 66(??):581–591, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301083.

Cheyney:1941:CAD

[Che41] Edward P. Cheyney. The connection of Alexander Dallas Bache with the University of Pennsylvania. Proceedings of the American Philosophical Society, 84(2):151–160, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984914.

Child:1916:ACO

[Chi16] C. M. Child. Age cycles and other periodicities in organisms. Proceedings of the American Philosophical Society, 55(5):330–339, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984048.

Chinard:1943:APS

[Chi43a] Gilbert Chinard. The American Philosophical Society and the world of science (1768–1800). Proceedings of the American Philosophical Society, 87(1):1–11, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984994.

Chinard:1943:JAP

[Chi43b] Gilbert Chinard. Jefferson and the American Philosophical Society. Proceedings of the American Philosophical Society, 87(3):263–276, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984873.

Chinard:1945:APS

[Chi45] Gilbert Chinard. The American Philosophical Society and the early history of forestry in America. Proceedings of the American Philosophical Society, 89(2):444–488, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985641.

Chinard:1947:ECT

[Chi47] Gilbert Chinard. Eighteenth Century theories on America as a human habitat. Proceedings of the American Philosophical Society, 91(1):27–57, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143122.

Chinard:1948:DFN

[Chi48] Gilbert Chinard. Dr. Franklin negotiates, December 1777. Proceedings of the American Philosophical Society, 92(2):101–104, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143406.

Chinard:1949:ASC

[Chi49] Gilbert Chinard. The American sketchbooks of Charles-Alexandre Lesueur. Proceedings of the American Philosophical Society, 93 (2):114–118, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143429.

Collins:1936:FTM

[CL36] R. Lee Collins and W. Gardner Lynn. Fossil turtles from Maryland. Proceedings of the American Philosophical Society, 76(2):151– 173, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984665.

Clarke:1906:SMC

[Cla06] F. W. Clarke. The statistical method in chemical geology. Proceedings of the American Philosophical Society, 45(182):14–32, January/ April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983676.

Clarke:1912:SGS

[Cla12] Frank Wigglesworth Clarke. Some geochemical statistics. Proceedings of the American Philosophical Society, 51(205):214–234, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984139.

Clarke:1922:AEP

[Cla22] John M. Clarke. The age of the Earth from the paleontological viewpoint. Proceedings of the American Philosophical Society, 61(4):272– 282, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984268.

Clark:1932:PCR

[Cla32] Victor S. Clark. Past crises in retrospect and in contemporary opinion. Proceedings of the American Philosophical Society, 71(3):73–83, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984554.

Clark:1945:EUS

[Cla45] Austin H. Clark. Echinoderms of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):295, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985334.

Cleland:1935:CTS

[Cle35] Ralph E. Cleland. Cyto-taxonomic studies on certain Oenotheras from California. Proceedings of the American Philosophical Society, 75(5):339–429, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984692.

Cleland:1936:HV

[Cle36] Ralph E. Cleland. Hugo de Vries. Proceedings of the American Philosophical Society, 76(2):248–250, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984672.

Cleland:1937:SRO

[Cle37] Ralph E. Cleland. Species relationships in Onagra. Proceedings of the American Philosophical Society, 77(4):477–542, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984775.

Clemence:1949:REP

[Cle49] G. M. Clemence. Relativity effects in planetary motion. Proceedings of the American Philosophical Society, 93(7):532–534, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143142.

Clay:1921:MJJ

[CM21] Albert T. Clay and James A. Montgomery. Morris Jastrow, Jr. Proceedings of the American Philosophical Society, 60(4):x–xxvi, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984522.

Cobb:1915:RWO

[Cob15] Margaret V. Cobb. Relationships of the white oaks of Eastern North America, with an introductory sketch of their phylogenetic history. Proceedings of the American Philosophical Society, 54(217):165–175, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983921.

Coghill:1926:GFN

[Cog26] G. E. Coghill. The growth of functional neurones and its relation to the development of behavior. Proceedings of the American Philosophical Society, 65(1):51–55, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984337.

Cohn:1914:NAM

[Coh14] Alfred E. Cohn. Newer aspects and methods in the study of the mechanism of the heart-beat. Proceedings of the American Philosophical Society, 53(215):287–302, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984078.

Cohn:1944:BBD

[Coh44] Edwin J. Cohn. Blood, blood derivatives, and blood substitutes. Proceedings of the American Philosophical Society, 88(3):159–173, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985409.

Cohen:1948:ERC

[Coh48] Benjamin V. Cohen. The evolving role of Congress in foreign affairs. Proceedings of the American Philosophical Society, 92(4):211–216, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143349.

Cole:1930:RTA

[Col30] Fay-Cooper Cole. Recent trends in anthropology. Proceedings of the American Philosophical Society, 69(1):385–390, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984372.

Colbert:1941:ADB

[Col41a] Leo Otis Colbert. Alexander Dallas Bache as Superintendent of United States Coast Survey 1843–1867. Proceedings of the American Philosophical Society, 84(2):173–180, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984916.

Coleman:1941:JBP

[Col41b] Arthur Prudden Coleman. John Bowring and the poetry of the Slavs. Proceedings of the American Philosophical Society, 84(3):431–459, May 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984959.

Cole:1943:CMW

[Col43a] Fay-Cooper Cole. Chronology in the Middle West. Proceedings of the American Philosophical Society, 86(2):299–303, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985107.

Collins:1943:EAB

[Col43b] Henry B. Collins, Jr. Eskimo archaeology and its bearing on the problem of Man’s antiquity in America. Proceedings of the American Philosophical Society, 86(2):220–235, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985100.

Colton:1943:RAH

[Col43c] Harold S. Colton. Reconstruction of Anasazi history. Proceedings of the American Philosophical Society, 86(2):264–269, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985105.

Comstock:1911:GP

[Com11] Daniel F. Comstock. The general principles. Proceedings of the American Philosophical Society, 50(200):321–332, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983926.

Compton:1922:ASI

[Com22] K. T. Compton. Arc spectra and ionization potentials in dissociated gases. Proceedings of the American Philosophical Society, 61(3):212– 226, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984229.

Compton:1931:AA

[Com31] Arthur H. Compton. Assault on atoms. Proceedings of the American Philosophical Society, 70(3):215–229, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984451.

Compton:1935:CCR

[Com35] Arthur H. Compton. The composition of cosmic rays. Proceedings of the American Philosophical Society, 75(4):251–274, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984627.

Compton:1940:SSA

[Com40] Arthur H. Compton. Science shaping American culture. Proceedings of the American Philosophical Society, 83(4):573–582, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984838.

Compton:1945:FCT

[Com45] Wilson Compton. Forest conservation: A task in engineering and in public and private cooperation. Proceedings of the American Philosophical Society, 89(2):423–427, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985638.

Compton:1946:FML

[Com46] Arthur H. Compton. Franklin Medal Lecture: Atomic energy as a human asset. Proceedings of the American Philosophical Society, 90(1):70–79, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301042.

Compton:1947:PSP

[Com47] Arthur H. Compton. The place of science in the program of UNESCO. Proceedings of the American Philosophical Society, 91(4):303–306, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143343.

Conklin:1902:EBT

[Con02] Edwin G. Conklin. The embryology of a brachiopod, Terebratulina Septentrionalis Couthouy. Proceedings of the American Philosophical Society, 41(168):41–76, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983525.

Conklin:1908:WKB

[Con08] Edwin G. Conklin. William Keith Brooks. Proceedings of the American Philosophical Society, 47(190):iii–x, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983841.

Conklin:1909:WDD

[Con09] Edwin G. Conklin. The world’s debt to Darwin. Proceedings of the American Philosophical Society, 48(193):xxxviii–lvii, September/ December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984066.

Conklin:1915:AW

[Con15] Edwin G. Conklin. August Weismann. Proceedings of the American Philosophical Society, 54(220):iii–xii, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984111.

Conklin:1934:PML

[Con34] Edwin G. Conklin. Penrose Memorial Lecture. A generation’s progress in the study of evolution. Proceedings of the American Philosophical Society, 74(2):137–158, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984738.

Conklin:1935:TS

[Con35] Edwin G. Conklin. Theobald Smith. Proceedings of the American Philosophical Society, 75(4):333–335, 1935. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984635.

Conklin:1936:RCR

[Con36] Edwin G. Conklin. Report of the committee on research. Proceedings of the American Philosophical Society, 76(3):379–389, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984550.

Conklin:1937:RCR

[Con37a] Edwin G. Conklin. Report of the Committee on Research, April 1936–April 1937. Proceedings of the American Philosophical Society, 77(4):621–629, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984783.

Conklin:1937:RTC

[Con37b] Edwin G. Conklin. Round table conference. Proceedings of the American Philosophical Society, 77(4):566–590, April 1937. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984778.

Conklin:1938:DDC

[Con38] Edwin G. Conklin. Disorientations of development in Crepidula Plana produced by low temperatures. Proceedings of the American Philosophical Society, 79(2):179–211, June 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984619.

Conklin:1940:CAP

[Con40] Edwin G. Conklin. Connection of the American Philosophical Society with our first national exploring expedition. Proceedings of the American Philosophical Society, 82(5):519–541, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984883.

Conklin:1941:ADB

[Con41] Edwin G. Conklin. Alexander Dallas Bache and his connection with the American Philosophical Society. Proceedings of the American Philosophical Society, 84(2):125–144, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984912.

Conklin:1942:IR

[Con42] Edwin G. Conklin. Introductory remarks. Proceedings of the American Philosophical Society, 86(1):1–2, September 1942. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985071.

Conklin:1943:IJB

[Con43] Edwin G. Conklin. Introduction to the Jefferson Bicentennial Program. Proceedings of the American Philosophical Society, 87(3):199– 200, July 14, 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984865.

Conant:1944:ALU

[Con44] James B. Conant. The advancement of learning in the United States in the post-war world. Proceedings of the American Philosophical Society, 87(4):291–298, January 29, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985040.

Conklin:1947:APS

[Con47] Edwin G. Conklin. The American Philosophical Society and international relations. Proceedings of the American Philosophical Society, 91(1):1–9, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143118.

Cooley:1940:EEP

[Coo40] Mary E. Cooley. The exploring expedition in the Pacific. Proceedings of the American Philosophical Society, 82(5):707–719, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984889.

Cooper:1946:PIU

[Coo46] John M. Cooper. Problems of international understanding. Proceedings of the American Philosophical Society, 90(4):314–317, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301028.

Cope:1948:CSM

[Cop48] Thomas D. Cope. Collecting source material about Charles Mason and Jeremiah Dixon. Proceedings of the American Philosophical Society, 92(2):111–114, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143409.

Cope:1949:DAW

[Cop49] Thomas D. Cope. Degrees along the West Line, the parallel between Maryland and Pennsylvania. Proceedings of the American Philosophical Society, 93(2):127–133, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143431.

Coulter:1916:ITS

[Cou16] John M. Coulter. Inheritance through spores. Proceedings of the American Philosophical Society, 55(5):344–347, 1916. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984050.

Coulter:1927:DBU

[Cou27] John M. Coulter. Development of botany in the United States. Proceedings of the American Philosophical Society, 66(??):309–318, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301071.

Court:1945:FAW

[Cou45a] Arnold Court. Field altitudes of West Base Parties, United States Antarctic Service Expedition, 1939–1941. Proceedings of the American Philosophical Society, 89(1):61–66, April 1945. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985314.

Court:1945:WOD

[Cou45b] Arnold Court. Weather observations during 1940–1941 at Little America III, Antarctica. Proceedings of the American Philosophical Society, 89(1):324–343, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985343.

Cox:1948:FRR

[Cox48] Isaac Joslin Cox. The Freeman Red River Expedition. Proceedings of the American Philosophical Society, 92(2):115–119, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143410.

Crandall:1907:GSF

[Cra07] Roderic Crandall. The geology of the San Francisco peninsula. Proceedings of the American Philosophical Society, 46(185):3–58, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983433.

Crane:1917:MVS

[Cra17] T. F. Crane. Mediæval sermon-books and stories and their study since 1883. Proceedings of the American Philosophical Society, 56(5): 369–402, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984027.

Craigie:1938:BVB

[Cra38] E. Horne Craigie. The blood vessels of the brain substance in some amphibians. Proceedings of the American Philosophical Society, 78(4):615–649, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984806.

Craigie:1939:VBT

[Cra39] E. Horne Craigie. Vascularity in the brains of tailed amphibians. I. Ambystoma tigrinum (Green). Proceedings of the American Philosophical Society, 81(1):21–27, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984963.

Craigie:1940:VBT

[Cra40] E. Horne Craigie. Vascularity in the brains of tailed amphibians. II. Necturus Maculosus Rafinesque. Proceedings of the American Philosophical Society, 82(4):395–410, May 20, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984877.

Cressman:1943:RRA

[Cre43] L. S. Cressman. Results of recent archaeological research in the Northern Great Basin Region of South Central Oregon. Proceedings of the American Philosophical Society, 86(2):236–246, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985101.

Crile:1912:PAR

[Cri12] George W. Crile. Phylogenetic association in relation to the emotions. Proceedings of the American Philosophical Society, 51(204):76–90, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984096.

Crile:1913:RBP

[Cri13] Geo. W. Crile. The relation between the physical state of brain cells and brain functions, — experimental and clinical. Proceedings of the American Philosophical Society, 52(210):397–406, May/August 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984084.

Crile:1914:KS

[Cri14] George W. Crile. The kinetic system. Proceedings of the American Philosophical Society, 53(215):263–286, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984077.

Crile:1921:ECT

[Cri21] G. W. Crile. An electro-chemical theory of normal and certain pathological processes. Proceedings of the American Philosophical Society, 60(4):546–552, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984519.

Crile:1930:ECI

[Cri30] George W. Crile. An electro-chemical interpretation of memory. Proceedings of the American Philosophical Society, 69(1):359–368, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984369.

Crickmay:1931:JHN

[Cri31] C. H. Crickmay. Jurassic history of North America: Its bearing on the development of continental structure. Proceedings of the American Philosophical Society, 70(1):15–102, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984235.

Crile:1933:OPI

[Cri33] George W. Crile. Orthogenesis and the power and infirmities of man. Proceedings of the American Philosophical Society, 72(4):245–254, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984689.

Crocker:1932:EEU

[Cro32] William Crocker. The effect of ethylene upon living organisms. Proceedings of the American Philosophical Society, 71(5):295–298, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984704.

Cross:1943:EAF

[Cro43] Dorothy Cross. The effect of the Abbott Farm on eastern chronology. Proceedings of the American Philosophical Society, 86(2):315–319, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985110.

Crosby:1949:EAC

[Cro49] Sumner McKnight Crosby. Excavations in the Abbey Church of St.- Denis 1948 the façade of Fulrad’s Church. Proceedings of the American Philosophical Society, 93(5):347–361, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143387.

Crile:1929:EIP

[CTR29] George W. Crile, Maria Telkes, and Amy F. Rowland. An experimental investigation of the physical nature of death. Proceedings of the American Philosophical Society, 68(2):69–81, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984307.

Curtis:1918:NSN

[Cur18] Heber D. Curtis. On the number of spiral nebulæ. Proceedings of the American Philosophical Society, 57(5):513–520, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984185.

Curti:1949:IRA

[Cur49] Merle Curti. The impact of the revolutions of 1848 on American thought. Proceedings of the American Philosophical Society, 93(3):209–215, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143467.

Cushman:1945:FUS

[Cus45] Joseph A. Cushman. Foraminifera of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):285–288, April 1945. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985330.

Cushman:1948:RFA

[Cus48] Robert E. Cushman. The repercussions of foreign affairs on the American tradition of civil liberty. Proceedings of the American Philosophical Society, 92(4):257–263, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143354.

Dallas:1905:EPR

[Dal05] W. L. Dallas. Enquiry into the pressure and rainfall conditions of the trades-monsoon area. Proceedings of the American Philosophical Society, 44(180):159–163, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983512.

Dale:1915:CMD

[Dal15] Nelson C. Dale. The Cambrian manganese deposits of Conception and Trinity Bays, Newfoundland. Proceedings of the American Philosophical Society, 54(220):371–456, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984110.

Daly:1925:RMB

[Dal25] Reginald A. Daly. Relation of mountain-building to igneous action. Proceedings of the American Philosophical Society, 64(2):283–307, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984195.

Daly:1946:OMT

[Dal46] Reginald A. Daly. Origin of the Moon and its topography. Proceedings of the American Philosophical Society, 90(2):104–119, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301051.

Dale:1947:FS

[Dal47] Henry Hallett Dale. The freedom of science. Proceedings of the American Philosophical Society, 91(1):64–72, February 1947. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143124.

Dana:1911:NCF

[Dan11] Charles E. Dana. Notes on cannon — Fourteenth and Fifteenth Centuries. Proceedings of the American Philosophical Society, 50 (199):147–167, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984031.

Darrach:1933:WWK

[Dar33] William Darrach. William Williams Keen promoter of useful knowledge. Proceedings of the American Philosophical Society, 72(3):115– 119, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984638.

Darrow:1938:NPP

[Dar38] Karl K. Darrow. Neutral particles in physics. Proceedings of the American Philosophical Society, 79(3):389–397, September 1938. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984856.

Darrow:1940:II

[Dar40a] Karl K. Darrow. Introduction to the ionosphere. Proceedings of the American Philosophical Society, 83(2):429–445, August 1940. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985019.

Darrow:1940:SNT

[Dar40b] Karl K. Darrow. Status of nuclear theory. Proceedings of the American Philosophical Society, 82(3):351–357, April 1940. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984934.

Darrow:1944:DSO

[Dar44a] Karl K. Darrow. Discussion on the symposium “organization, direction, and support of research”. Proceedings of the American Philosophical Society, 87(4):321–322, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985044.

Darrow:1944:E

[Dar44b] Karl K. Darrow. Entropy. Proceedings of the American Philosophical Society, 87(5):365–367, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985287.

Darrow:1947:GPS

[Dar47] Karl K. Darrow. Growth of the physical sciences and their applications in the United States. Proceedings of the American Philosophical Society, 91(1):17–21, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143120.

Davis:1902:SG

[Dav02] W. M. Davis. Systematic geography. Proceedings of the American Philosophical Society, 41(170):235–259, April 1902. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983555.

Davis:1906:WLE

[Dav06] William Morris Davis. Was Lewis Evans or Benjamin Franklin the first to recognize that our northeast storms come from the southwest? Proceedings of the American Philosophical Society, 45(183): 129–130, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983822.

Davenport:1908:DDM

[Dav08] Charles B. Davenport. Determination of dominance in Mendelian inheritance. Proceedings of the American Philosophical Society, 47 (188):59–63, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983798.

Davenport:1910:NVA

[Dav10a] C. B. Davenport. The new views about reversion. Proceedings of the American Philosophical Society, 49(196):291–296, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984121.

Davies:1910:TCH

[Dav10b] John Vipond Davies. The tunnel construction of the Hudson and Manhattan Railroad Company. Proceedings of the American Philosophical Society, 49(195):164–187, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983892.

Davis:1910:AGP

[Dav10c] William Morris Davis. Antarctic geology and polar climates. Proceedings of the American Philosophical Society, 49(195):200–202, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983896.

Davis:1915:TPS

[Dav15] Bradley Moore Davis. The test of a pure species of Œnothera. Proceedings of the American Philosophical Society, 54(218):226–245, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983936.

Davenport:1917:ERI

[Dav17] C. B. Davenport. The effects of race intermingling. Proceedings of the American Philosophical Society, 56(4):364–368, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983966.

Davis:1924:BON

[Dav24] Bradley Moore Davis. The behavior of “Œnothera Neo-Lamarckiana” in selfed line through seven generations. Proceedings of the American Philosophical Society, 63(3):239–278, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984327.

Davis:1926:HON

[Dav26] Bradley Moore Davis. The history of Œnothera Biennis Linnæus, “Œnothera Grandiflora” Solander, and “Œnothera Lamarckiana” of de Vries in England. Proceedings of the American Philosophical Society, 65(4):349–378, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984241.

Davis:1927:LEP

[Dav27] Bradley Moore Davis. Lamarck’s evening primrose (Oenothera lamarckiana Seringe) was a form of Oenothera grandiflora Solander. Proceedings of the American Philosophical Society, 66(??):319–355, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301072.

Davisson:1930:WPE

[Dav30] C. J. Davisson. The wave properties of electrons. Proceedings of the American Philosophical Society, 69(1):247–256, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984360.

Davenport:1931:IVM

[Dav31] C. B. Davenport. Individual vs. mass studies in child growth. Proceedings of the American Philosophical Society, 70(4):381–389, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984203.

Davis:1934:WON

[Dav34] Bradley Moore Davis. What is “Œnothera Hookeri” Torrey & Gray? Proceedings of the American Philosophical Society, 74(3):245–251, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984717.

Davenport:1935:PHC

[Dav35] Chas. B. Davenport. A provisional hypothesis of child development. Proceedings of the American Philosophical Society, 75(6):537–548, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984615.

Davidson:1936:SA

[Dav36] D. S. Davidson. The spearthrower in Australia. Proceedings of the American Philosophical Society, 76(4):445–483, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984597.

Davenport:1937:PGE

[Dav37a] C. B. Davenport. Postnatal growth of the external nose. Proceedings of the American Philosophical Society, 78(1):61–77, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984657.

Davis:1937:SSD

[Dav37b] Bradley Moore Davis. The segregation of sulfur and dwarf from crosses involving “Oenothera Franciscana” and certain hybrid derivatives. Proceedings of the American Philosophical Society, 77(2): 99–160, February 26, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984724.

Davidson:1938:EMA

[Dav38] D. Sutherland Davidson. An ethnic map of Australia. Proceedings of the American Philosophical Society, 79(4):649–679, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984944.

Davenport:1939:PND

[Dav39a] Charles Benedict Davenport. Post-natal development of the human outer nose. Proceedings of the American Philosophical Society, 80(2):175–316 + 318–355, January 31, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985128.

David:1939:NIN

[Dav39b] Charles Wendell David. Narratio de itinere navali peregrinorum hierosolymam tendentium et silviam capientium, A.D. 1189. Proceedings of the American Philosophical Society, 81(5):591–676, December 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985010.

Davenport:1940:PND

[Dav40] Charles Benedict Davenport. Post-natal development of the head. Proceedings of the American Philosophical Society, 83(1):1–215, July 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985119.

Davenport:1941:RB

[Dav41a] C. B. Davenport. Responsive bone. Proceedings of the American Philosophical Society, 84(1):65–70, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984851.

Davidson:1941:AAS

[Dav41b] Daniel Sutherland Davidson. Aboriginal Australian string figures. Proceedings of the American Philosophical Society, 84(6):763–901, August 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984876.

Davenport:1944:PDH

[Dav44] Charles Benedict Davenport. Postnatal development of the human extremities. Proceedings of the American Philosophical Society, 88(5):375–455, November 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985612.

Davidson:1949:DDW

[Dav49a] Daniel Sutherland Davidson. Disposal of the dead in Western Australia. Proceedings of the American Philosophical Society, 93(1):71– 97, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143273.

Davidson:1949:MCA

[Dav49b] Daniel Sutherland Davidson. Mourning-caps of the Australian Aborigines. Proceedings of the American Philosophical Society, 93(1): 57–70, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143272.

Ward:1923:ISC

[dCW23] Robert de C. Ward. The “Indian Summer” as a characteristic weather type of the Eastern United States. Proceedings of the American Philosophical Society, 62(2):48–56, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984509.

Deichmann:1945:OUS

[Dei45] Elisabeth Deichmann. An octocoral of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):294, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985333.

Delano:1934:EIN

[Del34] Frederic A. Delano. The economic implications of national planning. Proceedings of the American Philosophical Society, 74(1):21–28, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984582.

Deloria:1944:DTM

[Del44] Ella Cara Deloria. Dakota treatment of murderers. Proceedings of the American Philosophical Society, 88(5):368–371, November 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985610.

Delmage:1947:AIP

[Del47] Rutherford E. Delmage. The American idea of progress, 1750–1800. Proceedings of the American Philosophical Society, 91(4):307–314, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143344.

Dempster:1935:NMM

[Dem35] Arthur Jeffrey Dempster. New methods in mass spectroscopy. Proceedings of the American Philosophical Society, 75(8):755–767, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984592.

Dempster:1936:FEM

[Dem36] Arthur Jeffrey Dempster. Further experiments on the mass analysis of the chemical elements. Proceedings of the American Philosophical Society, 76(4):491–496, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984599.

Dercum:1927:OAA

[Der27] Francis X. Dercum. The origin and activities of the American Philosophical Society and an address on the dynamic factor in evolution. Proceedings of the American Philosophical Society, 66(??):19–45, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301056.

Dercum:1929:NTL

[Der29] Francis X. Dercum. On the nature of thought and its limitation. Proceedings of the American Philosophical Society, 68(4):275–302, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984341.

Dercum:1931:NLL

[Der31] Francis X. Dercum. Non-living and living matter. Proceedings of the American Philosophical Society, 70(1):1–13, 1931. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984234.

Detlefsen:1923:ELC

[Det23] J. A. Detlefsen. Are the effects of long-continued rotation in rats inherited? Proceedings of the American Philosophical Society, 62(5):

292–300, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984462.

Deterra:1937:CCA

[Det37] Hellmut Deterra. Cenozoic cycles in Asia and their bearing on human prehistory. Proceedings of the American Philosophical Society, 77(3):289–308, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984732.

Deierkauf-Holsboer:1947:VAH

[DH47] S. Wilma Deierkauf-Holsboer. Vie d’Alexandre Hardy, poète du roi: Quarante-deux documents inédits. Proceedings of the American Philosophical Society, 91(4):328–401, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143346.

DiGregorio:1903:SNC

[Di 03] Marchese Antonio Di Gregorio. On some names (chiefly Linnean) of animals and plants erroneously paired in synonymy. Proceedings of the American Philosophical Society, 42(173):263–264, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983738.

Dickson:1927:RLC

[Dic27] L. E. Dickson. Recent light on classic problems in the theory of numbers. Proceedings of the American Philosophical Society, 66(??):281– 286, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301069.

Dickinson:1943:OPP

[Dic43] John Dickinson. The old political philosophy and the new. Proceedings of the American Philosophical Society, 87(3):246–262, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984872.

Dill:1947:CPP

[Dil47] David Bruce Dill. Contributions of physiology to problems of industry. Proceedings of the American Philosophical Society, 91(3):281– 283, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143400.

Dinsmoor:1939:AOA

[Din39] William Bell Dinsmoor. Archæology and astronomy. Proceedings of the American Philosophical Society, 80(1):95–173, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984976.

Dinsmoor:1943:EAS

[Din43] William B. Dinsmoor. Early American studies of Mediterranean archaeology. Proceedings of the American Philosophical Society, 87(1): 70–104, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985002.

Dixon:1920:NTP

[Dix20] Roland B. Dixon. A new theory of Polynesian origins. Proceedings of the American Philosophical Society, 59(4):261–267, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984425.

Dixon:1930:RAO

[Dix30] R. B. Dixon. Recent archæological discoveries in the Philippines and their bearing on the prehistory of Eastern Asia. Proceedings of the American Philosophical Society, 69(1):225–229, 1930. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984358.

Dixon:1934:LVP

[Dix34] Roland B. Dixon. The long voyages of the Polynesians. Proceedings of the American Philosophical Society, 74(3):167–175, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984711.

Dvoichenko-Markoff:1947:BFA

[DM47] Eufrosina Dvoichenko-Markoff. Benjamin Franklin, the American Philosophical Society, and the Russian Academy of Science. Proceedings of the American Philosophical Society, 91(3):250–257, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143397.

Sajous:1927:EOF

[dMS27] Charles E. de M. Sajous. The endocrine organs as a foundation for scientific medicine. Proceedings of the American Philosophical Society, 66(??):47–65, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301057.

Sajous:1928:TRF

[dMS28] Charles E. de M. Sajous. Tissue respiration as the function of the internal secretions which science has sanctioned. Proceedings of the American Philosophical Society, 67(3):307–318, 1928. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984296.

Dodds:1933:RLG

[Dod33] Harold W. Dodds. The reorganization of local government. Proceedings of the American Philosophical Society, 72(2):67–75, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984743.

Doolittle:1903:ODS

[Doo03] Eric Doolittle. The orbit of the double star Σ 518. Proceedings of the American Philosophical Society, 42(173):170–179, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983730.

Doolittle:1910:SNF

[Doo10] C. L. Doolittle. Simon Newcomb, F.R.S., LL.D., D.C.L. Proceedings of the American Philosophical Society, 49(197):iii–xviii, October/ December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984092.

Doriot:1944:EP

[Dor44] Georges F. Doriot. Environmental protection. Proceedings of the American Philosophical Society, 88(3):196–203, September 1944. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985414.

Dorsey:1945:AMW

[Dor45] H. G. Dorsey, Jr. An Antarctic mountain weather station. Proceedings of the American Philosophical Society, 89(1):344–363, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985344.

Douglas:1901:RBS

[Dou01] James Douglas. Record of borings in the Sulphur Spring Valley, Arizona, and of agricultural experiments in the same locality. Proceedings of the American Philosophical Society, 40(167):161–164, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983759.

Douglass:1902:CLT

[Dou02] Earl Douglass. A Cretaceous and lower tertiary section in South Central Montana. Proceedings of the American Philosophical Society, 41(170):207–224, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983553.

Douglass:1922:STC

[Dou22] A. E. Douglass. Some topographic and climatic characters in the annual rings of the yellow pines and sequoias of the Southwest. Proceedings of the American Philosophical Society, 61(2):117–122, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984409.

Draper:1946:PHE

[Dra46] Warren F. Draper. Public health experiences in the European theatre of operations. Proceedings of the American Philosophical Society, 90(4):289–294, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301023.

Driver:1919:LNS

[Dri19] Chas. S. Driver. On the Luciopimelodinæ, a new subfamily of the South American Siluridæ. Proceedings of the American Philosophical Society, 58(7):448–456, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984214.

Dodge:1945:IFI

[DSA45] B. O. Dodge, Mary B. Schmitt, and Anita Appel. Inheritance of factors involved in one type of heterocaryotic vigor. Proceedings of the American Philosophical Society, 89(4):575–589, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985362.

deTerra:1942:MBK

[dT42] Hellmut de Terra. The megaliths of Bursahom, Kashmir, a new prehistoric civilization from India. Proceedings of the American Philosophical Society, 85(5):483–504, September 1942. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985068.

deTerra:1936:OUS

[dTdC36] H. de Terra and P. Teilhard de Chardin. Observations on the Upper Siwalik Formation and Later Pleistocene deposits in India. Proceedings of the American Philosophical Society, 76(6):791–822, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984572.

DuPont:1931:ALR

[Du 31] Francis I. Du Pont. The advantages of lessening radiation in the cylinders of internal combustion engines. Proceedings of the American Philosophical Society, 70(4):345–352, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984199.

DuVigneaud:1948:MMG

[Du 48] Vincent Du Vigneaud. Migration of the methyl group in the body. Proceedings of the American Philosophical Society, 92(3):127–135, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143580.

Duane:1916:SSN

[Dua16] Alexander Duane. Sight and signalling in the Navy. Proceedings of the American Philosophical Society, 55(5):400–414, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984054.

Duane:1922:RAP

[Dua22] William Duane. The radio-active point of view. Proceedings of the American Philosophical Society, 61(4):286–288, 1922. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984270.

Duane:1927:BF

[Dua27] Russell Duane. Benjamin Franklin. Proceedings of the American Philosophical Society, 66(??):729–735, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301091.

Dubois:1910:GJE

[Dub10] Patterson Dubois. The Great Japanese Embassy of 1860. A forgotten chapter in the history of international amity and commerce. Proceedings of the American Philosophical Society, 49(195):243–266, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983903.

Dubois:1914:SOP

[Dub14] Patterson Dubois. Some observations on the psychology of jurors and juries. Proceedings of the American Philosophical Society, 53(215): 307–322, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984080.

Dubos:1944:TSC

[Dub44] René J. Dubos. Trends in the study and control of infectious diseases. Proceedings of the American Philosophical Society, 88(3):208–213, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985416.

Dudley:1904:SPC

[Dud04] Chas. B. Dudley. A system of passenger car ventilation. Proceedings of the American Philosophical Society, 43(176):247–256, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983507.

Dugan:1915:SRO

[Dug15] Raymond S. Dugan. Some results from the observation of eclipsing variables. Proceedings of the American Philosophical Society, 54 (216):52–57, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983912.

Dumont:1948:HST

[Dum48] Paul-Emile Dumont. The horse-sacrifice in the Taittirīya-Brāhman.a: The eighth and ninth Prapāt.hakas of the Third Kān.d.a of the Taittirīya-Brāhman.a with translation. Proceedings of the American Philosophical Society, 92(6):447–503, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143199.

Dunham:1939:MIS

[Dun39] Theodore Dunham, Jr. The material of interstellar space. Proceedings of the American Philosophical Society, 81(2):277–293, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984986.

Durand:1917:MAA

[Dur17] W. F. Durand. Mechanical aspects of aëronautics. Proceedings of the American Philosophical Society, 56(3):170–188, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983945.

Dutilly:1944:ISL

[Dut44] Arthème Dutilly. An inexhaustible source of linguistic knowledge. Proceedings of the American Philosophical Society, 87(5):403–406, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985293.

deVries:1906:ESA

[dV06] Hugo de Vries. Elementary species in agriculture. Proceedings of the American Philosophical Society, 45(183):149–156, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983824.

Dyer:1944:RIW

[Dye44] R. E. Dyer. The role of immunization in wartime advances in medicine. Proceedings of the American Philosophical Society, 88 (3):182–188, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985412.

Eagleton:1930:CSI

[Eag30] Clyde Eagleton. The current status of international law. Proceedings of the American Philosophical Society, 69(1):203–215, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984356.

East:1915:ISC

[Eas15] E. M. East. An interpretation of sterility in certain plants. Proceedings of the American Philosophical Society, 54(216):70–72, January/ April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983915.

East:1940:DSS

[Eas40] E. M. East. The distribution of self-sterility in the flowering plants. Proceedings of the American Philosophical Society, 82(4):449–518, May 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984881.

Edgerton:1936:ISP

[Edg36] Franklin Edgerton. Indirect suggestion in poetry: A Hindu theory of literary æsthetics. Proceedings of the American Philosophical Society, 76(5):687–706, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984755.

Edgerton:1938:EI

[Edg38] Franklin Edgerton. Etymology and interpretation. Proceedings of the American Philosophical Society, 79(4):705–714, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984947.

Edgerton:1943:NEA

[Edg43] Franklin Edgerton. Notes on early American work in linguistics. Proceedings of the American Philosophical Society, 87(1):25–34, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984996.

Ehrenfeld:1916:JFF

[Ehr16] Frederick Ehrenfeld. Jointing as a fundamental factor in the degradation of the lithosphere. Proceedings of the American Philosophical Society, 55(5):363–399, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984053.

Eigenmann:1917:DSN

[Eig17a] Carl H. Eigenmann. Descriptions of sixteen new species of Pygidiidæ. Proceedings of the American Philosophical Society, 56(7):690– 703, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983889.

Eigenmann:1917:ENS

[Eig17b] Carl H. Eigenmann. Eighteen new species of fishes from Northwestern South America. Proceedings of the American Philosophical Society, 56 (7):673–689, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983888.

Eigenmann:1919:TPN

[Eig19] Carl H. Eigenmann. Trogloglanis Pattersoni a new blind fish from San Antonio, Texas. Proceedings of the American Philosophical Society, 58(6):397–400, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984264.

Eigenmann:1921:NOF

[Eig21a] Carl H. Eigenmann. The nature and origin of the fishes of the Pacific Slope of Ecuador, Peru and Chili. Proceedings of the American Philosophical Society, 60(4):503–523, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984516.

Eigenmann:1921:ODG

[Eig21b] Carl H. Eigenmann. The origin and distribution of the genera of the fishes of South America west of the Maracaibo, Orinoco, Amazon, and Titicaca basins. Proceedings of the American Philosophical Society, 60(1):1–6, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984221.

Eigenmann:1922:YFF

[Eig22] C. H. Eigenmann. Yellow fever and fishes. Proceedings of the American Philosophical Society, 61(3):204–211, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984228.

Eigenmann:1924:YFF

[Eig24] C. H. Eigenmann. Yellow fever and fishes in Colombia. Proceedings of the American Philosophical Society, 63(3):236–238, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984326.

Ekblaw:1940:AVD

[Ekb40] W. Elmer Ekblaw. The Arctic voyages and the discoveries of Dehaven, Kane and Hall. Proceedings of the American Philosophical Society, 82(5):877–887, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984896.

Eklund:1945:COR

[Ekl45] Carl R. Eklund. Condensed ornithology report, East Base, Palmer Land. Proceedings of the American Philosophical Society, 89(1):299– 304, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985339.

Ellsberg:1940:DJA

[Ell40] Edward Ellsberg. The drift of the “Jeannette” in the Arctic Sea. Proceedings of the American Philosophical Society, 82(5):889–896, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984897.

Elvehjem:1949:PDD

[Elv49] Conrad A. Elvehjem. Pellagra: A deficiency disease. Proceedings of the American Philosophical Society, 93(4):335–339, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143160.

Einstein:1949:DAE

[EM49] Albert Einstein and Robert A. Millikan. Dr. Albert Einstein and American colleagues, 1931. Proceedings of the American Philosophical Society, 93(7):543–545, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143144.

Emerson:1917:RTD

[Eme17] B. K. Emerson. Recurrent tetrahedral deformations and intercontinental torsions. Proceedings of the American Philosophical Society, 56 (6):445–472, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983968.

Emeneau:1937:ST

[Eme37] M. B. Emeneau. The songs of the Todas. Proceedings of the American Philosophical Society, 77(4):543–560, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984776.

Emeneau:1939:CT

[Eme39] M. B. Emeneau. The Christian Todas. Proceedings of the American Philosophical Society, 81(1):93–106, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984967.

Emeneau:1940:CIF

[Eme40] M. B. Emeneau. A classical Indian folk-tale as a reported modern event: The Brahman and the mongoose. Proceedings of the American Philosophical Society, 83(3):503–513, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985117.

Emmet:1903:CST

[Emm03] W. L. R. Emmet. The Curtis steam turbine. Proceedings of the American Philosophical Society, 42(172):68–84, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983643.

Emmet:1932:WMM

[Emm32] William Leroy Emmet. Why the markings on the Moon’s surface cannot be of volcanic origin. Proceedings of the American Philosophical Society, 71(5):285–294, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984703.

Evans:1937:QCA

[Eva37] Robley D. Evans. Quantum-counter amplifiers for gamma ray detection, and applications to studies in radium poisoning. Proceedings of the American Philosophical Society, 78(1):11–21, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984655.

Ewell:1907:PSS

[Ewe07] Marshall D. Ewell. A preliminary study of some modern micrometers. Proceedings of the American Philosophical Society, 46(185): 187–190, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983453.

Ewing:1938:MGM

[Ewi38] Maurice Ewing. Marine gravimetric methods and surveys. Proceedings of the American Philosophical Society, 79(1):47–70, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984679.

Ewing:1945:MUS

[Ewi45] H. E. Ewing. Mites of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1): 296, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985336.

Falk:1914:SFC

[Fal14] K. George Falk. Some further considerations in the development of the electron conception of valence. Proceedings of the American Philosophical Society, 53(213):25–30, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984128.

Falk:1946:RGA

[Fal46] I. S. Falk. Role of governmental agencies in a national health program. Proceedings of the American Philosophical Society, 90(4):265–270, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301019.

Fankhauser:1938:TNT

[Fan38] Gerhard Fankhauser. Triploidy in the newt, Triturus Viridescens. Proceedings of the American Philosophical Society, 79(4):715–730 + 732–739, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984948.

Farlow:1914:VSS

[Far14] William G. Farlow. The vegetation of the Sargasso Sea. Proceedings of the American Philosophical Society, 53(215):257–262, August/ December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984076.

Farabee:1917:SAI

[Far17] William Curtis Farabee. The South American Indian in his relation to geographic environment. Proceedings of the American Philosophical Society, 56(3):281–288, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983952.

Faris:1939:IUA

[Far39] Nabih Amin Faris. The Ih.yā’ ’Ulūm al-Dīn of al-Ghazzāli. Proceedings of the American Philosophical Society, 81(1):15–19, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984962.

Fennell:1901:NPC

[Fen01] C. A. M. Fennell. Notes on pure circulating decimals. Proceedings of the American Philosophical Society, 40(167):148–159, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983756.

Fenn:1948:PEA

[Fen48] Wallace O. Fenn. Physiology of exposures to abnormal concentrations of the respiratory gases. Proceedings of the American Philosophical Society, 92(3):144–154, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143583.

Fenton:1949:CMP

[Fen49a] William N. Fenton. Collecting materials for a political history of the six nations. Proceedings of the American Philosophical Society, 93(3):233–238, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143471.

Fenton:1949:SNT

[Fen49b] William N. Fenton. Seth Newhouse’s traditional history and constitution of the Iroquois Confederacy. Proceedings of the American Philosophical Society, 93(2):141–158, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143434.

Fernald:1942:SEB

[Fer42] M. L. Fernald. Some early botanists of the American Philosophical Society. Proceedings of the American Philosophical Society, 86(1):63–71, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985079.

Fermi:1946:DFC

[Fer46] Enrico Fermi. The development of the first chain reacting pile. Proceedings of the American Philosophical Society, 90(1):20–24, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301034.

Fetter:1943:EHP

[Fet43] Frank A. Fetter. The early history of political economy in the United States. Proceedings of the American Philosophical Society, 87(1):51– 60, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984999.

Fewkes:1932:ELN

[Few32] Vladimir J. Fewkes. Excavations in the Late Neolithic Fortress of Homolka in Bohemia: a preliminary report. Proceedings of the American Philosophical Society, 71(6):357–392, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984647.

Fewkes:1935:IDS

[Few35] Vladimir J. Fewkes. On the interpretation and dating of the site of “Belo Brdo” at Vinča in Yugoslavia. Proceedings of the American Philosophical Society, 75(7):651–672, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984812.

Fewkes:1937:NSY

[Few37] Vladimir J. Fewkes. Neolithic sites in the Yugoslavian portion of the Lower Danubian Valley. Proceedings of the American Philosophical Society, 78(2):329–406, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984541.

Fewkes:1944:CPM

[Few44] Vladimir J. Fewkes. Catawba pottery-making, with notes on Pamunkey pottery-making, Cherokee pottery-making, and coiling. Proceedings of the American Philosophical Society, 88(2):69–124, July 7, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985665.

Field:1938:IGS

[Fie38] Richard M. Field. The importance of geophysics to submarine geology. Proceedings of the American Philosophical Society, 79(1):1–7, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984674.

Finley:1924:WW

[Fin24] John H. Finley. Woodrow wilson. Proceedings of the American Philosophical Society, 63(3):iii–xi, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984329.

Fisher:1912:LMM

[Fis12] Sydney G. Fisher. The legendary and myth-making process in histories of the American Revolution. Proceedings of the American Philosophical Society, 51(204):53–75, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984095.

Fisher:1932:PDO

[Fis32] Irving Fisher. Progress and depressions — and our American dollar. Proceedings of the American Philosophical Society, 71(3):131–133, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984559.

Flanagan:1939:SPF

[Fla39] John C. Flanagan. A study of psychological factors related to fertility. Proceedings of the American Philosophical Society, 80(4):513–523, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985054.

Flexner:1930:A

[Fle30] Abraham Flexner. Address. Proceedings of the American Philosophical Society, 69(1):257–269, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984361.

Flexner:1935:WHW

[Fle35] Simon Flexner. William Henry Welch. Proceedings of the American Philosophical Society, 75(4):325–331, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984633.

Fleming:1938:TMO

[Fle38] J. A. Fleming. Terrestrial magnetism and oceanic structure. Proceedings of the American Philosophical Society, 79(1):109–125, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984682.

Fleming:1941:GWW

[Fle41] J. A. Fleming. Geomagnetism: World-wide and cosmic aspects with especial reference to early research in America. Proceedings of the American Philosophical Society, 84(2):263–298, 1941. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984923.

Fleming:1947:ISU

[Fle47] John A. Fleming. The international scientific unions. Proceedings of the American Philosophical Society, 91(1):121–125, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143133.

Foote:1913:FEV

[Foo13] Warren M. Foote. Factors in the exchange value of meteorites. Proceedings of the American Philosophical Society, 52(211):516– 542, September/October 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983974.

Foote:1948:DOF

[Foo48] Paul D. Foote. The discovery of an oil field. Proceedings of the American Philosophical Society, 92(1):15–25, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143621.

Ford:1936:ESD

[For36] William E. Ford. Edward Salisbury Dana. Proceedings of the American Philosophical Society, 76(2):237–241, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984668.

Ford:1943:AES

[For43] Guy Stanton Ford. America enters the scene. Proceedings of the American Philosophical Society, 87(2):121–125, August 16, 1943. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985087.

Fowler:1904:DFC

[Fow04] Henry W. Fowler. Description and figure of Coregonus Nelsonii bean. Proceedings of the American Philosophical Society, 43(178): 451–455, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983583.

Fowler:1940:FOW

[Fow40] Henry W. Fowler. The fishes obtained by the Wilkes Expedition, 1838–1842. Proceedings of the American Philosophical Society, 82(5):733–800, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984891.

Fox:1929:SOC

[Fox29] Herbert Fox. Some observations on comparative constitution in man and the lower mammals. Proceedings of the American Philosophical Society, 68(1):27–51, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984397.

Fox:1933:SNU

[Fox33] Herbert Fox. Some notes upon the nature, health and maintenance of the hyrax. Proceedings of the American Philosophical Society, 72(1):1–24, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984605.

Ferree:1918:LRE

[FR18] C. E. Ferree and G. Rand. Lighting in its relation to the eye. Proceedings of the American Philosophical Society, 57(5):440–478, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984181.

Francke:1908:MGS

[Fra08] Kuno Francke. Medieval German sculpture in the Germanic Museum of Harvard University. Proceedings of the American Philosophical Society, 47(190):636–643, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983833.

Francke:1927:CGR

[Fra27] Kuno Francke. Cosmopolitanism in German Romantic thought. Proceedings of the American Philosophical Society, 66(??):183–190, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301063.

Frank:1931:SEA

[Fra31a] Tenney Frank. Some economic aspects of Rome’s early law. Proceedings of the American Philosophical Society, 70(2):193–205, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984322.

Frankel:1931:LSL

[Fra31b] Lee K. Frankel. Lengthening the span of life. Proceedings of the American Philosophical Society, 70(3):241–254, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984453.

Frank:1944:PMC

[Fra44] Philipp Frank. The philosophical meaning of the Copernican Revolution. Proceedings of the American Philosophical Society, 87(5):381– 386, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985289.

Frazier:1945:AEC

[Fra45] Russell G. Frazier. Acclimatization and the effects of cold on the human body as observed at Little America III, on the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):249–255, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985327.

Friedmann:1945:BUS

[Fri45a] Herbert Friedmann. Birds of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):305–313, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985340.

Friedmann:1945:PBW

[Fri45b] Herbert Friedmann. Pelagic birds from the West Coast of South America. Proceedings of the American Philosophical Society, 89 (1):314–315, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985341.

Frost:1919:ROE

[Fro19] Edwin B. Frost. Results of observations of the eclipse by the expedition from the Yerkes Observatory. Proceedings of the American Philosophical Society, 58(4):282–288, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984406.

Fullerton:1909:IDM

[Ful09] George Stuart Fullerton. The influence of Darwin on the mental and moral sciences. Proceedings of the American Philosophical Society, 48(193):xxv–xxxvii, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984065.

Fuller:1936:ALR

[Ful36] L. L. Fuller. American legal realism. Proceedings of the American Philosophical Society, 76(2):191–235, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984667.

Furness:1916:OMC

[Fur16] William H. Furness. Observations on the mentality of chimpanzees and orang-utans. Proceedings of the American Philosophical Society, 55(3):281–290, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984118.

Gallup:1942:HIP

[Gal42] George Gallup. How important is public opinion in time of war? Proceedings of the American Philosophical Society, 85(5):440–444, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985063.

Gamble:1944:WRC

[Gam44] James L. Gamble. The water requirements of castaways. Proceedings of the American Philosophical Society, 88(3):151–158, September 8, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985408.

Gaposchkin:1938:SAD

[Gap38] Sergei I. Gaposchkin. A study of the absolute dimensions of eclipsing variables. Proceedings of the American Philosophical Society, 79 (3):327–340, September 30, 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984853.

Gaposchkin:1940:MRO

[Gap40] Sergei Gaposchkin. Masses, radii, and other absolute dimensions for 224 eclipsing variables. Proceedings of the American Philosophical Society, 82(3):291–322, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984931.

Gartlein:1941:AG

[Gar41] C. W. Gartlein. Aurora and geomagnetism. Proceedings of the American Philosophical Society, 84(2):299–307, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984924.

Gazin:1936:TNS

[Gaz36] C. Lewis Gazin. A Tæniodont skull from the Lower Eocene of Wyoming. Proceedings of the American Philosophical Society, 76(5): 597–612, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984749.

Gerould:1943:GSV

[Ger43] John H. Gerould. Genetic and seasonal variations of orange wingcolor in “Colias” butterflies. Proceedings of the American Philosophical Society, 86(3):405–438, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985029.

Gesell:1941:GBF

[Ges41] Arnold Gesell. The genesis of behavior form in fetus and infant. The growth of the mind from the standpoint of developmental morphology. Proceedings of the American Philosophical Society, 84(4):471– 488, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984952.

Goetz:1938:VCP

[GG38] Alexander Goetz and S. Scott Goetz. Vitrification and crystallization of protophyta at low temperatures. Proceedings of the American Philosophical Society, 79(3):361–388, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984855.

Gregory:1939:EMC

[GH39] William K. Gregory and Milo Hellman. On the evolution and major classification of the civets (Viverridae) and allied fossil and recent carnivora: A phylogenetic study of the skull and dentition. Proceedings of the American Philosophical Society, 81(3):309–392, August 31, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984861.

Gilmore:1942:PFP

[Gil42] Charles W. Gilmore. Paleocene faunas of the Polecat Bench Formation, Park County, Wyoming. Part II. Lizards. Proceedings of the American Philosophical Society, 85(2):159–167, January 1942. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985036.

Gish:1941:TER

[Gis41] O. H. Gish. Terrestrial electricity in relation to geomagnetism. Proceedings of the American Philosophical Society, 84(2):187–204, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984918.

Glance:1945:CUS

[Gla45] Grace Glance. Collembola of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):295, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985335.

Goddard:1912:HFM

[God12] Henry H. Goddard. Heredity of feeble-mindedness. Proceedings of the American Philosophical Society, 51(205):173–177, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984134.

Goldenweiser:1924:RPE

[Gol24] Alexander Goldenweiser. Are the races potentially equal? Proceedings of the American Philosophical Society, 63(2):215–221, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984475.

Goodspeed:1903:PFS

[Goo03] Arthur W. Goodspeed. On the properties of the field surrounding a Crookes tube. Proceedings of the American Philosophical Society, 42 (172):96–103, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983645.

Goodwin:1904:EC

[Goo04] Joseph H. Goodwin. Electrolytic calcium. Proceedings of the American Philosophical Society, 43(178):381–392, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983579.

Goodale:1909:IDN

[Goo09] George Lincoln Goodale. The influence of Darwin on the natural sciences. Proceedings of the American Philosophical Society, 48(193): xv–xxiv, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984064.

Goodspeed:1931:LWM

[Goo31] Arthur W. Goodspeed. Leslie William Miller. Proceedings of the American Philosophical Society, 70(5):399–401, 1931. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984494.

Goodrich:1948:PSA

[Goo48] Carter Goodrich. Public spirit and American improvements. Proceedings of the American Philosophical Society, 92(4):305–309, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143360.

Gould:1940:GA

[Gou40] Laurence M. Gould. Glaciers of Antarctica. Proceedings of the American Philosophical Society, 82(5):835–876, June 1940. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984895.

Gowen:1934:SAG

[Gow34] John W. Gowen. Some aspects of genetic constitution in relation to pathology. Proceedings of the American Philosophical Society, 74(3):215–226, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984715.

Graves:1928:PSR

[Gra28] Frank Pierrepont Graves. The present status of research in education. Proceedings of the American Philosophical Society, 67(2):175– 185, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984278.

Gregory:1927:OMA

[Gre27] William K. Gregory. The origin of man from the anthropoid stem: When and where? Proceedings of the American Philosophical Society, 66(??):439–463, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301078.

Gregory:1928:UPM

[Gre28a] William K. Gregory. The upright posture of man: A review of its origin and evolution. Proceedings of the American Philosophical Society, 67(4):339–377, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984249.

Gregory:1928:WAM

[Gre28b] William K. Gregory. Were the ancestors of man primitive brachiators? Proceedings of the American Philosophical Society, 67(2):129– 150, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984275.

Greene:1929:SEV

[Gre29] Evarts B. Greene. Some educational values of the American Revolution. Proceedings of the American Philosophical Society, 68(3): 185–194, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984207.

Greber:1934:ASS

[Grè34a] Jacques Grèber. Æsthetic and sociological aspects of city and regional planning. Proceedings of the American Philosophical Society, 74(1): 15–19, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984581.

Gregory:1934:HCT

[Gre34b] William King Gregory. A half century of trituberculy: the Cope– Osborn theory of dental evolution with a revised summary of molar evolution from fish to man. Proceedings of the American Philosophical Society, 73(4):169–317, April 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984748.

Gregory:1935:FOP

[Gre35] William K. Gregory. Further observations on the pectoral girdle and fin of Sauripterus Taylori Hall, a crossopterygian fish from the Upper Devonian of Pennsylvania, with special reference to the origin of the pentadactylate extremities of Tetrapoda. Proceedings of the American Philosophical Society, 75(7):673–690, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984813.

Gregory:1936:HFS

[Gre36a] William K. Gregory. Habitus factors in the skeleton of fossil and recent mammals. Proceedings of the American Philosophical Society, 76 (4):429–444, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984596.

Gregory:1936:HFO

[Gre36b] William King Gregory. Henry Fairfield Osborn. Proceedings of the American Philosophical Society, 76(3):395–408, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984553.

Gregg:1944:CMR

[Gre44] Alan Gregg. A critique of medical research. Proceedings of the American Philosophical Society, 87(4):313–320, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985043.

Grote:1900:DP

[Gro00a] A. Radcliffe Grote. The descent of the Pierids. Proceedings of the American Philosophical Society, 39(161):4–67, January 1900. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983541.

Grote:1900:HSG

[Gro00b] A. Radcliffe Grote. Historical sketch of Gortyna and allied genera. Proceedings of the American Philosophical Society, 39(162):346–357, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983665.

Grote:1902:ROS

[Gro02a] A. Radcliffe Grote. Results obtained from a search for the type of Noctua Linn., and conclusions as to types of Huebnerian noctuid genera represented in the North American fauna. Proceedings of the American Philosophical Society, 41(168):4–20, January 1902. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983518.

Grote:1902:SCT

[Gro02b] A. Radcliffe Grote. A specialized cocoon of Telea Polyphemus. Proceedings of the American Philosophical Society, 41(171):401– 406, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983487.

Guyer:1923:GCS

[Guy23] M. F. Guyer. The germ-cell and serological influences. Proceedings of the American Philosophical Society, 62(5):274–291, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984461.

Haehl:1904:MDS

[HA04] H. L. Haehl and Ralph Arnold. The Miocene diabase of the Santa Cruz Mountains in San Mateo County, California. Proceedings of the American Philosophical Society, 43(175):15–53, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983413.

Hale:1914:SMP

[Hal14] George E. Hale. Solar magnetic phenomena. Proceedings of the American Philosophical Society, 53(214):251–255, June/July 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984056.

Hall:1924:VRM

[Hal24] H. N. Hall. Are the various races of man potentially equal? Proceedings of the American Philosophical Society, 63(2):208–214, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984474.

Hallborg:1941:CSW

[Hal41] H. E. Hallborg. Correlations of short wave radio transmission across the Atlantic with magnetic conditions. Proceedings of the American Philosophical Society, 84(2):323–337, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984926.

Halecki:1943:PSD

[Hal43] Oskar Halecki. The problem of self-determinism. Proceedings of the American Philosophical Society, 87(2):194–198, August 1943. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985097.

Hall:1948:SFP

[Hal48] Vernon Hall, Jr. The Scaliger family papers. Proceedings of the American Philosophical Society, 92(2):120–123, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143411.

Halstead:1949:GST

[Hal49] Frank G. Halstead. The genesis and speed of the telegraph codes. Proceedings of the American Philosophical Society, 93(5):448–458, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143392.

Hammett:1929:CSE

[Ham29] Frederick S. Hammett. The chemical stimulus essential for growth by increase in cell number. Proceedings of the American Philosophical Society, 68(2):151–161, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984315.

Hammett:1930:NCR

[Ham30] Frederick S. Hammett. The natural chemical regulation of growth by increase in cell number. Proceedings of the American Philosophical Society, 69(1):217–223, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984357.

Hanson:1923:MAR

[Han23] Frank Blair Hanson. Modifications in the albino rat following treatment with alcohol fumes and X-rays; and the problem of their inheritance. Proceedings of the American Philosophical Society, 62(5): 301–310, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984463.

Haney:1944:PPP

[Han44] John L. Haney. Of the people, by the people, for the people. Proceedings of the American Philosophical Society, 88(5):359–367, November 7, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985609.

Harshberger:1906:GKS

[Har06] John W. Harshberger. A grass-killing slime mould. Proceedings of the American Philosophical Society, 45(184):271–273, October/ December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983380.

Harshberger:1907:TCM

[Har07] John W. Harshberger. Taxonomic charts of the monocotyledons and the dicotyledons. Proceedings of the American Philosophical Society, 46(187):313–321, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983470.

Harshberger:1908:CLS

[Har08] John W. Harshberger. The comparative leaf structure of the sand dune plants of Bermuda. Proceedings of the American Philosophical Society, 47(188):97–110, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983803.

Harshberger:1909:CLS

[Har09] John W. Harshberger. The comparative leaf structure of the strand plants of New Jersey. Proceedings of the American Philosophical Society, 48(191):72–89, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983814.

Harshberger:1911:HII

[Har11] John W. Harshberger. An hydrometric investigation of the influence of sea water on the distribution of salt marsh and estuarine plants. Proceedings of the American Philosophical Society, 50(201): 457–496, September/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984071.

Harshberger:1916:OVS

[Har16] John W. Harshberger. The origin and vegetation of salt marsh pools. Proceedings of the American Philosophical Society, 55(6):481–484, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983982.

Harper:1918:ORI

[Har18] R. A. Harper. Organization, reproduction and inheritance in pediastrum. Proceedings of the American Philosophical Society, 57(5): 375–439, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984180.

Harshberger:1922:EMS

[Har22] John W. Harshberger. Ecologic and morphologic study of the clovers (Trifolium). Proceedings of the American Philosophical Society, 61 (2):136–150, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984411.

Harshberger:1925:NPI

[Har25] John W. Harshberger. Notes on the Portuguese insectivorous plant, “Drosophyllum Lusitanicum”. Proceedings of the American Philosophical Society, 64(1):51–54, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984302.

Harshberger:1926:MGM

[Har26] John W. Harshberger. Mediterranean Garigue and Macchia. Proceedings of the American Philosophical Society, 65(1):56–63, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984338.

Harshberger:1928:TVC

[Har28] John W. Harshberger. Tundra vegetation of Central Alaska directly under the Arctic Circle. Proceedings of the American Philosophical Society, 67(3):215–234, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984291.

Harshberger:1929:VCJ

[Har29a] John W. Harshberger. The vegetation of Campos De Jordão, Brazil. Proceedings of the American Philosophical Society, 68(2):83–92, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984308.

Harshberger:1929:VST

[Har29b] John W. Harshberger. The vegetation of the screes, or talus slopes of Western North America. Proceedings of the American Philosophical Society, 68(1):13–25, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984396.

Harvey:1932:EBR

[Har32] E. Newton Harvey. The evolution of bioluminescence and its relation to cell respiration. Proceedings of the American Philosophical Society, 71(4):135–141, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984785.

Harrington:1933:FEC

[Har33] Elmer A. Harrington. Further experiments on the continuous generation of heat in certain silicates. Proceedings of the American Philosophical Society, 72(5):333–349, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984697.

Harvey:1940:BFV

[Har40] E. Newton Harvey. Benjamin Franklin’s views on the phosphorescence of the sea. Proceedings of the American Philosophical Society, 83(2):341–348, August 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985015.

Hart:1948:BIS

[Har48a] Clyde W. Hart. Bias in interviewing in studies of opinions, attitudes, and consumer wants. Proceedings of the American Philosophical Society, 92(5):399–404, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143052.

Harvey:1948:MWH

[Har48b] E. Newton Harvey. The mechanism of wounding by high velocity missiles. Proceedings of the American Philosophical Society, 92(4):294–304, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143359.

Hastings:1900:DLI

[Has00] W. G. Hastings. The development of law as illustrated by the decisions relating to the police power of the state. Proceedings of the American Philosophical Society, 39(163):359–554, September 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983818.

Hatcher:1902:OOM

[Hat02] J. B. Hatcher. Origin of the Oligocene and Miocene deposits of the Great Plains. Proceedings of the American Philosophical Society, 41(169):113–131, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983527.

Haupt:1901:MIO

[Hau01] Lewis M. Haupt. Methods of improving ocean bars. Proceedings of the American Philosophical Society, 40(166):62–85, July 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983572.

Haupt:1903:REA

[Hau03] Lewis M. Haupt. Reaction as an efficient agent in procuring deeper navigable channels in the improvement of rivers and harbors. Proceedings of the American Philosophical Society, 42(173):199–211, April/ May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983733.

Haupt:1904:MRP

[Hau04] Lewis M. Haupt. The Mississippi River problem. Proceedings of the American Philosophical Society, 43(175):71–96, January 1904. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983417.

Haupt:1905:EW

[Hau05] Lewis M. Haupt. Emancipation of the waterways. Proceedings of the American Philosophical Society, 44(179):42–50, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983604.

Haupt:1907:TUS

[Hau07a] Lewis M. Haupt. Transportation in the United States. Proceedings of the American Philosophical Society, 46(185):171–181, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983451.

Haupt:1907:JW

[Hau07b] Paul Haupt. Jonah’s whale. Proceedings of the American Philosophical Society, 46(185):151–164, January/April 1907. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983449.

Haupt:1909:NW

[Hau09a] Lewis M. Haupt. The nation and the waterways. Proceedings of the American Philosophical Society, 48(191):51–64, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983810.

Haupt:1909:BBO

[Hau09b] Paul Haupt. The burning bush and the origin of Judaism. Proceedings of the American Philosophical Society, 48(193):354–369, September/ December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984060.

Haupt:1911:APA

[Hau11] Paul Haupt. An ancient protest against the curse on eve. Proceedings of the American Philosophical Society, 50(201):505–517, September/ December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984073.

Haupt:1920:G

[Hau20] Paul Haupt. Golgotha. Proceedings of the American Philosophical Society, 59(3):237–244, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984501.

Haupt:1921:TBS

[Hau21] Paul Haupt. Tobit’s blindness and Sara’s hysteria. Proceedings of the American Philosophical Society, 60(2):71–95, 1921. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984385.

Haupt:1922:MNA

[Hau22] Paul Haupt. Manna, nectar, and ambrosia. Proceedings of the American Philosophical Society, 61(3):227–236, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984230.

Haury:1943:PCM

[Hau43] Emil W. Haury. A possible Cochise–Mogollon–Hohokam sequence. Proceedings of the American Philosophical Society, 86(2):260–263, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985104.

Hawthorne:1945:EFA

[Haw45] Roger Hawthorne. Exploratory flights of Admiral Byrd (1940). Proceedings of the American Philosophical Society, 89(1):398a–398e, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985350.

Hay:1900:DSV

[Hay00a] O. P. Hay. Descriptions of some vertebrates of the carboniferous age. Proceedings of the American Philosophical Society, 39(161):96–123, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983550.

Hays:1900:CBD

[Hay00b] I. Minis Hays. A contribution to the bibliography of the declaration of independence. Proceedings of the American Philosophical Society, 39(161):69–78, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983544.

Hay:1903:EGT

[Hay03] O. P. Hay. On the existing genera of the Trionychidæ. Proceedings of the American Philosophical Society, 42(174):268–274, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983713.

Hay:1911:FSA

[Hay11] Oliver P. Hay. A fossil specimen of the alligator snapper (Macrochelys Temminckii) from Texas. Proceedings of the American Philosophical Society, 50(200):452–455, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983934.

Hayford:1915:EGS

[Hay15] John F. Hayford. The Earth from the geophysical standpoint. Proceedings of the American Philosophical Society, 54(219):298–308, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983876.

Hayes:1920:UNM

[Hay20a] H. C. Hayes. U. S. Navy MV type of hydrophone as an aid and safeguard to navigation. Proceedings of the American Philosophical Society, 59(5):371–404, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984174.

Hayes:1920:DS

[Hay20b] Harvey C. Hayes. Detection of submarines. Proceedings of the American Philosophical Society, 59(1):1–47, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984218.

Hayes:1924:SDF

[Hay24a] Harvey C. Hayes. The sonic depth finder. Proceedings of the American Philosophical Society, 63(1):134–151, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984445.

Hays:1924:AAP

[Hay24b] I. Minis Hays. On the authorship of the anonymous pamphlet published in London in 1760 entitled “The Interest of Great Britain Considered with Regard to Her Colonies and the Acquisition of Canada and Guadaloupe”. Proceedings of the American Philosophical Society, 63(1):1–9, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984437.

Hayes:1940:NTH

[Hay40] Carlton J. H. Hayes. The novelty of totalitarianism in the history of Western civilization. Proceedings of the American Philosophical Society, 82(1):91–102, February 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984908.

Heyl:1922:EIC

[HB22] Paul R. Heyl and Lyman J. Briggs. The Earth inductor compass. Proceedings of the American Philosophical Society, 61(1):15–32, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984431.

Hertlein:1925:SNS

[HC25] Leo G. Hertlein and Colin H. Crickmay. A summary of the nomenclature and stratigraphy of the marine tertiary of Oregon and Washington. Proceedings of the American Philosophical Society, 64(2):224– 282, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984194.

Ham:1940:SDR

[HC40] Thomas Hale Ham and William B. Castle. Studies on destruction of red blood cells: Relation of increased hypotonic fragility and of erythro-stasis to the mechanism of hemolysis in certain anemias. Proceedings of the American Philosophical Society, 82(4):411–419, May 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984878.

Heaton:1940:DT

[Hea40] Herbert Heaton. Discussion of totalitarianism. Proceedings of the American Philosophical Society, 82(1):89–90, February 1940. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984907.

Hebb:1942:EEL

[Heb42] D. O. Hebb. The effect of early and late brain injury upon test scores, and the nature of normal adult intelligence. Proceedings of the American Philosophical Society, 85(3):275–292, February 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985007.

Heck:1938:RES

[Hec38] N. H. Heck. The role of earthquakes and the seismic method in submarine geology. Proceedings of the American Philosophical Society, 79(1):97–108, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984681.

Heck:1941:MSU

[Hec41] N. H. Heck. The magnetic survey of the United States. Proceedings of the American Philosophical Society, 84(2):205–224, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984919.

Heiser:1918:ASP

[Hei18] Victor G. Heiser. American sanitation in the Philippines and its influence on the Orient. Proceedings of the American Philosophical Society, 57(1):60–68, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983882.

Heiser:1926:HWL

[Hei26] Victor G. Heiser. The health work of the league of nations. Proceedings of the American Philosophical Society, 65(5):1–9, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984282.

Heilbrunn:1930:SPR

[Hei30] L. V. Heilbrunn. The surface precipitation reaction of living cells. Proceedings of the American Philosophical Society, 69(1):295–301, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984364.

Heiser:1932:RPC

[Hei32] Victor G. Heiser. Recent progress in the control of leprosy. Proceedings of the American Philosophical Society, 71(4):167–171, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984787.

Heilbrunn:1934:AAS

[Hei34] L. V. Heilbrunn. The action of anæsthetics on living protoplasm. Proceedings of the American Philosophical Society, 74(3):159–165, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984710.

Heizer:1947:PSK

[Hei47] Robert F. Heizer. Petroglyphs from Southwestern Kodiak Island, Alaska. Proceedings of the American Philosophical Society, 91(3):284– 293, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143401.

Heizer:1949:PSE

[Hei49] Robert F. Heizer. Pottery from the Southern Eskimo Region. Proceedings of the American Philosophical Society, 93(1):48–56, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143271.

Hellman:1928:RCH

[Hel28] Milo Hellman. Racial characters in human dentition. Part I. A racial distribution of the Dryopithecus pattern and its modifications in the lower molar teeth of man. Proceedings of the American Philosophical Society, 67(2):157–174, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984277.

Henderson:1923:LSL

[Hen23] Lawrence T. Henderson. The life and services of Louis Pasteur. Proceedings of the American Philosophical Society, 62(1):iii–xiii, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984177.

Hencken:1948:PAT

[Hen48] Hugh Hencken. The prehistoric archaeology of Tangier Zone, Morocco. Proceedings of the American Philosophical Society, 92(4): 282–288, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143357.

Hepburn:1918:BSP

[Hep18] Joseph Samuel Hepburn. Biochemical studies of the pitcher liquor of Nepenthes. Proceedings of the American Philosophical Society, 57(2): 112–129, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984163.

Hertlein:1937:NSS

[Her37] Leo George Hertlein. A note on some species of marine mollusks occurring in both Polynesia and the Western Americas. Proceedings of the American Philosophical Society, 78(2):303–312, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984539.

Herzfeld:1940:TLA

[Her40] Karl F. Herzfeld. Theory of light absorption in simple aromatic compounds. Proceedings of the American Philosophical Society, 82 (3):359–367, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984935.

Hess:1926:NKP

[Hes26] Alfred F. Hess. The newer knowledge of the physiological action of ultra-violet rays. Proceedings of the American Philosophical Society, 65(3):202–206, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984490.

Hess:1938:GAI

[Hes38] Harry Hammond Hess. Gravity anomalies and island arc structure with particular reference to the West Indies. Proceedings of the American Philosophical Society, 79(1):71–96, April 1938. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984680.

Hewett:1904:HUR

[Hew04] Waterman T. Hewett. The historical use of the relative pronouns in English literature. Proceedings of the American Philosophical Society, 43(177):278–289, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983766.

Heyl:1910:CEC

[Hey10] Paul R. Heyl. The conversion of the energy of carbon into electrical energy on solution in iron. Proceedings of the American Philosophical Society, 49(194):49–51, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983990.

Heyl:1913:PNC

[Hey13] Paul R. Heyl. Platinum in North Carolina. Proceedings of the American Philosophical Society, 52(208):21–30, January/April 1913. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983996.

Heyl:1941:MU

[Hey41] Paul R. Heyl. Magnetism and its uses. Proceedings of the American Philosophical Society, 84(2):339–351, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984927.

Hornbeck:1941:PSE

[HH41] George Hornbeck and Irl Howell. Production of secondary electrons by electrons of energy between 0.7 and 2.6 MEV. Proceedings of the American Philosophical Society, 84(1):33–51, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984849.

Howe:1917:SSP

[HHHN17] H. E. Howe, H. L. Howes, Percy Hodge, and Edward L. Nichols. Spectral structure of the phosphorescence of certain sulphides. Proceedings of the American Philosophical Society, 56(3):258–274, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983950.

Hibben:1943:DSC

[Hib43] Frank C. Hibben. Discoveries in Sandia Cave and early horizons in the Southwest. Proceedings of the American Philosophical Society, 86(2):247–254, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985102.

Hill:1916:WO

[Hil16] David Jayne Hill. World organization. Proceedings of the American Philosophical Society, 55(4):322–329, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984105.

Hinrichs:1910:TAW

[Hin10] Gustavus Hinrichs. The true atomic weights of oxygen and silver. Proceedings of the American Philosophical Society, 49(197): 359–363, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984090.

Hinrichs:1911:AWV

[Hin11] Gustavus D. Hinrichs. The atomic weight of vanadium determined from the laboratory work of eighty years. Proceedings of the American Philosophical Society, 50(199):191–216, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984034.

Hinrichs:1913:TAW

[Hin13] Gustavus Hinrichs. The true atomic weight of bromine. Proceedings of the American Philosophical Society, 52(211):543–557, September/ October 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983975.

Hirst:1909:RAF

[Hir09] Barton C. Hirst. Richard Alexander Fullerton Penrose, M.D., LL.D. Proceedings of the American Philosophical Society, 48(193):lviii– lxi, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984067.

Hammond:1937:CSF

[HK37] Datus M. Hammond and Charles A. Kofoid. The continuity of structure and function in the neuromotor system of Euplotes patella during its life cycle. Proceedings of the American Philosophical Society, 77(2):207–218, February 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984727.

Hober:1938:ECS

[HM38] Rudolf Höber and Elinor Moore. Experiments concerning the secretory activity of the liver. Proceedings of the American Philosophical Society, 78(4):587–593, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984804.

Hillebrand:1914:HMP

[HMW14] W. F. Hillebrand, H. E. Merwin, and Fred E. Wright. Hewettite, metahewettite and pascoite, hydrous calcium vanadates. Proceedings of the American Philosophical Society, 53(213):31–54, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984129.

Hobbs:1909:EOS

[Hob09] William Herbert Hobbs. The evolution and the outlook of seismic geology. Proceedings of the American Philosophical Society, 48 (192):259–302, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984158.

Hobbs:1910:CII

[Hob10] Wm. Herbert Hobbs. Characteristics of the inland-ice of the Arctic regions. Proceedings of the American Philosophical Society, 49 (194):57–129, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983992.

Hobbs:1915:RGA

[Hob15] William Herbert Hobbs. The rôle of the glacial anticyclone in the air circulation of the globe. Proceedings of the American Philosophical Society, 54(218):185–225, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983935.

Hobbs:1921:FGA

[Hob21] William Herbert Hobbs. The fixed glacial anticyclone compared to the migrating anticyclone. Proceedings of the American Philosophical Society, 60(1):34–42, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984225.

Hobbs:1923:RMV

[Hob23] William Herbert Hobbs. The rate of movement in vertical earth adjustments connected with the growth of mountains. Proceedings of the American Philosophical Society, 62(2):63–73, 1923. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984511.

Hobbs:1940:DWL

[Hob40] William Herbert Hobbs. The discovery of Wilkes Land, Antarctica. Proceedings of the American Philosophical Society, 82(5):561–582, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984886.

Hobbs:1943:GAC

[Hob43] William Herbert Hobbs. The glacial anticyclone and the continental glaciers of North America. Proceedings of the American Philosophical Society, 86(3):368–402, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985027.

Hobbs:1944:MGS

[Hob44] William Herbert Hobbs. Mountain growth, a study of the Southwestern Pacific Region. Proceedings of the American Philosophical Society, 88(4):221–268, October 18, 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985450.

Hoffleit:1939:OS

[Hof39a] Dorrit Hoffleit. Observations of supernovæ. Proceedings of the American Philosophical Society, 81(2):265–276, June 1939. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984985.

Hoffleit:1939:NSS

[Hof39b] Dorrit Hoffleit. On the necessity for spectral sub-classification, with a discussion of preliminary results for 1300 Southern stars. Proceedings of the American Philosophical Society, 81(4):533–545, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984825.

Hoffmeister:1940:JDD

[Hof40] J. Edward Hoffmeister. James Dwight Dana’s studies of volcanoes and of coral islands. Proceedings of the American Philosophical Society, 82(5):721–732, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984890.

Hollon:1947:ZMP

[Hol47] W. Eugene Hollon. Zebulon Montgomery Pike and the Wilkinson– Burr Conspiracy. Proceedings of the American Philosophical Society, 91(5):447–456, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143425.

Hooton:1926:PSR

[Hoo26] Earnest A. Hooton. Progress in the study of race mixtures with special reference to work carried on at Harvard University. Proceedings of the American Philosophical Society, 65(4):312–325, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984239.

Hooton:1932:PRA

[Hoo32] Earnest A. Hooton. Preliminary remarks on the anthropology of the American criminal. Proceedings of the American Philosophical Society, 71(6):349–355, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984646.

Hooton:1935:DCR

[Hoo35a] Earnest A. Hooton. Development and correlation of research in physical anthropology at Harvard University. Proceedings of the American Philosophical Society, 75(6):499–516, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984611.

Hooton:1935:RBD

[Hoo35b] Earnest A. Hooton. Roland Burrage Dixon. Proceedings of the American Philosophical Society, 75(8):770–774, 1935. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984594.

Hooker:1938:OGM

[Hoo38] Davenport Hooker. The origin of the grasping movement in man. Proceedings of the American Philosophical Society, 79(4):597–606, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984941.

Hopkins:1910:MOH

[Hop10] E. Washburn Hopkins. Magic observances in the Hindu epic. Proceedings of the American Philosophical Society, 49(194):24–40, January/ June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983988.

Hopkins:1931:UON

[Hop31] B. Smith Hopkins. Untangling one of Nature’s puzzles. Proceedings of the American Philosophical Society, 70(2):207–213, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984323.

Hough:1932:FHC

[Hou32] Walter Hough. Fire and human civilization. Proceedings of the American Philosophical Society, 71(6):403–406, April 1932. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984649.

Houston:1939:EM

[Hou39] W. V. Houston. Electrons in metals. Proceedings of the American Philosophical Society, 81(4):525–532, September 1939. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984824.

Hovey:1909:ETC

[Hov09] Edmund Otis Hovey. Earthquakes: Their causes and effects. Proceedings of the American Philosophical Society, 48(192):235–258, April/ August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984157.

Hovland:1948:SC

[Hov48] Carl I. Hovland. Social communication. Proceedings of the American Philosophical Society, 92(5):371–375, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143048.

Howard:1936:EMA

[How36] Edgar B. Howard. Early man in America. Proceedings of the American Philosophical Society, 76(3):327–333, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984547.

Howland:1941:SDC

[How41] Ruth B. Howland. Structure and development of centrifuged eggs and early embryos of Drosophila Melanogaster. Proceedings of the American Philosophical Society, 84(5):605–616, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984842.

Howard:1943:FYP

[How43] Edgar B. Howard. Folsom and Yuma problems. Proceedings of the American Philosophical Society, 86(2):255–259, February 1943. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985103.

Hrdlicka:1921:PA

[Hrd21] Aleš Hrdlička. The peopling of Asia. Proceedings of the American Philosophical Society, 60(4):535–545, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984518.

Hrdlicka:1926:PMA

[Hrd26a] Aleš Hrdlička. The people of the main American cultures. Proceedings of the American Philosophical Society, 65(3):157–160, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984485.

Hrdlicka:1926:PE

[Hrd26b] Aleš Hrdlička. The peopling of the Earth. Proceedings of the American Philosophical Society, 65(3):150–156, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984484.

Hrdlicka:1929:MFL

[Hrd29] Aleš Hrdlička. Man’s future in the light of his past and present. Proceedings of the American Philosophical Society, 68(1):1–11, 1929.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984395.

Hrdlicka:1932:CMA

[Hrd32] Aleš Hrdlička. The coming of Man from Asia in the light of recent discoveries. Proceedings of the American Philosophical Society, 71(6):393–402, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984648.

Hrdlicka:1933:F

[Hrd33] Aleš Hrdlička. The forehead. Proceedings of the American Philosophical Society, 72(5):315–324, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984695.

Hrdlicka:1934:NV

[Hrd34] Aleš Hrdlička. Normal variation. Proceedings of the American Philosophical Society, 74(3):253–261, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984718.

Hrdlicka:1935:SRR

[Hrd35] Aleš Hrdlička. Some reflections regarding human heredity. Proceedings of the American Philosophical Society, 75(4):295–312, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984631.

Hrdlicka:1936:GDA

[Hrd36] Aleš Hrdlička. Growth during adult life. Proceedings of the American Philosophical Society, 76(6):847–897, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984574.

Hrdlicka:1937:HT

[Hrd37] Aleš Hrdlička. Human typogeny. Proceedings of the American Philosophical Society, 78(1):79–95, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984658.

Hrdlicka:1943:CHP

[Hrd43] Aleš Hrdlička. Contribution to the history of physical anthropology in the United States of America, with special reference to Philadelphia. Proceedings of the American Philosophical Society, 87(1):61–64, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985000.

Hatcher:1904:ACM

[HS04] J. B. Hatcher and T. W. Stanton. An attempt to correlate the marine with the non-marine formations of the Middle West. Proceedings of the American Philosophical Society, 43(178):341–365, October/ December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983576.

Hall:1905:SOC

[HS05] Roy D. Hall and Edgar F. Smith. Some observations on columbium. Proceedings of the American Philosophical Society, 44(180):177–212, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983515.

Harvey:1930:KBR

[HS30] E. Newton Harvey and Peter A. Snell. The kinetics of bioluminescent reactions of short duration. Proceedings of the American Philosophical Society, 69(1):303–308, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984365.

Haas:1946:ASP

[HS46] Otto Haas and George Gaylord Simpson. Analysis of some phylogenetic terms, with attempts at redefinition. Proceedings of the American Philosophical Society, 90(5):319–349, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301045.

Huebner:1932:UI

[Hue32] S. S. Huebner. Unemployment insurance. Proceedings of the American Philosophical Society, 71(2):49–71, March 1932. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984685.

Huebner:1934:EFS

[Hue34] S. S. Huebner. Economic functions of the stock market as a business agency and the proposed national securities exchange act of 1934. Proceedings of the American Philosophical Society, 74(1):77–92, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984585.

Humphreys:1936:BL

[Hum36] W. J. Humphreys. Ball lightning. Proceedings of the American Philosophical Society, 76(5):613–626, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984750.

Humphreys:1942:RPM

[Hum42] William J. Humphreys. A review of papers on meteorology and climatology published by the American Philosophical Society prior to the Twentieth Century. Proceedings of the American Philosophical Society, 86(1):29–33, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985075.

Huntington:1913:GHN

[Hun13] Ellsworth Huntington. Guatemala and the highest Native American civilization. Proceedings of the American Philosophical Society, 52(211):467–487, September/October 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983971.

Hunsaker:1917:EA

[Hun17] Jerome C. Hunsaker. Engineering aspects. Proceedings of the American Philosophical Society, 56(3):249–254, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983948.

Hunter:1932:AMA

[Hun32] Robert J. Hunter. The activities of members of the American Philosophical Society in the early history of the Philadelphia Almshouse (The Philadelphia General Hospital). Proceedings of the American Philosophical Society, 71(6):308–319, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984643.

Hunter:1944:HFS

[Hun44] Merlin H. Hunter. The harmonization of fiscal and social aims. Proceedings of the American Philosophical Society, 88(1):43–49, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985188.

Hunsaker:1947:ISC

[Hun47] Jerome C. Hunsaker. International scientific congresses. Proceedings of the American Philosophical Society, 91(1):126–128, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143134.

Hyde:1935:AH

[Hyd35] Walter Woodburn Hyde. The Alps in history. Proceedings of the American Philosophical Society, 75(6):431–442, 1935. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984608.

Iddings:1911:PP

[Idd11] Joseph P. Iddings. Problems in petrology. Proceedings of the American Philosophical Society, 50(199):286–300, May/August 1911. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984045.

Innis:1944:PEM

[Inn44] Harold A. Innis. Political economy in the modern state. Proceedings of the American Philosophical Society, 87(4):323–341, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985045.

Jones:1908:ASN

[JA08] Harry C. Jones and John A. Anderson. The absorption spectra of neodymium chloride and praseodymium chloride in water, methyl alcohol, ethyl alcohol and mixtures of these solvents. Twenty-first communication. Proceedings of the American Philosophical Society, 47(189):276–297, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983844.

Jackson:1929:TME

[Jac29] A. V. Williams Jackson. The tomb of the Moghul Emperor Bābur in Afghanistan. Proceedings of the American Philosophical Society, 68 (3):195–205, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984208.

Jacobs:1931:DPN

[Jac31a] M. H. Jacobs. Diffusion processes in non-living and living systems. Proceedings of the American Philosophical Society, 70(2):167–186, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984320.

Jacobs:1931:OHZ

[Jac31b] M. H. Jacobs. Osmotic hemolysis and zoölogical classification. Proceedings of the American Philosophical Society, 70(4):363–370, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984201.

Jackson:1942:EOE

[Jac42] Dugald C. Jackson. Engineering in our early history. (The American Philosophical Society and engineering from 1768 to 1876). Proceedings of the American Philosophical Society, 86(1):45–51, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985077.

Jastrow:1904:HST

[Jas04] Morris Jastrow, Jr. The Hamites and Semites in the tenth chapter of Genesis. Proceedings of the American Philosophical Society, 43(176):173–207, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983504.

Jastrow:1908:HAB

[Jas08a] Morris Jastrow, Jr. Hepatoscopy and astrology in Babylonia and Assyria. Proceedings of the American Philosophical Society, 47(190): 646–676, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983836.

Jastrow:1908:SNP

[Jas08b] Morris Jastrow, Jr. The sign and name for planet in Babylonian. Proceedings of the American Philosophical Society, 47(189): 141–156, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983842.

Jastrow:1916:SGM

[Jas16] Morris Jastrow, Jr. Sir Gaston Maspero. Proceedings of the American Philosophical Society, 55(8):III–XIII, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984012.

Jastrow:1921:JGR

[Jas21] Morris Jastrow, Jr. Joseph George Rosengarten. Proceedings of the American Philosophical Society, 60(4):iii–ix, 1921. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984521.

Jaszi:1949:DON

[Jás49] Oscar Jászi. Danubia: Old and new. Proceedings of the American Philosophical Society, 93(1):1–31, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143267.

Jenkins:1913:GRA

[JB13] Olaf Pitt Jenkins and J. C. Branner. Geology of the region about Natal, Rio Grande Do Norte, Brazil. Proceedings of the American Philosophical Society, 52(211):431–466, September/October 1913. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983970.

Jennings:1908:HVE

[Jen08] H. S. Jennings. Heredity, variation and evolution in Protozoa. II. Heredity and variation of size and form in Paramecium, with studies of growth, environmental action and selection. Proceedings of the American Philosophical Society, 47(190):393–546, September/ December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983831.

Jensen:1936:ERS

[Jen36] J. C. Jensen. Evaporation and rainfall studies in the Northwest Minnesota Lake Region. Proceedings of the American Philosophical Society, 76(6):747–759, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984564.

Jensen:1938:ERS

[Jen38] J. C. Jensen. Evaporation and rainfall studies in the Northwest Minnesota Lake Region. Proceedings of the American Philosophical Society, 78(4):651–670, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984807.

Jennings:1941:GPB

[Jen41] H. S. Jennings. Genetics of Paramecium Bursaria. II Selfdifferentiation and self-fertilization of clones. Proceedings of the American Philosophical Society, 85(1):25–48, November 1941. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985124.

Jenkins:1948:FRF

[Jen48] Charles F. Jenkins. Franklin returns from France — 1785. Proceedings of the American Philosophical Society, 92(6):417–432, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143197.

Jepsen:1930:SPP

[Jep30a] G. L. Jepsen. Stratigraphy and paleontology of the Paleocene of Northeastern Park County, Wyoming. Proceedings of the American Philosophical Society, 69(1):463–528, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984379.

Jepsen:1930:NVF

[Jep30b] Glenn L. Jepsen. New vertebrate fossils from the Lower Eocene of the Bighorn Basin, Wyoming. Proceedings of the American Philosophical Society, 69(1):117–131, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984351.

Jepsen:1932:TTW

[Jep32] G. L. Jepsen. Tubulodon Taylori, a Wind River Eocene tubulidentate from Wyoming. Proceedings of the American Philosophical Society, 71(5):255–274, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984701.

Jepsen:1933:AES

[Jep33] Glenn L. Jepsen. American Eusmiloid sabre-tooth cats of the Oligocene Epoch. Proceedings of the American Philosophical Society, 72(5):355–369, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984699.

Jepsen:1934:RAA

[Jep34] Glenn L. Jepsen. A revision of the American Apatemyidæ and the description of a new genus, “Sinclairella”, from the White River Oligocene of South Dakota. Proceedings of the American Philosophical Society, 74(4):287–305, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984796.

Jepsen:1937:PRP

[Jep37] Glenn L. Jepsen. A Paleocene rodent, “Paramys Atavus”. Proceedings of the American Philosophical Society, 78(2):291–301, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984538.

Jepsen:1940:PFP

[Jep40] Glenn L. Jepsen. Paleocene faunas of the Polecat Bench Formation, Park County, Wyoming: Part I. Proceedings of the American Philosophical Society, 83(2):217–340, August 15, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985014.

Jepsen:1949:SOF

[Jep49] Glenn L. Jepsen. Selection, “orthogenesis,” and the fossil record. Proceedings of the American Philosophical Society, 93(6):479–500, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143337.

Jewett:1941:ADB

[Jew41] Frank B. Jewett. Alexander Dallas Bache, a founder, First President and benefactor of the National Academy of Sciences. Proceedings of the American Philosophical Society, 84(2):181–186, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984917.

Joerg:1940:DPP

[Joe40] W. L. G. Joerg. Demonstration of the peninsularity of Palmer Land, Antarctica, through Ellsworth’s flight of 1935. Proceedings of the American Philosophical Society, 82(5):821–832, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984893.

Johnson:1920:GAA

[Joh20] Douglas Johnson. Geographic aspects of the Adriatic problem. Proceedings of the American Philosophical Society, 59(6):512–516, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984390.

Johnson:1924:NTP

[Joh24] Emory R. Johnson. The Nation’s transportation problem. Proceedings of the American Philosophical Society, 63(1):18–24, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984439.

Johnson:1925:NSR

[Joh25] Emory R. Johnson. Next steps in railroad consolidation. Proceedings of the American Philosophical Society, 64(1):78–82, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984305.

Johnson:1928:CBM

[Joh28] Emory R. Johnson. Can business be made a science? Proceedings of the American Philosophical Society, 67(1):95–103, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984258.

Johnson:1929:MSS

[Joh29a] Douglas Johnson. Mean sealevel studies in New York waters. Proceedings of the American Philosophical Society, 68(2):93–94, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984309.

Johnson:1929:SPR

[Joh29b] Emory R. Johnson. Some problems of railroad consolidation. Proceedings of the American Philosophical Society, 68(2):119–128, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984312.

Johnson:1930:SPR

[Joh30] Emory R. Johnson. Some phases of railroad consolidation. Proceedings of the American Philosophical Society, 69(1):369–377, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984370.

Johnson:1932:WRK

[Joh32a] Alba B. Johnson. A way for the railways to keep out after they are out. Proceedings of the American Philosophical Society, 71(1): 23–29, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984817.

Johnson:1932:RSS

[Joh32b] Emory R. Johnson. The railroad situation: Some suggestions as to the way out. Proceedings of the American Philosophical Society, 71 (1):1–22, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984816.

Johnson:1934:PGO

[Joh34] Emery R. Johnson. Problems of government ownership and operation of the railroads. Proceedings of the American Philosophical Society, 74(1):103–128, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984587.

Johnson:1936:WWA

[Joh36] Emory R. Johnson. William Wallace Atterbury. Proceedings of the American Philosophical Society, 76(2):246–247, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984671.

Johnson:1938:PCT

[Joh38] Emory R. Johnson. The Panama Canal tolls problem and its solution. Proceedings of the American Philosophical Society, 78(4):671– 674, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984808.

Johnson:1939:PAM

[Joh39] E. R. Fenimore Johnson. The preservation and abuses of motion picture film by scientific institutions. Proceedings of the American Philosophical Society, 81(1):79–92, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984966.

Johnston:1941:MWS

[Joh41] H. F. Johnston. Magnetic work at sea. Proceedings of the American Philosophical Society, 84(2):257–262, 1941. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984922.

Johnson:1943:PCR

[Joh43] Emory R. Johnson. Panama Canal revenues and finances. Proceedings of the American Philosophical Society, 87(2):175–188, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985095.

Jones:1911:JHV

[Jon11] Harry C. Jones. Jacobus Henricus Van’t Hoff. Proceedings of the American Philosophical Society, 50(202):iii–xii, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984004.

Jones:1937:SSR

[Jon37] Donald F. Jones. Somatic segregation in relation to atypical growth. Proceedings of the American Philosophical Society, 77(3):411–416, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984735.

Jones:1942:OCI

[Jon42] Howard Mumford Jones. Origins of the colonial idea in england. Proceedings of the American Philosophical Society, 85(5):448–465, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985065.

Jones:1946:CIA

[Jon46] Howard Mumford Jones. The colonial impulse: An analysis of the “promotion” literature of colonization. Proceedings of the American Philosophical Society, 90(2):131–161, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301053.

Jordan:1906:HH

[Jor06] David Starr Jordan. The human harvest. Proceedings of the American Philosophical Society, 45(182):54–69, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983679.

Jones:1940:AVM

[JR40] Creighton C. Jones and Arthur Ruark. Apparatus for viewing and measuring stereoscopic cloud chamber photographs. Proceedings of the American Philosophical Society, 82(3):253–274, April 1940. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984929.

Jones:1909:ASV

[JS09] Harry C. Jones and W. W. Strong. The absorption spectra of various potassium, uranyl, uranous and neodymium salts in solution and the effect of temperature on the absorption spectra of certain colored salts in solution. Proceedings of the American Philosophical Society, 48(192):194–234, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984156.

Johnson:1927:SLS

[JW27] Douglas Johnson and Elizabeth Winter. Sea-level surfaces and the problem of coastal subsidence. Proceedings of the American Philosophical Society, 66(??):465–496, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301079.

Kaempffert:1931:CWP

[Kae31] Waldemar Kaempffert. Communication and world peace. Proceedings of the American Philosophical Society, 70(3):273–284, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984456.

Kaempffert:1939:TAR

[Kae39] Waldemar B. Kaempffert. Technological advance in relation to population trends. Proceedings of the American Philosophical Society, 80(4):559–568, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985057.

Kaylor:1945:PCC

[Kay45] Joseph F. Kaylor. Public control of cutting practices on private timberlands. Proceedings of the American Philosophical Society, 89(2):428–433, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985639.

Korff:1945:RCR

[KBC45] S. A. Korff, Dana K. Bailey, and Eric T. Clarke. Report on cosmicray observations made on the United States Antarctic Service Expedition, 1939–1941. Proceedings of the American Philosophical Society, 89(1):316–323, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985342.

Keasbey:1902:CE

[Kea02] Lindley M. Keasbey. A classification of economies. Proceedings of the American Philosophical Society, 41(169):146–153, April 1902. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983533.

Keen:1911:HCL

[Kee11] William W. Keen. Henry Charles Lea. Proceedings of the American Philosophical Society, 50(202):iii–iv, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984002.

Keedy:1934:PEA

[Kee34] Edwin R. Keedy. Preliminary examination of accused persons in england. Proceedings of the American Philosophical Society, 73(2): 103–116, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984759.

Keefer:1944:PSP

[Kee44] Chester S. Keefer. The present status of penicillin in the treatment of infections. Proceedings of the American Philosophical Society, 88(3):174–176, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985410.

Keefer:1948:STI

[Kee48] Chester S. Keefer. Streptomycin in the treatment of infections. Proceedings of the American Philosophical Society, 92(1):51–54, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143627.

Keller:1908:NSC

[Kel08] Harry F. Keller. Notes on some Chilean copper minerals. Proceedings of the American Philosophical Society, 47(188):79–85, January/ April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983800.

Keller:1909:NVC

[Kel09] Harry F. Keller. On a new variety of Chrysocolla from Chile. Proceedings of the American Philosophical Society, 48(191):65–66, January/ April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983811.

Kemmerer:1932:GGS

[Kem32] Edwin Walter Kemmerer. Gold and the gold standard. Proceedings of the American Philosophical Society, 71(3):85–104, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984555.

Kennelly:1909:LRB

[Ken09] A. E. Kennelly. The linear resistance between parallel conducting cylinders in a medium of uniform conductivity. Proceedings of the American Philosophical Society, 48(192):142–165, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984152.

Kennelly:1924:SPS

[Ken24a] A. E. Kennelly. Some properties of simple electric conducting networks. Proceedings of the American Philosophical Society, 63(2):171– 189, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984472.

Kent:1924:SIL

[Ken24b] Roland G. Kent. The scientist and an international language. Proceedings of the American Philosophical Society, 63(1):162–170, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984447.

Kennelly:1925:IEN

[Ken25] A. E. Kennelly. Identical electrical networks in series. Proceedings of the American Philosophical Society, 64(1):1–16, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984299.

Kennelly:1926:ESP

[Ken26] A. E. Kennelly. Equiangular spiral polygons as presenting themselves in electrical engineering. Proceedings of the American Philosophical Society, 65(4):326–348, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984240.

Kennelly:1928:MDC

[Ken28] A. E. Kennelly. A method for determining the constants of an electromagnetic oscillograph. Proceedings of the American Philosophical Society, 67(2):123–128, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984274.

Kennelly:1929:GLT

[Ken29] Arthur E. Kennelly. Gudermannians and Lambertians with their respective addition theorems. Proceedings of the American Philosophical Society, 68(3):175–184, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984206.

Kennelly:1930:PCI

[Ken30] A. E. Kennelly. Parallel connections of identical nets. Proceedings of the American Philosophical Society, 69(1):161–173, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984353.

Kennelly:1931:RVU

[Ken31] Arthur E. Kennelly. Rationalised versus unrationalised practical electromagnetic units. Proceedings of the American Philosophical Society, 70(2):103–119, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984317.

Kennelly:1933:MDC

[Ken33] Arthur E. Kennelly. Magnetic definitions from the circuital standpoint. Proceedings of the American Philosophical Society, 72(1):39– 55, January 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984607.

Kennelly:1935:MIP

[Ken35a] Arthur E. Kennelly. Michael Idvorsky Pupin. Proceedings of the American Philosophical Society, 75(4):335–338, 1935. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984636.

Kennelly:1935:PRG

[Ken35b] Arthur E. Kennelly. Proposed reforms of the Gregorian calendar. Proceedings of the American Philosophical Society, 75(1):71–110, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984723.

Kennelly:1936:MFE

[Ken36] A. E. Kennelly. Magnetic formulæ expressed in the M.K.S. system of units. Proceedings of the American Philosophical Society, 76(3):343– 377, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984549.

Keniston:1948:CRI

[Ken48] Hayward Keniston. Cultural relations and international understanding. Proceedings of the American Philosophical Society, 92(1):37–40, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143623.

Keppel:1937:REP

[Kep37] Frederick P. Keppel. The responsibility of endowments in the promotion of knowledge. Proceedings of the American Philosophical Society, 77(4):591–603, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984779.

Kofoid:1938:NSA

[KH38] Charles A. Kofoid and Datus M. Hammond. The neuromotor system of Anoplophrya Lumbrici, a degenerate ciliate, from the earthworm Octolasium Cyaneum. Proceedings of the American Philosophical Society, 78(4):595–614, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984805.

Kidder:1940:LB

[Kid40] Alfred V. Kidder. Looking backward. Proceedings of the American Philosophical Society, 83(4):527–537, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984833.

Kidder:1948:AAA

[Kid48] Alfred V. Kidder. American anthropology and archaeology. Proceedings of the American Philosophical Society, 92(2):126, May 1948.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143413.

Kienholz:1926:EAS

[Kie26] Raymond Kienholz. An ecological–anatomical study of beach vegetation in the Philippines. Proceedings of the American Philosophical Society, 65(5):58–100, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984288.

Kimball:1943:JA

[Kim43] Fiske Kimball. Jefferson and the arts. Proceedings of the American Philosophical Society, 87(3):238–245, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984871.

Kimball:1944:LPJ

[Kim44] Fiske Kimball. The life portraits of Jefferson and their replicas. Proceedings of the American Philosophical Society, 88(6):497–534, December 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985240.

Kiser:1939:VIA

[Kis39] Clyde V. Kiser. Voluntary and involuntary aspects of childlessness. Proceedings of the American Philosophical Society, 80(4):525–539, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985055.

Keller:1904:DRA

[KM04] Harry F. Keller and Philip Maas. Dimethyl racemic acid: Its synthesis and derivatives. Proceedings of the American Philosophical Society, 43(176):105–112, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983494.

Kennelly:1920:TPE

[KN20] A. E. Kennelly and U. Nabeshima. The transient process of establishing a steadily alternating current on a long line, from laboratory measurements on an artificial line. Proceedings of the American Philosophical Society, 59(5):325–370, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984173.

Kneeland:1937:FVU

[Kne37] Yale Kneeland, Jr. Filterable viruses in upper respiratory infection. Proceedings of the American Philosophical Society, 77(4):467–471, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984773.

Knowles:1940:TCI

[Kno40] Nathaniel Knowles. The torture of captives by the Indians of Eastern North America. Proceedings of the American Philosophical Society, 82(2):151–225, March 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985013.

Knowles:1945:GSPa

[Kno45a] Paul H. Knowles. Geology of Southern Palmer Peninsula, Antarctica. Proceedings of the American Philosophical Society, 89(1):132– 145, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985318.

Knowles:1945:GSPb

[Kno45b] Paul H. Knowles. Glaciology of Southern Palmer Peninsula, Antarctica. Proceedings of the American Philosophical Society, 89(1):174– 176, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985323.

Koenig:1903:APC

[Koe03] George A. Koenig. On artificial production of crystallized domeykite, algodonite, argentodomeykite and stibiodomeykite. Proceedings of the American Philosophical Society, 42(173):219–237, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983735.

Kohn:1940:TPW

[Koh40] Hans Kohn. The totalitarian philosophy of war. Proceedings of the American Philosophical Society, 82(1):57–72, February 1940. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984904.

Kohler:1941:NA

[Köh41] Wolfgang Köhler. On the nature of associations. Proceedings of the American Philosophical Society, 84(4):489–502, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984953.

Koht:1943:SNP

[Koh43] Halvdan Koht. The small nations in the post-war world. Proceedings of the American Philosophical Society, 87(2):152–160, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985092.

Kopal:1941:DLD

[Kop41] Zdeněk Kopal. Determination of limb darkening in eclipsing binaries from color-index observations. Proceedings of the American Philosophical Society, 84(5):565–587, July 30, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984840.

Kopal:1942:TLC

[Kop42] Zdeněk Kopal. Theoretical light curves of close eclipsing systems. Proceedings of the American Philosophical Society, 85(4):399–431, June 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984992.

Kopal:1943:AML

[Kop43a] Zdeněk Kopal. An application of the method of least squares to the adjustment of photometric elements of eclipsing binaries. Proceedings of the American Philosophical Society, 86(3):342–350, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985024.

Kopal:1943:ERU

[Kop43b] Zdeněk Kopal. The effects of reflection upon light and velocity curves of close binary systems. Proceedings of the American Philosophical Society, 86(3):351–363, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985025.

Kopal:1944:STF

[Kop44] Zdeněk Kopal. Some tables to facilitate computation of elements of eclipsing binaries. Proceedings of the American Philosophical Society, 88(2):145–149, July 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985668.

Kopal:1945:OTA

[Kop45a] Zdeněk Kopal. An outline of the theory of atmospheric eclipses. Proceedings of the American Philosophical Society, 89(4):590–600, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985363.

Kopal:1945:TVC

[Kop45b] Zdeněk Kopal. Theoretical velocity curves of close binary systems. Proceedings of the American Philosophical Society, 89(3):517–530, October 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985432.

Korff:1941:PNC

[Kor41] S. A. Korff. The production of neutrons by the cosmic radiation. Proceedings of the American Philosophical Society, 84(5):589–603, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984841.

Kotok:1945:ARM

[Kot45] E. I. Kotok. America’s role in meeting world timber needs. Proceedings of the American Philosophical Society, 89(2):418–422, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985637.

Kofoid:1940:NSO

[KR40] Charles A. Kofoid and Lauren E. Rosenberg. The neuromotor system of “Opisthonecta Henneguyi” (Fauré–Fremiet). Proceedings of the American Philosophical Society, 82(4):421–436, May 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984879.

Kraemer:1902:CP

[Kra02] Henry Kraemer. On the continuity of protoplasm. Proceedings of the American Philosophical Society, 41(169):174–180, April 1902. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983536.

Kraemer:1904:ONC

[Kra04] Henry Kraemer. The origin and nature of color in plants. Proceedings of the American Philosophical Society, 43(177):257–277, April/ September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983765.

Kraemer:1905:OAC

[Kra05] Henry Kraemer. The oligodynamic action of copper foil on certain intestinal organisms. Proceedings of the American Philosophical Society, 44(179):51–65, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983605.

Kraemer:1906:DSA

[Kra06] Henry Kraemer. Dilute sulphuric acid as a fungicide. Proceedings of the American Philosophical Society, 45(183):157–163, May/ September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983825.

Kramm:1910:SCC

[Kra10] H. E. Kramm. Serpentines of the Central Coast Ranges of California. Proceedings of the American Philosophical Society, 49(196): 315–349, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984124.

Krautheimer:1941:PVT

[Kra41] Richard Krautheimer. S. Pietro in Vincoli and the tripartite transept in the Early Christian Basilica. Proceedings of the American Philosophical Society, 84(3):353–429, May 31, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984958.

Kramer:1942:SLP

[Kra42] S. N. Kramer. Sumerian literature; a preliminary survey of the oldest literature in the world. Proceedings of the American Philosophical Society, 85(3):293–323, February 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985008.

Kramer:1946:HSN

[Kra46] Samuel Noah Kramer. Heroes of Sumer: A new heroic age in world history and literature. Proceedings of the American Philosophical Society, 90(2):120–130, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301052.

Krause:1947:HAR

[Kra47] Ernst H. Krause. High altitude research with V-2 rockets. Proceedings of the American Philosophical Society, 91(5):430–446, December 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143424.

Kollock:1905:URA

[KS05] Lily G. Kollock and Edgar F. Smith. The use of the rotating anode and mercury cathode in electro-analysis. Proceedings of the American Philosophical Society, 44(180):137–150, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983510.

Kollock:1906:ESA

[KS06] Lily G. Kollock and Edgar F. Smith. The effect of sulphuric acid on the deposition of metals when using a mercury cathode and rotating anode. Proceedings of the American Philosophical Society, 45(184): 255–267, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983378.

Kollock:1907:NRE

[KS07] Lily G. Kollock and Edgar F. Smith. New results in electrolysis. Proceedings of the American Philosophical Society, 46(187): 341–345, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983473.

Kennelly:1914:IAP

[KS14] A. E. Kennelly and H. S. Sanborn. The influence of atmospheric pressure upon the forced thermal convection from small electrically heated platinum wires. Proceedings of the American Philosophical Society, 53(213):55–77, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984130.

Kunz:1919:PEP

[KS19] Jakob Kunz and Joel Stebbins. Photo-electric photometry of the 1918 eclipse. Proceedings of the American Philosophical Society, 58(4):269– 271, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984404.

Kennelly:1927:AIS

[KS27] A. E. Kennelly and R. P. Siskind. The acoustic impedance of straight cylindrical air columns in brass tubes. Proceedings of the American Philosophical Society, 66(??):89–133, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301060.

Kennelly:1915:EVS

[KT15] A. E. Kennelly and H. O. Taylor. Explorations over the vibrating surfaces of telephonic diaphragms under simple impressed tones. Proceedings of the American Philosophical Society, 54(217):96–136, May/ July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983917.

Kennelly:1916:SPV

[KT16] A. E. Kennelly and H. O. Taylor. Some properties of vibrating telephone diaphragms. Proceedings of the American Philosophical Society, 55(6):415–460, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983979.

Kennelly:1908:HTC

[KU08] A. E. Kennelly and Walter L. Upson. The humming telephone, a contribution to the theoretical and practical analysis of its behavior. Proceedings of the American Philosophical Society, 47(189): 329–365, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983847.

Kunkel:1943:PWB

[Kun43] L. O. Kunkel. Potato witches’-broom transmission by dodder and cure by heat. Proceedings of the American Philosophical Society, 86(3):470–475, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985033.

Kurath:1934:LAN

[Kur34] Hans Kurath. The linguistic atlas of new england. Proceedings of the American Philosophical Society, 74(3):227–243, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984716.

Kuznets:1944:TNI

[Kuz44] Simon Kuznets. Taxes and national income. Proceedings of the American Philosophical Society, 88(1):10–21, June 1944. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985184.

Kuznets:1948:FER

[Kuz48] Simon Kuznets. Foreign economic relations of the United States and their impact upon the domestic economy. A review of long term trends. Proceedings of the American Philosophical Society, 92(4):228–243, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143351.

Kennelly:1919:ACP

[KV19] A. E. Kennelly and Edy Velander. Alternating-current planevector potentiometer measurements at telephonic frequencies. Proceedings of the American Philosophical Society, 58(2):97–132, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984529.

Keen:1926:AHC

[KW26] William W. Keen and Woodrow Wilson. Address by Hampton L. Carson in commemoration of I. Minis Hays. Proceedings of the American Philosophical Society, 65(1):iii–xxxii, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984333.

Kohler:1944:FAE

[KW44] Wolfgang Köhler and Hans Wallach. Figural after-effects. An investigation of visual processes. Proceedings of the American Philosophical Society, 88(4):269–357, October 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985451.

Kyte:1949:RLA

[Kyt49] George W. Kyte. Robert Liston and Anglo–American cooperation, 1796–1800. Proceedings of the American Philosophical Society, 93(3):259–266, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143476.

Lambert:1903:NAM

[Lam03] P. A. Lambert. New applications of Maclaurin’s series in the solution of equations and in the expansion of functions. Proceedings of the American Philosophical Society, 42(172):85–95, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983644.

Lambert:1904:EAF

[Lam04] Preston A. Lambert. Expansions of algebraic functions at singular points. Proceedings of the American Philosophical Society, 43(176):164–172, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983503.

Lambert:1905:SLC

[Lam05] P. A. Lambert. The straight line concept. Proceedings of the American Philosophical Society, 44(179):82–89, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983609.

Lamberton:1907:NWS

[Lam07] William A. Lamberton. The narratives of the walking on the sea. Proceedings of the American Philosophical Society, 46(185):80–86, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983441.

Lambert:1908:SAE

[Lam08] Preston A. Lambert. Solution of algebraic equations in infinite series. Proceedings of the American Philosophical Society, 47(188): 111–135, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983804.

Lambert:1911:SLD

[Lam11] Preston A. Lambert. On the solution of linear differential equations of successive approximations. Proceedings of the American Philosophical Society, 50(199):274–285, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984044.

Lampland:1919:LOE

[Lam19] C. O. Lampland. The Lowell Observatory eclipse observations, June 8, 1918. Prominences and coronal arches. Proceedings of the American Philosophical Society, 58(4):259–264, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984402.

Landreth:1906:CPV

[Lan06a] Burnet Landreth. A case of persistent vitality in seeds. Proceedings of the American Philosophical Society, 45(182):5–8, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983671.

Landreth:1906:NA

[Lan06b] Burnet Landreth. The new agriculture. Proceedings of the American Philosophical Society, 45(183):166–178, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983827.

Landis:1941:PSM

[Lan41] Carney Landis. Psychoanalysis and scientific method. Proceedings of the American Philosophical Society, 84(4):515–525, June 1941. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984955.

Langmuir:1946:WCA

[Lan46] Irving Langmuir. World control of atomic energy. Proceedings of the American Philosophical Society, 90(1):65–69, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301041.

Lane:1948:SSH

[Lan48a] Frederic C. Lane. The social sciences and the humanities. Proceedings of the American Philosophical Society, 92(5):356–362, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143046.

Langer:1948:SIP

[Lan48b] William L. Langer. Scholarship and the intelligence problem. Proceedings of the American Philosophical Society, 92(1):43–45, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143625.

Langmuir:1948:GPS

[Lan48c] Irving Langmuir. The growth of particles in smokes and clouds and the production of snow from supercooled clouds. Proceedings of the American Philosophical Society, 92(3):167–185, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143585.

Lashley:1941:CNP

[Las41] Karl S. Lashley. Coalescence of neurology and psychology. Proceedings of the American Philosophical Society, 84(4):461–470, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984951.

Lazarsfeld:1948:UPS

[Laz48] Paul F. Lazarsfeld. The use of panels in social research. Proceedings of the American Philosophical Society, 92(5):405–410, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143053.

Lane:1938:RTD

[LBD^{+}38] Alfred C. Lane, William Bowie, Roy E. Dickerson, E. A. Eckhardt, C. L. Garner, W. H. Hobbs, W. J. Humphreys, Chester R. Longwell, H. N. Russell, Paul S. Smith, W. T. Thom, Jr., and T. W. Vaughan. Round table discussion. Proceedings of the American Philosophical Society, 79(1):127–144, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984683.

LecomteDuNouy:1944:PT

[Lec44] P. Lecomte Du Noüy. Physiological time. Proceedings of the American Philosophical Society, 87(5):435–437, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985297.

Lee:1940:PTN

[Lee40] Hugh J. Lee. Peary’s transections of North Greenland, 1892–1895. Proceedings of the American Philosophical Society, 82(5):921–934, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984899.

Leith:1947:MRT

[Lei47] Charles Kenneth Leith. Mineral resources in their international relations. Proceedings of the American Philosophical Society, 91(1):83–87, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143128.

Leland:1932:ACL

[Lel32] Waldo G. Leland. The American Council of Learned Societies and its relation to humanistic studies. Proceedings of the American Philosophical Society, 71(4):179–189, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984789.

Leland:1937:MPR

[Lel37] Waldo G. Leland. Methods of promoting research, from the point of view of societies, academies, and councils. Proceedings of the American Philosophical Society, 77(4):611–616, April 1937. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984781.

Leland:1946:BAU

[Lel46] Waldo G. Leland. The background and antecedents of UNESCO. Proceedings of the American Philosophical Society, 90(4):295–299, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301024.

Lenher:1926:QDS

[Len26] Victor Lenher. The quantitative determination of selenium and tellurium. Proceedings of the American Philosophical Society, 65(1): 33–50, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984336.

Leverett:1926:PGS

[Lev26] Frank Leverett. The Pleistocene glacial stages: Were there more than four? Proceedings of the American Philosophical Society, 65(2):105– 118, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984418.

Levi:1944:NNS

[Lev44] Doro Levi. The novel of Ninus and Semiramis. Proceedings of the American Philosophical Society, 87(5):420–428, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985295.

Lewis:1942:IBR

[Lew42] William Draper Lewis. An International Bill of Rights. Proceedings of the American Philosophical Society, 85(5):445–447, September 1942.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985064.

Lewis:1943:LHS

[Lew43] T. M. N. Lewis. Late horizons in the Southeast. Proceedings of the American Philosophical Society, 86(2):304–312, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985108.

Lewis:1945:SEH

[Lew45] William Draper Lewis. The statement of essential human rights by representatives of the principal cultures of the world. Proceedings of the American Philosophical Society, 89(3):489–494, October 11, 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985429.

Lark-Horovitz:1936:IDE

[LHYH36] K. Lark-Horovitz, H. J. Yearian, and J. D. Howe. Intensity distribution in electron diffraction patterns. Proceedings of the American Philosophical Society, 76(6):766–770, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984566.

Likert:1948:OSG

[Lik48] Rensis Likert. Opinion studies and government policy. Proceedings of the American Philosophical Society, 92(5):341–350, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143044.

Lindsay:1918:SSI

[Lin18] Samuel McCune Lindsay. Soldiers’ and sailors’ insurance. Proceedings of the American Philosophical Society, 57(6):632–648, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984467.

Lingelbach:1927:DCF

[Lin27] William E. Lingelbach. Democracy and the control of foreign affairs. Proceedings of the American Philosophical Society, 66(??):655–679, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301087.

Lindgren:1933:MRA

[Lin33] Waldemar Lindgren. Memorial of Richard Alexander Fullerton Penrose, Jr. December 17, 1863–July 31, 1931. A tribute to his life and achievements. Proceedings of the American Philosophical Society, 72 (3):101–114, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984637.

Lingelbach:1938:NVA

[Lin38] William E. Lingelbach. Neutrality versus alliances. Belgium and the revolution in international politics. Proceedings of the American Philosophical Society, 79(4):607–636, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984942.

Lincoln:1940:GAT

[Lin40] Elsie Lincoln. Growth in Aeshna Tuberculifera (Odonata). Proceedings of the American Philosophical Society, 83(5):589–605, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985130.

Lingelbach:1948:BFP

[Lin48] William E. Lingelbach. B. Franklin, Printer — new source materials. Proceedings of the American Philosophical Society, 92(2):79–100, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143405.

Lingelbach:1949:OPR

[Lin49] William E. Lingelbach. Old Philadelphia: Redevelopment and conservation. Proceedings of the American Philosophical Society, 93(2):179– 207, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143438.

Livingston:1913:CAU

[Liv13] Burton Edward Livingston. Climatic areas of the United States as related to plant growth. Proceedings of the American Philosophical Society, 52(209):257–275, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983871.

Livingston:1940:SMF

[LL40] Burton E. Livingston and Stuart B. Le Compte, Jr. Soil-moisture fluctuation under a lawn, as indicated by absorption from porousporcelain irrigator cones with continuous operation. Proceedings of the American Philosophical Society, 82(3):227–251, April 30, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984928.

Lockhart:1945:ATD

[Loc45] Ernest E. Lockhart. Antarctic trail diet. Proceedings of the American Philosophical Society, 89(1):235–248, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985326.

Loeb:1908:TGT

[Loe08] Leo Loeb. Tumor growth and tissue growth. Proceedings of the American Philosophical Society, 47(188):3–12, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983789.

Loeb:1911:CCM

[Loe11] Leo Loeb. The cyclic changes in the mammalian ovary. Proceedings of the American Philosophical Society, 50(199):228–234, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984037.

Loewy:1949:TPS

[Loe49] Ariel G. Loewy. A theory of protoplasmic streaming. Proceedings of the American Philosophical Society, 93(4):326–329, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143158.

Long:1948:DTC

[Lon48] Esmond R. Long. The decline of tuberculosis as the chief cause of death. Proceedings of the American Philosophical Society, 92(3):139– 143, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143582.

Lorentz:1906:PNE

[Lor06] H. A. Lorentz. On positive and negative electrons. Proceedings of the American Philosophical Society, 45(183):103–109, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983819.

Lotka:1939:CPP

[Lot39] Alfred J. Lotka. Contact points of population study with related branches of science. Proceedings of the American Philosophical Society, 80(4):601–626, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985060.

Loudon:1943:STP

[Lou43] Alexander Loudon. Some thoughts on post-war planning. Proceedings of the American Philosophical Society, 87(2):139–146, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985090.

Lovett:1909:CGP

[Lov09] Edgar Odell Lovett. On certain generalizations of the problem of three bodies. Proceedings of the American Philosophical Society, 48 (191):111–118, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983816.

Lovett:1927:OU

[Lov27] Edgar Odell Lovett. Our universities. Proceedings of the American Philosophical Society, 66(??):723–727, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301090.

Lovejoy:1938:HI

[Lov38] Arthur O. Lovejoy. The historiography of ideas. Proceedings of the American Philosophical Society, 78(4):529–543, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984800.

Loveday:1943:ET

[Lov43] Alexander Loveday. The economics of transition. Proceedings of the American Philosophical Society, 87(2):189–193, August 1943. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985096.

Lowell:1900:MGE

[Low00] Percival Lowell. Mars on glacial epochs. Proceedings of the American Philosophical Society, 39(164):641–664, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983782.

Lowell:1901:ESS

[Low01] Percival Lowell. Explanation of the supposed signals from Mars of December 7 and 8, 1900. Proceedings of the American Philosophical Society, 40(167):166–176, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983762.

Lowell:1902:A

[Low02] Percival Lowell. Areography. Proceedings of the American Philosophical Society, 41(170):225–234, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983554.

Lowell:1903:CM

[Low03] Percival Lowell. The cartouches of Mars. Proceedings of the American Philosophical Society, 42(174):353–377, May/December 1903. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983719.

Lowell:1911:SPR

[Low11] Percival Lowell. Spectroscopic proof of the repulsion by the Sun of gaseous molecules in the tail of Halley’s Comet. Proceedings of the American Philosophical Society, 50(199):254–260, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984040.

Lowdermilk:1945:WWN

[Low45] W. C. Lowdermilk. World-wide needs of woods as a land conservation crop. Proceedings of the American Philosophical Society, 89(2):408– 417, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985636.

Luce:1918:BHS

[Luc18] Stephen Bleecker Luce. A brief history of the study of Greek vasepainting. Proceedings of the American Philosophical Society, 57(7): 649–668, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984526.

Luening:1940:M

[Lue40] Otto Luening. Music. Proceedings of the American Philosophical Society, 83(4):569–572, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984837.

Lundell:1933:ADM

[Lun33] Cyrus Longworth Lundell. Archeological discoveries in the Maya area. Proceedings of the American Philosophical Society, 72(3):147– 179, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984641.

Lutz:1926:FSB

[Lut26] Bertha Lutz. The flora of the Serra da Bocaina. Proceedings of the American Philosophical Society, 65(5):27–43, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984286.

Lutz:1927:PCS

[Lut27] Adolpho Lutz. Problems in the colonization and settlement of tropical South America. Proceedings of the American Philosophical Society, 66(??):135–157, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301061.

Lutz:1948:IFD

[Lut48] Friedrich A. Lutz. The influence of foreign developments on the American economy. Proceedings of the American Philosophical Society, 92(4):244–248, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143352.

Lydenberg:1948:AL

[Lyd48] Harry Miller Lydenberg. American linguistics. Proceedings of the American Philosophical Society, 92(2):124–126, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143412.

Lyman:1912:NJV

[Lym12] Benjamin Smith Lyman. The nature of the Japanese verb, so-called. Proceedings of the American Philosophical Society, 51(204):91–102, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984097.

Lyman:1915:PRA

[Lym15] Benjamin Smith Lyman. A practical rational alphabet. Proceedings of the American Philosophical Society, 54(220):359–370, October/ December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984109.

Lyman:1918:BN

[Lym18] Benjamin Smith Lyman. Brief notes. Proceedings of the American Philosophical Society, 57(6):627–631, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984466.

Mackay:1936:OTS

[MA36] Elizabeth Mackay and Charles E. Allen. The offspring of a triploid sporophyte. Proceedings of the American Philosophical Society, 76(6): 781–790, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984571.

Mabery:1903:RCO

[Mab03] Charles F. Mabery. A résumé of the composition and occurrence of petroleum. Proceedings of the American Philosophical Society, 42 (172):36–54, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983641.

Mackenzie:1902:SEP

[Mac02] A. Stanley Mackenzie. On some equations pertaining to the propagation of heat in an infinite medium. Proceedings of the American Philosophical Society, 41(169):181–199, April 1902. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983537.

MacDougal:1919:HG

[Mac19] D. T. MacDougal. Hydration and growth. Proceedings of the American Philosophical Society, 58(6):346–372, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984260.

MacDougal:1921:ABS

[Mac21a] D. T. MacDougal. The action of bases and salts on biocolloids and cell-masses. Proceedings of the American Philosophical Society, 60(1): 15–30, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984223.

MacDougal:1921:GT

[Mac21b] D. T. MacDougal. Growth in trees. Proceedings of the American Philosophical Society, 60(1):7–14, 1921. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984222.

MacDougal:1922:PAL

[Mac22] D. T. MacDougal. The probable action of lipoids in growth. Proceedings of the American Philosophical Society, 61(1):33–52, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984432.

MacDougal:1923:PIV

[Mac23] D. T. MacDougal. Permeability and the increase of volume of contents of living and of artificial cells. Proceedings of the American Philosophical Society, 62(2):1–25, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984507.

MacDougal:1924:AAM

[Mac24] D. T. MacDougal. The arrangement and action of material in the plasmatic layers and cell-walls of plants. Proceedings of the American Philosophical Society, 63(1):76–93, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984443.

MacDougal:1925:AEP

[Mac25] D. T. MacDougal. Absorption and exudation pressures of sap in plants. Proceedings of the American Philosophical Society, 64(2):102– 130, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984192.

MacDougal:1928:SRP

[Mac28] D. T. MacDougal. Substances regulating the passage of material into and out of plant cells: The lipoids. Proceedings of the American Philosophical Society, 67(1):33–45, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984254.

MacCurdy:1929:OWP

[Mac29] George Grant MacCurdy. Old-world prehistory in retrospect and prospect. Proceedings of the American Philosophical Society, 68(2): 95–106, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984310.

MacCurdy:1930:RPF

[Mac30a] George Grant MacCurdy. Recent progress in the field of old world prehistory. Proceedings of the American Philosophical Society, 69(1): 175–192, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984354.

MacDougal:1930:LGP

[Mac30b] Daniel T. MacDougal. Lengthened growth periods and continuous growth. Proceedings of the American Philosophical Society, 69(1): 329–345, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984367.

MacDougal:1932:PST

[Mac32] D. T. MacDougal. The pneumatic system of trees. Proceedings of the American Philosophical Society, 71(5):299–307, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984705.

MacCurdy:1933:PRN

[Mac33a] George Grant MacCurdy. Prehistoric research in the near east. Proceedings of the American Philosophical Society, 72(3):121–135, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984639.

MacDougal:1933:RLS

[Mac33b] D. T. MacDougal. The relation of leaf-surface to wood formation in pine trees. Proceedings of the American Philosophical Society, 72(5):325–332, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984696.

MacCurdy:1934:NLP

[Mac34] George Grant MacCurdy. New light on prehistoric Man in Asia. Proceedings of the American Philosophical Society, 74(3):185–191, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984713.

MacCurdy:1936:PMP

[Mac36a] George Grant MacCurdy. Prehistoric man in Palestine. Proceedings of the American Philosophical Society, 76(4):523–541, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984601.

Macdonald:1936:SSC

[Mac36b] Roderick Macdonald. A study of symmetry in the Centrechinoidea, based on behavior, with special reference to Lytechinus Variegatus; including a short discussion of linguistic difficulties in describing biological phenomena. Proceedings of the American Philosophical Society, 76(1):87–123, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984709.

MacDougal:1936:CPS

[Mac36c] D. T. MacDougal. The communication of the pneumatic systems of trees with the atmosphere. Proceedings of the American Philosophical Society, 76(6):823–845, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984573.

MacDougal:1939:FGC

[Mac39] D. T. MacDougal. Features of growth-control in trees. Proceedings of the American Philosophical Society, 81(3):421–445, August 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984863.

MacIver:1946:ICS

[Mac46] Robert Morrison MacIver. Intellectual cooperation in the social sciences. Proceedings of the American Philosophical Society, 90 (4):309–313, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301027.

Macelwane:1947:G

[Mac47] James B. Macelwane. Geophysics. Proceedings of the American Philosophical Society, 91(1):79–82, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143127.

Macmahon:1948:IPG

[Mac48] Arthur W. Macmahon. International policy and governmental structure (chiefly executive). Proceedings of the American Philosophical Society, 92(4):217–227, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143350.

Magie:1907:ATS

[Mag07] William Francis Magie. The association theory of solutions. Proceedings of the American Philosophical Society, 46(185):138–145, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983447.

Magie:1910:PNM

[Mag10] William Francis Magie. Physical notes on Meteor Crater, Arizona. Proceedings of the American Philosophical Society, 49(194): 41–48, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983989.

Magie:1912:TRS

[Mag12] William Francis Magie. Thermal relations of solutions. Proceedings of the American Philosophical Society, 51(205):235–239, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984140.

Magie:1937:HM

[Mag37] William F. Magie. Howard McClenahan. Proceedings of the American Philosophical Society, 77(2):219–221, February 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984728.

Magill:1949:NFT

[Mag49] Roswell Magill. The never finished tasks of federal tax revision. Proceedings of the American Philosophical Society, 93(4):283–289, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143154.

Mains:1934:GCO

[Mai34] Edwin B. Mains. The genera Cordyceps and Ophiocordyceps in Michigan. Proceedings of the American Philosophical Society, 74(4):263– 271, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984794.

Malone:1939:SP

[Mal39] Dumas Malone. The scholar and the public. Proceedings of the American Philosophical Society, 80(1):25–36, January 1939. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984970.

Malone:1949:TB

[Mal49] Kemp Malone. The text of “Beowulf”. Proceedings of the American Philosophical Society, 93(3):239–243, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143472.

Marshall:1905:SAP

[Mar05] Margaret E. Marshall. A study of the anatomy of Phalænoptilus, Ridgway. Proceedings of the American Philosophical Society, 44(181): 213–240, August/December 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983744.

Marvin:1917:CA

[Mar17] Charles Frederick Marvin. Cleveland Abbe, 1838–1916. Proceedings of the American Philosophical Society, 56(1):ix–xix, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983963.

Margolis:1928:TCG

[Mar28] Max L. Margolis. Textual criticism of the Greek Old Testament. Proceedings of the American Philosophical Society, 67(2):187–197, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984279.

Marx:1940:TP

[Mar40] Fritz Morstein Marx. Totalitarian politics. Proceedings of the American Philosophical Society, 82(1):1–38, February 1940. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984902.

Marquis:1948:SMH

[Mar48] Donald G. Marquis. Scientific methodology in human relations. Proceedings of the American Philosophical Society, 92(6):411–416, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143196.

Mason:1904:RTC

[Mas04] Otis T. Mason. The ripening of thoughts in common. Proceedings of the American Philosophical Society, 43(176):148–155, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983498.

Mason:1907:NMA

[Mas07] William Pitt Mason. Note as to the measurement of the action of water upon metals. Proceedings of the American Philosophical Society, 46(185):110–116, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983444.

Mason:1909:PWS

[Mas09] William Pitt Mason. The purification of water supplies by the use of hypochlorites. Proceedings of the American Philosophical Society, 48(191):67–68, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983812.

Mathews:1900:DSA

[Mat00a] R. H. Mathews. Divisions of the South Australian Aborigines. Proceedings of the American Philosophical Society, 39(161):78–91 + 93, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983545.

Mathews:1900:NTW

[Mat00b] R. H. Mathews. Native tribes of Western Australia. Proceedings of the American Philosophical Society, 39(161):123–125, January 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983551.

Mathews:1900:OOC

[Mat00c] R. H. Mathews. The origin, organization and ceremonies of the Australian Aborigines. Proceedings of the American Philosophical Society, 39(164):556–578, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983776.

Mathews:1900:PRI

[Mat00d] R. H. Mathews. Phallic rites and initiation ceremonies of the South Australian Aborigines. Proceedings of the American Philosophical Society, 39(164):622–638, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983778.

Mathews:1901:ARP

[Mat01a] R. H. Mathews. Aboriginal rock pictures in Queensland. Proceedings of the American Philosophical Society, 40(165):57–58, January 1901.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983429.

Mathews:1901:GL

[Mat01b] R. H. Mathews. The Gundungurra language. Proceedings of the American Philosophical Society, 40(167):140–148, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983755.

Mathews:1903:LNE

[Mat03a] R. H. Mathews. Languages of the New England Aborigines New South Wales. Proceedings of the American Philosophical Society, 42 (173):249–263, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983737.

Mathews:1903:SAL

[Mat03b] R. H. Mathews. Some Aboriginal languages of Queensland and Victoria. Proceedings of the American Philosophical Society, 42 (173):179–188, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983731.

Mathews:1904:NTV

[Mat04] R. H. Mathews. The native tribes of Victoria: Their languages and customs. Proceedings of the American Philosophical Society, 43(175):54–70, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983416.

Mathews:1905:SAW

[Mat05a] R. H. Mathews. Sociology of the Aborigines of Western Australia. Proceedings of the American Philosophical Society, 44(179): 32–35, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983600.

Matthew:1905:NOS

[Mat05b] W. D. Matthew. Notes on the osteology of Sinopa, a primitive member of the Hyænodontidæ. Proceedings of the American Philosophical Society, 44(179):69–72, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983607.

Mathews:1907:ALC

[Mat07a] R. H. Mathews. The Arran’da language, Central Australia. Proceedings of the American Philosophical Society, 46(187):322–339, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983471.

Mathews:1907:CDN

[Mat07b] R. H. Mathews. Correction: Divisions of North Australian tribes. Proceedings of the American Philosophical Society, 46(187):368, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
983479.

Mathews:1907:LBT

[Mat07c] R. H. Mathews. Language of the Birdhawal tribe, in Gippsland, Victoria. Proceedings of the American Philosophical Society, 46(187): 346–359, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983474.

Mathews:1907:LST

[Mat07d] R. H. Mathews. Languages of some tribes of Western Australia. Proceedings of the American Philosophical Society, 46(187):361– 368, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983478.

Mathews:1909:CSU

[Mat09a] R. H. Mathews. Ceremonial stones used by the Australian Aborigines. Proceedings of the American Philosophical Society, 48(191):1–7, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983807.

Mathews:1909:FNC

[Mat09b] R. H. Mathews. Further notes on ceremonial stones, Australia. Proceedings of the American Philosophical Society, 48(193):460– 462, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984062.

Mathews:1909:SBC

[Mat09c] R. H. Mathews. Some burial customs of the Australian Aborigines. Proceedings of the American Philosophical Society, 48(192): 313–318, April/August 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984160.

Mathews:1910:FNB

[Mat10] R. H. Mathews. Further notes on burial customs, Australia. Proceedings of the American Philosophical Society, 49(196):297–306, August/ September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984122.

Matthew:1919:RDF

[Mat19] W. D. Matthew. Recent discoveries of fossil vertebrates in the West Indies and their bearing on the origin of the Antillean fauna. Proceedings of the American Philosophical Society, 58(3):161–181, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984468.

Mayor:1919:DOC

[May19] Alfred Goldsborough Mayor. Detecting ocean currents by observing their hydrogen-ion concentration. Proceedings of the American Philosophical Society, 58(2):150–160, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984531.

Mayor:1920:RTS

[May20] Alfred Goldsborough Mayor. The reefs of Tutuila, Samoa, in their relation to coral reef theories. Proceedings of the American Philosophical Society, 59(3):224–236, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984500.

Mayor:1921:RAA

[May21] Alfred Goldsborough Mayor. Rose Atoll, American Samoa. Proceedings of the American Philosophical Society, 60(2):62–70, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984384.

Mayor:1943:EAP

[May43] A. Hyatt Mayor. Early American painters in England. Proceedings of the American Philosophical Society, 87(1):105–109, July 1943.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985003.

Mayr:1949:SS

[May49] Ernst Mayr. Speciation and selection. Proceedings of the American Philosophical Society, 93(6):514–519, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143339.

Marshall:1915:NFN

[MB15] J. T. W. Marshall and H. W. Banks. A new form of nephelometer. Proceedings of the American Philosophical Society, 54(217):176–184, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983922.

McCay:1904:TA

[McC04] Leroy W. McCay. Trisulphoxyarsenic acid. Proceedings of the American Philosophical Society, 43(176):112–115, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983495.

McClellan:1905:UFP

[McC05] William McClellan. On the use of the falling plate oscillograph as a phase meter. Proceedings of the American Philosophical Society, 44(180):166–176, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983514.

McCollum:1919:RDP

[McC19] E. V. McCollum. The relation of the diet to pellagra. Proceedings of the American Philosophical Society, 58(1):41–54, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984479.

McClung:1926:MNG

[McC26] C. E. McClung. “Martinichthys” — a new genus of Cretaceous fish from Kansas, with descriptions of six new species. Proceedings of the American Philosophical Society, 65(5):20–26, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984285.

McClung:1927:DBR

[McC27] C. E. McClung. The development of biological research in America. Proceedings of the American Philosophical Society, 66(??):67–78, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301058.

McComb:1941:GOI

[McC41] H. E. McComb. Geomagnetic observatories and instruments. Proceedings of the American Philosophical Society, 84(2):239–255, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984921.

McCoy:1945:RAA

[McC45] James C. McCoy. Report on aircraft activities at West Base, Antarctica. Proceedings of the American Philosophical Society, 89(1):387– 398, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985349.

McDiarmid:1915:VST

[McD15] R. J. McDiarmid. The variable stars TV, TW, TX Cassiopeiæ and T Leonis Minoris. Proceedings of the American Philosophical Society, 54(216):66–69, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983914.

McEwen:1938:MDO

[McE38] George F. McEwen. Modern dynamical oceanography: An achievement of applications to ocean observations of principles of mechanics and heat. Proceedings of the American Philosophical Society, 79(2):145–166, June 30, 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984617.

McGregor:1943:BEA

[McG43] John C. McGregor. Burial of an early American magician. Proceedings of the American Philosophical Society, 86(2):270–298, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985106.

McKee:1940:PAF

[McK40] Kenneth N. McKee. The popularity of the “American” on the French stage during the Revolution. Proceedings of the American Philosophical Society, 83(3):479–491, September 1940. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985115.

McMaster:1923:SGB

[McM23] Philip D. McMaster. The significance of the gall bladder. Proceedings of the American Philosophical Society, 62(3):185–189, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984506.

McMath:1938:RSS

[McM38] Robert R. McMath. Recent studies in solar phenomena. Proceedings of the American Philosophical Society, 79(4):475–498, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984938.

McNish:1941:SFM

[McN41] A. G. McNish. The significance of fossil magnetism. Proceedings of the American Philosophical Society, 84(2):225–237, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984920.

Mudd:1937:LBA

[MCPL37] Stuart Mudd, E. J. Czarnetzky, Horace Pettit, and David Lackman. Labile bacterial antigens and methods for their preparation and preservation. Proceedings of the American Philosophical Society, 77(4):463–466, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984772.

Mead:1926:PP

[Mea26] Charles W. Mead. The prehistoric Peruvians. Proceedings of the American Philosophical Society, 65(3):141–149, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984483.

Menzel:1939:PSA

[Men39] Donald H. Menzel. Problems of the solar atmosphere. Proceedings of the American Philosophical Society, 81(2):107–124, June 1939. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984978.

Mendelsohn:1940:BVC

[Men40] Charles J. Mendelsohn. Blaise de Vigenère and the “chiffre carré”. Proceedings of the American Philosophical Society, 82(2):103–129, March 22, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985011.

Merriman:1903:PLW

[Mer03] Mansfield Merriman. The principle of least work in mechanics and its use in investigations regarding the ether of space. Proceedings of the American Philosophical Society, 42(173):162–165, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983728.

Merriman:1905:RBE

[Mer05] Mansfield Merriman. The relation between the economic depth of a bridge truss and the depth that gives greatest stiffness. Proceedings of the American Philosophical Society, 44(180):164–165, May/ July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983513.

Merrill:1926:PCK

[Mer26] George P. Merrill. The present condition of knowledge on the composition of meteorites. Proceedings of the American Philosophical Society, 65(2):119–130, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984419.

Merriam:1927:OSS

[Mer27] John C. Merriam. Our sister societies. Proceedings of the American Philosophical Society, 66(??):737–738, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301092.

Merrill:1933:LHB

[Mer33] E. D. Merrill. Loureiro and his botanical work. Proceedings of the American Philosophical Society, 72(4):229–239, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984687.

Merriam:1934:CEC

[Mer34] John C. Merriam. Conservation and evolution in a changing social program. Proceedings of the American Philosophical Society, 73(5):351–370, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984626.

Merrill:1936:PBO

[Mer36] E. D. Merrill. Palisot de Beauvois as an overlooked American botanist. Proceedings of the American Philosophical Society, 76(6): 899–920, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984575.

Merrill:1937:SCO

[Mer37] E. D. Merrill. On the significance of certain oriental plant names in relation to introduced species. Proceedings of the American Philosophical Society, 78(1):111–146, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984661.

Merrill:1942:GOR

[Mer42] E. D. Merrill. A generally overlooked Rafinesque paper. Proceedings of the American Philosophical Society, 86(1):72–90, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985080.

Merrill:1943:RPS

[Mer43] E. D. Merrill. Rafinesque’s publications from the standpoint of world botany. Proceedings of the American Philosophical Society, 87(1):110– 119, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985004.

Metzger:1905:FHC

[Met05] John A. Metzger. The Filipino; his customs and character. Proceedings of the American Philosophical Society, 44(179):6–31, January/ April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983599.

McMath:1949:REI

[MG49] Robert R. McMath and Leo Goldberg. Recent exploration of the infra-red solar spectrum at the McMath–Hulbert Observatory. Proceedings of the American Philosophical Society, 93(5):362–372, November 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143388.

Munroe:1920:PDT

[MH20] Charles E. Munroe and Spencer P. Howell. Products of detonation of TNT. Proceedings of the American Philosophical Society, 59(3):194– 223, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984499.

McMath:1931:PRA

[MHM31] Francis C. McMath, Henry S. Hulbert, and Robert R. McMath. Preliminary results on the application of the motion picture camera to celestial photography. Proceedings of the American Philosophical Society, 70(4):371–379, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984202.

Michelson:1906:FA

[Mic06] A. A. Michelson. Form analysis. Proceedings of the American Philosophical Society, 45(183):110–116, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983820.

Michelson:1915:RPT

[Mic15] A. A. Michelson. The ruling and performance of a ten inch diffraction grating. Proceedings of the American Philosophical Society, 54 (217):137–142, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983918.

Miller:1907:GWG

[Mil07] G. A. Miller. The groups which are generated by two operators of orders two and four respectively whose commutator is of order two. Proceedings of the American Philosophical Society, 46(185):146–150, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983448.

Miller:1910:GGT

[Mil10a] G. A. Miller. Groups generated by two operators each of which transforms the square of the other into a power of itself. Proceedings of the American Philosophical Society, 49(195):238–242, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983902.

Miller:1910:SRB

[Mil10b] G. A. Miller. Some relations between substitution group properties and abstract groups. Proceedings of the American Philosophical Society, 49(196):307–314, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984123.

Miller:1911:TSL

[Mil11] G. A. Miller. On the totality of the substitutions on n letters which are commutative with every substitution of a given group on the same letters. Proceedings of the American Philosophical Society, 50 (198):139–146, January/April 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983957.

Miller:1919:SOE

[Mil19a] John A. Miller. The Sproul Observatory Eclipse Expedition, June 8, 1918. Coronal arches and streamers. Proceedings of the American Philosophical Society, 58(4):272–281, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984405.

Millikan:1919:SSA

[Mil19b] Robert A. Millikan. Some scientific aspects of the meteorological work of the United States Army. Proceedings of the American Philosophical Society, 58(2):133–149, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984530.

Millikan:1926:LFY

[Mil26] R. A. Millikan. The last fifteen years of physics. Proceedings of the American Philosophical Society, 65(2):68–78, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984413.

Millikan:1927:RDS

[Mil27] R. A. Millikan. Recent developments in spectroscopy. Proceedings of the American Philosophical Society, 66(??):211–230, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301065.

Miller:1930:ISS

[Mil30] John A. Miller. An introductory statement to the symposium. Proceedings of the American Philosophical Society, 69(1):397–399, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984374.

Miller:1935:CSJ

[Mil35] John A. Miller. Charles E. St. John. Proceedings of the American Philosophical Society, 75(8):769–770, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984593.

Milliman:1946:RBG

[Mil46] W. A. Milliman. The relationships between governmental and private responsibilities for national health. Proceedings of the American Philosophical Society, 90(4):280–288, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301022.

Millikan:1947:IMS

[Mil47] Robert A. Millikan. The interchange of men of science. Proceedings of the American Philosophical Society, 91(1):129–132, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143135.

Mims:1912:DVU

[Mim12] Stewart L. Mims. The diary of a voyage to the United States, by Moreau de Saint-Mery. Proceedings of the American Philosophical Society, 51(205):242–251, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984142.

Mitchell:1919:FS

[Mit19] S. A. Mitchell. The flash spectrum. Proceedings of the American Philosophical Society, 58(4):265–268, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984403.

Mitchell:1926:FSO

[Mit26] S. A. Mitchell. The flash spectrum observed at the 1925 eclipse. Proceedings of the American Philosophical Society, 65(2):99–104, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984417.

Mitchell:1928:DS

[Mit28] S. A. Mitchell. The distances of the stars. Proceedings of the American Philosophical Society, 67(3):267–286, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984293.

Mitchell:1930:HCE

[Mit30] S. A. Mitchell. Heights in the chromosphere from eclipse spectra. Proceedings of the American Philosophical Society, 69(1):433–439, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984377.

Mitchell:1932:SDR

[Mit32] S. A. Mitchell. Spectroscopic discoveries at the recent total eclipse. Proceedings of the American Philosophical Society, 71(6):343–348, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984645.

Mitchell:1938:AEE

[Mit38] S. A. Mitchell. With an astronomer on an eclipse expedition. Proceedings of the American Philosophical Society, 79(3):341–360, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984854.

Mitchell:1942:ADE

[Mit42] Samuel Alfred Mitchell. Astronomy during the early years of the American Philosophical Society. Proceedings of the American Philosophical Society, 86(1):13–21, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985073.

Millikan:1940:EDI

[MN40] Robert A. Millikan and H. Victor Neher. Energy distribution of incoming cosmic-ray particles. Proceedings of the American Philosophical Society, 83(2):409–428, August 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985018.

Montgomery:1904:MSF

[Mon04a] Thomas H. Montgomery. The morphological superiority of the female sex. Proceedings of the American Philosophical Society, 43(178):

365–380, October/December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983577.

Montgomery:1904:MFR

[Mon04b] Thomas H. Montgomery, Jr. The main facts in regard to the cellular basis of heredity. Proceedings of the American Philosophical Society, 43(175):5–14, January 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983411.

Montgomery:1908:MEO

[Mon08] Thos. H. Montgomery, Jr. On the morphology of the excretory organs of Metazoa: A critical review. Proceedings of the American Philosophical Society, 47(190):547–635, September/December 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983832.

Montgomery:1928:EFC

[Mon28] James A. Montgomery. An enactment of fundamental constitutional law in Old South Arabia. Proceedings of the American Philosophical Society, 67(3):207–213, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984290.

Moore:1908:MPL

[Moo08] George T. Moore. The most primitive living representative of the ancestors of the plant kingdom. Proceedings of the American Philosophical Society, 47(188):91–96, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983802.

Moore:1912:CW

[Moo12] John Bassett Moore. Contraband of war. Proceedings of the American Philosophical Society, 51(203):18–49, January/March 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983978.

Moore:1916:O

[Moo16] John Bassett Moore. Outline. Proceedings of the American Philosophical Society, 55(4):291–296, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984101.

Moon:1936:MPS

[Moo36] Virgil H. Moon. The mechanism and pathology of shock. Proceedings of the American Philosophical Society, 76(5):653–662, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984753.

Moore:1947:WMS

[Moo47] J. Percy Moore. William Maclure — scientist and humanitarian. Proceedings of the American Philosophical Society, 91(3):234–249, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143396.

Morris:1900:HH

[Mor00] J. Cheston Morris. Henry Hartshorne. Proceedings of the American Philosophical Society, 39(164):i–xii, October/December 1900. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983785.

Morris:1909:JWS

[Mor09] Harrison S. Morris. Joseph Wharton, Sc.D., LL.D. Proceedings of the American Philosophical Society, 48(193):lxxi–lxxvi, September/ December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984069.

Morley:1910:OWG

[Mor10] Edward W. Morley. Oliver Wolcott Gibbs, 1822–1908. Proceedings of the American Philosophical Society, 49(197):xix–xxxi, October/ December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984093.

Morgan:1915:CHM

[Mor15] T. H. Morgan. The constitution of the hereditary material. Proceedings of the American Philosophical Society, 54(217):143–153, May/ July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983919.

Morse:1926:MUA

[Mor26] Withrow Morse. The metabolism of uric acid. Proceedings of the American Philosophical Society, 65(3):196–201, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984489.

Morris:1932:ACM

[Mor32] Harrison S. Morris. The art called modern. Proceedings of the American Philosophical Society, 71(4):173–178, April 1932. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984788.

Morris:1935:RAT

[Mor35] Harrison S. Morris. The revolt against taste. Proceedings of the American Philosophical Society, 75(4):275–280, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984628.

Morey:1936:EAO

[Mor36] Charles R. Morey. The excavation of Antioch-on-the-Orontes. Proceedings of the American Philosophical Society, 76(5):637–651, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984752.

Morris:1937:I

[Mor37] Harrison S. Morris. Illusions. Proceedings of the American Philosophical Society, 78(2):287–290, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984537.

Morley:1943:AOI

[Mor43a] Sylvanus Griswold Morley. Archæological investigations of the Carnegie Institution of Washington in the Maya area of Middle America, during the past twenty-eight years. Proceedings of the American Philosophical Society, 86(2):205–219, February 1943. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985099.

Morris:1943:JL

[Mor43b] Roland S. Morris. Jefferson as a lawyer. Proceedings of the American Philosophical Society, 87(3):211–215, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984867.

Morse:1949:ENS

[Mor49] Marston Morse. Equilibria in nature: Stable and unstable. Proceedings of the American Philosophical Society, 93(3):222–225, June 1949.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143469.

Mozley:1937:BSS

[Moz37] Alan Mozley. A biological study of the sub-Arctic mollusca. Proceedings of the American Philosophical Society, 78(1):147–189, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984662.

Montagu:1944:EAA

[MP44] M. F. Ashley Montagu and C. Bernard Peterson. The earliest account of the association of human artifacts with fossil mammals in North America. Proceedings of the American Philosophical Society, 87(5):407–419, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985294.

Muller:1926:FSP

[MPG26] John Hughes Müller, Edwin Frederic Pike, and A. Kenneth Graham. Further studies on the preparation of metallic germanium and the volatility of the metal in hydrogen and in vacuo. Proceedings of the American Philosophical Society, 65(1):15–32, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984335.

Miller:1920:PFS

[MPS20] John A. Miller, John H. Pitman, and Hannah B. Steele. The parallaxes of fifty stars (second list) determined at Sproul Observatory. Proceedings of the American Philosophical Society, 59(2):85– 149, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984316.

MacDougal:1917:GI

[MS17] D. T. MacDougal and H. A. Spoehr. Growth and imbibition. Proceedings of the American Philosophical Society, 56(4):289–352, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983964.

MacDougal:1920:CCB

[MS20] D. T. MacDougal and H. A. Spoehr. The components and colloidal behavior of plant protoplasm. Proceedings of the American Philosophical Society, 59(3):150–170, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984495.

Millikan:1931:RTD

[MS31] Robert A. Millikan and W. F. G. Swann. Round table discussion. Proceedings of the American Philosophical Society, 70(3):309–316, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984459.

Mudd:1947:ROP

[Mud47] Stuart Mudd. Recent observations on programs for medicine and national health in the USSR. Proceedings of the American Philosophical Society, 91(2):181–188, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143603.

Muller:1926:FUM

[Mül26a] John Hughes Müller. The fluogermanates of the univalent metals. Proceedings of the American Philosophical Society, 65(5):44–57, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984287.

Muller:1926:FSA

[Mül26b] John Hughes Müller. Further studies on the allotrophy of Germanic oxide. Proceedings of the American Philosophical Society, 65(3):183– 195, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984488.

Muller:1936:ISS

[Mül36] John H. Müller. Ionic silver from silver hydroxide as a therapeutic agent. Proceedings of the American Philosophical Society, 76(3):303– 325, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984546.

Muller:1949:DMC

[Mul49] H. J. Muller. The Darwinian and modern conceptions of natural selection. Proceedings of the American Philosophical Society, 93(6):459–470, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143335.

Mumford:1940:LF

[Mum40] Lewis Mumford. Looking forward. Proceedings of the American Philosophical Society, 83(4):538–547, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984834.

Munroe:1909:DGC

[Mun09] Charles E. Munroe. The detonation of gun cotton. Proceedings of the American Philosophical Society, 48(191):69–71, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983813.

Munroe:1910:PEM

[Mun10] Charles E. Munroe. The propagation of explosions in mixtures of petroleum vapor with air in tubes. Proceedings of the American Philosophical Society, 49(195):203–206, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983897.

Munro:1911:CLT

[Mun11] Dana C. Munro. The cost of living in the twelfth century. Proceedings of the American Philosophical Society, 50(201):497–504, September/ December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984072.

Munro:1916:PC

[Mun16] Dana C. Munro. The popes and the crusades. Proceedings of the American Philosophical Society, 55(5):348–356, 1916. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984051.

Murray:1944:FRP

[Mur44] Elsie Murray. French refugees of 1793 in Pennsylvania. Proceedings of the American Philosophical Society, 87(5):387–393, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985290.

McPhedran:1925:AAS

[MW25] F. Maurice McPhedran and Charles N. Weyl. The application of automatic synchronization of X-ray exposures to the study of intra-thoracic movements. Proceedings of the American Philosophical Society, 64(2):89–94, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984190.

Merriam:1937:MIM

[MW37] John C. Merriam and Warren Weaver. The most important methods of promoting research. Proceedings of the American Philosophical Society, 77(4):605–610, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984780.

Myers:1942:QSP

[Mye42] Earl H. Myers. A quantitative study of the productivity of the foraminifera in the sea. Proceedings of the American Philosophical Society, 85(4):325–342, June 8, 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984988.

Myers:1943:LAF

[Mye43] Earl H. Myers. Life activities of foraminifera in relation to marine ecology. Proceedings of the American Philosophical Society, 86(3):439–458, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985030.

Neugebauer:1948:AOT

[Neu48] Otto Neugebauer. The astronomical origin of the theory of conic sections. Proceedings of the American Philosophical Society, 92(3):136– 138, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143581.

Newbold:1918:SDS

[New18] Wm. Romaine Newbold. The Syriac dialogue “Socrates.” A study in Syrian philosophy. Proceedings of the American Philosophical Society, 57(2):99–111, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984162.

Newcombe:1940:SPC

[New40] Curtis L. Newcombe. Studies on the phosphorus content of the estuarine waters of Chesapeake Bay. Proceedings of the American Philosophical Society, 83(5):621–630, October 1940. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985132.

Newcomer:1944:FSL

[New44] Mabel Newcomer. The federal, state, and local tax structure after the War. Proceedings of the American Philosophical Society, 88(1):50–54, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985189.

Newell:1948:MMP

[New48] Irwin M. Newell. Marine molluscan provinces of Western North America: A critique and a new analysis. Proceedings of the American Philosophical Society, 92(3):155–166, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143584.

Nichols:1902:SOA

[Nic02] M. Louise Nichols. The spermatogenesis of Oniscus Asellus Linn. With especial reference to the history of the chromatin. Proceedings of the American Philosophical Society, 41(168):77–112, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983526.

Nichols:1910:ETP

[Nic10] Edward L. Nichols. The effects of temperature on phosphorescence and fluorescence. Proceedings of the American Philosophical Society, 49(196):267–280, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984119.

Nichols:1916:CPP

[Nic16] Edward L. Nichols. Color-photographs of the phosphorescence of certain metallic sulphides. Proceedings of the American Philosophical Society, 55(6):494–502, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983984.

Nichols:1944:WRS

[Nic44a] Roy F. Nichols. War and research in social science. Proceedings of the American Philosophical Society, 87(4):361–364, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985048.

Nicolson:1944:MLS

[Nic44b] Marjorie H. Nicolson. Merchants of light: Scholarship in arts and letters. Proceedings of the American Philosophical Society, 87(4):352–360, January 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985047.

Nichols:1947:ADC

[Nic47] Roy F. Nichols. American democracy and the Civil War. Proceedings of the American Philosophical Society, 91(2):143–149, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143600.

Nichols:1948:PDI

[Nic48] Jeannette P. Nichols. Political distortions of international monetary relations. Proceedings of the American Philosophical Society, 92(6):433–446, December 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143198.

Nipher:1906:EVE

[Nip06] Francis E. Nipher. The elimination of velocity effects in measuring pressures in a fluid stream. Proceedings of the American Philosophical Society, 45(182):77–81, January/April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983681.

Nipher:1911:DDE

[Nip11a] Francis E. Nipher. Disruptive discharges of electricity through flames. Proceedings of the American Philosophical Society, 50(200): 397–403, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983931.

Nipher:1911:OP

[Nip11b] Francis E. Nipher. An optical phenomenon. Proceedings of the American Philosophical Society, 50(199):316–320, May/August 1911. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984047.

Nipher:1913:MEE

[Nip13] Francis E. Nipher. Matter in its electrically explosive state. Proceedings of the American Philosophical Society, 52(209):283–286, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983873.

Nipher:1919:GRF

[Nip19] Francis E. Nipher. Graphical representation of functions of the nth degree. Proceedings of the American Philosophical Society, 58(3):236– 240, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984471.

Newcombe:1939:DPC

[NL39] Curtis L. Newcombe and Andrew G. Lang. The distribution of phosphates in the Chesapeake Bay. Proceedings of the American Philosophical Society, 81(3):393–420, August 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984862.

Nock:1942:RAA

[Noc42] Arthur Darby Nock. Religious attitudes of the Ancient Greeks. Proceedings of the American Philosophical Society, 85(5):472–482, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985067.

Nolan:1934:LAP

[Nol34] J. Bennett Nolan. Lafayette and the American Philosophical Society. Proceedings of the American Philosophical Society, 73(2):117– 126, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984760.

Nolan:1944:BFM

[Nol44] J. Bennett Nolan. Ben Franklin’s mortgage on the Daniel Boone farm. Proceedings of the American Philosophical Society, 87(5):394– 397, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985291.

Norris:1932:IBP

[Nor32] George W. Norris. Improvements in banking practice suggested by the present depression. Proceedings of the American Philosophical Society, 71(3):117–124, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984557.

Northrop:1941:CDA

[Nor41] John H. Northrop. Crystalline diphtheria antitoxin. Proceedings of the American Philosophical Society, 85(1):13–17, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985122.

Notestein:1939:IFP

[Not39] Frank W. Notestein. Intrinsic factors in population growth. Proceedings of the American Philosophical Society, 80(4):499–511, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985053.

Notestein:1943:SIP

[Not43] Frank W. Notestein. Some implications of population change for Post-War Europe. Proceedings of the American Philosophical Society, 87(2):165–174, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985094.

Noyes:1927:V

[Noy27] William Albert Noyes. Valence. Proceedings of the American Philosophical Society, 66(??):287–308, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301070.

Noyes:1947:UNE

[Noy47] W. Albert Noyes, Jr. The United Nations Educational, Scientific and Cultural Organization. Proceedings of the American Philosophical Society, 91(1):133–136, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143136.

Noyes:1914:VNA

[NP14] William A. Noyes and Ralph S. Potter. The valence of nitrogen in ammonium salts. Proceedings of the American Philosophical Society, 53(213):18–24, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984127.

Noble:1937:SFF

[NS37] G. K. Noble and A. Schmidt. The structure and function of the facial and labial pits of snakes. Proceedings of the American Philosophical Society, 77(3):263–288, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984731.

Nuttall:1926:ATP

[Nut26] Zelia Nuttall. The Aztecs and their predecessors in the valley of Mexico. Proceedings of the American Philosophical Society, 65(4): 245–255, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984236.

Nuttycombe:1938:ASG

[NW38] John W. Nuttycombe and Aubrey J. Waters. The American species of the genus Stenostomum. Proceedings of the American Philosophical Society, 79(2):213–301, June 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984620.

Osborn:1931:EEM

[OC31] Henry Fairfield Osborn and Edwin H. Colbert. The elephant enamel method of measuring Pleistocene time. Also stages in the succession of fossil man and stone age industries. Proceedings of the American Philosophical Society, 70(2):187–191, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984321.

Odgers:1941:BE

[Odg41] Merle M. Odgers. Bache as an educator. Proceedings of the American Philosophical Society, 84(2):161–171, 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984915.

Odgers:1943:EAP

[Odg43] Merle M. Odgers. Education and the American Philosophical Society. Proceedings of the American Philosophical Society, 87(1):12–24, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984995.

Ogburn:1907:CLO

[Ogb07] John H. Ogburn. Comparison of latitude observations at Sayre Observatory, South Bethlehem, and at Flower Observatory, Philadelphia, from September 30, 1904, to September 3, 1906. Proceedings of the American Philosophical Society, 46(185):165–170, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983450.

Oliver:1902:BCM

[Oli02] Charles A. Oliver. Blindness from congenital malformation of the skull. Proceedings of the American Philosophical Society, 41 (169):161–174, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983535.

Oliver:1904:RCS

[Oli04] Charles A. Oliver. Regulation of color-signals in marine and naval service. Proceedings of the American Philosophical Society, 43(176):207– 210, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983505.

Oliver:1905:PGS

[Oli05] Charles A. Oliver. A plea for governmental supervision of posts necessitating normal perception of color. Proceedings of the American Philosophical Society, 44(179):40–41, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983603.

Olivier:1933:HTD

[Oli33] Charles P. Olivier. Heights and train-drifts of Leonid meteors of 1932. Proceedings of the American Philosophical Society, 72(4):215– 227, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984686.

Olivier:1935:PPS

[Oli35] Charles P. Olivier. Problems presented by a study of fireballs recently observed in America. Proceedings of the American Philosophical Society, 75(6):483–498, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984610.

Olivier:1936:RMT

[Oli36] Charles P. Olivier. Research on meteor trains. Proceedings of the American Philosophical Society, 76(6):772–773, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984568.

Olivier:1942:LEM

[Oli42] Charles P. Olivier. Long enduring meteor trains. Proceedings of the American Philosophical Society, 85(2):93–135, January 24, 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985034.

Olivier:1947:LEM

[Oli47] Charles P. Olivier. Long enduring meteor trains: Second paper. Proceedings of the American Philosophical Society, 91(4):315–327, October 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143345.

Olschki:1941:WCS

[Ols41] Leonardo Olschki. What Columbus saw on landing in the West Indies. Proceedings of the American Philosophical Society, 84(5):633– 659, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984844.

Osborn:1919:CRS

[OM19] Henry Fairfield Osborn and Charles Craig Mook. Characters and restoration of the sauropod genus Camarasaurus Cope. From type material in the Cope Collection in the American Museum of Natural History. Proceedings of the American Philosophical Society, 58(6): 386–396, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984263.

Oppenheimer:1946:AW

[Opp46] J. R. Oppenheimer. Atomic weapons. Proceedings of the American Philosophical Society, 90(1):7–10, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www. jstor.org/stable/3301032. Symposium on Atomic Energy and Its Implications (January 1946).

Ortmann:1902:GDF

[Ort02] A. E. Ortmann. The geographical distribution of freshwater decapods and its bearing upon ancient geography. Proceedings of the American Philosophical Society, 41(171):267–400, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983485.

Ortmann:1905:MAS

[Ort05] A. E. Ortmann. The mutual affinities of the species of the genus Cambarus, and their dispersal over the United States. Proceedings of the American Philosophical Society, 44(180):91–136, May/July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983509.

Ortmann:1909:DFW

[Ort09] A. E. Ortmann. The destruction of the fresh-water fauna in Western Pennsylvania. Proceedings of the American Philosophical Society, 48 (191):90–110, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983815.

Ortmann:1913:ADI

[Ort13] Arnold E. Ortmann. The Alleghenian Divide, and its influence upon the freshwater fauna. Proceedings of the American Philosophical Society, 52(210):287–390, May/August 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984083.

Ortmann:1918:NFM

[Ort18] A. E. Ortmann. The Nayades (freshwater mussels) of the upper Tennessee Drainage. With notes on synonymy and distribution. Proceedings of the American Philosophical Society, 57(6):521–626, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984465.

Ortmann:1920:CSS

[Ort20] A. E. Ortmann. Correlation of shape and station in fresh-water mussels (Naiades). Proceedings of the American Philosophical Society, 59 (4):268–312, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984426.

Osborn:1904:RAO

[Osb04] Henry F. Osborn. Recent advances in our knowledge of the evolution of the horse. Proceedings of the American Philosophical Society, 43(176):156–157, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983499.

Osborn:1923:EAM

[Osb23] Henry Fairfield Osborn. The explorations of the American Museum of Natural History in China and Mongolia. Proceedings of the American Philosophical Society, 62(3):90–94, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984503.

Osborn:1925:FCE

[Osb25] Henry Fairfield Osborn. Final conclusions on the evolution, phylogeny, and classification of the Proboscidea. Proceedings of the American Philosophical Society, 64(1):17–35, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984300.

Osborn:1927:RDR

[Osb27] Henry Fairfield Osborn. Recent discoveries relating to the origin and antiquity of man. Proceedings of the American Philosophical Society, 66(??):373–389, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301074.

Osborn:1928:PSP

[Osb28] Henry Fairfield Osborn. Present status of the problem of human ancestry. Proceedings of the American Philosophical Society, 67(2): 151–155, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984276.

Osborn:1934:TND

[Osb34] Henry Fairfield Osborn. The thirty-nine distinct lines of Proboscidean descent, and their migration into all parts of the world except Australia. Proceedings of the American Philosophical Society, 74(4):273– 285, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984795.

Osgood:1945:A

[Osg45] Charles G. Osgood. Areopagitica-1644. Proceedings of the American Philosophical Society, 89(3):495–498, October 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985430.

Osterhout:1916:DTA

[Ost16] W. J. V. Osterhout. A dynamical theory of antagonism. Proceedings of the American Philosophical Society, 55(7):533–553, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983924.

Overholser:1946:MH

[Ove46] Winfred Overholser. Mental hygiene. Proceedings of the American Philosophical Society, 90(4):259–264, September 1946. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3301018.

Packard:1903:HCA

[Pac03] Alpheus S. Packard. Hints on the classification of the Arthropoda; the group a polyphyletic one. Proceedings of the American Philosophical Society, 42(173):142–161, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983727.

Packard:1904:ODS

[Pac04a] Alpheus S. Packard. Opisthenogenesis, or the development of segments, median tubercles and markings “a tergo”. Proceedings of the American Philosophical Society, 43(177):289–294, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983767.

Packard:1904:OMO

[Pac04b] Alpheus S. Packard. The origin of the markings of organisms (pœcilogenesis) due to the physical rather than to the biological environment; with criticisms of the Bates–Müller hypotheses. Proceedings of the American Philosophical Society, 43(178):393–450, October/ December 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983581.

Packard:1935:JCW

[Pac35] Francis R. Packard. James Cornelius Wilson. Proceedings of the American Philosophical Society, 75(4):331–333, 1935. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984634.

Packard:1942:MAP

[Pac42] Francis R. Packard. Medicine and the American Philosophical Society. Proceedings of the American Philosophical Society, 86(1):91–102, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985081.

Pace:1946:APS

[Pac46] Antonio Pace. The American Philosophical Society and Italy. Proceedings of the American Philosophical Society, 90(5):387–421, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301047.

Palmer:1928:FRC

[Pal28] R. H. Palmer. Fossil and recent corals and coral reefs of Western Mexico. Three new species. Proceedings of the American Philosophical Society, 67(1):21–31, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984253.

Parker:1911:OSP

[Par11] G. H. Parker. The origin and significance of the primitive nervous system. Proceedings of the American Philosophical Society, 50 (199):217–225, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984035.

Parker:1915:PAI

[Par15] George H. Parker. The problem of adaptation as illustrated by the fur seals of the Pribilof Islands. Proceedings of the American Philosophical Society, 54(216):1–6, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983905.

Parker:1916:TNM

[Par16] G. H. Parker. Types of neuromuscular mechanism in sea-anemones. Proceedings of the American Philosophical Society, 55(5):340–343, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984049.

Parker:1918:CSS

[Par18] G. H. Parker. A critical survey of the sense of hearing in fishes. Proceedings of the American Philosophical Society, 57(2):69–98, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984161.

Parker:1920:PR

[Par20] G. H. Parker. The phosphorescence of Renilla. Proceedings of the American Philosophical Society, 59(3):171–175, 1920. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984496.

Parker:1922:RRI

[Par22] G. H. Parker. The relations of the retinal image to animal reactions. Proceedings of the American Philosophical Society, 61(2):107– 116, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984408.

Parker:1934:NAA

[Par34] G. H. Parker. Neurohumors as activating agents for fish melanophores. Proceedings of the American Philosophical Society, 74(3):177–184, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984712.

Parker:1935:DPC

[Par35] G. H. Parker. The disappearance of primary caudal bands in the tail of fundulus and its relation to the neurohumoral hypothesis. Proceedings of the American Philosophical Society, 75(1):1–10, May 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984721.

Parker:1937:ICC

[Par37] G. H. Parker. Integumentary color changes of elasmobranch fishes especially of Mustelus. Proceedings of the American Philosophical Society, 77(3):223–247, March 31, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984729.

Parker:1938:MRB

[Par38] G. H. Parker. Melanophore responses and blood supply (vasomotor changes). Proceedings of the American Philosophical Society, 78(3):513–527, January 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984766.

Parr:1939:PBM

[Par39] A. E. Parr. On the possibility of a biological mechanism controlling the occurrence of the oxygen minimum layer in the sea. Proceedings of the American Philosophical Society, 80(1):49–56, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984972.

Parker:1940:NCC

[Par40] G. H. Parker. On the neurohumors of the color changes in catfishes and on fats and oils as protective agents for such substances. Proceedings of the American Philosophical Society, 83(2):379–408, August 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985017.

Parker:1941:MAM

[Par41] G. H. Parker. The method of activation of melanophores and the limitations of melanophore responses in the catfish Ameiurus. Proceedings of the American Philosophical Society, 85(1):18–24, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985123.

Parker:1944:TFC

[Par44] G. H. Parker. The time factor in chromatophore responses. Proceedings of the American Philosophical Society, 87(5):429–434, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985296.

Passel:1945:SRS

[Pas45] Charles F. Passel. Sedimentary rocks of the Southern Edsel Ford Ranges, Marie Byrd Land, Antarctica. Proceedings of the American Philosophical Society, 89(1):123–131, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985317.

Patterson:1901:FFL

[Pat01] C. Stuart Patterson. Frederick Fraley, LL.D. Proceedings of the American Philosophical Society, 40(167):i–ix, December 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983763.

Patterson:1903:PT

[Pat03] C. Stuart Patterson. The problem of the trusts. Proceedings of the American Philosophical Society, 42(172):15–27, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983639.

Paton:1913:CSF

[Pat13] Stewart Paton. The correlation of structure and function in the development of the nervous system. Proceedings of the American Philosophical Society, 52(211):488–494, September/October 1913. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983972.

Patten:1916:CFE

[Pat16] William Patten. Coöperation as a factor in evolution. Proceedings of the American Philosophical Society, 55(7):503–532, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983923.

Patterson:1928:SEI

[Pat28] Ernest Minor Patterson. Some economic implications of America’s changing world status. Proceedings of the American Philosophical Society, 67(3):199–205, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984289.

Patterson:1931:EAM

[Pat31] Ernest Minor Patterson. Economic adjustment in a machine age. Proceedings of the American Philosophical Society, 70(3):263–272, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984455.

Patterson:1932:IFB

[Pat32] Ernest Minor Patterson. International factors in the business depression. Proceedings of the American Philosophical Society, 71(3):105– 115, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984556.

Patterson:1934:COT

[Pat34a] Bryan Patterson. A contribution to the osteology of “Titanoides” and the relationships of the amblypoda. Proceedings of the American Philosophical Society, 73(2):71–101, 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984758.

Patterson:1934:FCP

[Pat34b] Ernest Minor Patterson. Factors controlling prices, domestic and international. Proceedings of the American Philosophical Society, 74(1): 63–75, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984584.

Patterson:1935:SCO

[Pat35] Bryan Patterson. Second contribution to the osteology and affinities of the Paleocene amblypod Titanoides. Proceedings of the American Philosophical Society, 75(2):143–162, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984533.

Paul:1936:SP

[Pau36] Henry N. Paul. Shakespeare in Philadelphia. Proceedings of the American Philosophical Society, 76(6):719–729, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984561.

Pearl:1916:ECA

[Pea16] Raymond Pearl. On the effect of continued administration of certain poisons to the domestic fowl, with special reference to the progeny. Proceedings of the American Philosophical Society, 55(3):243–258, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984115.

Pearl:1917:SRD

[Pea17] Raymond Pearl. The sex ratio in the domestic fowl. Proceedings of the American Philosophical Society, 56(5):416–436, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984030.

Pearl:1919:RCS

[Pea19] Raymond Pearl. The relative contribution of the staple commodities to the national food consumption. Proceedings of the American Philosophical Society, 58(3):182–222, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984469.

Pease:1921:MSD

[Pea21] F. G. Pease. Measurement of star diameters by the interferometer method. Proceedings of the American Philosophical Society, 60(4): 524–534, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984517.

Pearl:1924:CPG

[Pea24] Raymond Pearl. The curve of population growth. Proceedings of the American Philosophical Society, 63(1):10–17, 1924. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984438.

Penck:1927:GAE

[Pen27] Albrecht Penck. Geography among the Earth sciences. Proceedings of the American Philosophical Society, 66(??):621–644, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301085.

Pennell:1942:BSB

[Pen42] Francis W. Pennell. Benjamin Smith Barton as naturalist. Proceedings of the American Philosophical Society, 86(1):108–122, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985083.

Perkins:1940:ALA

[Per40] Earle B. Perkins. Animal life in the Antarctic. Proceedings of the American Philosophical Society, 82(5):833–834, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984894.

Perkins:1945:BLA

[Per45] Jack E. Perkins. Biology at Little America III, the West Base of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):270–284, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985329.

Pettit:1903:JMW

[Pet03] Henry Pettit. Joseph Miller Wilson, A.M., C.E. Proceedings of the American Philosophical Society, 42(174):i–vi, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983723.

Payne-Gaposchkin:1939:VSP

[PG39] Cecilia Payne-Gaposchkin. Variable stars: A plan of study. Proceedings of the American Philosophical Society, 81(2):189–210, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984982.

Payne-Gaposchkin:1943:VSS

[PGG43] Cecilia Payne-Gaposchkin and Sergei Gaposchkin. Variable stars and the sources of stellar energy. Proceedings of the American Philosophical Society, 86(3):329–338, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985022.

Phelps:1939:NS

[Phe39] William Lyon Phelps. Notes on Shakespeare. Proceedings of the American Philosophical Society, 81(4):573–579, September 1939. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984828.

Phelps:1940:MNS

[Phe40] William Lyon Phelps. More notes on Shakespeare. Proceedings of the American Philosophical Society, 83(3):493–502, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985116.

Phillips:1903:RP

[Phi03] Everett F. Phillips. A review of parthenogenesis. Proceedings of the American Philosophical Society, 42(174):275–345, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983715.

Phillips:1904:RAO

[Phi04] Alexander H. Phillips. Radium in an American ore. Proceedings of the American Philosophical Society, 43(176):157–160, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983500.

Pickering:1906:IST

[Pic06] Edward C. Pickering. An international Southern telescope. Proceedings of the American Philosophical Society, 45(182):44–53, January/ April 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983678.

Pickering:1912:OP

[Pic12] Edward C. Pickering. The objective prism. Proceedings of the American Philosophical Society, 51(207):564–567, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984019.

Piggot:1938:TSU

[Pig38] Charles Snowden Piggot. The technique of securing undisturbed coresamples of the ocean bottom. Proceedings of the American Philosophical Society, 79(1):35–40, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984677.

Plaskett:1930:SRG

[Pla30] J. S. Plaskett. The structure and rotation of the galaxy. Proceedings of the American Philosophical Society, 69(1):401–417, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984375.

Poole:1933:SIA

[Poo33] Dewitt Clinton Poole. Structural improvements in the administration of foreign affairs. Proceedings of the American Philosophical Society, 72(2):77–86, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984744.

Prescott:1904:RC

[Pre04] Albert B. Prescott. The rôle of carbon. Proceedings of the American Philosophical Society, 43(176):102–104, April 1904. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983493.

Prince:1902:MDT

[Pri02] J. Dyneley Prince. A modern Delaware tale. Proceedings of the American Philosophical Society, 41(168):20–34, January 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983519.

Prince:1903:DAS

[Pri03] J. Dyneley Prince. Dying American speech — echoes from Connecticut. Proceedings of the American Philosophical Society, 42(174): 346–352, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983717.

Prince:1914:MPL

[Pri14] J. Dyneley Prince. The morphology of the Passamaquoddy language of Maine. Proceedings of the American Philosophical Society, 53 (213):92–117, January/May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984132.

Prince:1915:PVS

[Pri15] J. Dyneley Prince. The pronouns and verbs of Sumerian. Proceedings of the American Philosophical Society, 54(216):27–46, January/ April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983910.

Prince:1916:ROS

[Pri16] J. Dyneley Prince. A rare Old Slavonic religious manual. Proceedings of the American Philosophical Society, 55(5):357–362, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984052.

Prince:1917:NTB

[Pri17] J. Dyneley Prince. The names Troyan and Boyan in Old Russian. Proceedings of the American Philosophical Society, 56(2):152–160, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984172.

Prince:1919:TMO

[Pri19] J. Dyneley Prince. Tatar material in Old Russian. Proceedings of the American Philosophical Society, 58(1):74–88, 1919. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984481.

Prince:1920:SC

[Pri20] J. Dyneley Prince. Slav and Celt. Proceedings of the American Philosophical Society, 59(3):184–193, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984498.

Prince:1922:FL

[Pri22] J. Dyneley Prince. The Faeroe language. Proceedings of the American Philosophical Society, 61(2):151–166, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984412.

Prince:1924:DDB

[Pri24] John Dyneley Prince. The Danish dialect of Bornholm. Proceedings of the American Philosophical Society, 63(2):190–207, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984473.

Prince:1927:IJ

[Pri27] John Dyneley Prince. The idiom of Jutland. Proceedings of the American Philosophical Society, 66(??):191–209, ???? 1927. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3301064.

Prince:1931:GGS

[Pri31] John Dyneley Prince. The Gottschee Germans of Slovenia. Proceedings of the American Philosophical Society, 70(4):391–398, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984204.

Prince:1934:TRM

[Pri34] J. Dyneley Prince. Turkish remains in modern Belgrade. Proceedings of the American Philosophical Society, 73(5):331–340, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984624.

Prince:1935:SLM

[Pri35] J. Dyneley Prince. Slavonic loan material in Hungarian. Proceedings of the American Philosophical Society, 75(7):591–603, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984810.

Pincus:1940:CBM

[PS40] Gregory Pincus and Herbert Shapiro. The comparative behavior of mammalian eggs in vivo and in vitro. VII. Further studies on the activation of rabbit eggs. Proceedings of the American Philosophical Society, 83(5):631–647, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985133.

Putnam:1947:RCC

[Put47] B. H. Putnam. Records of Courts of Common Law, especially of the sessions of the Justices of the Peace. Sources for the economic history of England in the Fourteenth and Fifteenth Centuries. Proceedings of the American Philosophical Society, 91(3):258–273, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143398.

Quinn:1945:STS

[Qui45] David Beers Quinn. Sir Thomas Smith (1513–1577) and the beginnings of English Colonial theory. Proceedings of the American Philosophical Society, 89(4):543–560, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985360.

Radcliffe:1934:SGA

[Rad34] George L. Radcliffe. Some governmental aspects of regional planning. Proceedings of the American Philosophical Society, 74(1):1–13, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984580.

Randall:1923:IRS

[Ran23] H. M. Randall. Infra-red spectra. Proceedings of the American Philosophical Society, 62(6):326–340, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984188.

Rand:1932:HC

[Ran32] E. K. Rand. The humanism of Cicero. Proceedings of the American Philosophical Society, 71(4):207–216, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984791.

Ravenel:1903:WAT

[Rav03] Mazyck P. Ravenel. The warfare against tuberculosis. Proceedings of the American Philosophical Society, 42(173):212–219, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983734.

Ravenel:1913:CTF

[Rav13] Mazyck P. Ravenel. The control of typhoid fever by vaccination. Proceedings of the American Philosophical Society, 52(209):226–233, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983868.

Reichert:1908:PRU

[RB08] Edward T. Reichert and Amos P. Brown. Preliminary report upon a crystallographic study of the hemoglobins: A contribution to the specificity of corresponding vital substances in different vertebrates. Proceedings of the American Philosophical Society, 47 (189):298–301, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983845.

Ruthven:1937:SRU

[RC37] Alexander G. Ruthven and Edward P. Cheyney. The support of research in the universities. Proceedings of the American Philosophical Society, 77(4):617–620, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984782.

Riddle:1945:PSH

[RD45] Oscar Riddle and Louis B. Dotti. Pituitary and sex hormones which increase plasma calcium in birds and mammals. Proceedings of the American Philosophical Society, 89(3):499–516, October 1945. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985431.

Rugh:1940:DER

[RE40] Roberts Rugh and Frank Exner. Developmental effects resulting from exposure to X-rays. II. Development of leopard frog eggs activated by bullfrog sperm. Proceedings of the American Philosophical Society, 83(5):607–619, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985131.

Read:1949:MLL

[Rea49] Conyers Read. More light on the London Naval Treaty of 1930. Proceedings of the American Philosophical Society, 93(4):290–308, September 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143155.

Reeder:1936:FHS

[Ree36] Robert P. Reeder. The first homes of the Supreme Court of the United States. Proceedings of the American Philosophical Society, 76(4):543– 596, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984602.

Reed:1938:IGH

[Ree38] George M. Reed. Influence of the growth of the host on oat smut development. Proceedings of the American Philosophical Society, 79(2):303–326, June 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984621.

Rehfuss:1916:NGS

[Reh16] Martin E. Rehfuss. The normal gastric secretion. Proceedings of the American Philosophical Society, 55(6):461–470, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983980.

Rehn:1940:CAN

[Reh40] James A. G. Rehn. Connection of the Academy of Natural Sciences of Philadelphia with our first national exploring expedition. Proceedings of the American Philosophical Society, 82(5):543–549, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984884.

Reid:1909:SN

[Rei09] Harry Fielding Reid. Seismological notes. Proceedings of the American Philosophical Society, 48(192):303–312, April/August 1909. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984159.

Reid:1911:IMR

[Rei11] Harry Fielding Reid. Isostasy and mountain ranges. Proceedings of the American Philosophical Society, 50(200):444–451, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983933.

Reid:1915:CIE

[Rei15] Harry Fielding Reid. Constitution of the interior of the Earth as indicated by seismological investigations. Proceedings of the American Philosophical Society, 54(219):290–297, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983875.

Reid:1920:DLW

[Rei20] Harry Fielding Reid. The distribution of land and water on the Earth. Proceedings of the American Philosophical Society, 59(4):313– 324, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984427.

Reichelderfer:1940:CWT

[Rei40] F. W. Reichelderfer. The contributions of Wilkes to terrestrial magnetism, gravity and meteorology. Proceedings of the American Philosophical Society, 82(5):583–600, June 1940. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984887.

Richards:1903:IOS

[Ric03] Theodore William Richards. The inclusion and occlusion of solvent in crystals. An insidious source of error in quantitative chemical investigation. Proceedings of the American Philosophical Society, 42 (172):28–36, January/April 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983640.

Richards:1904:SED

[Ric04] Theodore William Richards. Sources of error in the determination of the atomic weight of nitrogen. Proceedings of the American Philosophical Society, 43(176):116–122, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983496.

Richardson:1911:DEA

[Ric11] Owen W. Richardson. The dynamical effects of aggregates of electrons. Proceedings of the American Philosophical Society, 50 (200):347–365, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983928.

Richards:1912:AMC

[Ric12] Horace Clark Richards. An autocollimating mounting for a concave grating. Proceedings of the American Philosophical Society, 51(207): 554–563, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984018.

Richards:1917:SWR

[Ric17] Theodore W. Richards. Sir William Ramsay, K.C.B. Proceedings of the American Philosophical Society, 56(1):iii–viii, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983962.

Richards:1933:MFN

[Ric33] Horace G. Richards. Marine fossils from New Jersey indicating a mild interglacial stage. Proceedings of the American Philosophical Society, 72(3):181–214, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984642.

Richards:1937:LFM

[Ric37] Horace G. Richards. Land and freshwater mollusks from the Island of Cozumel, Mexico, and their bearing on the geological history of the region. Proceedings of the American Philosophical Society, 77(3):249– 262, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984730.

Richards:1938:LMI

[Ric38] Horace G. Richards. Land mollusks from the island of Roatan, Honduras. Proceedings of the American Philosophical Society, 79(2):167– 178, June 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984618.

Richards:1939:LMC

[Ric39] Horace G. Richards. Land mollusks from Corn Island, Nicaragua. Proceedings of the American Philosophical Society, 81(1):29–36, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984964.

Richards:1942:SEA

[Ric42] Horace C. Richards. Some early American physicists. Proceedings of the American Philosophical Society, 86(1):22–28, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985074.

Riddle:1924:SRL

[Rid24] Oscar Riddle. Sex in the right and left sides of the bird’s body. Proceedings of the American Philosophical Society, 63(1):152–161, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984446.

Riddle:1927:MCB

[Rid27] Oscar Riddle. Metabolic changes in the body of female pigeons at ovulation. Proceedings of the American Philosophical Society, 66(??):497– 509, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301080.

Riddle:1929:SCD

[Rid29] Oscar Riddle. The special contribution of developmental mechanics to the thought and purpose of the man of tomorrow. Proceedings of the American Philosophical Society, 68(2):107–117, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984311.

Riddle:1935:AIH

[Rid35] Oscar Riddle. Aspects and implications of the hormonal control of the maternal instinct. Proceedings of the American Philosophical Society, 75(6):521–525, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984613.

Ridenour:1948:NSH

[Rid48] Louis N. Ridenour. The natural sciences and human relations. Proceedings of the American Philosophical Society, 92(5):351–355, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143045.

Riesman:1930:GM

[Rie30] David Riesman. Glozel: A mystery. Proceedings of the American Philosophical Society, 69(1):193–202, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984355.

Ritchie:1943:RAN

[Rit43] William A. Ritchie. Recent advances in New York State and the Northeast. Proceedings of the American Philosophical Society, 86(2):313–314, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985109.

Rivers:1947:ED

[Riv47] Thomas M. Rivers. Epidemic diseases. Proceedings of the American Philosophical Society, 91(1):88–94, February 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143129.

Rugh:1943:EHP

[RM43] Roberts Rugh and Douglas A. Marsland. The effect of hydrostatic pressure upon the early development of the frog’s egg “(Rana Pipiens)”: I. Macroscopic observations. Proceedings of the American Philosophical Society, 86(3):459–466, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985031.

Robinson:1911:NH

[Rob11] James Harvey Robinson. The new history. Proceedings of the American Philosophical Society, 50(199):179–190, May/August 1911. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984033.

Robinson:1940:NMB

[Rob40] David M. Robinson. A new marble bust of Menander, wrongly called Vergil. Proceedings of the American Philosophical Society, 83(3):465–477, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985114.

Robertson:1949:PSR

[Rob49a] H. P. Robertson. On the present state of relativistic cosmology. Proceedings of the American Philosophical Society, 93(7):527–531, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143141.

Robinson:1949:NCM

[Rob49b] H. W. Robinson. A note on Charles Mason’s ancestry and his family. Proceedings of the American Philosophical Society, 93(2):134–136, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143432.

Roddy:1915:CSF

[Rod15] H. Justin Roddy. Concretions in streams formed by the agency of blue green algæ and related plants. Proceedings of the American Philosophical Society, 54(218):246–258, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983937.

Roelker:1948:FGC

[Roe48] William Greene Roelker. The Franklin–Greene correspondence. Proceedings of the American Philosophical Society, 92(3):186–193, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143586.

Rogers:1910:NSP

[Rog10] Austin F. Rogers. Notes on some pseudomorphs, petrifactions and alterations. Proceedings of the American Philosophical Society, 49 (194):17–23, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983987.

Rogers:1912:VLR

[Rog12] Austin F. Rogers. The validity of the law of rational indices, and the analogy between the fundamental laws of chemistry and crystallography. Proceedings of the American Philosophical Society, 51 (204):103–117, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984098.

Rogers:1913:NM

[Rog13] Austin F. Rogers. The nomenclature of minerals. Proceedings of the American Philosophical Society, 52(212):606–615, November/ December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984147.

Rolfe:1913:SHB

[Rol13] John C. Rolfe. Suetonius and his biographies. Proceedings of the American Philosophical Society, 52(209):206–225, April 1913. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983867.

Rolfe:1922:UDI

[Rol22] John C. Rolfe. The use of devices for indicating vowel length in latin. Proceedings of the American Philosophical Society, 61(1):80–98, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984435.

Rolfe:1930:VAT

[Rol30] John C. Rolfe. Vergil after two thousand years. Proceedings of the American Philosophical Society, 69(1):347–357, 1930. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984368.

Ronne:1945:MSS

[Ron45] Finn Ronne. The main southern sledge journey from East Base, Palmer Land, Antarctica. Proceedings of the American Philosophical Society, 89(1):13–22, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985312.

Rosengarten:1900:AHG

[Ros00] J. G. Rosengarten. American history from German Archives. Proceedings of the American Philosophical Society, 39(162):129–154, April 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983661.

Rosengarten:1901:FB

[Ros01] J. G. Rosengarten. Franklin’s bagatelles. Proceedings of the American Philosophical Society, 40(166):87–135, July 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983574.

Rosengarten:1903:FPA

[Ros03a] J. G. Rosengarten. The “Franklin Papers” in the American Philosophical Society. Proceedings of the American Philosophical Society, 42(173):165–170, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983729.

Rosengarten:1903:ECM

[Ros03b] Joseph G. Rosengarten. The Earl of Crawford’s Ms. history in the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 42(174):397–404, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983722.

Rosengarten:1907:EFM

[Ros07] J. G. Rosengarten. The early French members of the American Philosophical Society. Proceedings of the American Philosophical Society, 46(185):87–93, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983442.

Rosengarten:1911:MSM

[Ros11] Joseph G. Rosengarten. Moreau de Saint Mery and his French friends in the American Philosophical Society. Proceedings of the American Philosophical Society, 50(199):168–178, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984032.

Rosengarten:1913:CPP

[Ros13] J. G. Rosengarten. A counsel of perfection: A plan for an automatic collection and distribution of a state tax for higher education.

Proceedings of the American Philosophical Society, 52(209):243–256, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983870.

Rosendahl:1928:LTA

[Ros28] C. E. Rosendahl. Lighter-than-air machines. Proceedings of the American Philosophical Society, 67(4):319–334, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984247.

Rosenberg:1940:COH

[Ros40] Lauren E. Rosenberg. Conjugation in “Opisthonecta Henneguyi” a free swimming vorticellid. Proceedings of the American Philosophical Society, 82(4):437–448, May 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984880.

Rose:1947:RAA

[Ros47a] William C. Rose. The role of the amino acids in human nutrition. Proceedings of the American Philosophical Society, 91(1):112–116, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143131.

Rosenwald:1947:WIW

[Ros47b] Lessing J. Rosenwald. “What is it we have got to solve? Are the Jews a state or are they a religion?”. Proceedings of the American Philosophical Society, 91(3):274–280, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143399.

Rous:1912:ATR

[Rou12] Peyton Rous. An avian tumor in its relation to the tumor problem. Proceedings of the American Philosophical Society, 51(205):201–205, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984137.

Rowe:1914:USF

[Row14] L. S. Rowe. The United States as a factor in world politics. Proceedings of the American Philosophical Society, 53(213):78–91, January/ May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984131.

Rowe:1916:ADN

[Row16] L. S. Rowe. America as the defender of neutral rights. Proceedings of the American Philosophical Society, 55(3):259–266, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984116.

Rowe:1929:CEB

[Row29] Alan Rowe. A comparison of Egyptian and Babylonian civilizations and their influence on Palestine. Proceedings of the American Philosophical Society, 68(4):313–319, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984343.

Rowntree:1942:HRP

[Row42] Leonard G. Rowntree. The health of registrants and the President’s plan of rehabilitation. Proceedings of the American Philosophical Society, 85(4):343–348, June 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984989.

Royce:1906:PSQ

[Roy06] Josiah Royce. The present state of the question regarding the first principles of theoretical science. Proceedings of the American Philosophical Society, 45(182):82–102, January/April 1906. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983682.

Rommel:1906:IFL

[RP06] Geo. M. Rommel and E. F. Phillips. Inheritance in the female line of size of litter in Poland China sows. Proceedings of the American Philosophical Society, 45(184):245–254, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983377.

Riggs:1935:DSN

[RP35] Elmer S. Riggs and Bryan Patterson. Description of some notoungulates from the Casamayor (“Notostylops”) beds of Patagonia. Proceedings of the American Philosophical Society, 75(2):163–215, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984534.

Rude:1938:NMM

[Rud38] G. T. Rudé. New methods of marine surveying. Proceedings of the American Philosophical Society, 79(1):9–25, April 1938. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984675.

Rugh:1939:DER

[Rug39] Roberts Rugh. Developmental effects resulting from exposure to Xrays: I. Effect on the embryo of irradiation of frog sperm. Proceedings of the American Philosophical Society, 81(3):447–471, August 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984864.

Rugh:1941:ESR

[Rug41] Roberts Rugh. Experimental studies on the reproductive physiology of the male spring peeper, Hyla Crucifer. Proceedings of the American Philosophical Society, 84(5):617–632, July 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984843.

Rumbold:1916:PAI

[Rum16] Caroline Rumbold. Pathological anatomy of the injected trunks of chestnut trees. Proceedings of the American Philosophical Society, 55 (6):485–493, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983983.

Russell:1910:DRS

[Rus10] Henry Norris Russell. On the distances of red stars. (abstract). Proceedings of the American Philosophical Society, 49(195):230–231, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983900.

Russell:1912:RBS

[Rus12] Henry Norris Russell. Relations between the spectra and other characteristics of the stars. Proceedings of the American Philosophical Society, 51(207):569–579, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984021.

Russell:1930:UCD

[Rus30] Loris S. Russell. Upper Cretaceous dinosaur faunas of North America. Proceedings of the American Philosophical Society, 69(1):133– 159, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984352.

Russell:1939:SE

[Rus39] Henry Norris Russell. Stellar energy. Proceedings of the American Philosophical Society, 81(2):295–307, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984987.

Rusk:1946:CR

[Rus46] Howard A. Rusk. Convalescence and rehabilitation. Proceedings of the American Philosophical Society, 90(4):271–274, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301020.

Russell:1947:ARD

[Rus47] Henry Norris Russell. America’s role in the development of astronomy. Proceedings of the American Philosophical Society, 91(1):10–16, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143119.

Rosengarten:1907:BNA

[RW07] Joseph G. Rosengarten and W. W. Biographical notice of Albert Henry Smyth. Proceedings of the American Philosophical Society, 46 (187):i–x, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983483.

Reed:1909:VCL

[RW09] Hugh D. Reed and Albert H. Wright. The vertebrates of the Cayuga Lake Basin, N.Y. Proceedings of the American Philosophical Society, 48(193):370–459, September/December 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984061.

Sanders:1935:SFO

[San35] Henry A. Sanders. Some fragments of the oldest Beatty papyrus in the Michigan Collection. Proceedings of the American Philosophical Society, 75(4):313–324, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984632.

Sanders:1939:SMC

[San39] Henry A. Sanders. A soldier’s marriage certificate in diploma form. Proceedings of the American Philosophical Society, 81(4):581–590, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984829.

Sarton:1944:LP

[Sar44] George Sarton. Lagrange’s personality (1736–1813). Proceedings of the American Philosophical Society, 88(6):457–496, December 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985239.

Sauveur:1927:WSI

[Sau27] Albert Sauveur. The Widmanstättian structure of iron alloys. Proceedings of the American Philosophical Society, 66(??):267–280, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301068.

Saunders:1940:FAB

[Sau40] Harold E. Saunders. The flight of Admiral Byrd to the South Pole and the exploration of Marie Byrd Land. Proceedings of the American Philosophical Society, 82(5):801–820, June 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984892.

Sauer:1948:ECD

[Sau48] Carl O. Sauer. Environment and culture during the last deglaciation. Proceedings of the American Philosophical Society, 92(1):65–77, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143629.

Saul:1949:PP

[Sau49] Leon J. Saul. Preventive psychiatry. Proceedings of the American Philosophical Society, 93(4):330–334, September 1949. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143159.

Schlesinger:1913:POR

[Sch13a] Frank Schlesinger. On the prospect of obtaining radial velocities by means of the objective prism. Proceedings of the American Philosophical Society, 52(209):175–183, April 1913. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983864.

Schreiner:1913:ENT

[Sch13b] Oswald Schreiner. Elimination and neutralization of toxic soil substances. Proceedings of the American Philosophical Society, 52 (210):420–430, May/August 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984086.

Schlesinger:1914:UPD

[Sch14] Frank Schlesinger. The use of a photographic doublet in cataloguing the positions of stars. Proceedings of the American Philosophical Society, 53(215):303–306, August/December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984079.

Schlesinger:1915:VLT

[Sch15a] Frank Schlesinger. Variations of latitude: Their bearing upon our knowledge of the interior of the Earth. Proceedings of the American Philosophical Society, 54(220):351–358, October/December 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984108.

Schuchert:1915:CBS

[Sch15b] Charles Schuchert. The conditions of black shale deposition as illustrated by the Kupferschiefer and Lias of Germany. Proceedings of the American Philosophical Society, 54(218):259–269, August 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983938.

Schelling:1916:CFS

[Sch16] Felix E. Schelling. The common folk of Shakespeare. Proceedings of the American Philosophical Society, 55(6):471–480, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983981.

Schlesinger:1917:DNP

[Sch17] Frank Schlesinger. A description of a new photographic transit instrument. Proceedings of the American Philosophical Society, 56(7): 484–486, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983885.

Schinz:1929:CNP

[Sch29] Albert Schinz. The concept of nature in philosophy and literature; a consideration of recent discussions. Proceedings of the American Philosophical Society, 68(3):207–225, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984209.

Schelling:1931:RSO

[Sch31] Felix E. Schelling. The return to Shakespearean orthodoxy. Proceedings of the American Philosophical Society, 70(4):353–361, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984200.

Schramm:1935:PSP

[Sch35] J. R. Schramm. Publication as a scientific problem. Proceedings of the American Philosophical Society, 75(6):527–536, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984614.

Schlesinger:1936:EBF

[Sch36] Frank Schlesinger. Edwin Brant Frost, 1866–1935. Proceedings of the American Philosophical Society, 76(2):241–244, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984669.

Schramm:1939:CAS

[Sch39] J. R. Schramm. Cost analysis of scholarly periodical printing. Preliminary report of the committee on abstracting and documentation of scientific literature of the National Research Council. Proceedings of the American Philosophical Society, 80(1):1–24, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984969.

Schneirla:1944:SAA

[Sch44a] T. C. Schneirla. Studies on the army-ant behavior pattern. — nomadism in the swarm-raider “Eciton burchelli”. Proceedings of the American Philosophical Society, 87(5):438–457, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985298.

Schuyler:1944:WH

[Sch44b] Robert Livingston Schuyler. War and historiography. Proceedings of the American Philosophical Society, 87(4):342–351, January 1944.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985046.

Schmitt:1945:MZM

[Sch45a] Waldo L. Schmitt. Miscellaneous zoological material collected by the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):297, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985337.

Schultz:1945:FUS

[Sch45b] Leonard P. Schultz. Fishes of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):298, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985338.

Schuyler:1946:ASH

[Sch46] Robert Livingston Schuyler. The antiquaries and Sir Henry Spelman: An essay in historiography. Proceedings of the American Philosophical Society, 90(2):91–103, May 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301050.

Schuyler:1948:MGI

[Sch48] Robert Livingston Schuyler. Man’s greatest illusion. Proceedings of the American Philosophical Society, 92(1):46–50, March 1948. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143626.

Schmitt:1949:S

[Sch49] Bernadotte E. Schmitt. 1848: As seen from 1948. Proceedings of the American Philosophical Society, 93(3):216–221, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143468.

Scott:1923:HS

[Sco23] William B. Scott. Historical sketch. Proceedings of the American Philosophical Society, 62(5):270–273, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984460.

Scott:1927:DAP

[Sco27] William Berryman Scott. Development of American palæontology. Proceedings of the American Philosophical Society, 66(??):409–429, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301076.

Scott:1928:RPU

[Sco28] William Berryman Scott. Reports of the Princeton University Expeditions to Patagonia, 1896–1899. Proceedings of the American Philosophical Society, 67(2):119–122, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984273.

Scott:1930:CSF

[Sco30] William Berryman Scott. On certain Silurian fish-like forms from Norway. Proceedings of the American Philosophical Society, 69(1): 391–396, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984373.

Scott:1932:FPP

[Sco32] James Brown Scott. Franklin: Political philosopher. Proceedings of the American Philosophical Society, 71(4):217–223, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984792.

Scott:1937:A

[Sco37] William Berryman Scott. The astrapotheria. Proceedings of the American Philosophical Society, 77(3):309–393, March 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984733.

Smith:1904:AWT

[SE04] Edgar F. Smith and Franz F. Exner. The atomic weight of tungsten. Proceedings of the American Philosophical Society, 43(176):123–148, April 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983497.

Speck:1942:MNB

[SE42] Frank G. Speck and Loren C. Eiseley. Montagnais–Naskapi bands and family hunting districts of the Central and Southeastern Labrador Peninsula. Proceedings of the American Philosophical Society, 85 (2):215–242, January 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985039.

Sears:1948:PDC

[Sea48] Robert R. Sears. Personality development in contemporary culture. Proceedings of the American Philosophical Society, 92(5):363–370, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143047.

See:1906:CEM

[See06] T. J. J. See. The cause of earthquakes, mountain formation and kindred phenomena connected with the physics of the Earth. Proceedings of the American Philosophical Society, 45(184):274–414, October/ December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983382.

See:1907:NTE

[See07a] T. J. J. See. The new theory of earthquakes and mountain formation, as illustrated by processes now at work in the depths of the sea. Proceedings of the American Philosophical Society, 46(187): 369–415, October/December 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983480.

See:1907:TSC

[See07b] T. J. J. See. On the temperature, secular cooling and contraction of the Earth, and on the theory of earthquakes held by the ancients. Proceedings of the American Philosophical Society, 46(186): 191–299, April/September 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983739.

See:1908:FRP

[See08] T. J. J. See. Further researches on the physics of the Earth, and especially on the folding of mountain ranges and the uplift of plateaus and continents produced by movements of lava beneath the crust arising from the secular leakage of the ocean bottoms. Proceedings of the American Philosophical Society, 47(189):157–275, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983843.

See:1909:PHE

[See09] T. J. J. See. The past history of the Earth as inferred from the mode of formation of the solar system. Proceedings of the American Philosophical Society, 48(191):119–128, January/April 1909. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983817.

See:1910:EPA

[See10a] T. J. J. See. The existence of planets about the fixed stars. Proceedings of the American Philosophical Society, 49(195):222–229, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983899.

See:1910:FCO

[See10b] T. J. J. See. Further considerations on the origin of the zone of asteroids and on the capture of satellites. Proceedings of the American Philosophical Society, 49(197):351–358, October/December 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984089.

See:1910:RRR

[See10c] T. J. J. See. Results of recent researches in cosmical evolution. Proceedings of the American Philosophical Society, 49(195):207–221, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983898.

See:1911:ESS

[See11a] T. J. J. See. The extension of the solar system beyond Neptune, and the connection existing between planets and comets. Proceedings of the American Philosophical Society, 50(199):266–268, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984042.

See:1911:NC

[See11b] T. J. J. See. The new cosmogony. Proceedings of the American Philosophical Society, 50(199):261–265, May/August 1911. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984041.

See:1911:SEI

[See11c] T. J. J. See. The secular effects of the increase of the Sun’s mass upon the mean motions, major axes and eccentricities of the orbits of the planets. Proceedings of the American Philosophical Society, 50 (199):269–273, May/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984043.

See:1912:DDM

[See12a] T. J. J. See. Determination of the depth of the Milky Way. Proceedings of the American Philosophical Society, 51(203):1–17, January/ March 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983977.

See:1912:DTG

[See12b] T. J. J. See. Dynamical theory of the globular clusters and of the clustering power inferred by Herschel from the observed figures of sidereal systems of high order. Proceedings of the American Philosophical Society, 51(204):118–166, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984099.

See:1913:FCO

[See13] T. J. J. See. Further considerations on the origin of the Himalaya Mountains and the Plateau of Tibet. Proceedings of the American Philosophical Society, 52(211):495–515, September/October 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983973.

See:1915:ELT

[See15] T. J. J. See. The Euler–Laplace Theorem on the decrease of the eccentricity of the orbits of the heavenly bodies under the secular action of a resisting medium. Proceedings of the American Philosophical Society, 54(219):344–350, September 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983878.

Seiwell:1940:TVH

[Sei40] H. R. Seiwell. Time variability of hydrographic elements determining the dynamic situation in the Western North Atlantic. Proceedings of the American Philosophical Society, 82(3):369–394, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984936.

Seiwell:1942:AVO

[Sei42] H. R. Seiwell. An analysis of vertical oscillations in the Southern North Atlantic. Proceedings of the American Philosophical Society, 85(2):136–158, January 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985035.

Seligman:1944:PWP

[Sel44] Eustace Seligman. A post-war program for the taxation of corporations and their stockholders. Proceedings of the American Philosophical Society, 88(1):60–67, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985191.

Sellers:1948:CWP

[Sel48] Charles Coleman Sellers. Charles Willson Peale’s career as a painter. Proceedings of the American Philosophical Society, 92(2):105–106, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143407.

Senstius:1930:SWS

[Sen30] M. W. Senstius. Studies on weathering and soil formation in tropical high altitudes. Proceedings of the American Philosophical Society, 69 (1):45–97, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984349.

Setchell:1918:PAR

[Set18] William Albert Setchell. Parasitism among the red algæ. Proceedings of the American Philosophical Society, 57(2):155–172, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984166.

Setchell:1926:NVC

[Set26] William Albert Setchell. Nullipore versus coral in reef-formation. Proceedings of the American Philosophical Society, 65(2):136–140, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984421.

Shapley:1919:SCT

[Sha19] Harlow Shapley. Star clusters and their contribution to knowledge of the universe. Proceedings of the American Philosophical Society, 58 (6):337–345, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984259.

Shapley:1930:TUM

[Sha30] Harlow Shapley. Ten unsolved mysteries of the stellar universe. Proceedings of the American Philosophical Society, 69(1):13–17, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984347.

Shaaber:1935:FVS

[Sha35] M. A. Shaaber. The Furness Variorum Shakespeare. Proceedings of the American Philosophical Society, 75(4):281–285, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984629.

Shapiro:1939:SRA

[Sha39] H. L. Shapiro. Some reflections of an anthropologist on the future of our population. Proceedings of the American Philosophical Society, 80(4):587–600, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985059.

Shapley:1943:NTJ

[Sha43] Harlow Shapley. Notes on Thomas Jefferson as a natural philosopher. Proceedings of the American Philosophical Society, 87(3):234– 237, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984870.

Shapley:1947:AIC

[Sha47] Harlow Shapley. Astronomy and international cooperation. Proceedings of the American Philosophical Society, 91(1):73–74, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143125.

Shapley:1949:SCL

[Sha49] Harlow Shapley. The standard Cepheid light curve. Proceedings of the American Philosophical Society, 93(1):40–44, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143269.

Sherman:1918:FCS

[She18] H. C. Sherman. Food conservation from the standpoint of the chemistry of nutrition. Proceedings of the American Philosophical Society, 57(5):491–500, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984183.

Shirley:1945:PAP

[Shi45] Charles C. Shirley. Photographic accomplishments and photographic technique at West Base, Antarctica. Proceedings of the American Philosophical Society, 89(1):382–385, April 1945. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985347.

Shorey:1927:PHS

[Sho27] Paul Shorey. Platonism and the history of science. Proceedings of the American Philosophical Society, 66(??):159–182, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301062.

Shotwell:1938:IP

[Sho38] James T. Shotwell. International peace. Proceedings of the American Philosophical Society, 78(4):545–550, March 1938. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984801.

Shoemaker:1945:AUS

[Sho45] Clarence R. Shoemaker. Amphipoda of the United States Antarctic Service Expedition 1939–1941. Proceedings of the American Philosophical Society, 89(1):289–293, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985331.

Shotwell:1946:CAE

[Sho46] James T. Shotwell. The control of atomic energy under the charter. Proceedings of the American Philosophical Society, 90(1):59–64, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301040.

Shope:1948:UMD

[Sho48] Richard E. Shope. An unfamiliar mechanism of disease transmission. Proceedings of the American Philosophical Society, 92(4):289–293, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143358.

Shryock:1943:AHC

[Shr43] Richard H. Shryock. American historiography: A critical analysis and a program. Proceedings of the American Philosophical Society, 87(1): 35–46, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984997.

Shryock:1946:HAP

[Shr46] Richard Harrison Shryock. The health of the American people: An historical survey. Proceedings of the American Philosophical Society, 90(4):251–258, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301017.

Shryock:1947:TAM

[Shr47] Richard Harrison Shryock. Trends in American medical research during the Nineteenth Century. Proceedings of the American Philosophical Society, 91(1):58–63, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143123.

Shufeldt:1900:OSS

[Shu00a] R. W. Shufeldt. On the osteology of the striges. [Strigidæ and Bubonidæ]. Proceedings of the American Philosophical Society, 39 (164):665–722, October/December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983784.

Shufeldt:1900:OW

[Shu00b] R. W. Shufeldt. On the osteology of the woodpeckers. Proceedings of the American Philosophical Society, 39(164):578–622, October/ December 1900. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983777.

Shufeldt:1901:OCC

[Shu01] R. W. Shufeldt. The osteology of the cuckoos: [coccyges]. Proceedings of the American Philosophical Society, 40(165):4–51, January 1901. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983422.

Shull:1910:GAT

[Shu10] George Harrison Shull. Germinal analysis through hybridization. Proceedings of the American Philosophical Society, 49(196):281– 290, August/September 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984120.

Shull:1942:NMS

[Shu42] George Harrison Shull. New mutational segregations from “Oenothera” mut. “Erythrina” de Vries. Proceedings of the American Philosophical Society, 85(2):183–214, January 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985038.

Shultz:1944:PWS

[Shu44] William J. Shultz. Post-war social orientation of state tax systems. Proceedings of the American Philosophical Society, 88(1):55–59, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985190.

Sigerist:1946:PPM

[Sig46] Henry E. Sigerist. The place of the physician in modern society. Proceedings of the American Philosophical Society, 90(4):275–279, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301021.

Simpson:1940:RMB

[Sim40] George Gaylord Simpson. Review of the mammal-bearing tertiary of South America. Proceedings of the American Philosophical Society, 83(5):649–709, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985134.

Simpson:1942:BVP

[Sim42] George Gaylord Simpson. The beginnings of vertebrate paleontology in North America. Proceedings of the American Philosophical Society, 86(1):130–188, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985085.

Simpson:1943:NMB

[Sim43] George Gaylord Simpson. Notes on the mammal-bearing tertiary of South America. Proceedings of the American Philosophical Society, 86(3):403–404, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985028.

Sinclair:1905:MFS

[Sin05] Wm. J. Sinclair. The marsupial fauna of the Santa Cruz beds. Proceedings of the American Philosophical Society, 44(179):73–81, January/April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983608.

Sinclair:1908:SCT

[Sin08] William J. Sinclair. The Santa Cruz Typotheria. Proceedings of the American Philosophical Society, 47(188):64–78, January/April 1908.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983799.

Sinclair:1910:DBP

[Sin10a] William J. Sinclair. Dermal bones of Paramylodon from the asphaltum deposits of Rancho la Brea, near Los Angeles, California. Proceedings of the American Philosophical Society, 49(195):191–195, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983894.

Sinclair:1910:RSL

[Sin10b] William J. Sinclair. The restored skeleton of Leptauchenia Decora. Proceedings of the American Philosophical Society, 49(195):196–199, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983895.

Sinclair:1915:AFL

[Sin15] William J. Sinclair. Additions to the fauna of the Lower Pliocene Snake Creek Beds (results of the Princeton University 1914 Expedition to Nebraska). Proceedings of the American Philosophical Society, 54(217):73–95, May/July 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983916.

Sinclair:1921:EBB

[Sin21a] William J. Sinclair. Entelodonts from the Big Badlands of South Dakota in the Geological Museum of Princeton University. Investigation aided by a grant from the Marsh Fund of the National Academy of Sciences. Proceedings of the American Philosophical Society, 60(4): 467–495, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984514.

Sinclair:1921:NHT

[Sin21b] William J. Sinclair. A new Hoplophoneus from the Titanotherium beds. Investigation aided by a grant from the Marsh Fund of the National Academy of Sciences. Proceedings of the American Philosophical Society, 60(2):96–98, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984386.

Sinclair:1921:TOL

[Sin21c] William J. Sinclair. The “turtle-oreodon layer” or “red layer,” a contribution to the stratigraphy of the White River Oligocene (results of the Princeton University 1920 Expedition to South Dakota). Investigation aided by a grant from the Marsh Fund of the National Academy of Sciences. Proceedings of the American Philosophical Society, 60(3):456–466, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984423.

Sinclair:1922:HBB

[Sin22a] William J. Sinclair. Hyracodons from the Big Badlands of South Dakota. Proceedings of the American Philosophical Society, 61(1): 65–79, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984434.

Sinclair:1922:SEW

[Sin22b] William J. Sinclair. The small entelodonts of the White River Oligocene. Proceedings of the American Philosophical Society, 61(1): 53–64, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984433.

Sinclair:1924:FCZ

[Sin24] William J. Sinclair. The faunas of the concretionary zones of the Oreodon Beds, White River Oligocene. Proceedings of the American Philosophical Society, 63(1):94–133, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984444.

Sinclair:1925:MSN

[Sin25a] Wm. J. Sinclair. The mounted skeleton of a new Mesohippus from the Protoceras beds. Proceedings of the American Philosophical Society, 64(1):55–63, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984303.

Singer:1925:LRL

[Sin25b] Edgar A. Singer, Jr. Logic and the relation of life to mechanism. Proceedings of the American Philosophical Society, 64(2):95–101, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984191.

Sinclair:1928:ONF

[Sin28] William J. Sinclair. Omorhamphus, a new flightless bird from the Lower Eocene of Wyoming. Proceedings of the American Philosophical Society, 67(1):51–65, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984256.

Sioussat:1949:PTR

[Sio49] George L. Sioussat. The “Philosophical Transactions” of the Royal Society in the libraries of William Byrd of Westover, Benjamin Franklin, and the American Philosophical Society. Proceedings of the American Philosophical Society, 93(2):99–113, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143428.

Siple:1945:GPG

[Sip45a] Paul A. Siple. General principles governing selection of clothing for cold climates. Proceedings of the American Philosophical Society, 89(1):200–234, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985325.

Siple:1945:GEL

[Sip45b] Paul A. Siple. Geographical exploration from Little America III, the West Base of the United States Antarctic Service Expedition 1939– 1941. Proceedings of the American Philosophical Society, 89(1):23–60, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985313.

Sitterly:1939:CS

[Sit39] Charlotte M. Sitterly. The composition of the Sun. Proceedings of the American Philosophical Society, 81(2):125–133, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984979.

Sinclair:1927:SE

[SJ27] William J. Sinclair and G. L. Jepsen. The skull of Eusmilus. Proceedings of the American Philosophical Society, 66(??):391–407, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301075.

Sinclair:1929:MSP

[SJ29] William J. Sinclair and Glenn L. Jepsen. A mounted skeleton of Palaeonictis. Proceedings of the American Philosophical Society, 68 (3):163–173, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984205.

Schenck:1936:MMP

[SK36] Hubert G. Schenck and A. Myra Keen. Marine molluscan provinces of Western North America. Proceedings of the American Philosophical Society, 76(6):921–938, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984576.

Schenck:1937:IMC

[SK37] Hubert G. Schenck and A. Myra Keen. An index-method for comparing molluscan faunules. Proceedings of the American Philosophical Society, 77(2):161–182, February 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984725.

Sitterly:1943:TS

[SK43] Charlotte M. Sitterly and Arthur S. King. Thorium in the Sun. Proceedings of the American Philosophical Society, 86(3):339–341, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985023.

Skard:1946:UCL

[Ska46] Sigmund Skard. The use of color in literature: A survey of research. Proceedings of the American Philosophical Society, 90(3):163–249, July 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301043.

Slipher:1917:N

[Sli17] V. M. Slipher. Nebulæ. Proceedings of the American Philosophical Society, 56(5):403–409, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984028.

Slipher:1938:POL

[Sli38a] E. C. Slipher. The planets from observations at the Lowell Observatory. Proceedings of the American Philosophical Society, 79 (3):441–473, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984860.

Slipher:1938:TNP

[Sli38b] V. M. Slipher. The trans Neptunian planet search. Proceedings of the American Philosophical Society, 79(3):435–440, September 1938.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984859.

Slichter:1948:SBI

[Sli48] Sumner H. Slichter. Some basic issues raised by trade unions. Proceedings of the American Philosophical Society, 92(3):203–209, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143588.

Spoehr:1939:SDA

[SM39] H. A. Spoehr and Harold W. Milner. Starch dissolution and amylolytic activity in leaves. Proceedings of the American Philosophical Society, 81(1):37–78, May 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984965.

Smith:1903:EPA

[Smi03] Edgar F. Smith. Extract from the President’s address. Proceedings of the American Philosophical Society, 42(174):vii–xvii, May/December 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983724.

Smith:1905:OCT

[Smi05] Edgar F. Smith. Observations on columbium and tantalum. Proceedings of the American Philosophical Society, 44(180):151–158, May/ July 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983511.

Smith:1906:PMG

[Smi06] James Perrin Smith. The paragenesis of the minerals in the glaucophane-bearing rocks of California. Proceedings of the American Philosophical Society, 45(184):183–242, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983374.

Smith:1917:MOP

[Smi17] Erwin F. Smith. Mechanism of overgrowth in plants. Proceedings of the American Philosophical Society, 56(6):437–444, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983967.

Smith:1918:SEA

[Smi18] J. Russell Smith. Some economic aspects of the American food supply. Proceedings of the American Philosophical Society, 57(5):501– 512, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984184.

Smith:1926:BOB

[Smi26a] Edgar F. Smith. Behavior of organic bases toward soluble tungstates. Proceedings of the American Philosophical Society, 65(5):10–17, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984283.

Smith:1926:IBS

[Smi26b] Edgar F. Smith. An interesting behavior of soluble meta-tungstates. Proceedings of the American Philosophical Society, 65(5):18–19, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984284.

Smith:1934:PCS

[Smi34] Frank H. Smith. Prochromosomes and chromosome structure in “Impatiens”. Proceedings of the American Philosophical Society, 74(3):193–214, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984714.

Smith:1938:WCP

[Smi38] Marian W. Smith. The war complex of the Plains Indians. Proceedings of the American Philosophical Society, 78(3):425–464, January 31, 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984763.

Smyth:1913:WWG

[Smy13] Herbert Weir Smyth. William Watson Goodwin. Proceedings of the American Philosophical Society, 52(212):iii–ix, November/December 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984149.

Smyth:1927:GAR

[Smy27] C. H. Smyth, Jr. The genesis of alkaline rocks. Proceedings of the American Philosophical Society, 66(??):535–580, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3301082.

Smyth:1936:MRS

[Smy36] Charles P. Smyth. Molecular rotation in solids. Proceedings of the American Philosophical Society, 76(4):485–489, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984598.

Smyth:1946:FYA

[Smy46] H. D. Smyth. Fifty years of atomic physics. Proceedings of the American Philosophical Society, 90(1):1–6, January 1946. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3301031.

Shapley:1948:CVS

[SN48] Harlow Shapley and Virginia McKibben Nail. Cepheid variable stars in the Small Magellanic Cloud. Proceedings of the American Philosophical Society, 92(4):310–323, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143361.

Snyder:1902:NMT

[Sny02] Monroe B. Snyder. A new method of transiting stars. Proceedings of the American Philosophical Society, 41(169):200–206, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983538.

Snyder:1926:PSC

[Sny26a] Monroe B. Snyder. Prediction of the spectrum of the cosmic rays. Proceedings of the American Philosophical Society, 65(4):379–380, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984242.

Snyder:1926:UAVa

[Sny26b] Monroe B. Snyder. Universal atomic volcanism and the Millikan cosmic rays. Proceedings of the American Philosophical Society, 65 (3):161–169, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984486.

Snyder:1926:UAVb

[Sny26c] Monroe B. Snyder. Universal atomic volcanism and the ultimate atom. Proceedings of the American Philosophical Society, 65(3):170– 182, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984487.

Sanders:1937:TAM

[SO37] Henry A. Sanders and Johanna Ogden. The text of Acts in MS 146 of the University of Michigan. Proceedings of the American Philosophical Society, 77(1):1–97, 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984768.

Sonneborn:1938:MTP

[Son38] T. M. Sonneborn. Mating types in Paramecium Aurelia: Diverse conditions for mating in different stocks; occurrence, number and interrelations of the types. Proceedings of the American Philosophical Society, 79(3):411–434, September 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984858.

Soper:1916:GPR

[Sop16] Ralph H. Soper. The geology of Parahyba and Rio Grande Do Norte, Brazil. Proceedings of the American Philosophical Society, 55(1):1– 20, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984170.

Shearin:1942:EEC

[SP42] Paul E. Shearin and T. Eugene Pardue. Electron-electron collisions in the primary energy range from 1.3 to 2.6 million electron volts. Proceedings of the American Philosophical Society, 85(3):243–249, February 25, 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985005.

Siple:1945:MDA

[SP45] Paul A. Siple and Charles F. Passel. Measurements of dry atmospheric cooling in subfreezing temperatures. Proceedings of the American Philosophical Society, 89(1):177–199, April 1945. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985324.

Skogsberg:1946:HMB

[SP46] Tage Skogsberg and Austin Phelps. Hydrography of Monterey Bay, California. Thermal conditions, Part II (1934–1937). Proceedings of the American Philosophical Society, 90(5):350–386, December 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301046.

Sparrow:1937:SCI

[Spa37] F. K. Sparrow, Jr. Some Chytridiaceous inhabitants of submerged insect Exuviæ. Proceedings of the American Philosophical Society, 78(1):23–60, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984656.

Speck:1926:CPN

[Spe26] Frank G. Speck. Culture problems in Northeastern North America. Proceedings of the American Philosophical Society, 65(4):272– 311, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984238.

Speidel:1940:SLN

[Spe40] Carl Caskey Speidel. Studies of living nerves. VI. Effects of metrazol on tissues of frog tadpoles with special reference to the injury and recovery of individual nerve fibers. Proceedings of the American Philosophical Society, 83(2):349–378, August 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985016.

Speidel:1942:SLN

[Spe42] Carl Caskey Speidel. Studies of living nerves. VIII. Histories of nerve endings in frog tadpoles subjected to various injurious treatments. Proceedings of the American Philosophical Society, 85(2):168–182, January 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985037.

Spitzka:1908:ORI

[Spi08a] Edw. Anthony Spitzka. Observations regarding the infliction of the death penalty by electricity. Proceedings of the American Philosophical Society, 47(188):39–50, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983796.

Spitzka:1908:PNB

[Spi08b] Edw. Anthony Spitzka. Preliminary note on the brains of natives of the Andaman and Nicobar Islands. Proceedings of the American Philosophical Society, 47(188):51–58, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983797.

Spiegelman:1939:MRW

[Spi39] Mortimer Spiegelman. Mortality in relation to widowhood. Proceedings of the American Philosophical Society, 80(4):541–558, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985056.

Sprout:1948:TTR

[Spr48] Harold Sprout. Trends in the traditional relation between military and civilian. Proceedings of the American Philosophical Society, 92(4):264–270, October 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143355.

Squier:1917:PAA

[Squ17] George O. Squier. Physical aspects. The air. Proceedings of the American Philosophical Society, 56(3):168–169, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983944.

Seiwell:1938:SDP

[SS38] H. R. Seiwell and Gladys Eddy Seiwell. The sinking of decomposing plankton in sea water and its relationship to oxygen consumption and phosphorus liberation. Proceedings of the American Philosophical Society, 78(3):465–481, January 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984764.

Schultz:1945:NVP

[SS45] Adolph H. Schultz and William L. Straus, Jr. The numbers of vertebrae in primates. Proceedings of the American Philosophical Society, 89(4):601–626, December 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985364.

Stanton:1903:NFW

[Sta03] Timothy W. Stanton. A new fresh-water molluscan faunule from the Cretaceous of Montana. Proceedings of the American Philosophical Society, 42(173):188–199, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983732.

Stanley:1937:SBI

[Sta37] W. M. Stanley. Some biochemical investigations on the crystalline tobacco-mosaic virus proteins. Proceedings of the American Philosophical Society, 77(4):447–453, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984770.

Stakman:1947:IPP

[Sta47] Elvin Charles Stakman. International problems in plant disease control. Proceedings of the American Philosophical Society, 91(1):95–111, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143130.

Starr:1948:WUE

[Sta48] Mark Starr. Why union education?: Aims, history, and philosophy of the educational work of the International Ladies’ Garment Workers’ Union. Proceedings of the American Philosophical Society, 92(3):194– 202, July 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143587.

Stengel:1902:SPT

[Ste02] Alfred Stengel. Specific precipitins and their medico-legal value in distinguishing human and animal blood. Proceedings of the American Philosophical Society, 41(171):407–414, April/December 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983490.

Stevenson:1911:FCBa

[Ste11a] John J. Stevenson. The formation of coal beds. I. An historical summary of opinion from 1700 to the present time. Proceedings of the American Philosophical Society, 50(198):1–116, January/April 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983954.

Stevenson:1911:FCBb

[Ste11b] John J. Stevenson. The formation of coal beds. II. Some elementary problems. Proceedings of the American Philosophical Society, 50(202): 519–643, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984001.

Stevenson:1912:FCB

[Ste12] John J. Stevenson. The formation of coal beds. III. Proceedings of the American Philosophical Society, 51(207):423–553, October/ December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984017.

Stevenson:1913:FCB

[Ste13] John J. Stevenson. The formation of coal beds. IV. Proceedings of the American Philosophical Society, 52(208):31–162, January/April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983997.

Stevenson:1916:IFF

[Ste16] John J. Stevenson. Interrelations of the fossil fuels. I. Proceedings of the American Philosophical Society, 55(2):21–203, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984026.

Stevenson:1917:IFF

[Ste17] John J. Stevenson. Interrelations of the fossil fuels. II. Proceedings of the American Philosophical Society, 56(2):53–151, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984171.

Stevenson:1918:IFF

[Ste18] John J. Stevenson. Interrelations of the fossil fuels. III. Proceedings of the American Philosophical Society, 57(1):1–48, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983880.

Stevenson:1920:IFF

[Ste20] John J. Stevenson. Interrelations of the fossil fuels. IV. Proceedings of the American Philosophical Society, 59(6):405–511, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984389.

Stewart:1930:PS

[Ste30] John Q. Stewart. The physics of a star. Proceedings of the American Philosophical Society, 69(1):419–432, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984376.

Stewart:1934:PSR

[Ste34a] Duncan Stewart, Jr. The petrography of some rocks from South Victoria Land. Proceedings of the American Philosophical Society, 74(4):307–310, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984797.

Stewart:1934:UMC

[Ste34b] Duncan Stewart, Jr. The University of Michigan collections of Antarctic rocks and minerals. Proceedings of the American Philosophical Society, 74(4):311–317, August 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984798.

Stengel:1936:TM

[Ste36a] Alfred Stengel. Thomas McCrae. Proceedings of the American Philosophical Society, 76(2):245–246, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984670.

Stetson:1936:SRP

[Ste36b] Harlan T. Stetson. Second report of progress on research in cosmicterrestrial relations. Proceedings of the American Philosophical Society, 76(6):759–765, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984565.

Stetson:1937:CFS

[Ste37] Harlan True Stetson. Correlation of frequencies of seismic disturbances with the hour angle of the Moon. Proceedings of the American Philosophical Society, 78(2):411–424, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984543.

Stetson:1938:PSP

[Ste38] H. C. Stetson. Present status of the problem of submarine canyons. Proceedings of the American Philosophical Society, 79(1):27–33, April 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984676.

Stefansson:1940:TYP

[Ste40] Vilhjalmur Stefansson. A ten-year program of Arctic study. Proceedings of the American Philosophical Society, 82(5):897–919, June 1940.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984898.

Stewart:1945:AWA

[Ste45a] Duncan Stewart, Jr. Abstracts of works on Antarctic petrography. Proceedings of the American Philosophical Society, 89(1):152–159, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985321.

Stewart:1945:PSI

[Ste45b] Duncan Stewart, Jr. The petrography of some intrusive rocks from King Edward VII and Marie Byrd Lands, Antarctica. Proceedings of the American Philosophical Society, 89(1):148–151, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985320.

Stewart:1945:PRS

[Ste45c] Duncan Stewart, Jr. Preliminary report on some intrusives of the Melchior Islands, Antarctica. Proceedings of the American Philosophical Society, 89(1):146–147, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985319.

Stewart:1947:RMI

[Ste47] Duncan Stewart, Jr. Rocks of the Melchior Islands, Antarctica. Proceedings of the American Philosophical Society, 91(3):229–233, August 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143395.

Stephan:1948:SSO

[Ste48] Frederick F. Stephan. Sampling in studies of opinions, attitudes, and consumer wants. Proceedings of the American Philosophical Society, 92(5):387–398, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143051.

Stebbins:1949:RES

[Ste49] G. Ledyard Stebbins, Jr. Reality and efficacy of selection in plants. Proceedings of the American Philosophical Society, 93(6):501–513, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143338.

Stirling:1933:JS

[Sti33] Matthew W. Stirling. Jivaro shamanism. Proceedings of the American Philosophical Society, 72(3):137–145, 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984640.

Stoney:1903:DWA

[Sto03] G. Johnstone Stoney. On the dependence of what apparently takes place in nature upon what actually occurs in the universe of real existences. Proceedings of the American Philosophical Society, 42 (173):105–142, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983726.

Stockard:1912:CED

[Sto12] Charles R. Stockard. Is the control of embryonic development a practical problem? Proceedings of the American Philosophical Society, 51(205):191–200, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984136.

Stone:1918:APB

[Sto18] Witmer Stone. Amos Peaslee Brown. Proceedings of the American Philosophical Society, 57(7):iii–xv, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984527.

Stockard:1923:EMG

[Sto23] Charles R. Stockard. Experimental modification of the germ-plasm and its bearing on the inheritance of acquired characters. Proceedings of the American Philosophical Society, 62(5):311–325, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984464.

Stone:1946:HPA

[Sto46] Robert S. Stone. Health protection activities of the Plutonium Project. Proceedings of the American Philosophical Society, 90(1):11– 19, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3301033.

Stouffer:1948:SAS

[Sto48] Samuel A. Stouffer. Studying the attitudes of soldiers. Proceedings of the American Philosophical Society, 92(5):336–340, November 1948.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143043.

Straus:1936:TAV

[Str36] William L. Straus, Jr. The thoracic and abdominal viscera of primates, with special reference to the orang-utan. Proceedings of the American Philosophical Society, 76(1):1–85, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984708.

Struve:1939:SEA

[Str39] Otto Struve. Stars with extended atmospheres. Proceedings of the American Philosophical Society, 81(2):211–251, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984983.

Struve:1942:MCC

[Str42] Otto Struve. The mystery of Cor Caroli. Proceedings of the American Philosophical Society, 85(4):349–358, June 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984990.

Strand:1943:OMC

[Str43] K. Aa. Strand. The orbital motion of 61 Cygni. Proceedings of the American Philosophical Society, 86(3):364–367, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985026.

Strecker:1944:N

[Str44] Edward A. Strecker. Neuropsychiatry. Proceedings of the American Philosophical Society, 88(3):204–207, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985415.

Sverdrup:1947:NIA

[Sve47] Harald Ulrik Sverdrup. New international aspects of oceanography. Proceedings of the American Philosophical Society, 91(1):75–78, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143126.

Stebbins:1949:LDG

[SW49] Joel Stebbins and A. E. Whitford. The light from distant galaxies. Proceedings of the American Philosophical Society, 93(1):45–47, April 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143270.

Swann:1928:A

[Swa28] W. F. G. Swann. The atom. Proceedings of the American Philosophical Society, 67(3):235–266, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984292.

Swann:1935:PML

[Swa35] W. F. G. Swann. Penrose Memorial Lecture. Is the Universe running down? Proceedings of the American Philosophical Society, 75(3):217– 249, June 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984784.

Swann:1936:CEM

[Swa36] W. F. G. Swann. Charles Elwood Mendenhall. Proceedings of the American Philosophical Society, 76(3):391–394, 1936. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984551.

Swift:1923:CSH

[Swi23] Homer F. Swift. The clinical significance of the histological lesions of rheumatic fever. Proceedings of the American Philosophical Society, 62(3):175–184, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984505.

Taeuber:1939:ACP

[Tae39] Conrad Taeuber. Agriculture and current population trends. Proceedings of the American Philosophical Society, 80(4):477–489, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985051.

Tatlock:1938:GMM

[Tat38] J. S. P. Tatlock. Geoffrey of Monmouth’s motives for writing his “Historia”. Proceedings of the American Philosophical Society, 79 (4):695–703, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984946.

Taussig:1932:D

[Tau32] Frank W. Taussig. Discussion. Proceedings of the American Philosophical Society, 71(3):125–129, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984558.

Taube:1949:NTC

[Tau49] Mortimer Taube. New tools for the control and use of research materials. Proceedings of the American Philosophical Society, 93(3):248–252, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143474.

Taylor:1926:MCR

[Tay26] Hugh S. Taylor. The mechanism of chemical reactions. Proceedings of the American Philosophical Society, 65(2):90–98, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984416.

Taylor:1939:PDC

[Tay39] Carl C. Taylor. Prospective development of cultural patterns in rural America and their possible influence on population trends. Proceedings of the American Philosophical Society, 80(4):569–585, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985058.

Taylor:1940:FA

[Tay40] Francis H. Taylor. The fine arts. Proceedings of the American Philosophical Society, 83(4):559–567, September 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984836.

Taylor:1941:LMT

[Tay41] Hugh S. Taylor. Large molecules through atomic spectacles. Proceedings of the American Philosophical Society, 85(1):1–12, November 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985121.

Taylor:1944:ODS

[Tay44] Hugh S. Taylor. The organization, direction, and support of research in the physical sciences. Proceedings of the American Philosophical Society, 87(4):299–306, January 1944. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985041.

Taylor:1948:USS

[Tay48] Hugh S. Taylor. Uses of the separated stable isotopes of oxygen. Proceedings of the American Philosophical Society, 92(1):1–5, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143619.

Taylor:1933:TPS

[TE33] Hugh S. Taylor and Henry Eyring. Theoretical and practical studies of atomic and molecular forms of the hydrogen isotopes. Proceedings of the American Philosophical Society, 72(4):255–264, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984690.

Thomson:1907:PIC

[Tho07] Elihu Thomson. The progress of the Isthmian Canal. Proceedings of the American Philosophical Society, 46(185):124–137, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983446.

Thomson:1911:GFB

[Tho11] Elihu Thomson. George Frederick Barker, M.D., Sc.D., LL.D. Proceedings of the American Philosophical Society, 50(202):xiii–xxix, October/December 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984005.

Thomson:1930:WIT

[Tho30] Elihu Thomson. Why the 200-inch telescope. Proceedings of the American Philosophical Society, 69(1):1–12, 1930. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984346.

Thorndike:1935:PS

[Tho35] Edward L. Thorndike. The paradox of science. Proceedings of the American Philosophical Society, 75(4):287–294, 1935. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984630.

Thorndike:1937:TT

[Tho37a] Edward L. Thorndike. Trains of thought. Proceedings of the American Philosophical Society, 77(3):417–438, March 1937. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984736.

Thorndike:1937:TTS

[Tho37b] Edward L. Thorndike. Trains of thought as symptoms of interests and attitudes: An exploratory investigation. Proceedings of the American Philosophical Society, 77(3):439–445, March 1937. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984737.

Thompson:1939:SEF

[Tho39] Warren S. Thompson. The social environment as a factor in population growth. Proceedings of the American Philosophical Society, 80(4):491–498, February 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985052.

Thorndike:1941:MA

[Tho41] Edward L. Thorndike. Mental abilities. Proceedings of the American Philosophical Society, 84(4):503–513, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984954.

Thomson:1946:RGU

[Tho46] Charles A. Thomson. The role of government in UNESCO. Proceedings of the American Philosophical Society, 90(4):300–303, September 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3301025.

Thompson:1949:FFR

[Tho49] Stith Thompson. The future of folklore research in the United States. Proceedings of the American Philosophical Society, 93(3):244–247, June 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3143473.

Tittmann:1908:PDA

[Tit08] O. H. Tittmann. Progress of the demarcation of the Alaska boundary. Proceedings of the American Philosophical Society, 47(188): 86–90, January/April 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983801.

Titchener:1914:PBV

[Tit14a] E. B. Titchener. On “psychology as the behaviorist views it”. Proceedings of the American Philosophical Society, 53(213):1–17, January/ May 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984126.

Titchener:1914:PAH

[Tit14b] E. B. Titchener. On the production of an artifical hiss. Proceedings of the American Philosophical Society, 53(215):323–329, August/ December 1914. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984081.

Titchener:1916:ETS

[Tit16] E. B. Titchener. On ethnological tests of sensation and perception with special reference to tests of color vision and tactile discrimination described in the reports of the Cambridge Anthropological Expedition to Torres Straits. Proceedings of the American Philosophical Society, 55(3):204–236, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984113.

Tolman:1941:MLA

[Tol41] Edward Chace Tolman. Motivation, learning and adjustment. Proceedings of the American Philosophical Society, 84(4):543–563, June 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984957.

Tower:1913:TOU

[Tow13] Charlemagne Tower. The treaty obligations of the United States relating to the Panama Canal. Proceedings of the American Philosophical Society, 52(209):234–242, April 1913. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983869.

Tower:1915:RDN

[Tow15] Charlemagne Tower. The rights and duties of neutralized territory. Proceedings of the American Philosophical Society, 54(216): 18–26, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983909.

Tower:1916:JAI

[Tow16] Charlemagne Tower. The judicial aspect of international law. — Arbitration. Proceedings of the American Philosophical Society, 55(4): 297–304, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984102.

Titchener:1907:EIS

[TP07] Edward Bradford Titchener and William Henry Pyle. The effect of imperceptible shadows on the judgment of distance. Proceedings of the American Philosophical Society, 46(185):94–109, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983443.

Titchener:1908:AIS

[TP08] Edward Bradford Titchener and William Henry Pyle. On the afterimages of subliminally colored stimuli. Proceedings of the American Philosophical Society, 47(189):366–384, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983848.

Travis:1906:PCL

[Tra06] Charles Travis. Pyrite from Cornwall, Lebanon County, Pennsylvania. Proceedings of the American Philosophical Society, 45(183): 131–148, May/September 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983823.

Trelease:1910:SA

[Tre10] William Trelease. Species in agave. Proceedings of the American Philosophical Society, 49(195):232–237, July 1910. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983901.

Trelease:1911:DGN

[Tre11] William Trelease. The desert group Nolineæ. Proceedings of the American Philosophical Society, 50(200):404–443, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983932.

Trelease:1912:CBO

[Tre12] William Trelease. The classification of the black oaks. Proceedings of the American Philosophical Society, 51(204):167–171, April/June 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984100.

Trelease:1915:LFA

[Tre15] William Trelease. The large fruited American oaks. Proceedings of the American Philosophical Society, 54(216):7–11, January/April 1915. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983906.

Trelease:1916:TNT

[Tre16] William Trelease. Two new terms Cormophytaster and Xeniophyte axiomatically fundamental in botany. Proceedings of the American Philosophical Society, 55(3):237–242, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984114.

Trelease:1917:NAH

[Tre17] William Trelease. Naming American hybrid oaks. Proceedings of the American Philosophical Society, 56(1):44–52, 1917. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983961.

Trelease:1921:NGU

[Tre21] William Trelease. A natural group of unusual black oaks. Proceedings of the American Philosophical Society, 60(1):31–33, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984224.

Trelease:1926:SPP

[Tre26] William Trelease. Some puzzling pepper leaves. Proceedings of the American Philosophical Society, 65(1):64–67, 1926. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984339.

Trelease:1928:TNG

[Tre28a] William Trelease. Trianaeopiper, a new genus of Piperaceae. Proceedings of the American Philosophical Society, 67(1):47–50, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984255.

Trelease:1928:TAS

[Tre28b] William Trelease. Two additional species of Trianaeopiper. Proceedings of the American Philosophical Society, 67(4):335–338, 1928.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984248.

Trelease:1929:LGS

[Tre29] William Trelease. Lindeniopiper, a generic segregate from piper. Proceedings of the American Philosophical Society, 68(1):53–54, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984398.

Trelease:1930:GAP

[Tre30] William Trelease. The geography of American peppers. Proceedings of the American Philosophical Society, 69(1):309–327, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984366.

Trelease:1933:UCR

[Tre33] William Trelease. An unusual chimæra. Proceedings of the American Philosophical Society, 72(4):241–244, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984688.

Trelease:1934:NAN

[Tre34] William Trelease. New amplifications of the North American Piperaceæ. Proceedings of the American Philosophical Society, 73(5):327– 330, May 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984623.

Trelease:1935:PPS

[Tre35] William Trelease. The pedicellate peppers of South America. Proceedings of the American Philosophical Society, 75(8):691–716, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984590.

Trelease:1937:MPY

[Tre37] William Trelease. A Mexican puzzle “(yucca Howard-Smithiï)”. Proceedings of the American Philosophical Society, 78(1):97–99, October 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984659.

Trowbridge:1902:SGH

[Tro02] John Trowbridge. Spectra of gases at high temperatures. Proceedings of the American Philosophical Society, 41(169):138–140, April 1902.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983529.

Trowbridge:1910:ED

[Tro10] Augustus Trowbridge. The ether drift. Proceedings of the American Philosophical Society, 49(194):52–56, January/June 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983991.

True:1908:CC

[Tru08] Frederick W. True. On the classification of the Cetacea. Proceedings of the American Philosophical Society, 47(189):385–391, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983849.

True:1928:SNB

[Tru28] Rodney H. True. Some neglected botanical results of the Lewis and Clark Expedition. Proceedings of the American Philosophical Society, 67(1):1–19, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984252.

True:1929:SLJ

[Tru29] Rodney H. True. A sketch of the life of John Bradbury, including his unpublished correspondence with Thomas Jefferson. Proceedings of the American Philosophical Society, 68(2):133–150, 1929. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984314.

True:1930:TMI

[Tru30] Rodney H. True. The toxicity of molecules and ions. Proceedings of the American Philosophical Society, 69(1):231–245, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984359.

True:1936:TJG

[Tru36] Rodney H. True. Thomas Jefferson’s garden book. Proceedings of the American Philosophical Society, 76(6):939–945, 1936. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984577.

True:1937:FAM

[Tru37] Rodney H. True. François André Michaux, the botanist and explorer. Proceedings of the American Philosophical Society, 78(2):313–327, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984540.

Tucker:1922:EDV

[Tuc22] R. H. Tucker. The effect of diurnal variation of clock rates upon longitude work. Proceedings of the American Philosophical Society, 61 (2):123–135, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984410.

Tucker:1933:SPV

[Tuc33] R. H. Tucker. Standard photometric and visual scales of magnitude. Proceedings of the American Philosophical Society, 72(5):351–354, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984698.

Tomanek:1924:PP

[TW24] Ewald Tomanek and Edwin B. Wilson. Pneumonia in Pittsburgh. Proceedings of the American Philosophical Society, 63(3):279–316, 1924. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984328.

Tyzzer:1932:POC

[Tyz32] Ernest Edward Tyzzer. Problems and observations concerning the transmission of blackhead infection in turkeys. Proceedings of the American Philosophical Society, 71(6):407–410, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984650.

Tyzzer:1942:CSG

[Tyz42] Ernest Edward Tyzzer. A comparative study of Grahamellae, Haemobartonellae and Eperythrozoa in small mammals. Proceedings of the American Philosophical Society, 85(4):359–398, June 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984991.

Urey:1946:MOS

[Ure46] Harold C. Urey. Methods and objectives in the separation of isotopes. Proceedings of the American Philosophical Society, 90(1):30–35, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/3301036.

Vaillant:1943:ATC

[Vai43] George C. Vaillant. The Aztecs their cultural and historical position in Middle American archaeology. Proceedings of the American Philosophical Society, 86(2):320–322, February 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985111.

Vaughan:1912:PP

[Vau12] Victor C. Vaughan. The protein poison. Proceedings of the American Philosophical Society, 51(205):206–213, July 1912. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984138.

vanDyke:1932:W

[vD32] Paul van Dyke. Washington. Proceedings of the American Philosophical Society, 71(4):191–205, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984790.

vanDoren:1943:BAP

[vD43] Carl van Doren. The beginnings of the American Philosophical Society. Proceedings of the American Philosophical Society, 87(3):277– 289, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984874.

vandeKamp:1944:PSP

[vdK44] Peter van de Kamp. A photographic study of the proper motion of nearby stars. Proceedings of the American Philosophical Society, 88(5):372–374, November 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985611.

Veblen:1912:JHP

[Veb12] Oswald Veblen. Jules Henri Poincaré. Proceedings of the American Philosophical Society, 51(207):iii–ix, October/December 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984024.

vonIhering:1902:MFP

[vI02] H. von Ihering. On the molluscan fauna of the Patagonian Tertiary. Proceedings of the American Philosophical Society, 41(169):132–137, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983528.

Viner:1946:IAB

[Vin46] Jacob Viner. The implications of the atomic bomb for international relations. Proceedings of the American Philosophical Society, 90(1):53–58, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301039.

Voegelin:1949:DAI

[Voe49] C. F. Voegelin. A decade of American Indian linguistics studies. Proceedings of the American Philosophical Society, 93(2):137–140, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143433.

Volwiler:1938:HBA

[Vol38] A. T. Volwiler. Harrison, Blaine, and American foreign policy, 1889–1893. Proceedings of the American Philosophical Society, 79 (4):637–648, November 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984943.

Webster:1919:NME

[WA19] A. G. Webster and Mildred Allen. On a new (?) method in exterior ballistics. Proceedings of the American Philosophical Society, 58(6): 382–385, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984262.

Wade:1945:GRM

[Wad45a] F. Alton Wade. The geology of the Rockefeller Mountains, King Edward VII Land, Antarctica. Proceedings of the American Philosophical Society, 89(1):67–77, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985315.

Wade:1945:ISS

[Wad45b] F. Alton Wade. An introduction to the Symposium on Scientific Results of the United States Antarctic Service Expedition, 1939–1941. Proceedings of the American Philosophical Society, 89(1):1–3, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985310.

Wade:1945:PAR

[Wad45c] F. Alton Wade. The physical aspects of the Ross Shelf ice. Proceedings of the American Philosophical Society, 89(1):160–173, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985322.

Walcott:1926:SPL

[Wal26] Charles D. Walcott. Samuel Pierpont Langley and modern aviation. Proceedings of the American Philosophical Society, 65(2):79–82, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984414.

Wallace:1948:HEM

[Wal48] Paul A. W. Wallace. Henry Ernest Muhlenberg. Proceedings of the American Philosophical Society, 92(2):107–110, May 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143408.

Wallace:1949:TSN

[Wal49a] Anthony F. C. Wallace. The Tuscaroras: Sixth Nation of the Iroquois Confederacy. Proceedings of the American Philosophical Society, 93(2):159–165, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143435.

Wallace:1949:MRU

[Wal49b] Paul A. W. Wallace. The Muhlenbergs and the revolutionary underground. Proceedings of the American Philosophical Society, 93 (2):119–126, May 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143430.

Wanless:1922:LWR

[Wan22] Harold R. Wanless. Lithology of the White River sediments. Proceedings of the American Philosophical Society, 61(3):184–203, 1922. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984227.

Wanless:1923:SWR

[Wan23] Harold R. Wanless. The stratigraphy of the White River Beds of South Dakota. Proceedings of the American Philosophical Society, 62 (4):190–269, 1923. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984216.

Ward:1925:IFW

[War25] Robert DeC. Ward. The importance of field-work in the study of climates. Proceedings of the American Philosophical Society, 64(1): 64–77, 1925. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984304.

Ward:1926:PSL

[War26] Robert DeC. Ward. The present status of long-range weather forecasting. Proceedings of the American Philosophical Society, 65(1): 1–14, 1926. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984334.

Ward:1928:PGB

[War28] Robert DeC. Ward. A proposed guide-book to the world’s weather and climates. Proceedings of the American Philosophical Society, 67 (1):67–94, 1928. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984257.

Ward:1931:RVW

[War31] Robert DeC. Ward. The railroads versus the weather. Proceedings of the American Philosophical Society, 70(2):137–166, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984319.

Warren:1943:MS

[War43] Robert B. Warren. Money and sovereignty. Proceedings of the American Philosophical Society, 87(2):147–151, August 1943. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/985091.

Warren:1944:CSS

[War44] Robert Warren. The capitalist and the social structure. Proceedings of the American Philosophical Society, 88(1):1–9, June 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985183.

Warner:1945:SPS

[War45] Lawrence A. Warner. Structure and petrography of the Southern Edsel Ford Ranges, Antarctica. Proceedings of the American Philosophical Society, 89(1):78–122, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985316.

Warren:1948:AP

[War48] Robert B. Warren. An attempt at perspective. Proceedings of the American Philosophical Society, 92(4):271–281, October 1948. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143356.

Waterman:1939:EAV

[Wat39a] Thomas Tileston Waterman. English antecedents of Virginia architecture. Proceedings of the American Philosophical Society, 80(1):57–63, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984973.

Watson:1939:ADM

[Wat39b] Fletcher Watson. Apparent distribution of meteoroid orbits. Proceedings of the American Philosophical Society, 81(4):481–491, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984820.

Watson:1939:SFR

[Wat39c] Fletcher Watson. A study of fireball radiant positions. Proceedings of the American Philosophical Society, 81(4):473–479, September 25, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984819.

Watson:1939:STM

[Wat39d] Fletcher Watson. A study of telescopic meteors. Proceedings of the American Philosophical Society, 81(4):493–504, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984821.

Wever:1937:CSE

[WB37] Ernest Glen Wever and Charles W. Bray. A comparative study of the electrical responses of the ear. Proceedings of the American Philosophical Society, 78(2):407–410, December 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984542.

White:1943:CNP

[WB43] Philip R. White and Armin C. Braun. A cancerous neoplasm of plants produced by autonomous, bacteria-free crown-gall tissue. Proceedings of the American Philosophical Society, 86(3):467–469, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985032.

Wilson:1939:MSF

[WBAW39] Edwin B. Wilson, Constance Bennett, Margaret Allen, and Jane Worcester. Measles and scarlet fever in Providence, R.I., 1929–1934 with respect to age and size of family. Proceedings of the American Philosophical Society, 80(3):357–476, February 10, 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984949.

Weaver:1944:RMP

[Wea44] John E. Weaver. Recovery of midwestern prairies from drought. Proceedings of the American Philosophical Society, 88(2):125–131, July 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985666.

Webster:1917:DA

[Web17] Arthur Gordon Webster. Dynamical aspects. Proceedings of the American Philosophical Society, 56(3):161–167, 1917. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983943.

Webster:1919:SCB

[Web19] Arthur Gordon Webster. Some considerations on the ballistics of a gun of seventy-five-mile range. Proceedings of the American Philosophical Society, 58(6):373–381, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984261.

Weeks:1912:HEA

[Wee12] David Fairchild Weeks. The heredity of epilepsy analyzed by the Mendelian method. Proceedings of the American Philosophical Society, 51(205):178–190, July 1912. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984135.

Weiss:1941:AVR

[Wei41] Paul Weiss. Autonomous versus reflexogenous activity of the central nervous system. Proceedings of the American Philosophical Society, 84(1):53–64, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984850.

Wenrich:1935:HPR

[Wen35] D. H. Wenrich. Host-parasite relations between parasitic protozoa and their hosts. Proceedings of the American Philosophical Society, 75 (7):605–650, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984811.

Wenrich:1937:SIB

[Wen37] D. H. Wenrich. Studies on Iodamœba Bütschlii (Protozoa) with special reference to nuclear structure. Proceedings of the American Philosophical Society, 77(2):183–205, February 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984726.

Wertenbaker:1943:AGA

[Wer43] Thomas Jefferson Wertenbaker. American Georgian architecture. Proceedings of the American Philosophical Society, 87(1):65–69, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/985001.

Westerfield:1934:SC

[Wes34] Ray B. Westerfield. Stabilization of the currency. Proceedings of the American Philosophical Society, 74(1):29–62, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984583.

Weslager:1944:WCI

[Wes44] C. A. Weslager. Wynicaco — a Choptank Indian chief. Proceedings of the American Philosophical Society, 87(5):398–402, May 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985292.

Westermann:1948:FSG

[Wes48] William Linn Westermann. The freedmen and the slaves of God. Proceedings of the American Philosophical Society, 92(1):55–64, March 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143628.

Wetherill:1934:SSF

[Wet34] Samuel Price Wetherill. Some subjective factors affecting industrial recovery. Proceedings of the American Philosophical Society, 74(1):129–135, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984588.

Wetmore:1947:PDP

[Wet47] Alexander Wetmore. Publication and the distribution of publications. Proceedings of the American Philosophical Society, 91(1):117–120, February 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143132.

Weyl:1949:RTS

[Wey49] Hermann Weyl. Relativity theory as a stimulus in mathematical research. Proceedings of the American Philosophical Society, 93(7):535–541, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143143.

Wharton:1904:PP

[Wha04] Joseph Wharton. Palladium (Pd). Proceedings of the American Philosophical Society, 43(177):332–338, April/September 1904. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983769.

Wheeler:1918:SSA

[Whe18] William Morton Wheeler. A study of some ant larvæ, with a consideration of the origin and meaning of the social habit among insects. Proceedings of the American Philosophical Society, 57(4):293–343, 1918. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983940.

Wheeler:1919:PAS

[Whe19] William Morton Wheeler. The parasitic Aculeata, a study in evolution. Proceedings of the American Philosophical Society, 58(1):1–40, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984478.

Wheeler:1931:HBS

[Whe31] William Morton Wheeler. Hopes in the biological sciences. Proceedings of the American Philosophical Society, 70(3):231–239, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984452.

Wheeler:1946:PPE

[Whe46] John Archibald Wheeler. Problems and prospects in elementary particle research. Proceedings of the American Philosophical Society, 90(1):36–47, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301037.

Whitehead:1911:HVC

[Whi11] J. B. Whitehead. The high voltage corona in air. Proceedings of the American Philosophical Society, 50(200):374–396, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983930.

White:1917:SIP

[Whi17] Orland E. White. Studies of inheritance in Pisum. II. The present state of knowledge of heredity and variation in peas. Proceedings of the American Philosophical Society, 56(7):487–588, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983886.

Whitehead:1920:HVC

[Whi20] J. B. Whitehead. The high voltage corona in air. Proceedings of the American Philosophical Society, 59(4):245–260, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984424.

Whitney:1931:TMP

[Whi31] Willis R. Whitney. Technology and material progress. Proceedings of the American Philosophical Society, 70(3):255–262, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984454.

Whiting:1935:GBS

[Whi35] P. W. Whiting. Genic balance, sex-determination and selective fertilization in Hymenoptera. Proceedings of the American Philosophical Society, 75(6):517–520, 1935. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984612.

Whipple:1938:PMS

[Whi38] Fred L. Whipple. Photographic meteor studies, I. Proceedings of the American Philosophical Society, 79(4):499–548, November 1938.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984939.

Whipple:1939:PCO

[Whi39a] Fred L. Whipple. On the physical characteristics and origin of the supernovæ. Proceedings of the American Philosophical Society, 81(2):253–264, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984984.

Whiting:1939:MBC

[Whi39b] Anna R. Whiting. Mutant body colors in the parasitic wasp “Habrobracon juglandis” (Ashm.) and their behavior in multiple recessives and in mosaics. Proceedings of the American Philosophical Society, 80(1):65–85, January 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984974.

Whipple:1940:PMSa

[Whi40a] Fred L. Whipple. Photographic meteor studies, II. Non-linear trails. Proceedings of the American Philosophical Society, 82(3):275–290, April 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984930.

Whipple:1940:PMSb

[Whi40b] Fred L. Whipple. Photographic meteor studies. III. The Taurid shower. Proceedings of the American Philosophical Society, 83(5):711–745, October 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985135.

Whittlesey:1940:RTI

[Whi40c] C. R. Whittlesey. The relation of totalitarianism to international trade and finance. Proceedings of the American Philosophical Society, 82(1):73–76, February 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984905.

Whipple:1944:RAT

[Whi44] Allen O. Whipple. Recent advances in the treatment of wounds. Proceedings of the American Philosophical Society, 88(3):177–181, September 1944. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/985411.

Whipple:1947:PMS

[Whi47] Fred L. Whipple. Photographic meteor studies. IV. The geminid shower. Proceedings of the American Philosophical Society, 91(2):189– 200, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143604.

Wiener:1945:RAO

[Wie45] Murray A. Wiener. Results of auroral observations at West Base, Antarctica, April to September, 1940. Proceedings of the American Philosophical Society, 89(1):364–378, April 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/985345.

Wigner:1946:RR

[Wig46] Eugene P. Wigner. Resonance reactions. Proceedings of the American Philosophical Society, 90(1):25–29, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3301035.

Wigner:1949:IPT

[Wig49] Eugene P. Wigner. Invariance in physical theory. Proceedings of the American Philosophical Society, 93(7):521–526, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143140.

Wilson:1902:BAD

[Wil02] J. C. Wilson. A brief account of the disease known as Osteitis Deformans. Proceedings of the American Philosophical Society, 41 (169):143–145, April 1902. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983531.

Wiley:1905:EPM

[Wil05] H. W. Wiley. The effect of preservatives on metabolism. Proceedings of the American Philosophical Society, 44(179):66–68, January/ April 1905. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983606.

Wilder:1908:BR

[Wil08a] Burt G. Wilder. The brain of Rhinochimaera. Proceedings of the American Philosophical Society, 47(188):37–38, January/April 1908.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983795.

Wiley:1908:IPO

[Wil08b] Harvey W. Wiley. Influence of preservatives and other substances added to foods upon health and metabolism. Proceedings of the American Philosophical Society, 47(189):302–328, May/August 1908. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983846.

Wilder:1910:BAO

[Wil10] B. G. Wilder. A brain of about one-half the average weight from an intelligent white man. Proceedings of the American Philosophical Society, 49(195):188–190, July 1910. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983893.

Wilson:1911:CA

[Wil11] Harold A. Wilson. The constitution of the atom. Proceedings of the American Philosophical Society, 50(200):366–373, July/August 1911. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983929.

Wilson:1916:LA

[Wil16] George Grafton Wilson. Legislative aspects. Proceedings of the American Philosophical Society, 55(4):305–312, 1916. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984103.

Wilson:1917:TAE

[Wil17] Edwin Bidwell Wilson. Theory of an aëroplane encountering gusts, II. Proceedings of the American Philosophical Society, 56(3):212– 248, 1917. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983947.

Williams:1920:DKP

[Wil20] Edward H. Williams, Jr. The deep Kansan pondings in Pennsylvania and the deposits therein. Part One. Proceedings of the American Philosophical Society, 59(1):48–84, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/984219.

Wilson:1932:CWT

[Wil32] G. Lloyd Wilson. Coördination as a “way-out” of the transportation crisis. Proceedings of the American Philosophical Society, 71(1):31– 38, 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984818.

Wildt:1939:CP

[Wil39a] Rupert Wildt. The constitution of the planets. Proceedings of the American Philosophical Society, 81(2):135–152, June 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984980.

Williams:1939:BOM

[Wil39b] J. D. Williams. Binocular observations of 718 meteors. Proceedings of the American Philosophical Society, 81(4):505–520, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984822.

Williams:1939:SBM

[Wil39c] J. D. Williams. Systematic biases of meteor directions. Proceedings of the American Philosophical Society, 81(4):521–524, September 1939. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984823.

Willis:1942:AG

[Wil42a] Bailey Willis. American geology, 1850–1900. Proceedings of the American Philosophical Society, 86(1):34–44, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985076.

Wilson:1942:SSA

[Wil42b] M. L. Wilson. Survey of scientific agriculture. Proceedings of the American Philosophical Society, 86(1):52–62, September 1942. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985078.

Williams:1943:PPW

[Wil43a] John H. Williams. Problems of post-war international monetary stabilization. Proceedings of the American Philosophical Society, 87(2):133–138, August 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985089.

Wilson:1943:TJF

[Wil43b] M. L. Wilson. Thomas Jefferson — farmer. Proceedings of the American Philosophical Society, 87(3):216–222, July 1943. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984868.

Willits:1946:SAA

[Wil46] Joseph H. Willits. Social adjustments to atomic energy. Proceedings of the American Philosophical Society, 90(1):48–52, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3301038.

Winlock:1932:ETD

[Win32] H. E. Winlock. Excavations at the Temple of Deir el Bahri, 1921– 1931. Proceedings of the American Philosophical Society, 71(6):321– 341, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/984644.

Winlock:1940:OAE

[Win40] H. E. Winlock. The origin of the Ancient Egyptian calendar. Proceedings of the American Philosophical Society, 83(3):447–464, September 10, 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/985113.

Winslow:1945:WNE

[Win45] Carlile P. Winslow. Wood in the national economy. Proceedings of the American Philosophical Society, 89(2):403–407, July 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/985635.

Wirth:1948:RRC

[Wir48] Louis Wirth. Research in racial and cultural relations. Proceedings of the American Philosophical Society, 92(5):381–386, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143050.

Wissler:1927:ACA

[Wis27] Clark Wissler. Age changes in anthropological characters in childhood and adult life. Proceedings of the American Philosophical Society, 66(??):431–438, ???? 1927. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3301077.

Wissler:1942:AIA

[Wis42] Clark Wissler. The American Indian and the American Philosophical Society. Proceedings of the American Philosophical Society, 86(1):189–204, September 1942. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/985086.

Wolf:1932:EAN

[Wol32] S. K. Wolf. Engineering aspects of noise studies. Proceedings of the American Philosophical Society, 71(5):275–283, April 1932. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984702.

Woodbridge:1930:CCA

[Woo30] F. J. E. Woodbridge. A comment on college administration. Proceedings of the American Philosophical Society, 69(1):281–294, 1930. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984363.

Woods:1931:USS

[Woo31] Arthur Woods. Unemployment and its social significance. Proceedings of the American Philosophical Society, 70(3):285–296, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984457.

Woody:1940:PTE

[Woo40] Thomas Woody. Principles of totalitarian education. Proceedings of the American Philosophical Society, 82(1):39–55, February 1940. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984903.

Whipple:1941:TSS

[WPG41] Fred L. Whipple and Cecilia Payne-Gaposchkin. Theoretical synthesis of supernova spectra. Proceedings of the American Philosophical Society, 84(1):1–32, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984848.

Wiener:1947:PLC

[WPL47] Philip P. Wiener, Charles Sanders Peirce, and Samuel P. Langley. The Peirce–Langley correspondence and Peirce’s manuscript on Hume and the laws of nature (at the Smithsonian Institution). Proceedings of the American Philosophical Society, 91(2):201–228, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143605.

Wright:1903:CP

[Wri03] Fred Eugene Wright. Crystallographic properties. Proceedings of the American Philosophical Society, 42(173):237–249, April/May 1903. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983736.

Wright:1919:PLS

[Wri19] Fred. E. Wright. Polarized light in the study of ores and metals. Proceedings of the American Philosophical Society, 58(7):401–447, 1919. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984213.

Wright:1920:CAR

[Wri20] W. H. Wright. Certain aspects of recent spectroscopic observations of the gaseous nebulæ which appear to establish a relationship between them and the stars. Proceedings of the American Philosophical Society, 59(6):517–531, 1920. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/984391.

Wright:1921:CFR

[Wri21] Quincy Wright. The control of the foreign relations of the United States: The relative rights, duties, and responsibilities of the President, of the Senate and the House, and of the Judiciary, in theory and in practice. Proceedings of the American Philosophical Society, 60(3): 99–455, 1921. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984422.

Wright:1943:TJC

[Wri43] Louis B. Wright. Thomas Jefferson and the classics. Proceedings of the American Philosophical Society, 87(3):223–233, July 1943. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984869.

Wright:1949:PSE

[Wri49] Sewall Wright. Population structure in evolution. Proceedings of the American Philosophical Society, 93(6):471–478, December 1949. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143336.

Wherry:1906:URA

[WS06] Edgar T. Wherry and Edgar F. Smith. The use of a rotating anode in the electrolytic precipitation of uranium and molybdenum. Proceedings of the American Philosophical Society, 45(184): 268–270, October/December 1906. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983379.

Wiley:1907:PSA

[WS07] H. W. Wiley and Herman Schreiber. Production of synthetic alcohol. Proceedings of the American Philosophical Society, 46(185): 117–123, January/April 1907. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983445.

Wyckoff:1937:USV

[Wyc37] Ralph W. G. Wyckoff. The ultracentrifugal study of virus proteins. Proceedings of the American Philosophical Society, 77(4):455–462, April 1937. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/984771.

Yerkes:1941:PD

[Yer41] Robert M. Yerkes. Psychology and defense. Proceedings of the American Philosophical Society, 84(4):527–542, June 1941. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/984956.

Young:1933:GOG

[You33] James T. Young. The growing overhead in government. Proceedings of the American Philosophical Society, 72(2):57–66, April 1933. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984742.

Young:1934:LPR

[You34] James T. Young. Liberty, property and recovery. Proceedings of the American Philosophical Society, 74(1):93–101, July 1934. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/984586.

Young:1947:TRR

[You47] Donald Young. Techniques of race relations. Proceedings of the American Philosophical Society, 91(2):150–161, April 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143601.

Young:1948:LFD

[You48] Donald Young. Limiting factors in the development of the social sciences. Proceedings of the American Philosophical Society, 92(5):325–335, November 1948. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3143042.

Zeleny:1931:CWG

[Zel31] John Zeleny. The changes which gaseous ions undergo with time. Proceedings of the American Philosophical Society, 70(2):121–135, 1931. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/984318.

Zworykin:1947:ERA

[ZF47] V. K. Zworykin and L. E. Flory. An electronic reading aid for the blind. Proceedings of the American Philosophical Society, 91(2):139– 142, April 1947. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143599.

Zirkle:1941:NSB

[Zir41] Conway Zirkle. Natural selection before the “origin of species”. Proceedings of the American Philosophical Society, 84(1):71–123, April 1941. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/984852.

Zon:1945:FRS

[Zon45] Raphael Zon. Forests in relation to soil and water. Proceedings of the American Philosophical Society, 89(2):399–402, July 18, 1945. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/985634.